### v1

# Personality Structured Interview for Large Language Model Simulation in Personality Research

Abstract
Although psychometrics researchers have recently explored the use of large language models (LLMs) as proxies for human participants, LLMs often fail to generate heterogeneous data with human-like diversity, which diminishes their value in advancing social science research. To address these challenges, we explored the potential of the theory-informed Personality Structured Interview (PSI) as a tool for simulating human responses in personality research. In this approach, the simulation is grounded in nuanced real-human interview transcripts that target the personality construct of interest. We have provided a growing set of 357 structured interview transcripts from a representative sample [^1] Dataset and code will be shared later., each containing an individual’s response to 32 open-ended questions carefully designed to gather theory-based personality evidence. Additionally, grounded in psychometric research, we have summarized an evaluation framework to systematically validate LLM-generated psychometric data. Results from three experiments demonstrate that well-designed structured interviews could improve human-like heterogeneity in LLM-simulated personality data and predict personality-related behavioral outcomes (i.e., organizational citizenship behaviors and counterproductive work behavior). We further discuss the role of theory-informed structured interviews in LLM-based simulation and outline a general framework for designing structured interviews to simulate human-like data for psychometric research.
Personality Structured Interview for Large Language Model Simulation in Personality Research
Pengda Wang1, Huiqi Zou3, Hanjie Chen2, Tianjun Sun1, Ziang Xiao3, and Frederick L. Oswald1 1Department of Psychological Sciences, Rice University 2Department of Computer Science, Rice University 3Department of Computer Science, Johns Hopkins University {pw32,hc86,ts110,foswald}@rice.edu; {hzou11,ziang.xiao}@jhu.edu
**footnotetext: Co-corresponding authors.
1 Introduction
Personality research is an important field in psychology for understanding how individual traits shape significant life outcomes and trajectories, including career success, mental health, and overall well-being (e.g., Judge et al., 2002; Roberts et al., 2007; Robins et al., 2002). For instance, measures of conscientiousness have been consistently and positively correlated with job performance and academic success (Barrick and Mount, 1991), whereas measures of neuroticism are found to be strongly and positively correlated with mental health issues, such as anxiety and depression (John et al., 2008).
However, studying personality requires understanding habits of thoughts, attitudes, and behaviors across diverse contexts and cultures, which is often a research challenge. For example, some research questions involve rare or extreme cases, making it difficult to gather sufficient samples (e.g., Lynam and Widiger, 2001). Others require longitudinal designs to track personality development or dynamic social interactions (e.g., Damian et al., 2019; Roberts et al., 2006). Additionally, certain experimental manipulations, like inducing stress or altering life circumstances, may be feasible or unethical for causal insights Fisher (2021). If a method could accurately simulate the human personality distributions, they could supplement and accelerate personality research by offering a scalable, cost-effective approach to supplement traditional data collection and experimentation (Messeri and Crockett, 2024).
Several prior studies have supported this possibility, showing that LLMs can generate responses for personality scales that reflect personality traits resembling those of humans (e.g., Lee et al. 2024, Huang et al. 2023). However, studies found such methods face significant limitations in capturing individual differences at the item level (e.g., Wang et al. 2024a). Information at the item level provides a more granular understanding of how traits manifest in specific individuals and situations rather than relying solely on broad trait averages. However, without additional empirical information, items confound these sources of variance. Moreover, key psychometric challenges remain in modeling LLM responses, as they may not accurately reflect the intended personality construct.
To address this gap, we explored the potential of the Personality Structured Interview (PSI), which employs theory-informed questions to elicit information embedded in human narrative responses that are relevant to the target personality constructs of interest in LLM simulations. We believe this sophisticated linguistic approach can inform and enhance the measurement of human-like traits in heterogeneous LLM-simulated personality data, potentially leading to a greater understanding of personality.
Grounded in psychometric research (e.g., Cronbach, 1951; Cronbach and Meehl, 1955), we have summarized a framework for evaluating psychometric data that fully considers its hierarchical structure, in which observed responses (item level) are mapped to latent traits (domain level). This evaluation framework encompasses a range of analyses, from overall descriptive statistics to a more in-depth examination of psychometric performance.
Based on our evaluation framework, we conducted three experiments. Those three experiments evaluated our method’s effectiveness in replicating individual personality traits, simulating human-like personality distribution, and demonstrating personality-related behaviors with respect to established personality theory. We found our PSI method improves human-like diversity in personality traits simulation and recovers nuances in personality-related behaviors.
Figure 1: Overview of the Development of the Personality Structured Interview and Experimental Implementation in This Study
Note: For an overview of the personality structured interview development process, refer to §3 and Appendix B. The structured interview development framework is detailed in Appendix C. Information on data collection and the human sample is provided in §4 and Appendix D. The LLM simulation process is described in §4 and Appendix F. Details on the experimental setup and results of the three experiments can be found in §4, §5, and Appendix I. For the psychometric data evaluation framework, see Appendix H.
In summary, our paper offers three contributions:
  • A theory-informed LLM-based simulation method (PSI) for personality research and a development framework (see Appendix C).
  • A growing dataset, now with 357 structured interview transcripts.
  • An evaluation framework for LLM-simulated psychometric data that is grounded in psychological theories (see Appendix H).
We further discuss the potential of PSI methods in advancing research (see Appendix C). We believe the framework behind PSI could be generalized to other psychological constructs and other psychometric data to develop theory-informed simulation methods for human understanding.
2 Personality Research and Related Works
Here we provide a concise overview of key personality concepts and discuss related works.
MacKinnon (1944) proposed two definitions of personality. One emphasizes internal factors like temperament and interpersonal strategies that drive consistent behavior across time, situations, and cultures. The other focuses on interpersonal characteristics as perceived by others, linking personality to reputation. The former highlights internal drives, while the latter centers on external behaviors and social perception. Together, they underscore personality’s role in shaping thought patterns and behaviors in social interactions (Hogan et al., 1996).
Personality encodes rich and complex information in language and text (Goldberg, 1990; Saucier and Goldberg, 2001). In fact, the Five-Factor Model (FFM) of personality is extensively researched (e.g., Costa and McCrae, 2008; John, 1999; McCrae and Costa Jr, 1997); it is directly based on the lexical hypothesis, which posits that individual differences that are important in human interactions (e.g., have survival value across cultures) are often encoded in some or all languages of the world. The five factors are openness, conscientiousness, extraversion, agreeableness, and neuroticism (OCEAN). From Galton’s (1884) early research to Allport and Odbert’s (1936) systematic organization and factor analysis of lexical terms, and through further developments by Norman (1963) and Goldberg (1990), the FFM theory has progressively evolved (see Appendix A for further details on the structure of personality). In other words, because language and text contain rich and complex information about personality, LLMs may capture and model such encoding by learning from vast amounts of training data.
There are many studies to-date on LLMs that focus on personality, where researchers aim to benchmark their psychological profiles (Lee et al., 2024; Li et al., 2024; Pellert et al., 2023), investigate their ability to simulate personality (Huang et al., 2023; Jiang et al., 2023; Serapio-García et al., 2023), and examine whether LLM agents align with character personality settings and how these settings influence their generated dialogues (Shao et al., 2023; Wang et al., 2024b; Xu et al., 2024).
Currently, common methods for LLM simulation of human personality distributions have several limitations. The Persona-Chat dataset (Persona method) by Zhang et al. (2018) was originally designed to enhance the engagement of chit-chat models by increasing personalization. Therefore, its focus is more on enhancing personalization, rather than capturing personality traits. The adjective-based categorization method (Shape method) proposed by Serapio-García et al. (2023) effectively simulates profile-specific patterns of high/low standing on personality dimensions. However, this method presents a challenge in simulating human personality distributions, as restricting each case to a single dimension fails to capture the inherently multidimensional and hierarchical nature of personality data (e.g., Kachur et al., 2020).
To address the aforementioned limitations of personality-simulated data, we explore the use of theory-informed PSI transcription to simulate personality data.
3 Personality Structured Interview
The theory-informed structured interview mentioned in this study differs from previous interview approaches used in LLM-based simulations, such as the Life Interview  (Park et al., 2024), which broadly inquires about various aspects of an individual’s life. In contrast, the theory-informed structured interview is specifically designed to target the personality constructs that we aim to model.
A theory-informed approach is essential not only for improving the accuracy of data simulation but also for enhancing interpretability. Our goal is to generate data that not only aligns with actual data at the personality scale level but also exhibits theoretically coherent personality-related behaviors. By grounding LLM-generated data in theory, we ensure that the simulation goes beyond reproducing statistical properties to capturing meaningful psychological patterns. This enhances interpretability, facilitates transparent comparisons with human responses, and preserves the psychological validity of the constructs.
Learning from psychometric theories, we summarized a framework for developing structured interviews, which can serve as a template for creating interviews designed to collect targeted information for simulating specific types of psychometric data (see Appendix C). We present a PSI example designed to capture personality-related information, with relevant personality theories and question design detailed in §3.1.
We want to further emphasize that the “structured interview questions” mentioned here differ from those described by Wang et al. (2024b); Ran et al. (2024), who refers to converting personality scale items into open-ended questions. The main issue with their approach is that, while it attempts to rewrite items as open-ended questions to elicit richer responses, it may fail to achieve the intended effect. For instance, in their example, “Values artistic, aesthetic experiences” is rewritten as “Do you value artistic, aesthetic experiences?” Although this question appears open-ended, it tends to prompt a simple “yes” or “no” response and function as if it is close-ended, limiting the exploration of the respondent’s deeper thoughts or reasoning (Trull et al., 1998).
Moreover, such questions differ from the behavioral or situational questions typically found in structured interviews. Structured interviews usually involve more specific scenarios or tasks that encourage respondents to describe past experiences or hypothetical reactions (Campion et al., 1997). For example, to explore a respondent’s views on artistic and aesthetic experiences, a more effective question might be, “Please describe a moment when you felt deeply inspired by an artistic or aesthetic experience?” This encourages detailed, contextual responses beyond a simple “yes” or “no.”
3.1 PSI Questions Design
How can we design questions more effectively and obtain textual information with greater depth in measuring personality? McAdams’ three levels of personality theory and the corresponding, theory-informed interviews offer a valuable direction (e.g., McAdams, 1995, 1996, 2001).
According to McAdams’s theory, personality can be divided into three levels: traits (broad personality characteristics), personal concerns (goals, values, strategies), and narrative identity (the story individuals tell themselves about their lives). The latter two levels, in particular, reveal how individuals act in specific situations and developmental stages and how they organize and make sense of their life experiences.
The key to obtaining more meaningful textual information lies in guiding individuals toward higher levels of self-expression. Rather than simply recording mundane daily events, we should encourage reflection on pivotal life moments, significant relationships, and future aspirations. For example, effective questions could be: “Can you describe an event that changed the trajectory of your life?” or “Tell me about a moment you are most proud of.” By designing questions like these, we can elicit deeper, more meaningful narratives, enriching the personality information.
We adapted and modified McAdams (1995, 1996, 2001) theory-based interview and narrative identity approach, incorporating elements from the structured interview of the FFM (SIFFM; Trull et al., 1998). The initial draft of the question pool was prepared and revised by subject matter experts (SMEs) through discussion. The SMEs included two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology. Subsequently, we conducted pilot testing with six undergraduate research assistants from a personality research lab.
As a result, the PSI, comprising 32 questions, was developed, as shown in Table 3 in Appendix B.
3.2 PSI Dataset
Using this set of questions, we designed a chatbot to conduct structured interviews, facilitating data collection (Institutional Review Board approval was obtained). Because data collection is ongoing, we used responses collected through the end of December 2024. After excluding incomplete responses and those failing attention checks, the final sample consisted of 357 participants. On average, participants were in their early 30s (M = 33.30 years, SD = 13.06), with 52.40% identifying as men, and 44.80% as women.
To ensure a demographically diverse and broadly representative sample, we recruited participants from both undergraduate and working adult populations, ensuring a diverse range of backgrounds in terms of educational, employment, and life experience. Undergraduate participants were recruited from a large public university in the Midwest and received research credit for their participation. Working adult participants were sourced from two widely recognized, high-quality crowdsourcing platforms—Prolific [^2] https://www.prolific.com/ and CloudResearch Connect [^3] https://connect.cloudresearch.com/—which are known for their diverse and demographically representative participant pools. All working adult participants were compensated for their time. The average interview duration was 34 minutes.
In addition to completing the structured interview, participants provided demographic information and responded to both the personality scale and the personality-related behaviors scale (see Appendix D for detailed descriptions of data collection and examples of the dataset). Responses to PSI were subsequently incorporated into the prompt to generate simulated data (see Appendix F for the specific prompts used in the simulation process).
4 Experimental Setup
We designed three primary experiments to verify the effectiveness of the PSI method and its advantages over other methods: (1) Response Similarity (§4.2): evaluating the degree of similarity between responses generated by LLMs based on the PSI method and actual corresponding human responses on personality scales; (2) Human Personality Distribution Simulation (§4.3): comparing the PSI method with other methods in simulating human personality distributions; (3) Personality-Related Behavioral Performance (§4.4): testing the models’ performance using the PSI method to simulate personality-related behaviors. General settings shared by all experiments are described in §4.1, while specific differences are detailed in their respective sections.
4.1 General Settings
Psychological Scale
We used Big Five Inventory-2 (BFI-2; Soto and John, 2017): The BFI-2 is designed to capture three core facets of each of the FFM personality factors. Each facet is measured by two positively worded items and two negatively worded items, resulting in 60 items in total. Human respondents and LLMs were instructed to indicate the degree to which they agree with each item on a 5-point scale (1 = “Strongly disagree”, 2 = “Somewhat disagree”, 3 = “Neither agree nor disagree”, 4 = “Somewhat agree”, 5 = “Strongly agree”). Specific scale items and scoring criteria can be found in Table 5, Table 6 and Table 7 in Appendix E .
LLMs
We tested two LLMs, encompassing both closed-source and open-source models: GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024), and Llama 3 (llama-3-70b-instruct) (AI@Meta, 2024). We include these models to allow for cross-validation and comparative analysis, ensuring that the findings are robust and not specific to a single model.
GPT-4o and Llama 3 are among the current leading models, making them ideal for assessing the latest advancements in LLM performance. To ensure reproducibility, the temperature of the LLMs was set to zero to generate deterministic responses.
4.2 Response Similarity
The main purpose of this experiment is to assess the similarity between responses generated by LLMs based on the PSI method and human self-reports.
Metrics
Because the data generated by LLMs correspond on a one-to-one basis with data from human respondents, we used the Mean Absolute Error (MAE) and Pearson correlation coefficient ($r$) to examine the strength of the similarity between them. Smaller MAE and higher $r$ indicate greater similarity. For the $r$ calculation formula, see Appendix G.  Park et al. (2024) employed a Life Interview simulation, which is currently the most information-rich method for personality simulation. It is also one of the few papers that provides a direct one-to-one performance comparison with human samples. The reported MAE and $r$ with corresponding human samples are likely the best values available so far. We used this as a standard for comparison.
If data from human respondents are regarded as the gold-standard criterion, this test can also be considered an examination of criterion-related validity, which assesses the correlation between a measure and an external standard (Cronbach and Meehl, 1955).
4.3 Human Personality Distribution Simulation
The main purpose of this experiment is to investigate further the differences among various methods of simulating human personality distributions. We compared the PSI method, the Persona method—which relies on dialogue information (Zhang et al., 2018)—and the Shape method, which uses adjective-based dimensional categorization (Serapio-García et al., 2023), to observe differences when simulating a normal human sample. The specific descriptions of the Persona method and Shape method are shown in Appendix G.
Human Sample Criterion
The human samples used for this experiment were collected as part of a broader project related to personality assessment through Prolific and received Institutional Review Board approval. Respondents were instructed to complete a set of demographic questions, the BFI-2, and a set of criterion measures. The respondents were compensated with \$3.75 for their participation. In total, 1,559 respondents provided valid responses. On average, participants were in their early 40s (M = 42.29 years, SD = 11.79), 50.80% identifying as men, 49.20% identifying as women.
Metrics
To quantify the degree of similarity between human and LLM responses, we leveraged multiple metrics. At the domain and facet levels, we compared the mean (M; sample mean of each domain and facet), standard deviation (SD; sample standard deviation of each domain and facet), Cronbach’s alpha, and correlations among scores. At the item level, we also compared the mean and standard deviation of item scores (sample mean and standard deviation of each item).
Specifically, we also used MAE and $r$ to quantify similarities in these metrics. Here, however, MAE and $r$ are based on sample-level analysis rather than the one-to-one correspondence seen in the previous experiment. Our focus is on whether, when simulating an overall human sample, similarities can be observed across various measurement levels—including items, facets, and domains.
Additionally, separately for human and LLM responses, we fitted a three-factor confirmatory factor analysis (CFA; Jöreskog, 1969) model (TFM; where facets are “factors”) to the responses to each domain of the BFI-2. Aside from the facet structure of each domain, we also used the facet scores as indicators and fitted the FFM. Model fit, standardized factor loadings, and latent correlations among the facets were also compared between human and LLM responses. Tucker’s congruence coefficient (TCC; Tucker, 1951) and MAE of factor loadings were used to quantify the similarity between factor solutions from human and LLM responses. For the CFA model, model fit information, and the TCC calculation formula, see Appendix G. This evaluation framework is adaptable and can be used to analyze the fidelity of other simulated psychometric data (see Appendix H for further details).
4.4 Personality-Related Behavioral Performance
The primary aim of this experiment is to explore whether LLMs, when assigned specific personality settings, exhibit behaviors theoretically aligned with those personalities. This is an exploratory study, as LLMs generate responses based on statistical probabilities derived from the training corpus (Yang et al., 2024), making it uncertain whether assigning personality settings (through the PSI data) will influence the model to act consistently with that personality. Fortunately, the data we have collected through the PSI method includes LLM ratings of human behaviors, providing a valuable basis to test this hypothesis.
Personality-Related Behavior
We collected self-report data on two classic types of workplace behaviors: organizational citizenship behavior (OCB) and counterproductive work behavior (CWB). Both types of behaviors are widely supported by research as being related to personality (e.g., Organ and Ryan, 1995; Berry et al., 2007).
These data were collected using the scale developed by Spector and Fox (2010) and prompted the LLM to respond to the same questions (for a detailed description of Spector and Fox’s (2010) measures, see Appendix G; for specific prompt details, refer to Appendix F).
Metrics
We primarily focus on the $r$ between personality domains from different data sources and OCB, as well as CWB. We anticipate that the $r$ between self-reported personality domains and OCB/CWB in LLM simulation will closely align with those observed in the human participants.
5 Experimental Results
Here, we present the results of the three experiments mentioned above: Response Similarity (§5.1), Human Personality Distribution Simulation (§5.2), and Personality-Related Behavioral Performance (§ 5.3).
5.1 Response Similarity Results
Table 1 provides a clear illustration of the similarity between the personality generated using the PSI method and the corresponding human self-reported personality data. The results show that regardless of the model used, the average correlation is around .5, suggesting a moderately strong positive relationship and a significant association between the two variables. In particular, compared to the Life Interview method (Park et al., 2024), the PSI method consistently outperforms or matches the Life Interview approach across nearly all metrics, as indicated by its similar or lower MAE and similar or higher $r$.
Notably, the Life Interview method relies on up to two hours of interview data, while the PSI method achieves comparable results using interview data consisting of just 32 questions, with an average duration of about 34 minutes. This further underscores the advantage of the PSI method in effectively leveraging LLMs to simulate personality profiles.
Domain MAE $r$
PSI GPT-4o PSI Llama3 Life Interview PSI GPT-4o PSI Llama3 Life Interview
Extraversion 0.58 0.75 0.72 .64 .43 .45
Agreeableness 0.53 0.56 0.60 .41 .36 .35
Conscientiousness 0.59 0.66 0.63 .46 .40 .52
Neuroticism 0.63 0.59 0.75 .63 .57 .68
Openness 0.80 1.27 0.62 .43 .39 .39
Table 1: MAE and $r$ of Human Response and LLM Response Across PSI GPT-4o, PSI Llama3, and Life Interview (see Park et al., 2024, page 35 Table 3) for BFI-2 Each Personality Domain
Note: n = 357 for PSI method; n = 1,052 for Life Interview method. The bold text highlights the aspects where the PSI method performs better than the Life Interview method.
5.2 Human Personality Distribution Simulation Results
Descriptive Statistics
We compared the means and standard deviations of the different methods at three levels: personality item, facet, and domain. MAE and $r$ for both the means and standard deviations are shown in Table 2. The MAE for the means reflected the average difference between the LLM responses and the human responses at each level, whereas the MAE for the standard deviations revealed the difference in variability between the two datasets. The $r$ further illustrated the linear relationship between the two datasets, with values closer to one indicating a stronger correlation. Detailed means and standard deviations for human responses and LLM responses at the item, facet, and domain levels are provided in Tables 1213141516, and 17 in Appendix I.1.
Table 2 shows that the PSI method demonstrates better performance in simulating human samples compared to the Persona and Shape methods at the item level. Although at the facet and domain level, some results from other methods slightly outperform PSI (e.g., Shape GPT-4o in M), the process of aggregating scores from item to facet to domain levels reduces the impact of extreme values, smoothing them out at higher levels. Therefore, the PSI method’s advantage at the more granular item level becomes more important.
Additionally, PSI consistently demonstrates statistically significant positive correlations with the SD of human samples across items, facets, and domains, whereas the other methods exhibit either negative or non-significant correlations. Simulating the SD is a challenge for the other methods. Although the Shape method increases response variability and introduces more variance, it still struggles to replicate the true variance observed in human samples. The PSI method proposed in this paper addresses this issue to a certain extent.
Model Item Level Facet Level Domain Level
MAE $r$ MAE $r$ MAE $r$
M SD M SD M SD M SD M SD M SD
Persona
GPT-4o 0.48 0.44 .50 -.13 0.42 0.36 .64 -.24 0.42 0.30 .69 -.64
Llama3 0.51 0.26 .57 .02 0.33 0.15 .80 .31 0.31 0.06 .80 .76
Shape
GPT-4o 0.58 0.16 .38 -.06 0.50 0.18 .82 -.41 0.46 0.21 .84 -.77
Llama3 0.62 0.25 .45 -.26 0.52 0.38 .78 -.64 0.45 0.31 .82 -.65
PSI
GPT-4o 0.48 0.34 .68 .27 0.40 0.22 .71 .49 0.38 0.15 .67 .69
Llama3 0.66 0.36 .50 .14 0.54 0.27 .57 .19 0.49 0.21 .54 .51
Table 2: MAE and $r$ for BFI-2 Human Responses and Different Methods LLM Responses at Item, Facet, and Domain Levels
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. The bold text indicates the best-performing data in that column.
Psychometric Performance
Psychometric performance includes multiple components; here, we primarily present model fit and structural validity (i.e., factor loadings). For other aspects, such as scale reliability and discriminant validity, please refer to Appendix I.1.
Model Fit: For both the human sample criterion and different methods of LLM responses, the TFM was fitted to each BFI-2 domain, and the FFM was fitted to all the data. TFM fit information is shown in Table 18, and FFM fit information can be seen in Table 19 in Appendix I.1.
From Table 18, it can be observed that the model fit indices of the Persona method and the PSI method are relatively similar to those of the human sample, while the Shape method performs slightly worse. Table 19 reflects a similar trend, where the Shape method shows notable discrepancies compared to the human sample in the RMSEA and SRMR model fit indices (see Appendix G for the explanation of RMSEA, SRMR, and other model fit indices).
Structural Validity: TCC was used to evaluate the similarity of factor loadings between human responses and LLM responses. TCC results for the TFM of each BFI-2 domain are shown in Table 20, while results for the FFM are shown in Table 21 in Appendix I.1.
The values of the TCC are generally high and thus supportive of factor-loading correspondence, with all methods showing strong alignment with the human sample (a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity; Lorenzo-Seva and Ten Berge, 2006). However, it is important to note that the TCC focuses on overall profile similarity, largely ignoring differences in absolute values (which might be too lenient). Therefore, we also need to examine the results of the specific standardized factor loadings.
Tables 22 and 23 in Appendix I.1 present the standardized factor loadings. From these tables, it is clear that the PSI method outperforms the Shape method in terms of similarity to human samples, with its factor loadings more closely matching human data. However, its results are relatively close to those of the Persona method, yet there are clearly differences in some factor loadings compared to the human sample.
It is important to note that the highest factor loadings do not necessarily indicate the best performance, as we are evaluating the similarity between the factor loadings of the human data and the LLM responses. Higher factor loadings suggest stronger correlations between the factor and the items, but this is not always ideal. It indicates that LLMs may struggle to make refined distinctions among items within a facet, meaning they may have difficulty determining which items are more or less related to the factor, unlike human respondents, who exhibit greater differentiation.
Tables 24 and 25 in Appendix I.1 present the specific inter-factor correlations. Aligning with our previous findings, the PSI method generally performs better than the other two methods when simulating human personality distributions, especially at a more fine-grained level. For instance, in the TFM results, other methods encountered anomalies, such as inter-factor correlations exceeding one (accompanied by a warning message when fitting the model), but the PSI method generally performed well.
The other results (e.g., scale reliability and discriminant validity) can be found in Appendix I.1. Overall, the results produced by the PSI method generally outperform those of the Persona method and the Shape method. However, there is still a certain gap compared to the human sample.
5.3 Personality-Related Behavioral Performance Results
Table 28 presents the correlations between the personality dimensions and OCB/CWB reported by PSI GPT-4o, with a comparison to human self-reported data. The relevant results for PSI Llama3 are provided in Table 29 in Appendix I.2, while Figures 3 and Figure 4 display the complete correlation matrices for GPT-4o and Llama3, respectively.
The results demonstrate that the correlations simulated by LLMs between OCB/CWB and various personality dimensions closely resemble the patterns seen in human self-reported data, except for the openness domain. When human reports show positive, negative, or no correlation, the LLM simulations generally exhibit consistent trends in the corresponding directions.
However, the correlations generated by LLM simulations are often higher, likely because the model primarily relies on the “typical” or “idealized” knowledge structures absorbed from its training corpus during simulation, consistently repeating and amplifying such associations. In contrast, human self-reported data contains more random noise, leading to relatively weaker correlations.
6 Conclusion
In summary, this study proposes a novel method for LLM to simulate human personality distributions—PSI and provides a comprehensive explanation of its development process and corresponding dataset.
Through three experiments, we validated PSI’s effectiveness. Results show that PSI matches or even outperforms two-hour interviews (Park et al., 2024) in personality simulation and surpasses existing methods in terms of both item-level accuracy and overall psychometric indicators. One experiment also examined PSI’s ability to simulate personality-related behavior, revealing that although LLMs approach human-like performance, they still exhibit idealized tendencies.
In summary, a theory-informed structured interview like PSI can better simulate human-like psychometric data. We explore the role of theory-informed structured interviews in simulation in greater detail and examine their potential to advance research, both of which are elaborated upon in Appendix C.
Limitations
This study has several limitations. First, we explored the potential of PSI in advancing research. Our single experiment demonstrates that theory-informed questions can effectively elicit information embedded in human narrative responses—–particularly those relevant to the target personality construct—–allowing LLMs to simulate the corresponding behavior. However, additional experiments are needed to investigate its broader applications.
Second, the personality assessment in this paper relies on self-reported personality scales, which can be viewed as a limitation of the study. However, these self-report scales are specifically designed for humans to target personality traits across individuals, in ways that cannot be directly located within LLMs themselves (see Appendix C for a detailed discussion on psychometrics). Furthermore, self-report scales are used often because respondents have privileged access to their own personality, and even with their well-known limitations (e.g., faking or social desirability in self-reports; see Appendix I.3 for the social desirability analysis), they remain valuable to evaluate the personality traits exhibited by LLMs. However, because the primary goal of this study is to have the LLM simulate human responses and behaviors that correspond to extracted personality information, it is reasonable to assess LLMs using methods designed to measure human personality traits.
Moreover, this paper tests only two types of LLMs, which may somewhat limit its scope, given the array of LLMs that are now available. However, the conclusions drawn from our two models were highly consistent, and we have reason to believe that the advantages of the PSI method are generally applicable across different LLMs. Furthermore, we designed three different experiments and conducted comprehensive psychometric analyses of various methods within LLMs, ensuring greater rigor and reliability of the results.
Ethical Statement
We hereby confirm that all authors of this study are aware of the provided ACL Code of Ethics and comply with the Code of Conduct.
Human Sample Data
This paper discusses the comparison between multiple sets of human data and results generated by LLMs. The collection of human data strictly adhered to relevant ethical guidelines and received approval from the Institutional Review Board. To ensure fair treatment of participants and proper recognition of their contributions, reasonable compensation or course credit (for student participants) was provided.
Throughout the research process, we placed a strong emphasis on transparency and openness, also ensuring that all participants signed informed consent forms. Additionally, to safeguard data privacy and uphold ethical standards, the publicly available dataset underwent rigorous screening to exclude any personally identifiable information and was shared only with explicit public consent.
Appendix A Appendix: The Structure of Personality
A key question in personality-related LLM research pertains to personality structure: What is the nature and breadth of the personality traits we want to simulate the human personality distributions? In the research literature, personality structures often emerge from applying factor analysis to individuals’ responses to a large number of personality-relevant items. This approach is what has been used to identify the five personaltiy factors in the FFM. Moreover, personality is better understood in terms of one’s continuous standing on each of multiple dimensions rather than as static types or profiles. Research data clearly supports this view (Wilmot, 2015; Wilmot et al., 2019). Dividing individuals into limited categories (e.g., 16 types in MBTI) artificially segments continuous dimensions into discrete units, which may overlook important individual differences (Ones and Wiernik, 2018).
Another question is whether we should incorporate personality traits and structures beyond the FFM to achieve a more comprehensive understanding of personality, given that additional personality variables and alternative structures have been proposed over the years. It turns out that there can be conceptual overlap among these models (e.g., Hough et al., 2015).
For example, a meta-analysis by Joseph and Newman (2010) revealed that emotional intelligence (EI) shows statistically and practically significant relationships with neuroticism and extroversion within the FFM. In fact, when controlling for personality variables, the unique contribution of EI almost disappears. Similarly, Credé et al. (2017) conducted a meta-analysis on grit and found that its core components can largely be explained by conscientiousness, with little added predictive validity beyond that. Moreover, although the HEXACO structure introduces an honesty-humility dimension, the remaining five dimensions align closely with the FFM structure (Lee and Ashton, 2004, 2006). Research by Cutler and Condon (2023) further indicates that nearly all personality semantic information can be classified within the FFM structure.
These studies collectively suggest that although further subdivision of personality structures might provide new perspectives, it often leads to conceptual redundancy and measurement complexity without necessarily enhancing predictive validity or theoretical value. Rather than pursuing a wide range of personality frameworks, we will delve deeper into the FFM structure, which is very widely accepted by personality psychologists.
Appendix B Appendix: Personality Structured Interview Questions
Table 3 presents the final set of 32 questions that form the basis of our personality structured interview. This framework was developed by adapting and modifying McAdams’s life history interview and narrative identity approach (McAdams, 1995, 1996, 2001), while also incorporating components from the Structured Interview of the Five-Factor Model (SIFFM; Trull et al., 1998).
The initial draft of the question pool was created and refined through collaborative discussions among subject matter experts (SMEs). The SME team consisted of two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology. To ensure the quality and clarity of the questions, pilot testing was conducted with six undergraduate research assistants from a personality research lab. This iterative process of development and feedback led to the construction of the personality structured interview.
The development process is similar to the development of psychological tests or scales (psychometric). We have provided more details on psychometrics and the development framework in Appendix C.
# Questions
1 To get us started, where are you from? Where did you grow up and what was the place like?
2 Thinking back, what kind of student were you in school?
3 Did you have a teacher or teachers that were influential? If so, why? What were they like?
4 What was your favorite subject in school, and why?
5 What was your least favorite subject in school, and why?
6 Still thinking back, who were your heroes when you were young and why?
7 When you were little, what did you want to be when you grew up? And why?
8 What were your dreams and plans when you graduated from high school? What made you have those dreams or plans?
9 If you had complete freedom, what would your dream job be, and why?
10 How have your dreams and goals changed throughout your life?
11 Shifting gears to your childhood, how would you describe the personalities of people in the family you grew up in? For example, what were your parents and/or siblings like?
12 How are you similar or different from your parents and/or siblings?
13 How do you think your similarities and/or differences influenced your relationship with them?
14 What was the best part of your childhood?
15 What do you think were the worst parts of your childhood?
16 Switching gears a little bit, what was your first paid job? How old were you then? (If this is not applicable to you, then please put ‘NA’)
17 What other jobs have you had? (If this is not applicable to you, then please put ‘NA’)
18 What do you do now for a living? And why did you choose it?
19 Please describe your typical work day.
20 What is the best and worst part of your current work?
21 Did you serve in the military? Please tell us about that experience, what was the best and worst part of it?
22 Moving on, what are your adult friendships like?
23 How are your adult friendships different from your childhood friendships?
24 What are your strongest qualities as a friend? In other words, what makes you a great friend to have?
25 What about your weakest qualities in friendships? In other words, what do you struggle with when you are trying to be a friend to someone?
26 Moving onto more general questions, when thinking about your life in general, what are you most proud of?
27 What hobbies or other interests do you have?
28 What things frighten you now?
29 What were some things that frightened you most as a child?
30 What are the three biggest news events that have occurred in your lifetime?
31 If you had the power to solve one and only one problem in the world, what would it be, and why?
32 Tell me about a time when you did not know if you would make it. How did you overcome that challenge?
Table 3: Structured Interview Questions
Appendix C Appendix: Psychometrics and Structured Interview Development Framework
Psychometrics is a field of psychology dedicated to the theory and practice of psychological measurement. It primarily focuses on quantifying psychological traits, behaviors, and abilities through systematic testing and analysis. Psychological traits, such as personality dimensions, cognitive abilities, and emotional states, are inherently abstract constructs that cannot be measured directly. Therefore, psychometricians rely on tools like surveys, questionnaires, scales, or structured interviews to infer these traits through observable indicators or responses.
Measuring Psychological Traits
To measure a psychological trait, psychometricians typically operationalize the trait by identifying observable behaviors or self-identities that correlate with the underlying construct. For example, extraversion can be measured by assessing behaviors such as sociability, assertiveness, and enjoyment of social interactions, or by examining identities such as seeing oneself as an outgoing person and believing that one thrives in social situations.
These behaviors/identities are translated into measurable items (for scale; e.g., “I enjoy being the center of attention”) or questions (for interview; e.g., “What are your strongest qualities as a friend? In other words, what makes you a great friend to have?”). The challenge lies in ensuring that these items/questions accurately and consistently capture the construct across different populations and contexts.
Theory-Informed Structured Interview for LLM Data Simulation
Theory-informed structured interviews are the most suitable method for enabling LLMs to simulate psychometric data. These interviews are specifically designed to capture the constructs underlying the targeted psychometric measures, ensuring that the simulated data aligns with the intended psychological construct. By extracting textual information that directly reflects the target construct, theory-informed structured interviews facilitate the representation of heterogeneous data while preserving a high degree of human diversity, thereby enhancing the validity and applicability of the simulated psychometric data.
Moreover, since the information is extracted based on theoretical foundations, it also provides a certain level of interpretability for the LLM’s simulation. This not only allows the generated data to be compared with theoretical expectations but also increases its potential for practical applications.
The Potential for Advancing Research
A theory-informed structured interview transcript-based simulation can generate data more effectively by focusing on the target construct. Ideally, the simulated data should reproduce the same constructs reflected in real-world data and simulate behaviors associated with these constructs.
Take personality as an example—if simulated data can accurately replicate real-world personality constructs, it enables research that would be difficult to conduct in reality, such as developing contextualized personality assessment tools and exploring new personality theories through multi-agent simulations.
Developing Contextualized Personality Assessment Tools: Traditional personality assessments mainly rely on standardized questionnaires or laboratory tasks, which often fail to adequately simulate real-world social contexts. By using theory-informed structured interview transcript-based simulations, we can generate more fine-grained and context-sensitive individual response data. For instance, we can simulate various occupational scenarios (such as crisis management, teamwork, or remote work) and analyze how different personality traits manifest in these contexts. This approach not only aids in developing measurement tools tailored to specific applications but also enhances ecological validity, allowing for more accurate assessments of personality across different situations.
Exploring New Personality Theories through Multi-Agent Simulation: If simulated data can accurately reflect real-world personality constructs, we can leverage multi-agent interactive systems to simulate individuals’ behavioral patterns and observe how different personality traits evolve in group dynamics. For example, virtual agents with distinct personality traits can be placed in cooperative tasks, competitive environments, or social interactions, enabling researchers to test whether existing personality theories effectively predict these interaction patterns. Additionally, this approach can uncover new personality dynamics, such as whether certain personality trait combinations produce unexpected group effects or whether behavior in specific situations deviates from traditional theoretical predictions.
However, these assumptions are based on an ideal premise—that simulated data can successfully reflect the same constructs as real-world data. A theory-informed structured interview undoubtedly offers a promising pathway in this regard, warranting further in-depth exploration.
Structured Interview Development Framework
Developing a structured interview for LLM to simulate data involves a series of carefully designed steps to ensure that the resulting test reliably and validly measures the construct of interest. Table 4 outlines the framework.
# Steps
1 Identify behaviors/perceptions that represent the construct or define the domain.
2 Prepare a set of structured interview specifications—structured interview blueprint.
3 Build an initial question pool.
4 Have questions reviewed by substantive experts (and revised as necessary).
5 Hold preliminary question tryouts.
6 Determine statistical properties of questions (and eliminate poor questions or revise as necessary).
7 Field-test the structured interview on a large representative sample of the intended examinee population.
8 Design and conduct reliability and validity studies for the final form of the structured interview.
9 Develop guidelines for administration, and interpretation of the structured interview.
Table 4: Structured Interview Development Framework
Identify behaviors/perceptions that represent the construct or define the domain: Clearly defining the construct is essential to developing relevant structured interview questions. The construct should be operationalized by identifying specific behaviors or attributes that indicate its presence. In other words, you need to find a theory to guide you on how to measure the target constructs. This step may involve reviewing literature, conducting expert interviews, or organizing focus groups to understand the various dimensions and observable characteristics of the construct.
Prepare a set of structured interview specifications—structured interview blueprint: A blueprint outlines the structured interview’s structure and content, specifying how questions will be distributed across the construct’s dimensions or components. It typically includes information on the number of questions per domain, questions content.
Build an initial question pool: In this step, an extensive list of questions is created to cover the full range of the construct. Question wording should be clear, concise, and relevant to the target population. It is common practice to generate more questions than needed to ensure that poorly performing questions can be removed later without compromising the structured interview.
Have questions reviewed by substantive experts (and revised as necessary): SMEs review the question pool for content accuracy, relevance, clarity, and bias. Experts assess whether the questions align with the construct’s definition and whether any important aspects are missing. Feedback from experts helps refine the wording, remove ambiguous questions, and identify questions with potential cultural or gender biases.
Hold preliminary question tryouts: Before large-scale testing, questions are piloted on a small group of individuals representative of the target population. This stage helps identify any immediate issues with question comprehension, response format, or instructions. It can also include cognitive interviews where participants are asked to explain their thought processes when answering questions. The feedback from this stage informs further revisions, ensuring that questions are clear and easy to understand.
Determine statistical properties of questions (and eliminate poor questions or revise as necessary): Question performance is assessed through statistical analyses to evaluate difficulty, discrimination, and internal consistency. Standardized scoring criteria, such as Behaviorally Anchored Rating Scales (BARS), can be used for obtaining question scores. Then, methods like correlations, factor analysis, and item response theory (IRT) can help determine how effectively each question measures the intended construct.
Questions that demonstrate poor psychometric properties—such as low discrimination or high measurement error—are either revised or removed. For example, questions with low correlations with the overall score or incorrect factor loadings may be eliminated.
Field-test the structured interview on a large representative sample of the intended examinee population: The revised item set is administered to a large, representative sample to gather comprehensive data on the scale’s performance. This step ensures that the sample reflects the population for which the test is intended, which is critical for generalizability. Statistical analyses are conducted to refine the test further. This process may involve removing redundant items, assessing dimensionality, and ensuring that items work well across demographic subgroups.
Design and conduct reliability and validity studies for the final form of the structured interview: To ensure the structured interview is psychometrically sound, various reliability and validity studies are conducted: Reliability studies measure the structured interview’s consistency and stability, including internal consistency (e.g., Cronbach’s alpha), test-retest reliability, and inter-rater reliability (if applicable). Validity studies assess whether the structured interview measures what it is intended to measure. This includes: Content validity (the extent to which items cover the construct); Construct validity (e.g., convergent and discriminant validity); Criterion-related validity (e.g., predictive or concurrent validity). These studies provide evidence that the structured interview is both reliable and valid for its intended purpose.
Develop guidelines for administration and interpretation of the structured interview: The final step involves creating a comprehensive, structured interview manual that includes instructions for structured interview administration, scoring procedures, and guidelines for interpreting results. This manual ensures consistency in the structured interview’s use across different settings and helps minimize errors in administration and scoring. Guidelines for interpreting scores may include norms, cutoff points, and descriptions of what various score ranges indicate.
Appendix D Appendix: Personality Structured Interview Dataset and Data Collection
Institutional Review Board approval was obtained. The data was collected through an online questionnaire, followed by an online structured interview.
Since we are still actively collecting data, we will share a part of the dataset that does not include any personal privacy information and has received permission for data sharing. Each example is composed of the following characteristics:
  1. 1.
    Gender: The gender of the participant.
  2. 2.
    Race: The racial background of the participant.
  3. 3.
    English: Whether English is the participant’s first language.
  4. 4.
    Age: The participant’s age.
  5. 5.
    Weight: The participant’s weight.
  6. 6.
    Height: The participant’s height.
  7. 7.
    OCB1–OCB10: Self-reported Organizational Citizenship Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 10 for specific item details.
  8. 8.
    CWB1–CWB10: Self-reported Counterproductive Work Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 11 for specific item details.
  9. 9.
    Q1–Q32: Participant responses to each personality structured interview questions, see Table 3 for specific question details.
  10. 10.
    Item1–Item60: Responses to each item of the BFI-2. Refer to Appendix E for item descriptions and scoring guidelines.
Appendix E Appendix: BFI-2 Scale
Instructions: Here are a number of characteristics that may or may not apply to you. For example, do you agree that you are someone who likes to spend time with others? Please write a number next to each statement to indicate the extent to which you agree or disagree with that statement.
Scales:
# Statement
1 I am someone who is outgoing, sociable.
2 I am someone who is compassionate, has a soft heart.
3 I am someone who tends to be disorganized.
4 I am someone who is relaxed, handles stress well.
5 I am someone who has few artistic interests.
6 I am someone who has an assertive personality.
7 I am someone who is respectful, treats others with respect.
8 I am someone who tends to be lazy.
9 I am someone who stays optimistic after experiencing a setback.
10 I am someone who is curious about many different things.
11 I am someone who rarely feels excited or eager.
12 I am someone who tends to find fault with others.
13 I am someone who is dependable, steady.
14 I am someone who is moody, has up and down mood swings.
15 I am someone who is inventive, finds clever ways to do things.
16 I am someone who tends to be quiet.
17 I am someone who feels little sympathy for others.
18 I am someone who is systematic, likes to keep things in order.
19 I am someone who can be tense.
20 I am someone who is fascinated by art, music, or literature.
21 I am someone who is dominant, acts as a leader.
22 I am someone who starts arguments with others.
23 I am someone who has difficulty getting started on tasks.
24 I am someone who feels secure, comfortable with self.
25 I am someone who avoids intellectual, philosophical discussions.
26 I am someone who is less active than other people.
27 I am someone who has a forgiving nature.
28 I am someone who can be somewhat careless.
29 I am someone who is emotionally stable, not easily upset.
30 I am someone who has little creativity.
31 I am someone who is sometimes shy, introverted.
32 I am someone who is helpful and unselfish with others.
33 I am someone who keeps things neat and tidy.
34 I am someone who worries a lot.
35 I am someone who values art and beauty.
36 I am someone who finds it hard to influence people.
37 I am someone who is sometimes rude to others.
38 I am someone who is efficient, gets things done.
39 I am someone who often feels sad.
40 I am someone who is complex, a deep thinker.
41 I am someone who is full of energy.
42 I am someone who is suspicious of others’ intentions.
43 I am someone who is reliable, can always be counted on.
44 I am someone who keeps their emotions under control.
45 I am someone who has difficulty imagining things.
46 I am someone who is talkative.
47 I am someone who can be cold and uncaring.
48 I am someone who leaves a mess, doesn’t clean up.
49 I am someone who rarely feels anxious or afraid.
50 I am someone who thinks poetry and plays are boring.
51 I am someone who prefers to have others take charge.
52 I am someone who is polite, courteous to others.
53 I am someone who is persistent, works until the task is finished.
54 I am someone who tends to feel depressed, blue.
55 I am someone who has little interest in abstract ideas.
56 I am someone who shows a lot of enthusiasm.
57 I am someone who assumes the best about people.
58 I am someone who sometimes behaves irresponsibly.
59 I am someone who is temperamental, gets emotional easily.
60 I am someone who is original, comes up with new ideas.
Table 5: BFI-2 Scale
Scoring: Reverse-keyed items are denoted by “R.”
In psychological measurement and scale development, reverse coding is a common and essential technique. Its primary purpose is to ensure that the scoring direction of all items remains consistent, thereby improving the reliability of the measurement results and the accuracy of their interpretation.
A balanced arrangement of positive and negative items helps reduce response biases, such as consistency effects or response patterns where participants select the same option repeatedly. These tendencies can mask the respondent’s true attitudes or behavioral traits, leading to distorted measurement outcomes. By alternating positive and negative statements and applying reverse coding to the appropriate items, this issue can be effectively mitigated.
Reverse coding can also help reduce the influence of social desirability bias. When all items are presented in the same direction, participants may easily guess the purpose of the test and provide responses that align with perceived expectations. By mixing positive and negative statements and applying reverse coding, the scale can disrupt this pattern, making it harder for participants to determine the “correct” answers, thus providing a more genuine reflection of their inner states or attitudes.
Domain Level Item Numbers
Extraversion 1, 6, 11R, 16R, 21, 26R, 31R, 36R, 41, 46, 51R, 56
Agreeableness 2, 7, 12R, 17R, 22R, 27, 32, 37R, 42R, 47R, 52, 57
Conscientiousness 3R, 8R, 13, 18, 23R, 28R, 33, 38, 43, 48R, 53, 58R
Neuroticism 4R, 9R, 14, 19, 24R, 29R, 34, 39, 44R, 49R, 54, 59
Openness 5R, 10, 15, 20, 25R, 30R, 35, 40, 45R, 50R, 55R, 60
Table 6: BFI-2 Domain Level with Item Numbers
Facet Level Item Numbers
Sociability 1, 16R, 31R, 46
Assertiveness 6, 21, 36R, 51R
Energy Level 11R, 26R, 41, 56
Compassion 2, 17R, 32, 47R
Respectfulness 7, 22R, 37R, 52
Trust 12R, 27, 42R, 57
Organization 3R, 18, 33, 48R
Productiveness 8R, 23R, 38, 53
Responsibility 13, 28R, 43, 58R
Anxiety 4R, 19, 34, 49R
Depression 9R, 24R, 39, 54
Emotional Volatility 14, 29R, 44R, 59
Intellectual Curiosity 10, 25R, 40, 55R
Aesthetic Sensitivity 5R, 20, 35, 50R
Creative Imagination 15, 30R, 45R, 60
Table 7: BFI-2 Facet Level with Item Numbers
Appendix F Appendix: Prompts List
Personality Scale Prompt Format
Table 8 presents the prompts used to generate LLM responses for the selected personality test on a Likert scale, where personality_description denotes the personality prompt.
In the PSI method, the personality prompt integrates both the interview question and the interviewee’s corresponding response, as collected in the PSI dataset (see Appendix D). For the Persona and Shape methods, more detailed descriptions of the personality prompts can be found in Appendix G.
The test_item corresponds to each individual item in the BFI-2 scale.
For the following task, respond in a way that matches this description: {personality_description}.
Considering the statement, please indicate the extent to which you agree or disagree on a scale from 1 to 5 (where 1 = “disagree strongly”, 2 = “disagree a little”, 3 = “neither agree nor disagree”, 4 = “agree a little”, and 5 = “agree strongly”): {test_item}.
Table 8: Prompt format for gathering LLMs’ responses to BFI-2 scale.
Behavioral Question Prompt Format
The complete prompt format for eliciting LLMs’ responses to personality-related behavioral questions is defined in Table 9.
The questions_and_responses refers to the transcript of a structured human personality interview, while question_list comprises statements evaluating OCB and CWB, as detailed in Table 10 and Table 11 in Appendix G. The model is instructed to rate each statement using a frequency scale from one to five.
Task: Simulate an individual’s behavior and predict their responses to a series of work-related questions.
Data to Analyze: {questions_and_responses}
Instructions:
1. Simulate the Individual: Use the provided data to simulate the individual’s personality, behavior, and work habits.
2. Predict Responses:
- Based on the provided data, simulate the individual and answer the question.
- Recall that in the past year, how often have you done the following things? (1 = Never, 2 = Once or twice, 3 = Once or twice per month, 4 = Once or twice per week, 5 = Every day)
Questions: {question_list}
Output Format:
Question Score: [Provide the numerical score here.]
Table 9: Prompt format for gathering LLMs’ responses to personality-related behavioral questions.
Appendix G Appendix: Additional Experiment Settings
Pearson Correlation Coefficient
The correlation calculation formula is as follows:
$r=\frac{\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}}$
$X_{i}$ and $Y_{i}$ represent the data values for LLMs and human. The symbols $\overline{X}$ and $\overline{Y}$ represent the means of variables $X$ and $Y$, respectively. The numerator, $\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})$, represents the covariance between $X$ and $Y$. The denominator, $\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}$, standardizes the result, constraining the value of $r$ to range between -1 and 1.
When $r$ is close to 1, it indicates a strong positive correlation between the two variables; when $r$ is close to -1, it indicates a strong negative correlation; and when $r$ is close to 0, it indicates no significant linear relationship between the two variables.
Persona Method
The Persona method is based on the Persona-Chat dataset constructed by Zhang et al. (2018). The dataset consists of persona descriptions, and each is made up of five short sentences containing demographic information collected through Amazon Mechanical Turk crowdsourcing. To avoid sentence similarity or repetition, these persona descriptions were required to be rewritten (e.g., changing “I am very shy” to “I am not a social person”). Zhang et al. (2018) demonstrated through machine learning model validation and human evaluations that such persona descriptions provide an effective method to enhance personalization. Currently, incorporating personal profiles into prompts is widely used in research related to LLM agents (Park et al., 2023; Wang et al., 2023; Xi et al., 2023).
In the current study, we treated each persona description as an individual entity (i.e., a single subject) and randomly selected 300 persona descriptions from the dataset. One example is, “I wear a lot of leather. I have boots I always wear. I sleep in late during the day. I listen to metal music. I have black spiky hair.”
Shape Method
The Shape method is based on the work of (Serapio-García et al., 2023), who introduced a prompting approach to shape synthetic personality in LLMs along desired dimensions. The researchers expanded upon Goldberg (1990) lexical hypothesis, expanding his list of 70 bipolar adjectives (Goldberg, 1992) to include 104 trait adjectives. Additionally, they employed linguistic qualifiers commonly used in Likert-type scales (Likert, 1932), such as “a bit,” “very,” and “extremely,” to set target levels for each adjective. This resulted in a fine-grained prompting method with nine levels: 1. extremely low adjective; 2. very low adjective; 3. low adjective; 4. a bit low adjective; 5. neither low adjective nor high adjective; 6. a bit high adjective; 7. high adjective; 8. very high adjective; 9. extremely high adjective.
In our study, each prompt involves five randomly selected adjective markers from a specific personality domain. These markers are positioned after a consistent linguistic qualifier to set the prompt at one of nine intensity levels. For example, one prompt is: “You are extremely friendly, extremely energetic, extremely assertive, extremely bold, and extremely active.” These are five positively-keyed adjectives targeting extraversion. In this case, the prompt seeks to create a highly sociable and dynamic personality profile, which might result in responses characterized by enthusiasm, confidence, and proactivity. We also randomly select 300 prompts here.
CFA Model
The basic form of the CFA model is:
$y=\Lambda\eta+\epsilon$
where $y$ represents the vector of observed variables; $\Lambda$ is the factor loading matrix (i.e., the loadings of each observed variable on the latent factors); $\eta$ is the vector of latent factors; $\epsilon$ is the vector of error terms, with the assumption that the error terms have a mean of zero and are mutually independent (Jöreskog, 1969).
Latent factors are variables that are not directly observed but are inferred from other variables that are observed (measured). In the context of a CFA model, latent factors represent underlying constructs or traits that are believed to influence the observed variables. For example, in psychology, a latent factor might represent a construct like “intelligence” or “anxiety,” which cannot be measured directly but can be estimated through related observed behaviors or responses on a test.
Model Fit Information
Chi-Square Test, $\mathbf{\chi^{2}}$: The chi-square test is used to measure the difference between the observed covariance matrix and the factor model’s fitted covariance matrix.
$\chi^{2}=(N-1)\times F_{\text{ML}}$
where $N$ is the sample size; $F_{\text{ML}}$ is the value of the fit function under maximum likelihood estimation. A smaller chi-square value indicates a better model fit. However, with large samples, the chi-square value tends to be large, so other fit indices are usually the primary reference.
Degrees of Freedom, df: The df represent the relationship between model parameters and observed variables:
$df=\frac{p(p+1)}{2}-q$
where $p$ is the number of observed variables, and $q$ is the number of model parameters.
Degrees of freedom reflect the amount of independent information available in a statistical model to estimate its parameters. They are calculated as the number of information points provided by the data minus the number of model parameters, representing the extent to which the model can adjust freely. Therefore, a higher degree of freedom indicates fewer parameters in the model and fewer constraints on the data. With more degrees of freedom, the model has fewer restrictions, though the fitting difficulty may increase. Too few degrees of freedom may lead to overfitting, while too many can result in underfitting.
Comparative Fit Index, CFI: The CFI is used to compare the goodness of fit of a model with a baseline model (usually an independent model).
$\text{CFI}=1-\frac{\max(\chi^{2}-\textit{df},0)}{\max(\chi_{\text{null}}^{2}-\textit{df}_{\text{null}},0)}$
where $\chi^{2}$ and df are the chi-square value and degrees of freedom of the target model; $\chi_{\textit{null}}^{2}$ and $\textit{df}_{\text{null}}$ are the chi-square value and degrees of freedom of the baseline (independent) model.
Tucker-Lewis Index, TLI: TLI, also known as the Non-Normed Fit Index (NNFI), considers model complexity.
$\text{TLI}=\frac{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-\left(\chi^{2}/\textit{df}\right)}{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-1}$
This value ranges from 0 to 1, with a value typically greater than.90 indicating good model fit.
Root Mean Square Error of Approximation, RMSEA: The RMSEA quantifies the error per degree of freedom in a model, with smaller values indicating better model fit.
$\text{RMSEA}=\frac{\chi^{2}-df}{df(N-1)}$
where $\chi^{2}$ is the chi-square value of the model; df is the degrees of freedom; N is the sample size.
Standardized Root Mean Square Residual, SRMR: The SRMR measures the discrepancy between model-predicted values and actual observed values, calculated as:
$\text{SRMR}=\sqrt{\frac{\sum_{i=1}^{p}\sum_{j=1}^{p}\left(s_{ij}-\hat{s}_{ij}\right)^{2}}{\frac{p(p+1)}{2}}}$
where $s_{ij}$ is an element in the observed covariance matrix; $\hat{s}_{ij}$ is an element in the model-fitted covariance matrix.
Based on Hu and Bentler (1999), CFI/TLI $\geq$ .95, RMSEA $\leq$ .06, and SRMR $\leq$ .08 are considered good fit thresholds.
Tucker’s Congruence Coefficient
Tucker’s congruence coefficient, also known as the coefficient of congruence, is typically used to assess the similarity between two-factor structures in factor analysis (Tucker, 1951). The formula is given by:
$\phi=\frac{\sum_{i=1}^{n}a_{i}b_{i}}{\sqrt{\sum_{i=1}^{n}a_{i}^{2}\cdot\sum_{i=1}^{n}b_{i}^{2}}}$
where $a_{i}$ and $b_{i}$ are the loadings of the $i$-th factor for two different factor solutions (or different samples or methods); $n$ is the total number of factors.
The coefficient ranges from -1 to 1, where values close to 1 indicate high similarity (congruence) between the factor solutions, values close to 0 indicate low similarity, and negative values indicate a dissimilar or inverse relationship. According to Lorenzo-Seva and Ten Berge (2006), a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity. However, this is a relatively lenient criterion; specific differences still need to be determined based on the factor loadings.
Behavior Variable Measures
Organizational Citizenship Behavior (OCB): OCB was measured using ten items from Spector and Fox (2010) to assess extra-role behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: “In the past year, how often have you helped new employees get oriented to the job?”. Internal consistency was Cronbach’s alpha = .83.
Counterproductive Work Behavior (CWB): CWB was measured using ten items from Spector and Fox (2010), designed to assess harmful workplace behaviors. Items were rated on a frequency scale ranging from 1 (never) to 5 (every day). Example item: “In the past year, how often have you ignored someone at work?”. Internal consistency was Cronbach’s alpha = .86.
The detailed scale information for both can be found in Table 10 and Table 11 below.
Instructions: Recall that in the past year, how often have you done the following things (1 = Never, 2 = Once or twice, 3 = Once or twice per month, 4 = Once or twice per week, 5 = Every day)?
# Statement
1 Took time to advise, coach, or mentor a co-worker.
2 Helped co-worker learn new skills or shared job knowledge.
3 Helped new employees get oriented to the job.
4 Lent a compassionate ear when someone had a work problem.
5 Offered suggestions to improve how work is done.
6 Helped a co-worker who had too much to do.
7 Volunteered for extra work assignments.
8 Worked weekends or other days off to complete a project or task.
9 Volunteered to attend meetings or work on committees on own time.
10 Gave up meal and other breaks to complete work.
Table 10: OCB Scale
# Statement
1 Purposely wasted your employer’s materials/supplies.
2 Complained about insignificant things at work.
3 Told people outside the job what a lousy place you work for.
4 Came to work late without permission.
5 Stayed home from work and said you were sick when you weren’t.
6 Insulted someone about their job performance.
7 Made fun of someone’s personal life.
8 Ignored someone at work.
9 Started an argument with someone at work.
10 Insulted or made fun of someone at work.
Table 11: CWB Scale
Appendix H Appendix: Psychometric Data Evaluation Framework
Here, we present the evaluation framework used to assess the fidelity of simulated psychometric data (i.e., how well it aligns with human data).
The evaluation is conducted at different levels. For example, for personality, it includes Item, Facet, and Domain levels. Generally, the structure of psychometric data is hierarchical, where observed responses (item level) map onto latent traits (domain level). Our evaluation framework incorporates both descriptive statistics and psychometric performance metrics to ensure a comprehensive evaluation (see Figure 2). Below, we outline each component and the rationale for its inclusion.
Descriptive Statistics
Descriptive statistics are primarily used to summarize, outline, and present the basic features of data. They help researchers understand the distribution and fundamental trends of the data without interpreting or measuring specific psychological constructs. Do not forget that the evaluation of psychometric data is hierarchical, we need to evaluate on item and domain level.
Mean (M): The mean represents the central tendency of the responses, reflecting the average score across individuals.
Standard Deviation (SD): The standard deviation captures the dispersion of responses, indicating how much variation exists within the data.
Both M and SD can be quantified for similarity to human distribution using MAE and $r$. However, these metrics provide a more summarized level of comparison; we also need to examine performance on specific items and the domain.
Distribution Shape: The distribution shape describes the overall pattern of how responses are spread across the scale. It provides insight into whether the data follows a normal distribution or exhibits skewness and kurtosis.
Skewness measures the asymmetry of the distribution. A positive skew indicates a longer right tail (more low scores with a few high scores), while a negative skew indicates a longer left tail (more high scores with a few low scores).
Kurtosis captures the “tailedness” of the distribution. High kurtosis (leptokurtic) suggests heavy tails with more extreme values, while low kurtosis (platykurtic) indicates a flatter distribution with fewer extreme values.
Psychometric Performance
Psychometric performance is primarily used to evaluate the measurement quality of psychological constructs, ensuring that assessments accurately and reliably capture the intended traits.
Structural Validity: Structural validity refers to the extent to which the internal structure of a measurement instrument aligns with the theoretical construct it is intended to assess. It can be assessed through model fit, factor loadings, and inter-factor correlations.
The model fit indices provide an overall assessment of how well the proposed structure aligns with the observed data. Ideally, the simulated psychometric data will have a similar model fit compared with the human data. Common indices include $\mathbf{\chi^{2}}$, CFI, TLI, RMSEA, and SRMR.
Factor loadings indicate the extent to which each item represents the intended construct, while inter-factor correlations reveal relationships between latent variables. The simulated psychometric data should also resemble human data in these two aspects. We can use TCC as a summarized level of comparison; however, it is typically considered too lenient, so we need to examine factor loadings and inter-factor correlation values more closely.
Scale Reliability: Scale reliability assesses the internal consistency of items measuring the same construct. Cronbach’s alpha is a widely used reliability coefficient, The simulated psychometric data should also resemble human data in this aspect.
Discriminant Validity: Discriminant validity ensures that distinct constructs are not excessively correlated. It can be examined by calculating the mean absolute correlation.
Figure 2: Psychometric Data Evaluation Framework
Appendix I Appendix: Additional Analyses and Results
I.1 Additional Method Comparison Results
Additional Descriptive Statistics Results
Here we show the detailed means and standard deviations for human responses and LLM responses at the item, facet, and domain levels, see Tables 1213141516, and 17.
No. Item Content Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
1 I am someone who is outgoing, sociable 3.03 3.41 2.77 2.76 2.45 2.67 2.14
2 I am someone who is compassionate, has a soft heart 4.34 3.51 3.87 3.10 3.17 4.05 3.91
3 I am someone who tends to be disorganized 3.62 3.39 3.77 3.32 3.31 3.34 3.24
4 I am someone who is relaxed, handles stress well 2.56 3.20 2.94 3.47 3.32 2.91 2.96
5 I am someone who has few artistic interests 3.68 3.72 3.80 3.27 2.92 2.75 2.73
6 I am someone who has an assertive personality 2.96 3.08 3.07 2.82 2.81 2.57 2.25
7 I am someone who is respectful, treats others with respect 4.62 3.38 4.43 3.48 3.62 4.16 4.22
8 I am someone who tends to be lazy 3.62 4.00 4.23 3.51 3.44 3.54 3.75
9 I am someone who stays optimistic after a setback 2.44 2.74 2.29 2.95 2.92 2.18 2.29
10 I am someone who is curious about many different things 4.43 3.89 3.71 3.02 3.15 3.60 3.05
11 I am someone who rarely feels excited or eager 3.54 4.30 4.18 3.21 2.99 2.98 2.40
12 I am someone who tends to find fault with others 3.34 3.29 4.31 3.47 3.46 3.58 3.71
13 I am someone who is dependable, steady 4.33 3.29 3.70 2.93 3.11 4.12 3.96
14 I am someone who is moody, has up and down mood swings 2.56 3.01 2.52 2.72 2.60 3.04 2.31
15 I am someone who is inventive, finds clever ways to do things 3.90 3.44 3.24 3.06 2.91 3.05 2.37
16 I am someone who tends to be quiet 2.35 3.26 3.44 3.02 3.25 2.17 2.51
17 I am someone who feels little sympathy for others 3.71 3.73 4.47 3.72 3.54 4.22 2.88
18 I am someone who is systematic, likes to keep things in order 3.96 3.00 2.46 2.94 2.59 3.18 2.44
19 I am someone who can be tense 3.29 3.23 3.18 3.12 3.04 3.66 2.73
20 I am someone who is fascinated by art, music, or literature 4.06 3.61 2.99 3.15 2.90 3.43 2.81
21 I am someone who has an assertive personality (dominant) 2.90 2.94 2.46 2.77 2.53 2.42 2.07
22 I am someone who starts arguments with others 4.23 3.65 4.56 3.65 3.70 4.44 4.35
23 I am someone who has difficulty getting started on tasks 3.31 3.29 3.85 3.16 3.23 3.10 2.79
24 I am someone who feels secure, comfortable with self 2.26 2.75 2.24 3.14 2.65 2.65 2.70
25 I am someone who avoids intellectual, philosophical discussions 3.93 3.40 3.36 3.23 3.13 2.85 2.57
26 I am someone who is less active than other people 3.35 3.80 3.86 3.36 3.14 2.65 2.67
27 I am someone who has a forgiving nature 3.78 3.05 3.48 3.05 3.28 3.51 3.59
28 I am someone who can be somewhat careless 3.56 3.30 3.19 3.16 3.00 3.06 3.15
29 I am someone who is emotionally stable, not easily upset 2.44 3.16 2.66 3.17 2.87 2.70 2.67
30 I am someone who has little creativity 3.95 4.10 4.36 3.59 3.42 3.37 2.85
31 I am someone who is sometimes shy, introverted 2.35 3.28 3.43 3.11 3.34 2.14 2.70
32 I am someone who is helpful and unselfish with others 4.18 3.32 3.88 3.27 3.24 4.06 4.12
33 I am someone who keeps things neat and tidy 3.72 2.99 2.47 3.07 2.63 3.04 2.70
34 I am someone who worries a lot 3.30 2.96 2.59 2.79 2.50 3.64 3.04
35 I am someone who values art and beauty 4.16 3.65 3.23 3.14 3.06 3.38 2.87
36 I am someone who finds it hard to influence people 3.09 3.32 3.44 3.25 2.99 2.74 2.25
37 I am someone who is sometimes rude to others 3.79 3.24 4.43 3.50 3.63 3.70 4.10
38 I am someone who is efficient, gets things done 4.23 3.31 3.45 3.07 3.12 3.69 3.07
39 I am someone who often feels sad 2.71 2.86 2.27 2.74 2.42 3.24 2.69
40 I am someone who is complex, a deep thinker 4.02 3.64 3.59 3.19 3.14 3.34 2.74
41 I am someone who is full of energy 3.14 3.45 3.39 2.99 2.86 2.81 2.41
42 I am someone who is suspicious of others’ intentions 2.75 3.14 3.91 3.27 3.40 3.42 3.60
43 I am someone who is reliable, can always be counted on 4.32 3.28 3.76 3.12 3.18 4.06 3.75
44 I am someone who keeps their emotions under control 2.22 3.10 2.92 2.98 2.90 2.50 2.34
45 I am someone who has difficulty imagining things 4.17 4.06 4.08 3.57 3.43 3.21 2.14
46 I am someone who is talkative 2.88 3.11 2.53 2.87 2.50 3.31 2.36
47 I am someone who can be cold and uncaring 3.84 3.95 4.44 3.60 3.47 4.08 4.19
48 I am someone who leaves a mess, doesn’t clean up 4.15 3.47 4.25 3.39 3.25 3.32 4.03
49 I am someone who rarely feels anxious or afraid 3.44 3.20 3.12 2.85 2.71 3.48 3.66
50 I am someone who thinks poetry and plays are boring 3.66 3.54 3.85 3.44 3.15 2.90 2.29
51 I am someone who prefers to have others take charge 2.96 3.54 4.33 3.39 3.20 3.52 3.87
52 I am someone who is polite, courteous to others 4.50 3.28 4.13 3.33 3.56 3.97 4.38
53 I am someone who is persistent, works until the task is finished 4.27 3.56 3.87 3.01 3.13 4.07 3.75
54 I am someone who tends to feel depressed, blue 2.65 2.70 2.01 2.60 2.36 3.25 2.78
55 I am someone who has little interest in abstract ideas 3.83 3.25 3.40 3.08 2.93 2.53 2.15
56 I am someone who shows a lot of enthusiasm 3.35 3.86 4.08 2.81 2.81 2.70 2.58
57 I am someone who assumes the best about people 3.37 3.12 3.40 3.04 3.13 3.46 3.64
58 I am someone who sometimes behaves irresponsibly 3.64 3.11 3.44 2.90 3.22 3.13 3.55
59 I am someone who is temperamental, gets emotional easily 2.53 3.02 2.33 2.80 2.66 2.96 2.04
60 I am someone who is original, comes up with new ideas 3.84 3.28 3.25 2.96 2.88 3.05 2.40
Table 12: Item Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
No. Item Content Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
1 I am someone who is outgoing, sociable 1.40 0.87 1.31 1.22 1.37 0.98 0.78
2 I am someone who is compassionate, has a soft heart 0.90 0.84 1.12 1.22 1.44 0.93 1.03
3 I am someone who tends to be disorganized 1.36 0.78 0.95 1.17 1.46 0.76 0.96
4 I am someone who is relaxed, handles stress well 1.30 0.79 1.16 1.16 1.39 0.92 0.96
5 I am someone who has few artistic interests 1.33 1.04 1.31 1.25 1.32 1.39 1.11
6 I am someone who has an assertive personality 1.38 0.79 1.18 1.18 1.37 0.76 0.73
7 I am someone who is respectful, treats others with respect 0.61 0.89 0.92 1.23 1.42 0.75 0.78
8 I am someone who tends to be lazy 1.31 0.87 1.04 1.28 1.50 0.81 0.97
9 I am someone who stays optimistic after experiencing a setback 1.24 0.84 0.98 1.34 1.54 0.89 0.93
10 I am someone who is curious about many different things 0.81 0.89 1.12 1.37 1.55 0.98 0.91
11 I am someone who rarely feels excited or eager 1.26 0.84 0.75 1.33 1.41 0.93 0.82
12 I am someone who tends to find fault with others 1.32 0.59 0.69 1.26 1.48 0.83 0.70
13 I am someone who is dependable, steady 0.86 0.85 1.12 1.33 1.59 0.79 0.75
14 I am someone who is moody, has up and down mood swings 1.36 0.56 1.09 1.27 1.41 0.76 0.93
15 I am someone who is inventive, finds clever ways to do things 1.03 0.69 1.14 1.21 1.44 0.63 0.74
16 I am someone who tends to be quiet 1.34 0.84 1.12 1.12 1.27 1.02 0.97
17 I am someone who feels little sympathy for others 1.45 0.93 0.79 1.27 1.54 0.68 1.00
18 I am someone who is systematic, likes to keep things in order 1.06 0.88 1.09 1.16 1.37 0.94 0.81
19 I am someone who can be tense 1.27 0.75 1.08 1.22 1.20 0.95 0.95
20 I am someone who is fascinated by art, music, or literature 1.17 0.89 1.44 1.18 1.36 1.18 1.11
21 I am someone who is dominant, acts as a leader 1.35 0.74 1.15 1.13 1.49 0.70 0.61
22 I am someone who starts arguments with others 1.05 0.72 0.77 1.24 1.54 0.72 0.60
23 I am someone who has difficulty getting started on tasks 1.39 0.69 0.86 1.26 1.49 0.74 0.69
24 I am someone who feels secure, comfortable with self 1.27 0.88 1.08 1.42 1.49 0.96 0.92
25 I am someone who avoids intellectual, philosophical discussions 1.21 0.79 1.36 1.25 1.56 1.10 0.99
26 I am someone who is less active than other people 1.31 0.99 1.04 1.24 1.31 0.98 0.83
27 I am someone who has a forgiving nature 1.20 0.57 0.90 1.20 1.37 0.85 0.78
28 I am someone who can be somewhat careless 1.25 0.83 1.11 1.18 1.38 0.68 1.00
29 I am someone who is emotionally stable, not easily upset 1.27 0.70 1.01 1.31 1.47 1.06 1.09
30 I am someone who has little creativity 1.20 0.76 0.69 1.22 1.32 0.83 1.02
31 I am someone who is sometimes shy, introverted 1.36 0.87 1.11 1.07 1.30 1.04 0.99
32 I am someone who is helpful and unselfish with others 0.88 0.80 1.05 1.21 1.50 0.73 0.72
33 I am someone who keeps things neat and tidy 1.24 0.64 1.00 1.09 1.35 0.73 0.71
34 I am someone who worries a lot 1.47 0.70 1.08 1.23 1.33 1.18 1.07
35 I am someone who values art and beauty 1.07 0.82 1.30 1.18 1.35 1.00 1.06
36 I am someone who finds it hard to influence people 1.20 0.69 1.04 1.26 1.37 0.67 0.67
37 I am someone who is sometimes rude to others 1.25 0.66 0.91 1.31 1.49 0.77 1.00
38 I am someone who is efficient, gets things done 0.88 0.82 1.12 1.25 1.46 0.84 0.91
39 I am someone who often feels sad 1.43 0.68 0.98 1.27 1.25 0.99 0.91
40 I am someone who is complex, a deep thinker 1.08 0.72 1.06 1.06 1.26 0.83 0.99
41 I am someone who is full of energy 1.30 0.89 1.24 1.23 1.35 0.85 0.76
42 I am someone who is suspicious of others’ intentions 1.30 0.58 0.83 1.26 1.49 1.00 0.68
43 I am someone who is reliable, can always be counted on 0.88 0.93 1.09 1.32 1.59 0.80 0.78
44 I am someone who keeps their emotions under control 1.14 0.64 1.08 1.22 1.44 0.77 1.04
45 I am someone who has difficulty imagining things 1.06 0.77 0.53 1.25 1.27 0.95 0.54
46 I am someone who is talkative 1.38 0.70 1.11 1.04 1.28 1.09 0.92
47 I am someone who can be cold and uncaring 1.23 0.80 0.87 1.35 1.49 0.78 0.98
48 I am someone who leaves a mess, doesn’t clean up 1.12 0.71 0.95 1.26 1.64 0.55 0.85
49 I am someone who rarely feels anxious or afraid 1.37 0.83 1.12 1.15 1.37 1.02 0.89
50 I am someone who thinks poetry and plays are boring 1.37 0.75 1.31 1.38 1.53 0.87 0.68
51 I am someone who prefers to have others take charge 1.28 0.69 0.76 1.27 1.57 0.70 0.51
52 I am someone who is polite, courteous to others 0.70 0.65 0.97 1.14 1.44 0.75 0.79
53 I am someone who is persistent, works until the task is finished 0.91 0.87 1.13 1.26 1.48 0.80 0.94
54 I am someone who tends to feel depressed, blue 1.46 0.79 1.07 1.31 1.37 1.08 1.00
55 I am someone who has little interest in abstract ideas 1.22 0.96 1.08 1.22 1.36 1.17 0.63
56 I am someone who shows a lot of enthusiasm 1.22 0.85 1.09 1.34 1.51 0.83 0.94
57 I am someone who assumes the best about people 1.28 0.78 0.98 1.27 1.47 0.92 0.74
58 I am someone who sometimes behaves irresponsibly 1.30 0.74 1.13 1.14 1.40 0.70 0.95
59 I am someone who is temperamental, gets emotional easily 1.33 0.64 1.08 1.24 1.52 1.00 0.87
60 I am someone who is original, comes up with new ideas 1.07 0.65 1.13 1.22 1.41 0.68 0.75
Table 13: Item Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Facet Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Sociability 2.65 3.26 3.04 2.94 2.89 2.57 2.43
Assertiveness 2.97 3.22 3.33 3.06 2.88 2.81 2.61
Energy Level 3.35 3.85 3.88 3.09 2.95 2.79 2.52
Compassion 4.02 3.63 4.16 3.42 3.36 4.10 3.78
Respectfulness 4.29 3.39 4.39 3.49 3.63 4.07 4.26
Trust 3.31 3.15 3.78 3.21 3.32 3.49 3.64
Organization 3.86 3.21 3.24 3.18 2.94 3.22 3.10
Productiveness 3.86 3.54 3.85 3.19 3.23 3.60 3.34
Responsibility 3.96 3.25 3.52 3.03 3.13 3.59 3.60
Anxiety 3.15 3.15 2.96 3.06 2.89 3.42 3.10
Depression 2.51 2.76 2.20 2.86 2.59 2.83 2.61
Emotional Volatility 2.44 3.07 2.61 2.92 2.76 2.80 2.34
Intellectual Curiosity 4.05 3.55 3.51 3.13 3.09 3.08 2.63
Aesthetic Sensitivity 3.89 3.63 3.47 3.25 3.01 3.12 2.67
Creative Imagination 3.96 3.72 3.73 3.29 3.16 3.17 2.44
Table 14: Facet Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Facet Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Sociability 1.17 0.69 0.98 0.95 1.00 0.91 0.74
Assertiveness 1.04 0.55 0.82 1.10 1.22 0.57 0.42
Energy Level 0.95 0.75 0.88 1.12 1.19 0.77 0.64
Compassion 0.81 0.69 0.81 1.18 1.36 0.69 0.76
Respectfulness 0.70 0.58 0.74 1.13 1.33 0.61 0.68
Trust 1.00 0.44 0.67 1.17 1.34 0.77 0.62
Organization 1.01 0.57 0.79 0.98 1.18 0.63 0.68
Productiveness 0.91 0.65 0.87 1.12 1.29 0.68 0.68
Responsibility 0.84 0.59 0.95 1.10 1.26 0.64 0.74
Anxiety 1.11 0.54 0.89 0.95 1.04 0.88 0.75
Depression 1.14 0.56 0.89 1.16 1.26 0.87 0.81
Emotional Volatility 1.10 0.43 0.92 1.06 1.25 0.76 0.87
Intellectual Curiosity 0.83 0.68 0.96 1.04 1.24 0.89 0.70
Aesthetic Sensitivity 1.01 0.75 1.09 1.12 1.26 0.96 0.83
Creative Imagination 0.88 0.57 0.74 1.05 1.17 0.67 0.60
Table 15: Facet Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Domain Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion 2.99 3.44 3.42 3.03 2.90 2.72 2.52
Agreeableness 3.87 3.39 4.11 3.37 3.43 3.89 3.89
Conscientiousness 3.89 3.33 3.54 3.13 3.10 3.47 3.35
Neuroticism 2.70 2.99 2.59 2.94 2.75 3.02 2.68
Openness 3.97 3.63 3.57 3.22 3.09 3.12 2.58
Table 16: Domain Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Domain Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion 0.85 0.57 0.81 0.95 1.00 0.67 0.52
Agreeableness 0.69 0.51 0.70 1.11 1.28 0.62 0.62
Conscientiousness 0.80 0.53 0.79 1.00 1.15 0.59 0.64
Neuroticism 1.02 0.43 0.81 0.97 1.09 0.77 0.72
Openness 0.76 0.58 0.78 1.03 1.15 0.72 0.59
Table 17: Domain Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Additional Psychometric Performance Results
Here we show the TFM fit information in Table 18; FFM fit information in Table 19; TCC results for the TFM of each BFI-2 domain in Table 20; TCC results for the FFM in Table 21; specific standardized factor loading results in Tables 22 and 23; specific inter-factor correlation results in Tables 24 and 25.
Domain Model $\mathbf{\chi^{2}}$ df CFI TLI RMSEA SRMR
Extraversion Human 993.189 51 .892 .861 .109 .057
Persona
GPT-4o 261.099 51 .886 .852 .117 .058
Llama3 316.715 51 .885 .851 .132 .060
Shape
GPT-4o 663.947 51 .823 .771 .200 .101
Llama3 934.002 51 .684 .592 .241 .132
PSI
GPT-4o 329.341 51 .892 .860 .135 .057
Llama3 332.944 51 .810 .755 .136 .092
Agreeableness Human 904.640 51 .875 .838 .104 .060
Persona
GPT-4o 210.863 51 .900 .870 .102 .062
Llama3 616.199 51 .798 .739 .192 .088
Shape
GPT-4o 525.799 51 .909 .882 .176 .039
Llama3 874.914 51 .840 .793 .233 .070
PSI
GPT-4o 177.494 51 .952 .937 .091 .049
Llama3 255.684 51 .923 .900 .116 .050
Conscientiousness Human 1041.784 51 .897 .867 .112 .058
Persona
GPT-4o 385.339 51 .780 .716 .148 .088
Llama3 705.359 51 .768 .699 .207 .109
Shape
GPT-4o 705.340 51 .826 .775 .207 .069
Llama3 1502.901 51 .642 .536 .310 .196
PSI
GPT-4o 348.206 51 .887 .854 .139 .052
Llama3 427.908 51 .834 .785 .157 .069
Neuroticism Human 929.871 51 .931 .911 .105 .052
Persona
GPT-4o 304.611 51 .756 .684 .129 .091
Llama3 346.207 51 .884 .850 .139 .059
Shape
GPT-4o 960.437 51 .720 .638 .244 .099
Llama3 808.305 51 .772 .705 .224 .101
PSI
GPT-4o 448.933 51 .876 .840 .161 .069
Llama3 448.871 51 .854 .811 .161 .067
Openness Human 909.210 51 .899 .870 .104 .064
Persona
GPT-4o 226.649 51 .910 .883 .107 .057
Llama3 256.032 51 .904 .876 .116 .077
Shape
GPT-4o 502.104 51 .878 .842 .172 .055
Llama3 569.466 51 .861 .820 .185 .057
PSI
GPT-4o 310.479 51 .904 .876 .130 .071
Llama3 352.224 51 .854 .811 .140 .114
Table 18: Model Fits for BFI-2 Three-Factor Models of Each Domain
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Model $\mathbf{\chi^{2}}$ df CFI TLI RMSEA SRMR
Human 1747.433 80 .852 .806 .116 .080
Persona
GPT-4o 551.459 80 .819 .762 .140 .097
Llama3 954.848 80 .780 .711 .191 .120
Shape
GPT-4o 2132.208 80 .711 .621 .292 .143
Llama3 1988.584 80 .717 .629 .283 .149
PSI
GPT-4o 899.057 80 .769 .697 .185 .119
Llama3 802.352 80 .774 .703 .173 .114
Table 19: Model Fits for BFI-2 Five-Factor Model
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Domain Facet Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion Sociability 1.00 .99 1.00 .99 1.00 .99
Assertiveness .97 .98 .99 .97 .98 .97
Energy Level .98 .98 .97 .96 .98 .98
Agreeableness Compassion .95 .96 .97 .95 .97 1.00
Respectfulness 1.00 1.00 1.00 1.00 .98 .99
Trust .98 .99 1.00 .99 1.00 .99
Conscientiousness Organization .99 1.00 .99 .95 .99 1.00
Productiveness .99 .99 .99 .99 1.00 .97
Responsibility .99 .98 .99 .96 1.00 .99
Neuroticism Anxiety .99 .99 .98 .99 1.00 1.00
Depression .97 .99 .99 .99 1.00 1.00
Emotional Volatility .97 1.00 .99 .99 1.00 1.00
Openness Intellectual Curiosity 1.00 1.00 .99 .98 1.00 .98
Aesthetic Sensitivity .99 .98 .99 .98 .99 .97
Creative Imagination 1.00 .98 1.00 1.00 .99 .96
Table 20: Tucker’s Congruence Coefficient for BFI-2 Three-Factor Models of Each Domain
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Domain Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion 1.00 .99 1.00 1.00 1.00 1.00
Agreeableness 1.00 1.00 .99 .99 1.00 1.00
Conscientiousness .99 .98 .99 .99 .99 .99
Neuroticism .99 1.00 .99 1.00 1.00 1.00
Openness 1.00 .99 1.00 .99 1.00 .99
Table 21: Tucker’s Congruence Coefficient for BFI-2 Five-Factor Model
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Domain Facet=~Item Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion Sociability=~item1 .80 .77 .72 .85 .53 .87 .63
Sociability=~item16 .82 .79 .88 .74 .75 .84 .88
Sociability=~item31 .82 .79 .87 .73 .80 .91 .76
Sociability=~item46 .76 .74 .64 .85 .57 .76 .67
Assertiveness=~item6 .76 .66 .84 .94 .85 .81 .69
Assertiveness=~item21 .89 .61 .73 .90 .69 .80 .69
Assertiveness=~item36 .53 .73 .72 .84 .88 .76 .44
Assertiveness=~item51 .69 .59 .54 .80 .73 .55 .27
Energy Level=~item11 .45 .82 .73 .91 .89 .82 .66
Energy Level=~item26 .54 .69 .75 .72 .78 .74 .47
Energy Level=~item41 .83 .81 .87 .82 .76 .87 .71
Energy Level=~item56 .77 .82 .78 .88 .77 .80 .84
Agreeableness Compassion=~item2 .73 .76 .75 .90 .74 .84 .83
Compassion=~item17 .34 .84 .86 .93 .97 .79 .46
Compassion=~item32 .62 .74 .74 .87 .86 .80 .79
Compassion=~item47 .75 .67 .85 .96 .98 .89 .86
Respectfulness=~item7 .71 .82 .80 .93 .84 .90 .91
Respectfulness=~item22 .60 .54 .68 .80 .87 .43 .48
Respectfulness=~item37 .70 .74 .79 .94 .94 .80 .82
Respectfulness=~item52 .70 .70 .80 .83 .81 .85 .86
Trust=~item12 .69 .63 .35 .57 .95 .95 .78
Trust=~item27 .68 .60 .80 .84 .84 .79 .84
Trust=~item42 .67 .35 .63 .95 .95 .78 .83
Trust=~item57 .78 .64 .82 .90 .85 .88 .89
Conscientiousness Organization=~item3 .86 .64 .75 .75 .48 .83 .83
Organization=~item18 .70 .62 .66 .73 .90 .83 .65
Organization=~item33 .86 .63 .68 .95 .55 .74 .74
Organization=~item48 .74 .72 .72 .68 .55 .80 .79
Productiveness=~item8 .66 .67 .68 .64 .67 .73 .43
Productiveness=~item23 .86 .71 .65 .81 .70 .78 .43
Productiveness=~item38 .87 .72 .88 .82 .85 .85 .52
Productiveness=~item53 .65 .47 .66 .80 .63 .81 .82
Responsibility=~item13 .72 .67 .94 .92 .97 .79 .83
Responsibility=~item28 .77 .58 .58 .68 .46 .59 .71
Responsibility=~item43 .66 .51 .74 .75 .70 .68 .57
Responsibility=~item58 .80 .59 .80 .75 .87 .71 .74
Neuroticism Anxiety=~item4 .60 .71 .72 .68 .62 .80 .79
Anxiety=~item19 .63 .65 .61 .65 .72 .78 .43
Anxiety=~item34 .65 .47 .81 .73 .79 .59 .65
Anxiety=~item49 .74 .71 .84 .85 .70 .92 .89
Depression=~item9 .66 .67 .60 .81 .77 .88 .44
Depression=~item24 .86 .82 .59 .81 .78 .92 .91
Depression=~item39 .87 .86 .65 .47 .86 .95 .95
Depression=~item54 .65 .69 .88 .92 .95 .79 .85
Emotional Volatility=~item14 .72 .77 .72 .59 .88 .84 .86
Emotional Volatility=~item29 .77 .66 .51 .74 .75 .70 .68
Emotional Volatility=~item44 .66 .64 .27 .68 .75 .60 .67
Emotional Volatility=~item59 .80 .67 .72 .83 .79 .65 .85
Openness Intellectual Curiosity=~item10 .60 .90 .81 .84 .91 .92 .89
Intellectual Curiosity=~item25 .63 .91 .67 .88 .89 .95 .94
Intellectual Curiosity=~item40 .65 .80 .59 .70 .57 .85 .85
Intellectual Curiosity=~item55 .74 .77 .81 .84 .89 .88 .56
Aesthetic Sensitivity=~item5 .66 .66 .84 .85 .91 .71 .74
Aesthetic Sensitivity=~item20 .86 .77 .81 .65 .83 .79 .85
Aesthetic Sensitivity=~item35 .87 .59 .70 .84 .81 .92 .88
Aesthetic Sensitivity=~item50 .65 .63 .69 .41 .91 .63 .32
Creative Imagination=~item15 .72 .77 .40 .71 .81 .95 .79
Creative Imagination=~item30 .77 .80 .59 .70 .81 .86 .82
Creative Imagination=~item45 .66 .84 .68 .49 .77 .92 .35
Creative Imagination=~item60 .80 .86 .70 .95 .81 .84 .95
Table 22: Standardized Factor Loadings for BFI-2 Three-Factor Models of Each Domain with Human Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
Domain=~Facet Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
E=~Sociability .66 .76 .85 .91 .87 .81 .78
E=~Assertiveness .59 .67 .86 .77 .73 .71 .68
E=~Energy Level .77 .88 .86 .89 .85 .92 .89
A=~Compassion .70 .82 .90 .96 .99 .82 .79
A=~Respectfulness .80 .90 .94 .90 .85 .88 .94
A=~Trust .65 .75 .87 .93 .95 .84 .84
C=~Organization .70 .87 .87 .94 .97 .87 .86
C=~Productiveness .89 .77 .75 .84 .81 .84 .82
C=~Responsibility .79 .82 .95 .93 .92 .88 .91
N=~Anxiety .84 .69 .75 .80 .86 .77 .77
N=~Depression .88 .91 .96 1.02 .96 .97 .87
N=~Emotional Volatility .85 .63 .79 .71 .80 .78 .85
O=~Intellectual Curiosity .74 .77 .66 .97 .93 .74 .85
O=~Aesthetic Sensitivity .71 .73 .65 .95 .96 .69 .59
O=~Creative Imagination .81 .91 .89 .90 .82 .96 .76
Table 23: Standardized Factor Loadings for BFI-2 Five-Factor Model
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
Domain Facet~~Facet Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion Sociability~~Assertiveness .59 .72 .89 .77 .91 .69 .70
Sociability~~Energy Level .64 .79 .84 .91 .80 .84 .85
Assertiveness~~Energy Level .47 .66 .82 .65 .61 .72 .65
Agreeableness Compassion~~Respectfulness .81 .90 .98 .92 .87 .85 .90
Compassion~~Trust .70 .83 .94 .94 .96 .79 .84
Respectfulness~~Trust .58 .95 .96 .93 .84 .85 .86
Conscientiousness Organization~~Productiveness .75 .86 .81 .97 .70 .86 .86
Organization~~Responsibility .70 1.01 .85 .99 .86 .89 .87
Productiveness~~Responsibility .86 .82 .87 .87 .89 .82 .87
Neuroticism Anxiety~~Depression .81 .88 .88 .95 .96 .86 .79
Anxiety~~Emotional Volatility .90 .87 .79 .87 .87 .92 .83
Depression~~Emotional Volatility .79 .84 .85 .89 .83 .86 .82
Openness Intellectual Curiosity~~Aesthetic Sensitivity .66 .68 .59 1.02 1.00 .51 .51
Intellectual Curiosity~~Creative Imagination .74 .85 .67 .99 .89 .78 .72
Aesthetic Sensitivity~~Creative Imagination .63 .79 .60 .93 .86 .77 .50
Table 24: Inter-factor Correlations for BFI-2 Three-Factor Models of Each Domain with Human Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
Domain~~Domain Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Extraversion~~Agreeableness .28 .26 .10 .59 .54 .41 .24
Extraversion~~Conscientiousness .56 .43 .26 .36 .31 .44 .34
Extraversion~~Neuroticism -.61 -.61 -.59 -.80 -.72 -.62 -.31
Extraversion~~Openness .35 .37 .51 .72 .71 .29 .48
Agreeableness~~Conscientiousness .47 .56 .77 .65 .65 .59 .63
Agreeableness~~Neuroticism -.43 -.40 -.54 -.66 -.73 -.44 -.47
Agreeableness~~Openness .27 .31 .25 .79 .89 .14 .14
Conscientiousness~~Neuroticism -.61 -.52 -.63 -.67 -.82 -.58 -.60
Conscientiousness~~Openness .17 .10 .16 .27 .45 .06 .18
Neuroticism~~Openness -.17 -.06 -.23 -.48 -.65 .14 .18
Table 25: Inter-factor Correlations for BFI-2 Five-Factor Models with Human Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
Scale Reliability: Facet level and domain level Cronbach’s alpha for different method LLM responses on BFI-2 and human responses are shown in Table 26.
It can be observed that the PSI method, compared to the Persona and Shape methods, performs closer to the results of the human sample in terms of Cronbach‘s alpha (with the number of data marked in italics and bold being the smallest). This indicates that the LLM personality data generated by the PSI method holds an advantage in consistency and reliability, enabling it to more accurately simulate the statistical characteristics of human samples.
Facet / Domain Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
Sociability .87 .85 .86 .87 .77 .91 .82
Assertiveness .81 .74 .80 .93 .86 .82 .57
Energy Level .73 .86 .86 .90 .87 .88 .76
Compassion .67 .84 .87 .95 .93 .90 .82
Respectfulness .74 .79 .85 .94 .93 .84 .86
Trust .80 .64 .79 .96 .94 .87 .88
Organization .87 .73 .80 .86 .82 .86 .82
Productiveness .80 .80 .85 .91 .90 .88 .76
Responsibility .77 .67 .88 .91 .86 .88 .86
Anxiety .84 .65 .81 .81 .79 .88 .78
Depression .86 .66 .88 .90 .91 .91 .88
Emotional Volatility .89 .60 .89 .87 .88 .89 .91
Intellectual Curiosity .75 .82 .85 .87 .89 .89 .79
Aesthetic Sensitivity .83 .87 .83 .92 .93 .87 .84
Creative Imagination .82 .80 .82 .88 .89 .88 .77
Extraversion .87 .90 .93 .95 .92 .93 .87
Agreeableness .85 .90 .94 .98 .97 .93 .93
Conscientiousness .90 .88 .93 .95 .94 .94 .92
Neuroticism .94 .83 .93 .94 .94 .95 .93
Openness .89 .91 .89 .96 .96 .92 .88
Table 26: Cronbach’s alpha for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
Discriminant Validity: The results for discriminant validity are shown in Table 27. We can further observe that, compared to the Persona and Shape methods, the PSI method demonstrates the closest performance to human samples in terms of discriminant validity. Specifically, the PSI method shows the closest mean of absolute values when examining its correlations with human samples.
Both higher and lower levels of external validity reveal the degree of differences between the methods and human samples. Higher external validity indicates that the Big Five factors in human samples are more distinctly differentiated from one another, while lower external validity suggests the opposite. Therefore, our focus here is on identifying the approach that most closely aligns with the performance of human samples.
Dimensions Human Persona Shape PSI
GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
E, A .17 .21 .12 .48 .40 .33 .21
E, C .35 .40 .33 .46 .41 .38 .30
E, N -.45 -.43 -.42 -.58 -.57 -.43 -.21
E, O .22 .23 .28 .66 .68 .13 .33
A, C .33 .47 .66 .67 .70 .49 .52
A, N -.36 -.30 -.54 -.64 -.71 -.29 -.37
A, O .20 .30 .31 .71 .78 .14 .14
C, N -.49 -.43 -.60 -.74 -.87 -.47 -.52
C, O .09 .09 .20 .29 .48 .02 .08
N, O -.11 .03 -.08 -.31 -.57 .32 .23
Mean of Absolute Values .28 .29 .35 .55 .62 .30 .29
Table 27: Domain Level Correlation Analysis for BFI-2 Human Responses and Different Methods LLM Responses
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis. E = Extraversion; A = Agreeableness; C = Conscientiousness; N = Neuroticism; O = Openness.
I.2 Additional Personality-Related Behavioral Performance Results
Table 28 presents the correlations between the personality dimensions and OCB/CWB reported by PSI GPT-4o, with a comparison to human self-reported data. The relevant results for PSI Llama3 are provided in Table 29, while Figures 3 and Figure 4 display the complete correlation matrices for GPT-4o and Llama3, respectively.
Domain OCB CWB
Human PSI Human PSI
Ext .42 .54 -.01 -.08
Agr .18 .36 -.30 -.54
Con .12 .45 -.35 -.47
Neu -.21 -.35 .23 .36
Ope .17 .02 -.17 .01
Table 28: Comparison of OCB and CWB Correlations with Personality Domains: Human vs. PSI GPT-4o
Note: n = 357 for human and PSI method.
Domain OCB CWB
Human PSI Human PSI
Ext .42 .48 -.01 .05
Agr .18 .46 -.30 -.45
Con .12 .47 -.35 -.41
Neu -.21 -.38 .23 .33
Ope .17 .11 -.17 .00
Table 29: Comparison of OCB and CWB Correlations with Personality Domains: Human vs. PSI Llama3
Note: n = 357 for human and PSI method.
Figure 3: Complete Correlation Matrices for Human and PSI GPT-4o
Note: n = 357 for human and PSI method.
Figure 4: Complete Correlation Matrices for Human and PSI Llama3
Note: n = 357 for human and PSI method.
I.3 The Influence of Social Desirability
In previous studies, researchers have observed some intriguing phenomena, specifically that LLMs tend to exhibit human-like social desirability bias when simulating human samples. This bias is reflected in the LLMs’ responses, which lean toward behaviors and traits that are socially approved.
Numerous studies support this finding. For example, Hilliard et al. (2024) noted that newer and larger-parameter LLMs display more diverse personality traits, including higher levels of agreeableness, emotional stability, and openness. Similarly, Salecha et al. (2024) found that LLMs exhibit human-like social desirability bias when generating simulated data.
In this section, using the data we have collected, we directly compare the results from human samples and LLM-simulated data. We examined the self-reported personality traits of human samples alongside the simulated personality traits from LLMs, analyzing how both were influenced by social desirability ratings. Specifically, we used social desirability ratings to predict and test the correlations and differences between the mean scores of human and LLM data across various items.
By analyzing the predictive relationships and correlations between item means and social desirability ratings, we aim to uncover how social desirability influences the characteristics and patterns in both human and LLM data. Furthermore, by comparing the differences in item means between human data and LLM-simulated data, we explore whether social desirability ratings affect human samples and model-simulated data in similar ways.
Social Desirability Rating
Social desirability ratings of the BFI-2 items were obtained from another ongoing study where 142 human resource practitioners were asked to rate how desirable each item was in general (1 = “Very undesirable,” 2 = “Undesirable,” 3 = “Slightly undesirable,” 4 = “Neither desirable nor undesirable,” 5 = “Slightly desirable,” 6 = “Desirable,” 7 = “Very desirable”).
If a certain measurement item receives a high score (6 or 7), it indicates that the trait is generally viewed as desirable or positive in society. On the other hand, a low score (1 or 2) suggests that the trait may be considered less ideal or inconsistent with social expectations. Some desirable traits may lead participants to lean toward choosing higher scores (due to social desirability bias) rather than responding based on their true circumstances.
Results
Table 30 presents the regression results predicting the self-reported personality in the human sample, the LLM-simulated personality, and the difference between human and LLM personalities, all based on the social desirability rating. Additionally, Figure 5 visually illustrates the trends through a regression plot.
It can be observed that when an item’s social desirability rating is neutral (4), most LLMs tend to generate neutral responses (rated 3 on the scale). This result indicates that LLMs have learned to associate neutral social expectations with neutral opinions, consistent with previous research findings (Salecha et al., 2024). Regarding the mean difference between the human responses and the LLM responses, most differences show a positive correlation with the social desirability rating, and when the social desirability rating is neutral (4), the mean difference in most data is approximately 0.
In Figure 5, the line graph clearly illustrates that as the social desirability rating of the item increases, the average scores of both human responses and LLM responses rise accordingly. Moreover, as the social desirability rating increases, the average difference between human and LLM responses gradually widens, whereas this difference narrows when the social desirability scores are lower.
This trend indicates that when social desirability scores are high, the mean score of human responses tends to be higher than that of LLM responses. Conversely, when social desirability scores are low, the average score of human responses falls below that of LLM responses. This suggests that human responses are more strongly influenced by social desirability than those of LLMs.
Variable Human Persona GPT-4o Shape GPT-4o PSI GPT-4o Persona Llama3 Shape Llama3 PSI Llama3
Intercept 1.15 (0.23) 1.90 (0.17) 2.33 (0.09) 2.13 (0.23) 0.71 (0.21) 2.25 (0.11) 2.08 (0.28)
Social desirability 0.46 (0.05) 0.23 (0.04) 0.12 (0.02) 0.23 (0.05) 0.47 (0.05) 0.14 (0.02) 0.20 (0.06)
$R^{2}$59 .39 .38 .27 .65 .37 .15
Predicted score at neutral point 2.99 2.82 2.81 3.05 2.59 2.81 2.88
Correlation .77 .63 .62 .52 .80 .61 .39
Mean Score Difference
Intercept - -0.76 (0.19) -1.18 (0.19) -0.99 (0.20) 0.44 (0.21) -1.11 (0.23) -0.93 (0.32)
Social desirability - 0.23 (0.04) 0.34 (0.04) 0.23 (0.04) -0.01 (0.05) 0.33 (0.05) 0.26 (0.07)
$R^{2}$34 .53 .32 .48 .42 .19
Predicted score at neutral point - 0.16 0.18 -0.07 0.40 0.21 0.11
Correlation - .58 .73 .56 -.03 .65 .43
Table 30: Regression Analysis and Correlation of Social Desirability Ratings
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
Figure 5: Regression Line Chart of Social Desirability Ratings
Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses. Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.

### v2

@@
-Personality Structured Interview for Large Language Model Simulation in Personality Research
+Generative Personality Simulation via Theory-Informed Structured Interview
 Abstract
-Although psychometrics researchers have recently explored the use of large language models (LLMs) as proxies for human participants, LLMs often fail to generate heterogeneous data with human-like diversity, which diminishes their value in advancing social science research.
-To address these challenges, we explored the potential of the theory-informed Personality Structured Interview (PSI) as a tool for simulating human responses in personality research.
-In this approach, the simulation is grounded in nuanced real-human interview transcripts that target the personality construct of interest.
-We have provided a growing set of 357 structured interview transcripts from a representative sample [^1] Dataset and code will be shared later., each containing an individual’s response to 32 open-ended questions carefully designed to gather theory-based personality evidence.
-Additionally, grounded in psychometric research, we have summarized an evaluation framework to systematically validate LLM-generated psychometric data.
-Results from three experiments demonstrate that well-designed structured interviews could improve human-like heterogeneity in LLM-simulated personality data and predict personality-related behavioral outcomes (i.e., organizational citizenship behaviors and counterproductive work behavior).
-We further discuss the role of theory-informed structured interviews in LLM-based simulation and outline a general framework for designing structured interviews to simulate human-like data for psychometric research.
-Personality Structured Interview for Large Language Model Simulation in Personality Research
-Pengda Wang1, Huiqi Zou3, Hanjie Chen2, Tianjun Sun1, Ziang Xiao3, and Frederick L.
-Oswald1 1Department of Psychological Sciences, Rice University 2Department of Computer Science, Rice University 3Department of Computer Science, Johns Hopkins University {pw32,hc86,ts110,foswald}@rice.edu; {hzou11,ziang.xiao}@jhu.edu
+Despite their potential as human proxies, LLMs often fail to generate heterogeneous data with human-like diversity, thereby diminishing their value in advancing social science research.
+To address this gap, we propose a novel method to incorporate psychological insights into LLM simulation through the Personality Structured Interview (PSI [^1] Code and data: https://github.com/isle-dev/PSI.).
+PSI leverages psychometric scale-development procedures to capture personality-related linguistic information from a formal psychological perspective.
+To systematically evaluate simulation fidelity, we developed a measurement theory grounded evaluation procedure that considers the latent construct nature of personality and evaluates its reliability, structural validity, and external validity.
+Results from three experiments demonstrate that PSI effectively improves human-like heterogeneity in LLM-simulated personality data and predicts personality-related behavioral outcomes.
+We further offer a theoretical framework for designing theory-informed structured interviews to enhance the reliability and effectiveness of LLMs in simulating human-like data for broader psychometric research.
+Generative Personality Simulation via Theory-Informed Structured Interview
+Pengda Wang1, Huiqi Zou3, Han Jiang4, Hanjie Chen2, Tianjun Sun1, Xiaoyuan Yi5, Ziang Xiao4, and Frederick L.
+Oswald1 1Department of Psychological Sciences & 2Department of Computer Science, Rice University 3Department of Electrical and Computer Engineering, Northeastern University 4Department of Computer Science, Johns Hopkins University 5Microsoft Research Asia {pw32,ts110,fo3}@rice.edu; {ziang.xiao}@jhu.edu
 **footnotetext: Co-corresponding authors.
 1 Introduction
-Personality research is an important field in psychology for understanding how individual traits shape significant life outcomes and trajectories, including career success, mental health, and overall well-being (e.g., Judge et al., 2002; Roberts et al., 2007; Robins et al., 2002).
-For instance, measures of conscientiousness have been consistently and positively correlated with job performance and academic success (Barrick and Mount, 1991), whereas measures of neuroticism are found to be strongly and positively correlated with mental health issues, such as anxiety and depression (John et al., 2008).
-However, studying personality requires understanding habits of thoughts, attitudes, and behaviors across diverse contexts and cultures, which is often a research challenge.
-For example, some research questions involve rare or extreme cases, making it difficult to gather sufficient samples (e.g., Lynam and Widiger, 2001).
-Others require longitudinal designs to track personality development or dynamic social interactions (e.g., Damian et al., 2019; Roberts et al., 2006).
-Additionally, certain experimental manipulations, like inducing stress or altering life circumstances, may be feasible or unethical for causal insights Fisher (2021).
-If a method could accurately simulate the human personality distributions, they could supplement and accelerate personality research by offering a scalable, cost-effective approach to supplement traditional data collection and experimentation (Messeri and Crockett, 2024).
-Several prior studies have supported this possibility, showing that LLMs can generate responses for personality scales that reflect personality traits resembling those of humans (e.g., Lee et al.
-2024, Huang et al.
-2023).
-However, studies found such methods face significant limitations in capturing individual differences at the item level (e.g., Wang et al.
-2024a).
-Information at the item level provides a more granular understanding of how traits manifest in specific individuals and situations rather than relying solely on broad trait averages.
-However, without additional empirical information, items confound these sources of variance.
-Moreover, key psychometric challenges remain in modeling LLM responses, as they may not accurately reflect the intended personality construct.
-To address this gap, we explored the potential of the Personality Structured Interview (PSI), which employs theory-informed questions to elicit information embedded in human narrative responses that are relevant to the target personality constructs of interest in LLM simulations.
-We believe this sophisticated linguistic approach can inform and enhance the measurement of human-like traits in heterogeneous LLM-simulated personality data, potentially leading to a greater understanding of personality.
-Grounded in psychometric research (e.g., Cronbach, 1951; Cronbach and Meehl, 1955), we have summarized a framework for evaluating psychometric data that fully considers its hierarchical structure, in which observed responses (item level) are mapped to latent traits (domain level).
-This evaluation framework encompasses a range of analyses, from overall descriptive statistics to a more in-depth examination of psychometric performance.
-Based on our evaluation framework, we conducted three experiments.
-Those three experiments evaluated our method’s effectiveness in replicating individual personality traits, simulating human-like personality distribution, and demonstrating personality-related behaviors with respect to established personality theory.
-We found our PSI method improves human-like diversity in personality traits simulation and recovers nuances in personality-related behaviors.
-Figure 1: Overview of the Development of the Personality Structured Interview and Experimental Implementation in This Study
-Note: For an overview of the personality structured interview development process, refer to §3 and Appendix B.
-The structured interview development framework is detailed in Appendix C.
-Information on data collection and the human sample is provided in §4 and Appendix D.
-The LLM simulation process is described in §4 and Appendix F.
-Details on the experimental setup and results of the three experiments can be found in §4, §5, and Appendix I.
-For the psychometric data evaluation framework, see Appendix H.
-In summary, our paper offers three contributions:
-  • A theory-informed LLM-based simulation method (PSI) for personality research and a development framework (see Appendix C).
-  • A growing dataset, now with 357 structured interview transcripts.
-  • An evaluation framework for LLM-simulated psychometric data that is grounded in psychological theories (see Appendix H).
-We further discuss the potential of PSI methods in advancing research (see Appendix C).
-We believe the framework behind PSI could be generalized to other psychological constructs and other psychometric data to develop theory-informed simulation methods for human understanding.
-2 Personality Research and Related Works
-Here we provide a concise overview of key personality concepts and discuss related works.
-MacKinnon (1944) proposed two definitions of personality.
+The discipline of personality psychology seeks to understand how individual differences shape significant life outcomes and trajectories.
+Decades of empirical work have demonstrated that personality traits predict a wide range of significant domains, including career success, mental and physical health, interpersonal relationships, and overall well-being (e.g., Judge et al., 2002; Roberts et al., 2007; Robins et al., 2002).
+Therefore, personality assessment has often been used to inform personnel selection and clinical interventions.
+Moreover, the relevance of personality research is rapidly expanding into the field of Artificial Intelligence (AI), where it informs the design of more adaptive, human-centered, and personalized systems.
+Figure 1: Overview of the development of the personality structured interview and experimental implementation.
+However, studying personality requires large amounts of psychological data, such as personality trait scores and behavioral indicators.
+These data collection demands pose substantial logistical and financial challenges for researchers.
+Moreover, certain types of data are particularly difficult to obtain.
+For example, studies focusing on individuals with elevated levels of psychopathy or narcissism often struggle to obtain adequate sample sizes due to the rarity and sensitivity of these traits (e.g., Lynam and Widiger, 2001).
+Other research questions require longitudinal designs to capture personality development or dynamic patterns of social interaction over time, further compounding the complexity and resource demands (e.g., Damian et al., 2019; Roberts et al., 2006).
+If a method could accurately simulate human personality distributions, it would accelerate personality research by offering a more scalable and cost-effective way to support data collection and experimentation (Messeri and Crockett, 2024).
+In this context, the current work investigates the potential of LLMs to simulate human-like responses in psychometric research.
+Although prior studies have shown such promise, indicating that LLMs can generate responses for personality scales that reflect personality traits resembling those of humans (e.g., Huang et al., 2023, Lee et al., 2024), those approaches exhibit notable limitations, particularly in their ability to capture individual differences at the item-level and to reproduce the heterogeneity observed in human responses (e.g., Wang et al., 2024).
+Item-level data provides a more granular understanding of how traits manifest in specific individuals rather than relying solely on broad trait averages.
+Furthermore, modeling response heterogeneity is crucial for reflecting the variability and complexity of real-world human behavior.
+Additionally, current evaluation practices are often misaligned with the standards of psychological research.
+For example, prior research has predominantly assessed personality simulation data at the aggregated level, neglecting item- and facet-level analyses as well as the psychometric properties of the data.
+The evaluation gap undermines psychologists’ confidence in the validity of such methods.
+To address these research gaps, we investigated the use of Personality Structured Interview (PSI) transcripts as a means of guiding LLMs in simulating personality data (see Figure 1).
+The PSI comprises theory-driven questions paired with responses, designed to elicit narrative-based information pertinent to the personality constructs targeted in LLM simulations.
+This theory-informed approach enables the expression of personality in a manner that enhances the human-like heterogeneity and representativeness of the simulated data.
+To bridge the evaluation gap, we developed a comprehensive framework for assessing personality simulation data grounded in measurement theory (e.g., Cronbach, 1951; Cronbach and Meehl, 1955).
+This evaluation framework explicitly accounts for the hierarchical structure of personality measurement, wherein observed responses at the item-level map onto latent traits at the domain-level.
+It further supports a wide range of analytic strategies, from basic descriptive statistics to more advanced assessments of psychometric validity and reliability.
+With this framework, we conducted three experiments to evaluate the efficacy of the PSI method in: (1) replicating individual-level personality data, (2) simulating distributions of personality approximate those observed in human populations, and (3) capturing personality-related behaviors aligned with established findings.
+Our findings demonstrate that the PSI approach enhances human-like variability of simulated personality data and effectively captures the patterns of personality-relevant behavior.
+In summary, our paper makes three key contributions: (1) a theory-informed LLM-based simulation method, PSI, for personality research, along with a framework for developing such interview protocols; (2) the release of a dataset comprising 357 structured interview transcripts; and (3) an evaluation framework for LLM-simulated psychometric data, grounded in measurement theory.
+We further discuss the potential of the PSI method and the dataset to advance AI and psychology research (see Appendix G).
+The interview question development framework underlying PSI can be generalized to capture information relevant to other constructs (e.g., value), thereby facilitating the development of more theory-aligned data simulations and enhancing the alignment of LLMs with humans.
+2 Related Works
+Personality Definition and Construct
+MacKinnon (1944) proposed two complementary definitions of personality.
 One emphasizes internal factors like temperament and interpersonal strategies that drive consistent behavior across time, situations, and cultures.
@@
 The former highlights internal drives, while the latter centers on external behaviors and social perception.
-Together, they underscore personality’s role in shaping thought patterns and behaviors in social interactions (Hogan et al., 1996).
-Personality encodes rich and complex information in language and text (Goldberg, 1990; Saucier and Goldberg, 2001).
-In fact, the Five-Factor Model (FFM) of personality is extensively researched (e.g., Costa and McCrae, 2008; John, 1999; McCrae and Costa Jr, 1997); it is directly based on the lexical hypothesis, which posits that individual differences that are important in human interactions (e.g., have survival value across cultures) are often encoded in some or all languages of the world.
+Together, these perspectives underscore personality’s role in shaping thought patterns and behaviors in social interactions (e.g., Hogan et al., 1996).
+Personality encodes rich and complex information in language and text (e.g., Goldberg, 1990; Saucier and Goldberg, 2001).
+The Five Factor Model (FFM) of personality is extensively researched (e.g., Costa and McCrae, 2008; John, 1999; McCrae and Costa Jr, 1997); it is based on the lexical hypothesis, which posits that individual differences that are important in human interactions (e.g., have survival value across cultures) tend to become encoded in language.
 The five factors are Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism (OCEAN).
-From Galton’s (1884) early research to Allport and Odbert’s (1936) systematic organization and factor analysis of lexical terms, and through further developments by Norman (1963) and Goldberg (1990), the FFM theory has progressively evolved (see Appendix A for further details on the structure of personality).
-In other words, because language and text contain rich and complex information about personality, LLMs may capture and model such encoding by learning from vast amounts of training data.
-There are many studies to-date on LLMs that focus on personality, where researchers aim to benchmark their psychological profiles (Lee et al., 2024; Li et al., 2024; Pellert et al., 2023), investigate their ability to simulate personality (Huang et al., 2023; Jiang et al., 2023; Serapio-García et al., 2023), and examine whether LLM agents align with character personality settings and how these settings influence their generated dialogues (Shao et al., 2023; Wang et al., 2024b; Xu et al., 2024).
-Currently, common methods for LLM simulation of human personality distributions have several limitations.
-The Persona-Chat dataset (Persona method) by Zhang et al.
-(2018) was originally designed to enhance the engagement of chit-chat models by increasing personalization.
-Therefore, its focus is more on enhancing personalization, rather than capturing personality traits.
-The adjective-based categorization method (Shape method) proposed by Serapio-García et al.
-(2023) effectively simulates profile-specific patterns of high/low standing on personality dimensions.
-However, this method presents a challenge in simulating human personality distributions, as restricting each case to a single dimension fails to capture the inherently multidimensional and hierarchical nature of personality data (e.g., Kachur et al., 2020).
-To address the aforementioned limitations of personality-simulated data, we explore the use of theory-informed PSI transcription to simulate personality data.
-3 Personality Structured Interview
-The theory-informed structured interview mentioned in this study differs from previous interview approaches used in LLM-based simulations, such as the Life Interview  (Park et al., 2024), which broadly inquires about various aspects of an individual’s life.
-In contrast, the theory-informed structured interview is specifically designed to target the personality constructs that we aim to model.
-A theory-informed approach is essential not only for improving the accuracy of data simulation but also for enhancing interpretability.
-Our goal is to generate data that not only aligns with actual data at the personality scale level but also exhibits theoretically coherent personality-related behaviors.
-By grounding LLM-generated data in theory, we ensure that the simulation goes beyond reproducing statistical properties to capturing meaningful psychological patterns.
-This enhances interpretability, facilitates transparent comparisons with human responses, and preserves the psychological validity of the constructs.
-Learning from psychometric theories, we summarized a framework for developing structured interviews, which can serve as a template for creating interviews designed to collect targeted information for simulating specific types of psychometric data (see Appendix C).
-We present a PSI example designed to capture personality-related information, with relevant personality theories and question design detailed in §3.1.
-We want to further emphasize that the “structured interview questions” mentioned here differ from those described by Wang et al.
-(2024b); Ran et al.
-(2024), who refers to converting personality scale items into open-ended questions.
-The main issue with their approach is that, while it attempts to rewrite items as open-ended questions to elicit richer responses, it may fail to achieve the intended effect.
-For instance, in their example, “Values artistic, aesthetic experiences” is rewritten as “Do you value artistic, aesthetic experiences?” Although this question appears open-ended, it tends to prompt a simple “yes” or “no” response and function as if it is close-ended, limiting the exploration of the respondent’s deeper thoughts or reasoning (Trull et al., 1998).
-Moreover, such questions differ from the behavioral or situational questions typically found in structured interviews.
-Structured interviews usually involve more specific scenarios or tasks that encourage respondents to describe past experiences or hypothetical reactions (Campion et al., 1997).
-For example, to explore a respondent’s views on artistic and aesthetic experiences, a more effective question might be, “Please describe a moment when you felt deeply inspired by an artistic or aesthetic experience?” This encourages detailed, contextual responses beyond a simple “yes” or “no.”
-3.1 PSI Questions Design
-How can we design questions more effectively and obtain textual information with greater depth in measuring personality?
-McAdams’ three levels of personality theory and the corresponding, theory-informed interviews offer a valuable direction (e.g., McAdams, 1995, 1996, 2001).
-According to McAdams’s theory, personality can be divided into three levels: traits (broad personality characteristics), personal concerns (goals, values, strategies), and narrative identity (the story individuals tell themselves about their lives).
-The latter two levels, in particular, reveal how individuals act in specific situations and developmental stages and how they organize and make sense of their life experiences.
-The key to obtaining more meaningful textual information lies in guiding individuals toward higher levels of self-expression.
-Rather than simply recording mundane daily events, we should encourage reflection on pivotal life moments, significant relationships, and future aspirations.
-For example, effective questions could be: “Can you describe an event that changed the trajectory of your life?” or “Tell me about a moment you are most proud of.” By designing questions like these, we can elicit deeper, more meaningful narratives, enriching the personality information.
-We adapted and modified McAdams (1995, 1996, 2001) theory-based interview and narrative identity approach, incorporating elements from the structured interview of the FFM (SIFFM; Trull et al., 1998).
-The initial draft of the question pool was prepared and revised by subject matter experts (SMEs) through discussion.
-The SMEs included two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology.
-Subsequently, we conducted pilot testing with six undergraduate research assistants from a personality research lab.
-As a result, the PSI, comprising 32 questions, was developed, as shown in Table 3 in Appendix B.
-3.2 PSI Dataset
-Using this set of questions, we designed a chatbot to conduct structured interviews, facilitating data collection (Institutional Review Board approval was obtained).
-Because data collection is ongoing, we used responses collected through the end of December 2024.
-After excluding incomplete responses and those failing attention checks, the final sample consisted of 357 participants.
-On average, participants were in their early 30s (M = 33.30 years, SD = 13.06), with 52.40% identifying as men, and 44.80% as women.
-To ensure a demographically diverse and broadly representative sample, we recruited participants from both undergraduate and working adult populations, ensuring a diverse range of backgrounds in terms of educational, employment, and life experience.
-Undergraduate participants were recruited from a large public university in the Midwest and received research credit for their participation.
-Working adult participants were sourced from two widely recognized, high-quality crowdsourcing platforms—Prolific [^2] https://www.prolific.com/ and CloudResearch Connect [^3] https://connect.cloudresearch.com/—which are known for their diverse and demographically representative participant pools.
+This theoretical foundation has evolved over decades, from Galton (1884)’s early work on trait descriptors, through Allport and Odbert (1936)’s lexical studies, and into more systematic factor-analytic approaches by scholars such as Norman (1963) and Goldberg (1990).
+These developments have shaped our modern understanding of personality structure (see Appendix F for further details).
+Given the linguistic basis of personality, LLMs trained on vast and diverse text corpora are well-equipped to model personality data from natural language use.
+Generative Personality Simulation
+Current approaches to simulating personality data with LLMs exhibit several notable limitations.
+The Persona-Chat dataset, introduced by Zhang et al.
+(2018), was originally designed to enhance chit-chat models through increased personalization.
+As such, its primary goal is to foster conversational engagement rather than to accurately model underlying personality constructs.
+Another approach, the adjective-based categorization method (Shape method) proposed by Serapio-García et al.
+(2023), enables the simulation of profile-specific patterns by targeting high or low expressions of individual personality dimensions.
+However, this method struggles to realistically capture personality distributions, as it constrains each case to a single trait dimension—oversimplifying the inherently multidimensional and hierarchical nature of personality structure (e.g., Kachur et al., 2020).
+To address these limitations, we explore the use of PSI as a means to generate more nuanced, ecologically valid, and representative simulations of personality data.
+3 Personality Structured Interview for LLM Simulation
+The PSI differs from earlier interview datasets, such as the Life Interview transcripts (Park et al., 2024), which ask broad, general questions about a person’s life.
+In contrast, the PSI is purposefully designed to elicit linguistic cues associated with personality constructs.
+Based on psychometric scale development methods, we developed a structured framework to guide the formulation of interview questions (see Appendix G).
+Simply rephrasing scale items into open-ended questions often fails to elicit meaningful elaboration.
+Converting an item ‘‘Values artistic, aesthetic experiences’’ into ‘‘Do you value artistic, aesthetic experiences?’’ typically yields a brief ‘‘yes’’ or ‘‘no’’ answer, offering little insight (Trull et al., 1998).
+Such questions also diverge from the behavioral prompts commonly used in structured interviews, which are more effective for eliciting rich, narrative data (e.g., Campion et al., 1997).
+To explore aesthetic values, asking ‘‘Describe a moment when you felt inspired by an artistic or aesthetic experience’’ invites deeper, more contextual responses.
+3.1 Personality Structured Interview Development Process
+The overall process of developing a structured interview designed to assess the target construct is detailed in Appendix G.
+Here, let $\mathcal{C}$ denote the personality construct of interest.
+We present a complete example, formulated as a functional chain, to illustrate how the final set of structured interview questions, $\mathcal{Q}_{\text{final}}$, is produced: $\mathcal{C}\xrightarrow{f_{1}}\mathcal{I}\xrightarrow{f_{2}}\mathcal{B}\xrightarrow{f_{3}}\mathcal{Q}_{1}\xrightarrow{f_{4}}\mathcal{Q}_{2}^{(0)}\xrightarrow{f_{5}~(\text{iterative})}(\mathcal{R},\mathcal{V})\xrightarrow{f_{6}}\mathcal{Q}_{\text{final}}$.
+In the original steps, we first need to identify the behavior/perception indicators $\mathcal{I}$ demonstrated by $\mathcal{C}$ ($\mathcal{C}\xrightarrow{f_{1}}\mathcal{I}$), and based on $\mathcal{I}$ to build the initial blueprint $\mathcal{B}$ for the structured interview questions ($\mathcal{I}\xrightarrow{f_{2}}\mathcal{B}$).
+Then, based on $\mathcal{B}$, we can generate the initial question pool $\mathcal{Q}_{1}$ ($\mathcal{B}\xrightarrow{f_{3}}\mathcal{Q}_{1}$).
+However, given the substantial body of research on personality assessment and constructs—such as narrative identity theory (McAdams, 1995, 1996, 2001) and the Structured Interview for the Five-Factor Model (SIFFM; Trull et al., 1998) [^2] McAdams’ three levels of personality (traits, personal concerns, and narrative identity) as well as other established personality structured interviews, highlight the importance of thoughtfully guiding individuals toward deeper self-expression when designing such interviews.
+Instead of focusing on routine events or superficial details, we should prompt reflection on pivotal moments, meaningful relationships, and future aspirations.
+For example, questions like ‘‘Can you describe an event that changed the trajectory of your life?’’ or ‘‘Tell me about a moment you are most proud of.’’ can elicit rich, narrative-driven insights into personality.—four subject matter experts ($\text{SMEs}=\{\text{Ph.D.}_{student1},\text{Ph.D.}_{student2},\text{Postdoc},\text{Professor}\}$), all specializing in Personality Psychology, adapted and modified these existing theory and interview questions accordingly to generate the initial question pool: $\mathcal{Q}_{1}$ ($\mathcal{Q}_{1}=\text{adapt}(\mathcal{Q}_{\text{prior}},\text{SMEs})$).
+Following this, we conducted pilot testing with six undergraduate research assistants affiliated with a personality research laboratory ($\mathcal{D}_{\text{pilot}}$).
+Based on their feedback, the interview questions were further revised by SMEs ($\mathcal{Q}_{2}=f_{4}(\mathcal{Q}_{1},\text{Stat}(\mathcal{D}_{\text{pilot}}))$).
+Then we conducted field testing with a convenience sample of university participants ($\mathcal{D}_{\text{field}}$).
+During field testing, the other three doctoral students in Industrial/Organizational and Personality Psychology served as independent observers: they reviewed the interview transcripts and provided observer ratings of participants’ personalities to validate the reliability ($\mathcal{R}$) and validity ($\mathcal{V}$) of the structured personality interview questions ($\mathcal{R}^{(t)},\mathcal{V}^{(t)}=f_{5}(\mathcal{Q}_{2}^{(t)},\mathcal{D}_{\text{field}})$).
+This is an iterative process—if the evaluated metrics do not meet the predefined thresholds $\tau_{\text{rel}}$ and $\tau_{\text{val}}$, the questions are further revised ($\mathcal{Q}_{2}^{(t+1)}=f_{\text{revise}}(\mathcal{Q}_{2}^{(t)},\mathcal{R}^{(t)},\mathcal{V}^{(t)})$); otherwise, the process produces $\mathcal{Q}_{\text{final}}$ ($\mathcal{R},\mathcal{V}\xrightarrow{f_{6}}\mathcal{Q}_{\text{final}}$).
+Our field testing results have shown that the average correlation between observer ratings and self-report ratings ($\mathcal{V}$) was 0.36, and the average inter-rater reliability (intraclass correlation; $\mathcal{R}$) was 0.76.
+These values are consistent with theoretical expectations—meta-analytic research suggests an average self–other rating correlation of 0.36 ($\tau_{\text{val}}$) for the Big Five traits (e.g., Connolly et al., 2007), and the inter-rater reliability obtained exceeds the commonly accepted threshold of 0.70 ($\tau_{\text{rel}}$).
+Detailed results from the validation phase are presented in Appendix B.
+As a result, $\mathcal{Q}_{\text{final}}$, comprising 32 carefully developed questions, was developed, as shown in Table 7 in Appendix B.
+3.2 Personality Structured Interview Dataset
+Data were collected online using the finalized set of interview questions, $\mathcal{Q}_{\text{final}}$ (Institutional Review Board approval was obtained; protocol number and institution are masked for blind review).
+Participants were sampled from both undergraduate and working adult populations.
+Undergraduate participants were recruited from a large public university in the U.S.
+Midwest and received course credit for their participation.
+Working adults were recruited via two widely recognized, high-quality crowdsourcing platforms—Prolific [^3] https://www.prolific.com/ and CloudResearch Connect [^4] https://connect.cloudresearch.com/—both known for providing access to diverse and demographically representative samples.
 All working adult participants were compensated for their time.
-The average interview duration was 34 minutes.
-In addition to completing the structured interview, participants provided demographic information and responded to both the personality scale and the personality-related behaviors scale (see Appendix D for detailed descriptions of data collection and examples of the dataset).
-Responses to PSI were subsequently incorporated into the prompt to generate simulated data (see Appendix F for the specific prompts used in the simulation process).
-4 Experimental Setup
-We designed three primary experiments to verify the effectiveness of the PSI method and its advantages over other methods: (1) Response Similarity (§4.2): evaluating the degree of similarity between responses generated by LLMs based on the PSI method and actual corresponding human responses on personality scales; (2) Human Personality Distribution Simulation (§4.3): comparing the PSI method with other methods in simulating human personality distributions; (3) Personality-Related Behavioral Performance (§4.4): testing the models’ performance using the PSI method to simulate personality-related behaviors.
-General settings shared by all experiments are described in §4.1, while specific differences are detailed in their respective sections.
+On average, each structured interview session lasted approximately 34 minutes.
+In addition to responding to $\mathcal{Q}{\text{final}}$, participants provided demographic information and completed a standardized personality inventory (BFI-2; Soto and John, 2017b), along with scales measuring personality-relevant behaviors (organizational citizenship behavior & counterproductive work behavior; Spector and Fox, 2010) (see Appendix C for detailed descriptions and examples).
+Their responses to $\mathcal{Q}{\text{final}}$ were subsequently incorporated into prompt templates used for generating simulated data (see Appendix D for the specific simulation prompts).
+After excluding incomplete responses and those that failed attention checks, the final analytic sample included 357 participants.
+The average age of the participants was 33.30 years, with a standard deviation of 13.06 years.
+In terms of gender identity, 52.40% identified as male and 44.80% as female.
+The racial composition of the sample aligned with national demographic trends, with 70.59% identifying as White, 10.92% as Black or African American, 9.24% as Hispanic/Latinx, and 7.84% as Asian, and 1.40% as Other.
+4 Experiments and Results
 4.1 General Settings
-Psychological Scale
-We used Big Five Inventory-2 (BFI-2; Soto and John, 2017): The BFI-2 is designed to capture three core facets of each of the FFM personality factors.
-Each facet is measured by two positively worded items and two negatively worded items, resulting in 60 items in total.
-Human respondents and LLMs were instructed to indicate the degree to which they agree with each item on a 5-point scale (1 = “Strongly disagree”, 2 = “Somewhat disagree”, 3 = “Neither agree nor disagree”, 4 = “Somewhat agree”, 5 = “Strongly agree”).
-Specific scale items and scoring criteria can be found in Table 5, Table 6 and Table 7 in Appendix E .
+Personality Scale
+We used the Big Five Inventory–2 (BFI-2; Soto and John, 2017b), a widely used and well-validated psychological measure specifically designed to assess three distinct facets within each of the five major personality domains of the FFM.
+Each facet is measured by four items, resulting in a total of 60 items.
+Both human respondents and LLMs were explicitly instructed to rate the extent to which they agreed with each item using a 5-point Likert scale (1 = “Strongly disagree”, 2 = “Somewhat disagree”, 3 = “Neither agree nor disagree”, 4 =“Somewhat agree”, 5 = “Strongly agree”).
+Detailed item content and scoring criteria are presented in Table 2-4 in Appendix A.
 LLMs
-We tested two LLMs, encompassing both closed-source and open-source models: GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024), and Llama 3 (llama-3-70b-instruct) (AI@Meta, 2024).
-We include these models to allow for cross-validation and comparative analysis, ensuring that the findings are robust and not specific to a single model.
-GPT-4o and Llama 3 are among the current leading models, making them ideal for assessing the latest advancements in LLM performance.
-To ensure reproducibility, the temperature of the LLMs was set to zero to generate deterministic responses.
+We evaluated seven widely used LLMs: Mistral-7B (Jiang et al., 2023), Gemma-2-9B and Gemma-2-27B (GemmaTeam, 2024), Llama3-8B and Llama3-70B (AI@Meta, 2024), GPT-4o-mini (OpenAI, 2024a), and GPT-4o (OpenAI, 2024b).
+To ensure reproducibility, all models were consistently tested at a temperature of zero.
+Detailed prompt descriptions are in Appendix D.
+Metrics
+The evaluation metrics vary across experiments and are introduced in detail within their respective sections (§4.2, §4.3, and §4.4).
 4.2 Response Similarity
-The main purpose of this experiment is to assess the similarity between responses generated by LLMs based on the PSI method and human self-reports.
+The main purpose of this experiment is to assess the similarity between LLM-simulated responses based on the PSI method and human self-reports.
 Metrics
-Because the data generated by LLMs correspond on a one-to-one basis with data from human respondents, we used the Mean Absolute Error (MAE) and Pearson correlation coefficient ($r$) to examine the strength of the similarity between them.
-Smaller MAE and higher $r$ indicate greater similarity.
-For the $r$ calculation formula, see Appendix G.
-Park et al.
-(2024) employed a Life Interview simulation, which is currently the most information-rich method for personality simulation.
-It is also one of the few papers that provides a direct one-to-one performance comparison with human samples.
-The reported MAE and $r$ with corresponding human samples are likely the best values available so far.
-We used this as a standard for comparison.
-If data from human respondents are regarded as the gold-standard criterion, this test can also be considered an examination of criterion-related validity, which assesses the correlation between a measure and an external standard (Cronbach and Meehl, 1955).
+To evaluate the degree of similarity between the outputs of LLMs and human respondents, we report the Mean Absolute Error (MAE) and the Pearson correlation coefficient ($r$).
+These two metrics are widely used in personality modeling and psychological assessment due to their ability to capture both absolute differences and linear relationships.
+Specifically, lower MAE values and higher $r$ values indicate a stronger alignment between the two data sources.
+The formula used for calculating $r$ is provided in Appendix E.
+Results
+Figure 2 illustrates the $r$ between PSI method LLM-simulated data and human self-reported data.
+Across models, average $r$ range from 0.43 to 0.52, indicating a statistically significant and moderately strong positive association between simulated and human responses.
+Figure 2: Correlation of human vs.
+LLM-simulated data across different LLMs and two methods.
+PSI achieves comparable correlations to Life Interview using only one-quarter of the time, demonstrating higher efficiency.
+For specific $r$ see Table 12 in Appendix I.1.
+We further compared our results with prior work by Park et al.
+(2024), who employed a Life Interview simulation—widely regarded as one of the most information-rich methods for simulating personality data.
+Their work is notable for offering one of the few direct, one-to-one performance comparisons between LLMs and human participants.
+Compared to the Life Interview approach, the PSI method demonstrates better performance across nearly all personality domains.
+When evaluated using the same backbone model, GPT-4o, PSI yields a lower MAE in four domains and a higher $r$ in three.
+Even where PSI shows slightly higher MAE or lower $r$, the differences are minor (e.g., MAE for Openness: 0.80 vs.
+0.62; $r$ for Neuroticism: 0.63 vs.
+0.68; and Conscientiousness: 0.46 vs.
+0.52).
+Notably, the Life Interview method described by Park et al.
+(2024) requires approximately two hours of interview time.
+In contrast, the PSI approach achieves comparable outcomes with an average duration of just 34 minutes, about one quarter of the time.
+This significant reduction in data collection time not only lessens participant burden but also conserves computational resources needed for simulation.
+These results highlight the efficiency of the PSI method and its effectiveness in leveraging LLMs to simulate personality data.
 4.3 Human Personality Distribution Simulation
-The main purpose of this experiment is to investigate further the differences among various methods of simulating human personality distributions.
-We compared the PSI method, the Persona method—which relies on dialogue information (Zhang et al., 2018)—and the Shape method, which uses adjective-based dimensional categorization (Serapio-García et al., 2023), to observe differences when simulating a normal human sample.
-The specific descriptions of the Persona method and Shape method are shown in Appendix G.
+The primary goal of this experiment is to further examine differences among prompt-based methods for simulating human personality distributions.
+Specifically, we compared the PSI method, the Persona method (which utilizes dialogue-based information; Zhang et al., 2018), and the Shape method (which employs adjective-based dimensional categorization; Serapio-García et al., 2023), to assess how each simulates a representative human sample.
 Human Sample Criterion
-The human samples used for this experiment were collected as part of a broader project related to personality assessment through Prolific and received Institutional Review Board approval.
-Respondents were instructed to complete a set of demographic questions, the BFI-2, and a set of criterion measures.
-The respondents were compensated with \$3.75 for their participation.
+The human samples used for this experiment were collected as part of a broader project related to personality assessment through Prolific (Institutional Review Board approval was obtained).
+Participants were instructed to complete a set of demographic questions, the BFI-2, and a set of criterion measures.
 In total, 1,559 respondents provided valid responses.
-On average, participants were in their early 40s (M = 42.29 years, SD = 11.79), 50.80% identifying as men, 49.20% identifying as women.
+The average age of the participants was 42.29 years, with a standard deviation of 11.79 years, 50.80% identifying as men, 49.20% identifying as women.
+Figure 3: HAI results for human vs.
+LLM-simulated sample across seven backbone LLMs and three methods.
+Each subplot shows results at a different granularity level.
+Higher HAIs (larger area) indicate closer alignment with the heterogeneity distribution in the human sample.
+PSI method performance far exceeds that of Persona and Shape.
 Metrics
-To quantify the degree of similarity between human and LLM responses, we leveraged multiple metrics.
-At the domain and facet levels, we compared the mean (M; sample mean of each domain and facet), standard deviation (SD; sample standard deviation of each domain and facet), Cronbach’s alpha, and correlations among scores.
-At the item level, we also compared the mean and standard deviation of item scores (sample mean and standard deviation of each item).
-Specifically, we also used MAE and $r$ to quantify similarities in these metrics.
-Here, however, MAE and $r$ are based on sample-level analysis rather than the one-to-one correspondence seen in the previous experiment.
-Our focus is on whether, when simulating an overall human sample, similarities can be observed across various measurement levels—including items, facets, and domains.
-Additionally, separately for human and LLM responses, we fitted a three-factor confirmatory factor analysis (CFA; Jöreskog, 1969) model (TFM; where facets are “factors”) to the responses to each domain of the BFI-2.
-Aside from the facet structure of each domain, we also used the facet scores as indicators and fitted the FFM.
-Model fit, standardized factor loadings, and latent correlations among the facets were also compared between human and LLM responses.
-Tucker’s congruence coefficient (TCC; Tucker, 1951) and MAE of factor loadings were used to quantify the similarity between factor solutions from human and LLM responses.
-For the CFA model, model fit information, and the TCC calculation formula, see Appendix G.
-This evaluation framework is adaptable and can be used to analyze the fidelity of other simulated psychometric data (see Appendix H for further details).
+To evaluate the similarity in distribution between the human sample and the LLM-simulated sample, we leveraged multiple metrics.
+We computed the sample means ($\mu$) and standard deviations ($\sigma$) at the domain, facet, and item levels.
+For the domain and facet levels, we also calculated Cronbach’s alpha coefficients to assess internal consistency, as well as inter-scale correlations to examine the relationships among constructs.
+We further computed the Heterogeneity Alignment Index (HAI [^5] The HAI can be understood as a profile correlation between different observations or subgroups of $\sigma$.
+It has been widely applied across various fields to evaluate the similarity between samples or individuals based on multidimensional features (e.g., Humbad et al., 2013; McCrae, 2008).) by correlating the $\sigma$ profiles at each level, providing a measure of similarity in the variability structure of personality data between the human and LLM samples.
+$\text{HAI}=$
+$\frac{\sum_{i=1}^{n}(\sigma_{\text{H},i}\!-\!\bar{\sigma}_{\text{H}})(\sigma_{\text{M},i}\!-\!\bar{\sigma}_{\text{M}})}{\sqrt{\sum_{i=1}^{n}(\sigma_{\text{H},i}\!-\!\bar{\sigma}_{\text{H}})^{2}}\!\cdot\!\sqrt{\sum_{i=1}^{n}(\sigma_{M,i}\!-\!\bar{\sigma}_{M})^{2}}},$
+where $\sigma_{\text{H},i}$ and $\sigma_{\text{M},i}$ denote the $\sigma$ of the $i$-th human and LLM sample, respectively.
+Additionally, we conducted a three-factor confirmatory factor analysis (CFA; Jöreskog, 1969) for both the human and LLM-simulated samples, separately within each BFI-2 domain.
+In these models, referred to as the Three-Factor Models (TFMs), facets were treated as latent factors.
+Beyond modeling the facet structure within each domain, we also employed the facet-level scores as observed indicators in higher-order CFA models to evaluate the broader FFM structure.
+We compared model fit indices, factor loadings, and inter-factor correlations across human and LLM-simulated samples.
+To quantify the personality construct structural similarity between the two sample types, we used Tucker’s congruence coefficient (TCC; Tucker, 1951) and the MAE of the factor loadings.
+Detailed descriptions of the CFA model specifications, fit indices, and the TCC formula are provided in Appendix E.
+This evaluation framework is adaptable and can be applied to assess the fidelity of other types of simulated psychometric data (see Appendix H).
+Descriptive Statistics Results
+Figure 3 presents the HAI results comparing human sample vs.
+LLM-simulated samples using seven different LLMs and three simulation methods (Persona, Shape, and PSI).
+The three radar plots represent HAI scores at varying levels of granularity (item, facet, and domain).
+Each axis corresponds to a different LLM, and higher HAI values (indicated by a larger area within the plot) reflect a closer alignment with the heterogeneity patterns observed in human responses.
+Across all three levels of granularity, the PSI method consistently outperforms both Persona and Shape, demonstrating substantially greater alignment with human-like heterogeneity.
+[…]
+[…]
+The above results reflect analyses of $\mu$ and $\sigma$ for both human and LLM-simulated samples across different levels of granularity.
+Detailed $\mu$ and $\sigma$ values are reported in Tables LABEL:tab:Mistral_dLABEL:tab:GPT-4o in Appendix I.2.
+Personality Data Distribution Visualization
+As shown in Figure 4, we visualized and compared personality data from the human sample and three LLM-based simulation methods across levels by projecting them into a two-dimensional space using principal component analysis (PCA).
+We fitted the two-dimensional PCA model using only the human sample, which served as the reference for subsequent projections of the simulated data.
+Figure 4: PCA comparison of human and GPT-4o simulated data across different levels and methods.
+PSI distributes closest to human; Persona centers with reduced variance; Shape disperses at the extremes.
+Among the three methods, PSI yielded the closest alignment with the human distribution, suggesting a stronger capacity to replicate the multivariate structure of real personality data.
+In contrast, the Persona method exhibited a centrally clustered distribution, indicating limited heterogeneity.
+The Shape method showed a more dispersed distribution, reflecting greater variability, as well as a larger divergence from the human pattern.
+Model Method SOCTCC $\uparrow$ ASSTCC $\uparrow$ ENETCC $\uparrow$ SOCMAE $\downarrow$ ASSMAE $\downarrow$ ENEMAE $\downarrow$
+Mistral-7B Persona $0.91$ $0.92$ $0.98$ $0.09$ $0.08$ $0.04$
+Shape $0.99$ $0.94$ $0.95$ $0.03$ $0.07$ $0.07$
+PSI (ours) $1.00$ $0.96$ $0.99$ $0.02$ $0.07$ $0.02$
+Gemma-2-9B Persona $1.00$ $0.97$ $1.00$ $0.01$ $0.05$ $0.02$
+Shape $0.97$ $0.86$ $0.97$ $0.04$ $0.10$ $0.07$
+PSI (ours) $1.00$ $0.97$ $0.86$ $0.01$ $0.05$ $0.09$
+Gemma-2-27B Persona $1.00$ $0.97$ $0.98$ $0.02$ $0.06$ $0.05$
+Shape $1.00$ $0.97$ $0.97$ $0.02$ $0.06$ $0.08$
+PSI (ours) $1.00$ $0.92$ $0.98$ $0.02$ $0.09$ $0.04$
+Llama3-8B Persona 0.78-0.78 $0.55$ $0.71$ $0.41$ $0.16$ $0.14$
+Shape 0.81-0.81 $0.79$ 0.80-0.80 $0.40$ $0.11$ $0.33$
+PSI (ours) $1.00$ $0.98$ $0.98$ $0.02$ $0.04$ $0.04$
+Llama3-70B Persona $0.99$ $0.98$ $0.98$ $0.03$ $0.05$ $0.05$
+Shape $0.99$ $0.97$ $0.96$ $0.05$ $0.05$ $0.06$
+PSI (ours) $1.00$ $0.95$ $0.99$ $0.03$ $0.09$ $0.03$
+GPT-4o-mini Persona $1.00$ $0.99$ $0.99$ $0.01$ $0.04$ $0.05$
+Shape $0.91$ $0.95$ $1.00$ $0.09$ $0.07$ $0.03$
+PSI (ours) $1.00$ $0.99$ $0.98$ $0.02$ $0.03$ $0.05$
+GPT-4o Persona $1.00$ $0.97$ $0.98$ $0.01$ $0.06$ $0.05$
+Shape $1.00$ $0.99$ $0.97$ $0.02$ $0.05$ $0.06$
+PSI (ours) $1.00$ $0.98$ $0.98$ $0.02$ $0.04$ $0.06$
+Table 1: TCC and MAE for TFM of Extraversion for human vs.
+LLM-simulated data across different Backbone Model + Method combinations.
+Best TCC and MAE are in bold; Higher TCC ($\uparrow$) and lower MAE ($\downarrow$) are better.
+Factor Analysis Results
+Factor analysis encompasses multiple components; here, we present model fit and structural validity results (e.g., fit indices and factor loadings).
+Additional aspects, such as scale reliability and discriminant validity, are reported in Appendix I.2.
+Model Fit: Model fit information for the TFMs and the FFM is presented in Table LABEL:tab:modelfit in Appendix I.2.
+For both TFMs and FFM, the model fit indices for the Persona and PSI methods are relatively comparable to those observed in the human sample, whereas the Shape method exhibits notably poorer fit.
+Detailed explanations of model fit indices are provided in Appendix E.
+Structural Validity: TCC and MAE are used to assess the overall similarity and discrepancy in factor loadings.
+Table 1 illustrates the TFM results for the Extraversion domain as an example, where SOC, ASS, and ENE represent the three facet-level factors.
+Following Lorenzo-Seva and Ten Berge (2006), a TCC above 0.95 indicates good similarity, while values between 0.85 and 0.94 suggest moderate similarity.
+TCC values were generally high, indicating good alignment in factor loadings between the LLM-simulated and human samples across most methods.
+However, notable exceptions were observed for the Persona and Shape methods when using Llama3-8B as the backbone model, where TCC values were negative, suggesting a failure to replicate the human factor structure.
+In terms of MAE, the PSI method generally yielded lower errors, demonstrating better performance in capturing the true factor loading patterns.
+Detailed values of TCC and MAE for all TFMs and FFM are provided in Tables LABEL:tcc_tfm_e33 in Appendix I.2.
+Standardized factor loadings for all TFMs and the FFM are shown in Tables LABEL:tab:fc_sociabilityLABEL:tab:fc_creative_imagination and Tables LABEL:tab:ffm_extraversionLABEL:tab:ffm_openness, respectively.
+These patterns are consistent with the example presented in Table 1.
+Tables LABEL:tab:inter_extraversion59 in Appendix I.2 report the inter-factor correlation.
+Consistent with our previous findings, the PSI method generally outperforms the other two methods.
+For example, in the TFM results, the Persona and Shape methods exhibited anomalies, such as inter-factor correlations exceeding one and accompanied by warning messages during model fitting.
+In contrast, the PSI method produced more stable and plausible results.
+Other results like scale reliability and discriminant validity are in Appendix I.2. Overall, the PSI method consistently yields better performance compared to the Persona and Shape methods.
 4.4 Personality-Related Behavioral Performance
-The primary aim of this experiment is to explore whether LLMs, when assigned specific personality settings, exhibit behaviors theoretically aligned with those personalities.
-This is an exploratory study, as LLMs generate responses based on statistical probabilities derived from the training corpus (Yang et al., 2024), making it uncertain whether assigning personality settings (through the PSI data) will influence the model to act consistently with that personality.
-Fortunately, the data we have collected through the PSI method includes LLM ratings of human behaviors, providing a valuable basis to test this hypothesis.
+This experiment aims to explore whether LLMs, when assigned specific personality settings, exhibit behaviors that align with theoretical expectations.
 Personality-Related Behavior
-We collected self-report data on two classic types of workplace behaviors: organizational citizenship behavior (OCB) and counterproductive work behavior (CWB).
-Both types of behaviors are widely supported by research as being related to personality (e.g., Organ and Ryan, 1995; Berry et al., 2007).
-These data were collected using the scale developed by Spector and Fox (2010) and prompted the LLM to respond to the same questions (for a detailed description of Spector and Fox’s (2010) measures, see Appendix G; for specific prompt details, refer to Appendix F).
+We examined well-studied workplace behaviors: organizational citizenship behavior (OCB) and counterproductive work behavior (CWB), both of which are closely linked to personality (e.g., Organ and Ryan, 1995; Berry et al., 2007).
+Human responses were collected using the scale developed by Spector and Fox (2010), and the same items were administered to the LLMs to generate simulated data.
+For details on the measurement, see Appendix A; for prompt construction, refer to Appendix D.
 Metrics
-We primarily focus on the $r$ between personality domains from different data sources and OCB, as well as CWB.
-We anticipate that the $r$ between self-reported personality domains and OCB/CWB in LLM simulation will closely align with those observed in the human participants.
-5 Experimental Results
-Here, we present the results of the three experiments mentioned above: Response Similarity (§5.1), Human Personality Distribution Simulation (§5.2), and Personality-Related Behavioral Performance (§ 5.3).
-5.1 Response Similarity Results
-Table 1 provides a clear illustration of the similarity between the personality generated using the PSI method and the corresponding human self-reported personality data.
-The results show that regardless of the model used, the average correlation is around .5, suggesting a moderately strong positive relationship and a significant association between the two variables.
-In particular, compared to the Life Interview method (Park et al., 2024), the PSI method consistently outperforms or matches the Life Interview approach across nearly all metrics, as indicated by its similar or lower MAE and similar or higher $r$.
-Notably, the Life Interview method relies on up to two hours of interview data, while the PSI method achieves comparable results using interview data consisting of just 32 questions, with an average duration of about 34 minutes.
-This further underscores the advantage of the PSI method in effectively leveraging LLMs to simulate personality profiles.
-Domain MAE $r$
-PSI GPT-4o PSI Llama3 Life Interview PSI GPT-4o PSI Llama3 Life Interview
-Extraversion 0.58 0.75 0.72 .64 .43 .45
-Agreeableness 0.53 0.56 0.60 .41 .36 .35
-Conscientiousness 0.59 0.66 0.63 .46 .40 .52
-Neuroticism 0.63 0.59 0.75 .63 .57 .68
-Openness 0.80 1.27 0.62 .43 .39 .39
-Table 1: MAE and $r$ of Human Response and LLM Response Across PSI GPT-4o, PSI Llama3, and Life Interview (see Park et al., 2024, page 35 Table 3) for BFI-2 Each Personality Domain
-Note: n = 357 for PSI method; n = 1,052 for Life Interview method. The bold text highlights the aspects where the PSI method performs better than the Life Interview method.
-5.2 Human Personality Distribution Simulation Results
-Descriptive Statistics
-We compared the means and standard deviations of the different methods at three levels: personality item, facet, and domain.
-MAE and $r$ for both the means and standard deviations are shown in Table 2.
-The MAE for the means reflected the average difference between the LLM responses and the human responses at each level, whereas the MAE for the standard deviations revealed the difference in variability between the two datasets.
-The $r$ further illustrated the linear relationship between the two datasets, with values closer to one indicating a stronger correlation.
-Detailed means and standard deviations for human responses and LLM responses at the item, facet, and domain levels are provided in Tables 1213141516, and 17 in Appendix I.1.
-Table 2 shows that the PSI method demonstrates better performance in simulating human samples compared to the Persona and Shape methods at the item level.
-Although at the facet and domain level, some results from other methods slightly outperform PSI (e.g., Shape GPT-4o in M), the process of aggregating scores from item to facet to domain levels reduces the impact of extreme values, smoothing them out at higher levels.
-Therefore, the PSI method’s advantage at the more granular item level becomes more important.
-Additionally, PSI consistently demonstrates statistically significant positive correlations with the SD of human samples across items, facets, and domains, whereas the other methods exhibit either negative or non-significant correlations.
-Simulating the SD is a challenge for the other methods.
-Although the Shape method increases response variability and introduces more variance, it still struggles to replicate the true variance observed in human samples.
-The PSI method proposed in this paper addresses this issue to a certain extent.
-Model Item Level Facet Level Domain Level
-MAE $r$ MAE $r$ MAE $r$
-M SD M SD M SD M SD M SD M SD
-Persona
-GPT-4o 0.48 0.44 .50 -.13 0.42 0.36 .64 -.24 0.42 0.30 .69 -.64
-Llama3 0.51 0.26 .57 .02 0.33 0.15 .80 .31 0.31 0.06 .80 .76
-Shape
-GPT-4o 0.58 0.16 .38 -.06 0.50 0.18 .82 -.41 0.46 0.21 .84 -.77
-Llama3 0.62 0.25 .45 -.26 0.52 0.38 .78 -.64 0.45 0.31 .82 -.65
-PSI
-GPT-4o 0.48 0.34 .68 .27 0.40 0.22 .71 .49 0.38 0.15 .67 .69
-Llama3 0.66 0.36 .50 .14 0.54 0.27 .57 .19 0.49 0.21 .54 .51
-Table 2: MAE and $r$ for BFI-2 Human Responses and Different Methods LLM Responses at Item, Facet, and Domain Levels
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-The bold text indicates the best-performing data in that column.
-Psychometric Performance
-Psychometric performance includes multiple components; here, we primarily present model fit and structural validity (i.e., factor loadings).
-For other aspects, such as scale reliability and discriminant validity, please refer to Appendix I.1.
-Model Fit: For both the human sample criterion and different methods of LLM responses, the TFM was fitted to each BFI-2 domain, and the FFM was fitted to all the data.
-TFM fit information is shown in Table 18, and FFM fit information can be seen in Table 19 in Appendix I.1.
-From Table 18, it can be observed that the model fit indices of the Persona method and the PSI method are relatively similar to those of the human sample, while the Shape method performs slightly worse.
-Table 19 reflects a similar trend, where the Shape method shows notable discrepancies compared to the human sample in the RMSEA and SRMR model fit indices (see Appendix G for the explanation of RMSEA, SRMR, and other model fit indices).
-Structural Validity: TCC was used to evaluate the similarity of factor loadings between human responses and LLM responses.
-TCC results for the TFM of each BFI-2 domain are shown in Table 20, while results for the FFM are shown in Table 21 in Appendix I.1.
-The values of the TCC are generally high and thus supportive of factor-loading correspondence, with all methods showing strong alignment with the human sample (a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity; Lorenzo-Seva and Ten Berge, 2006).
-However, it is important to note that the TCC focuses on overall profile similarity, largely ignoring differences in absolute values (which might be too lenient).
-Therefore, we also need to examine the results of the specific standardized factor loadings.
-Tables 22 and 23 in Appendix I.1 present the standardized factor loadings.
-From these tables, it is clear that the PSI method outperforms the Shape method in terms of similarity to human samples, with its factor loadings more closely matching human data.
-However, its results are relatively close to those of the Persona method, yet there are clearly differences in some factor loadings compared to the human sample.
-It is important to note that the highest factor loadings do not necessarily indicate the best performance, as we are evaluating the similarity between the factor loadings of the human data and the LLM responses.
-Higher factor loadings suggest stronger correlations between the factor and the items, but this is not always ideal.
-It indicates that LLMs may struggle to make refined distinctions among items within a facet, meaning they may have difficulty determining which items are more or less related to the factor, unlike human respondents, who exhibit greater differentiation.
-Tables 24 and 25 in Appendix I.1 present the specific inter-factor correlations.
-Aligning with our previous findings, the PSI method generally performs better than the other two methods when simulating human personality distributions, especially at a more fine-grained level.
-For instance, in the TFM results, other methods encountered anomalies, such as inter-factor correlations exceeding one (accompanied by a warning message when fitting the model), but the PSI method generally performed well.
-The other results (e.g., scale reliability and discriminant validity) can be found in Appendix I.1.
-Overall, the results produced by the PSI method generally outperform those of the Persona method and the Shape method.
-However, there is still a certain gap compared to the human sample.
-5.3 Personality-Related Behavioral Performance Results
-Table 28 presents the correlations between the personality dimensions and OCB/CWB reported by PSI GPT-4o, with a comparison to human self-reported data.
-The relevant results for PSI Llama3 are provided in Table 29 in Appendix I.2, while Figures 3 and Figure 4 display the complete correlation matrices for GPT-4o and Llama3, respectively.
-The results demonstrate that the correlations simulated by LLMs between OCB/CWB and various personality dimensions closely resemble the patterns seen in human self-reported data, except for the openness domain.
-When human reports show positive, negative, or no correlation, the LLM simulations generally exhibit consistent trends in the corresponding directions.
-However, the correlations generated by LLM simulations are often higher, likely because the model primarily relies on the “typical” or “idealized” knowledge structures absorbed from its training corpus during simulation, consistently repeating and amplifying such associations.
-In contrast, human self-reported data contains more random noise, leading to relatively weaker correlations.
+Our primary focus is on the $r$ between personality domains from different data sources and two behavioral outcomes: OCB and CWB.
+We expect that the $r$ between personality and OCB/CWB in the LLM-simulated data will closely mirror those observed in the human sample.
+Results
+Table 67 and 68 in Appendix I.3 reports the $r$ between personality domains and OCB/CWB in both human data and data simulated using the PSI method.
+The results demonstrate that correlations $r$ generated by PSI method largely reflect the patterns found in human data across most personality domains, with the exception of Openness.
+Generally, when human data exhibit positive, negative, or negligible $r$, the LLM-simulated results tend to align in the same direction.
+However, these simulated $r$ are often stronger in magnitude.
+This amplification may stem from the models’ reliance on “typical” or “idealized” knowledge structures internalized during training, which can reinforce and magnify common associations.
+In contrast, human responses are shaped by individual variability and random noise, leading to more diffuse patterns.
+Figure 5: Ablation test results for predicting Conscientiousness.
+The x-axis represents each question removed on its own, and the y-axis shows the $R^{2}$ value.
+The red dashed line indicates the baseline $R^{2}$ when all questions are included.
+4.5 Ablation Study on PSI Questions
+We further investigated the individual contributions of each PSI question to personality prediction through an ablation study.
+The PSI was designed as a highly integrated framework, with its questions intentionally co-developed and structurally interlinked to capture diverse facets of personality-related information.
+As a result, isolating individual questions is inherently challenging, as doing so could undermine the integrity of the overall design.
+Moreover, each question tends to tap into different aspects of personality, rendering them functionally complementary.
+This also implies that certain questions may be more effective than others in capturing information relevant to specific domains; some items may be more closely aligned with one domain than others.
+Given this design, we expected the differences in individual question effectiveness to be relatively minor overall.
+However, we also anticipated that the relative importance of each question would vary depending on the specific personality domain being predicted.
+We conducted an ablation study by leaving one question out each time and simulated personality traits based on the remaining 31 question–answer pairs (see Table 7 in Appendix B for all questions).
+We then calculated the coefficient of determination ($R^{2}$) using human self-reported personality scores as the reference, in order to assess the impact of each question on model performance.
+If $R^{2}$ increases after a particular question is removed, this suggests that the question may have negatively affected prediction—potentially introducing redundancy, noise, or irrelevant information that weakened overall performance.
+Conversely, if $R^{2}$ decreases after removal, it indicates that the question made a substantial contribution to prediction and was a key factor supporting model performance.
+As shown in the Figure 5, the removal of most questions leads to a slight decrease in the $R^{2}$, as one would expect, although the changes are minor and generally remain within a ±0.01 range.
+Notably, the relative importance of individual questions varies across different prediction domains (see Figure 7-10 in the Appendix I.4).
+This empirical finding supports our earlier hypothesis: different types of questions capture distinct facets of personality, offering complementary rather than redundant information.
+Furthermore, the relative importance of questions varies across target domains, indicating that different questions make distinct contributions to domain-specific predictions.
+Together, these findings underscore the meaningful contributions of each PSI question.
 5 Conclusion
-In summary, this study proposes a novel method for LLM to simulate human personality distributions—PSI and provides a comprehensive explanation of its development process and corresponding dataset.
-Through three experiments, we validated PSI’s effectiveness.
-Results show that PSI matches or even outperforms two-hour interviews (Park et al., 2024) in personality simulation and surpasses existing methods in terms of both item-level accuracy and overall psychometric indicators.
-One experiment also examined PSI’s ability to simulate personality-related behavior, revealing that although LLMs approach human-like performance, they still exhibit idealized tendencies.
-In summary, a theory-informed structured interview like PSI can better simulate human-like psychometric data.
-We explore the role of theory-informed structured interviews in simulation in greater detail and examine their potential to advance research, both of which are elaborated upon in Appendix C.
+This study introduces a novel method, PSI, for simulating human personality data with LLMs and provides a detailed account of its development and the associated dataset.
+Across three experiments, we evaluated the effectiveness of PSI.
+The results show that PSI performs well in simulating personality, yielding moderately strong and statistically significant correlations with human self-report data.
+Moreover, when modeling the distribution of personality data at the population level, PSI outperforms existing methods in both human-like heterogeneity and broader factor analysis results.
+Our experiments also assessed PSI’s ability to simulate personality-related behavior.
+Although LLMs approximate human-like patterns, they still exhibit idealized responses that deviate from the natural variability found in human behavior.
+In summary, theory-informed structured interviews such as PSI offer a more realistic and psychometrically grounded approach to simulating human-like data.
+We further discuss the utility of such data in simulation-based research and its potential to improve future psychometric studies in Appendix G.
 Limitations
-This study has several limitations.
-First, we explored the potential of PSI in advancing research.
-Our single experiment demonstrates that theory-informed questions can effectively elicit information embedded in human narrative responses—–particularly those relevant to the target personality construct—–allowing LLMs to simulate the corresponding behavior.
-However, additional experiments are needed to investigate its broader applications.
-Second, the personality assessment in this paper relies on self-reported personality scales, which can be viewed as a limitation of the study.
-However, these self-report scales are specifically designed for humans to target personality traits across individuals, in ways that cannot be directly located within LLMs themselves (see Appendix C for a detailed discussion on psychometrics).
-Furthermore, self-report scales are used often because respondents have privileged access to their own personality, and even with their well-known limitations (e.g., faking or social desirability in self-reports; see Appendix I.3 for the social desirability analysis), they remain valuable to evaluate the personality traits exhibited by LLMs.
-However, because the primary goal of this study is to have the LLM simulate human responses and behaviors that correspond to extracted personality information, it is reasonable to assess LLMs using methods designed to measure human personality traits.
-Moreover, this paper tests only two types of LLMs, which may somewhat limit its scope, given the array of LLMs that are now available.
-However, the conclusions drawn from our two models were highly consistent, and we have reason to believe that the advantages of the PSI method are generally applicable across different LLMs.
-Furthermore, we designed three different experiments and conducted comprehensive psychometric analyses of various methods within LLMs, ensuring greater rigor and reliability of the results.
+This study is not without limitations.
+First, our initial evidence suggests that theory-driven prompts can successfully elicit personality-relevant information embedded in human narratives, enabling LLMs to simulate corresponding psychometric data.
+However, further research is necessary to examine the robustness and generalizability of this approach across diverse contexts and constructs.
+Second, the personality assessment in this study is based on self-report measures originally designed for humans.
+These tools are intended to capture the stable, internal personality that individuals can access and report, which LLMs, by design, do not inherently possess (see Appendix G for a detailed discussion on psychometric considerations).
+Nevertheless, LLMs can simulate human-like responses that reflect particular personality patterns when guided by theory-informed prompts.
+Because the goal of this study is not to assert that LLMs have personality, but rather that they can simulate personality-driven behavior, it is reasonable to apply human-based assessment tools to evaluate the quality and fidelity of that simulation.
 Ethical Statement
-We hereby confirm that all authors of this study are aware of the provided ACL Code of Ethics and comply with the Code of Conduct.
-Human Sample Data
-This paper discusses the comparison between multiple sets of human data and results generated by LLMs.
-The collection of human data strictly adhered to relevant ethical guidelines and received approval from the Institutional Review Board.
-To ensure fair treatment of participants and proper recognition of their contributions, reasonable compensation or course credit (for student participants) was provided.
-Throughout the research process, we placed a strong emphasis on transparency and openness, also ensuring that all participants signed informed consent forms.
-Additionally, to safeguard data privacy and uphold ethical standards, the publicly available dataset underwent rigorous screening to exclude any personally identifiable information and was shared only with explicit public consent.
-Appendix A Appendix: The Structure of Personality
-A key question in personality-related LLM research pertains to personality structure: What is the nature and breadth of the personality traits we want to simulate the human personality distributions?
-In the research literature, personality structures often emerge from applying factor analysis to individuals’ responses to a large number of personality-relevant items.
-This approach is what has been used to identify the five personaltiy factors in the FFM.
-Moreover, personality is better understood in terms of one’s continuous standing on each of multiple dimensions rather than as static types or profiles.
-Research data clearly supports this view (Wilmot, 2015; Wilmot et al., 2019).
-Dividing individuals into limited categories (e.g., 16 types in MBTI) artificially segments continuous dimensions into discrete units, which may overlook important individual differences (Ones and Wiernik, 2018).
-Another question is whether we should incorporate personality traits and structures beyond the FFM to achieve a more comprehensive understanding of personality, given that additional personality variables and alternative structures have been proposed over the years.
-It turns out that there can be conceptual overlap among these models (e.g., Hough et al., 2015).
-For example, a meta-analysis by Joseph and Newman (2010) revealed that emotional intelligence (EI) shows statistically and practically significant relationships with neuroticism and extroversion within the FFM.
-In fact, when controlling for personality variables, the unique contribution of EI almost disappears.
-Similarly, Credé et al.
-(2017) conducted a meta-analysis on grit and found that its core components can largely be explained by conscientiousness, with little added predictive validity beyond that.
-Moreover, although the HEXACO structure introduces an honesty-humility dimension, the remaining five dimensions align closely with the FFM structure (Lee and Ashton, 2004, 2006).
-Research by Cutler and Condon (2023) further indicates that nearly all personality semantic information can be classified within the FFM structure.
-These studies collectively suggest that although further subdivision of personality structures might provide new perspectives, it often leads to conceptual redundancy and measurement complexity without necessarily enhancing predictive validity or theoretical value.
-Rather than pursuing a wide range of personality frameworks, we will delve deeper into the FFM structure, which is very widely accepted by personality psychologists.
-Appendix B Appendix: Personality Structured Interview Questions
-Table 3 presents the final set of 32 questions that form the basis of our personality structured interview.
-This framework was developed by adapting and modifying McAdams’s life history interview and narrative identity approach (McAdams, 1995, 1996, 2001), while also incorporating components from the Structured Interview of the Five-Factor Model (SIFFM; Trull et al., 1998).
-The initial draft of the question pool was created and refined through collaborative discussions among subject matter experts (SMEs).
-The SME team consisted of two doctoral students, one postdoctoral researcher, and one professor, all specializing in personality psychology.
-To ensure the quality and clarity of the questions, pilot testing was conducted with six undergraduate research assistants from a personality research lab.
-This iterative process of development and feedback led to the construction of the personality structured interview.
-The development process is similar to the development of psychological tests or scales (psychometric). We have provided more details on psychometrics and the development framework in Appendix C.
-# Questions
-1 To get us started, where are you from? Where did you grow up and what was the place like?
-2 Thinking back, what kind of student were you in school?
-3 Did you have a teacher or teachers that were influential? If so, why? What were they like?
-4 What was your favorite subject in school, and why?
-5 What was your least favorite subject in school, and why?
-6 Still thinking back, who were your heroes when you were young and why?
-7 When you were little, what did you want to be when you grew up? And why?
-8 What were your dreams and plans when you graduated from high school? What made you have those dreams or plans?
-9 If you had complete freedom, what would your dream job be, and why?
-10 How have your dreams and goals changed throughout your life?
-11 Shifting gears to your childhood, how would you describe the personalities of people in the family you grew up in? For example, what were your parents and/or siblings like?
-12 How are you similar or different from your parents and/or siblings?
-13 How do you think your similarities and/or differences influenced your relationship with them?
-14 What was the best part of your childhood?
-15 What do you think were the worst parts of your childhood?
-16 Switching gears a little bit, what was your first paid job? How old were you then? (If this is not applicable to you, then please put ‘NA’)
-17 What other jobs have you had? (If this is not applicable to you, then please put ‘NA’)
-18 What do you do now for a living? And why did you choose it?
-19 Please describe your typical work day.
-20 What is the best and worst part of your current work?
-21 Did you serve in the military? Please tell us about that experience, what was the best and worst part of it?
-22 Moving on, what are your adult friendships like?
-23 How are your adult friendships different from your childhood friendships?
-24 What are your strongest qualities as a friend? In other words, what makes you a great friend to have?
-25 What about your weakest qualities in friendships? In other words, what do you struggle with when you are trying to be a friend to someone?
-26 Moving onto more general questions, when thinking about your life in general, what are you most proud of?
-27 What hobbies or other interests do you have?
-28 What things frighten you now?
-29 What were some things that frightened you most as a child?
-30 What are the three biggest news events that have occurred in your lifetime?
-31 If you had the power to solve one and only one problem in the world, what would it be, and why?
-32 Tell me about a time when you did not know if you would make it. How did you overcome that challenge?
-Table 3: Structured Interview Questions
-Appendix C Appendix: Psychometrics and Structured Interview Development Framework
-Psychometrics is a field of psychology dedicated to the theory and practice of psychological measurement.
-It primarily focuses on quantifying psychological traits, behaviors, and abilities through systematic testing and analysis.
-Psychological traits, such as personality dimensions, cognitive abilities, and emotional states, are inherently abstract constructs that cannot be measured directly.
-Therefore, psychometricians rely on tools like surveys, questionnaires, scales, or structured interviews to infer these traits through observable indicators or responses.
-Measuring Psychological Traits
-To measure a psychological trait, psychometricians typically operationalize the trait by identifying observable behaviors or self-identities that correlate with the underlying construct.
-For example, extraversion can be measured by assessing behaviors such as sociability, assertiveness, and enjoyment of social interactions, or by examining identities such as seeing oneself as an outgoing person and believing that one thrives in social situations.
-These behaviors/identities are translated into measurable items (for scale; e.g., “I enjoy being the center of attention”) or questions (for interview; e.g., “What are your strongest qualities as a friend?
-In other words, what makes you a great friend to have?”).
-The challenge lies in ensuring that these items/questions accurately and consistently capture the construct across different populations and contexts.
-Theory-Informed Structured Interview for LLM Data Simulation
-Theory-informed structured interviews are the most suitable method for enabling LLMs to simulate psychometric data.
-These interviews are specifically designed to capture the constructs underlying the targeted psychometric measures, ensuring that the simulated data aligns with the intended psychological construct.
-By extracting textual information that directly reflects the target construct, theory-informed structured interviews facilitate the representation of heterogeneous data while preserving a high degree of human diversity, thereby enhancing the validity and applicability of the simulated psychometric data.
-Moreover, since the information is extracted based on theoretical foundations, it also provides a certain level of interpretability for the LLM’s simulation.
-This not only allows the generated data to be compared with theoretical expectations but also increases its potential for practical applications.
-The Potential for Advancing Research
-A theory-informed structured interview transcript-based simulation can generate data more effectively by focusing on the target construct.
-Ideally, the simulated data should reproduce the same constructs reflected in real-world data and simulate behaviors associated with these constructs.
-Take personality as an example—if simulated data can accurately replicate real-world personality constructs, it enables research that would be difficult to conduct in reality, such as developing contextualized personality assessment tools and exploring new personality theories through multi-agent simulations.
-Developing Contextualized Personality Assessment Tools: Traditional personality assessments mainly rely on standardized questionnaires or laboratory tasks, which often fail to adequately simulate real-world social contexts.
-By using theory-informed structured interview transcript-based simulations, we can generate more fine-grained and context-sensitive individual response data.
-For instance, we can simulate various occupational scenarios (such as crisis management, teamwork, or remote work) and analyze how different personality traits manifest in these contexts.
-This approach not only aids in developing measurement tools tailored to specific applications but also enhances ecological validity, allowing for more accurate assessments of personality across different situations.
-Exploring New Personality Theories through Multi-Agent Simulation: If simulated data can accurately reflect real-world personality constructs, we can leverage multi-agent interactive systems to simulate individuals’ behavioral patterns and observe how different personality traits evolve in group dynamics.
-For example, virtual agents with distinct personality traits can be placed in cooperative tasks, competitive environments, or social interactions, enabling researchers to test whether existing personality theories effectively predict these interaction patterns.
-Additionally, this approach can uncover new personality dynamics, such as whether certain personality trait combinations produce unexpected group effects or whether behavior in specific situations deviates from traditional theoretical predictions.
-However, these assumptions are based on an ideal premise—that simulated data can successfully reflect the same constructs as real-world data.
-A theory-informed structured interview undoubtedly offers a promising pathway in this regard, warranting further in-depth exploration.
-Structured Interview Development Framework
-Developing a structured interview for LLM to simulate data involves a series of carefully designed steps to ensure that the resulting test reliably and validly measures the construct of interest.
-Table 4 outlines the framework.
-# Steps
-1 Identify behaviors/perceptions that represent the construct or define the domain.
-2 Prepare a set of structured interview specifications—structured interview blueprint.
-3 Build an initial question pool.
-4 Have questions reviewed by substantive experts (and revised as necessary).
-5 Hold preliminary question tryouts.
-6 Determine statistical properties of questions (and eliminate poor questions or revise as necessary).
-7 Field-test the structured interview on a large representative sample of the intended examinee population.
-8 Design and conduct reliability and validity studies for the final form of the structured interview.
-9 Develop guidelines for administration, and interpretation of the structured interview.
-Table 4: Structured Interview Development Framework
-Identify behaviors/perceptions that represent the construct or define the domain: Clearly defining the construct is essential to developing relevant structured interview questions.
-The construct should be operationalized by identifying specific behaviors or attributes that indicate its presence.
-In other words, you need to find a theory to guide you on how to measure the target constructs.
-This step may involve reviewing literature, conducting expert interviews, or organizing focus groups to understand the various dimensions and observable characteristics of the construct.
-Prepare a set of structured interview specifications—structured interview blueprint: A blueprint outlines the structured interview’s structure and content, specifying how questions will be distributed across the construct’s dimensions or components.
-It typically includes information on the number of questions per domain, questions content.
-Build an initial question pool: In this step, an extensive list of questions is created to cover the full range of the construct.
-Question wording should be clear, concise, and relevant to the target population.
-It is common practice to generate more questions than needed to ensure that poorly performing questions can be removed later without compromising the structured interview.
-Have questions reviewed by substantive experts (and revised as necessary): SMEs review the question pool for content accuracy, relevance, clarity, and bias.
-Experts assess whether the questions align with the construct’s definition and whether any important aspects are missing.
-Feedback from experts helps refine the wording, remove ambiguous questions, and identify questions with potential cultural or gender biases.
-Hold preliminary question tryouts: Before large-scale testing, questions are piloted on a small group of individuals representative of the target population.
-This stage helps identify any immediate issues with question comprehension, response format, or instructions.
-It can also include cognitive interviews where participants are asked to explain their thought processes when answering questions.
-The feedback from this stage informs further revisions, ensuring that questions are clear and easy to understand.
-Determine statistical properties of questions (and eliminate poor questions or revise as necessary): Question performance is assessed through statistical analyses to evaluate difficulty, discrimination, and internal consistency.
-Standardized scoring criteria, such as Behaviorally Anchored Rating Scales (BARS), can be used for obtaining question scores.
-Then, methods like correlations, factor analysis, and item response theory (IRT) can help determine how effectively each question measures the intended construct.
-Questions that demonstrate poor psychometric properties—such as low discrimination or high measurement error—are either revised or removed.
-For example, questions with low correlations with the overall score or incorrect factor loadings may be eliminated.
-Field-test the structured interview on a large representative sample of the intended examinee population: The revised item set is administered to a large, representative sample to gather comprehensive data on the scale’s performance.
-This step ensures that the sample reflects the population for which the test is intended, which is critical for generalizability.
-Statistical analyses are conducted to refine the test further.
-This process may involve removing redundant items, assessing dimensionality, and ensuring that items work well across demographic subgroups.
-Design and conduct reliability and validity studies for the final form of the structured interview: To ensure the structured interview is psychometrically sound, various reliability and validity studies are conducted: Reliability studies measure the structured interview’s consistency and stability, including internal consistency (e.g., Cronbach’s alpha), test-retest reliability, and inter-rater reliability (if applicable).
-Validity studies assess whether the structured interview measures what it is intended to measure.
-This includes: Content validity (the extent to which items cover the construct); Construct validity (e.g., convergent and discriminant validity); Criterion-related validity (e.g., predictive or concurrent validity).
-These studies provide evidence that the structured interview is both reliable and valid for its intended purpose.
-Develop guidelines for administration and interpretation of the structured interview: The final step involves creating a comprehensive, structured interview manual that includes instructions for structured interview administration, scoring procedures, and guidelines for interpreting results.
-This manual ensures consistency in the structured interview’s use across different settings and helps minimize errors in administration and scoring.
-Guidelines for interpreting scores may include norms, cutoff points, and descriptions of what various score ranges indicate.
-Appendix D Appendix: Personality Structured Interview Dataset and Data Collection
-Institutional Review Board approval was obtained. The data was collected through an online questionnaire, followed by an online structured interview.
-Since we are still actively collecting data, we will share a part of the dataset that does not include any personal privacy information and has received permission for data sharing.
-Each example is composed of the following characteristics:
-  1. 1.
-    Gender: The gender of the participant.
-  2. 2.
-    Race: The racial background of the participant.
-  3. 3.
-    English: Whether English is the participant’s first language.
-  4. 4.
-    Age: The participant’s age.
-  5. 5.
-    Weight: The participant’s weight.
-  6. 6.
-    Height: The participant’s height.
-  7. 7.
-    OCB1–OCB10: Self-reported Organizational Citizenship Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 10 for specific item details.
-  8. 8.
-    CWB1–CWB10: Self-reported Counterproductive Work Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 11 for specific item details.
-  9. 9.
-    Q1–Q32: Participant responses to each personality structured interview questions, see Table 3 for specific question details.
-  10. 10.
-    Item1–Item60: Responses to each item of the BFI-2. Refer to Appendix E for item descriptions and scoring guidelines.
-Appendix E Appendix: BFI-2 Scale
+This paper presents a comparison between human data and LLM-simulated data by three different methods.
+All human data were collected in strict accordance with relevant ethical guidelines and were approved by the Institutional Review Board.
+Participants received either reasonable monetary compensation or course credit (in the case of student participants) to ensure fair treatment and appropriate recognition of their contributions.
+We placed a strong emphasis on transparency and ethical integrity throughout the research process.
+All participants provided informed consent prior to participation.
+Moreover, to protect privacy and maintain ethical standards, the publicly shared dataset was carefully screened to remove any personally identifiable information.
+Acknowledgments
+This research project is partially supported by the Science of Trustworthy AI Award from Schmidt Sciences, the AI2AI Award from Amazon, the Microsoft Accelerate Foundation Models Research (AFMR) grant program, and by the National Science Foundation (NSF) under Award Number 2522411.
+The content is solely the responsibility of the authors and does not necessarily represent the official views of the NSF.
+Pengda Wang: conceptualization (equal, lead), methodology (equal, lead), software (lead), formal analysis (lead), investigation (lead), data curation (lead), writing - original draft (lead), writing - review & editing (equal, lead), visualization (lead); Huiqi Zou: methodology (support), validation (support), formal analysis (support), investigation (support), data curation (support), writing - original draft (support), writing - review & editing (equal, support), visualization (support); Han Jiang: methodology (support), validation (support), investigation (support), data curation (support), writing - original draft (support), writing - review & editing (equal, support), visualization (support); Hanjie Chen: writing - review & editing (equal, support); Tianjun Sun: conceptualization (equal, lead), methodology (equal, lead), validation (support), formal analysis (support), investigation (support), resources (equal, lead), data curation (support), writing - original draft (support), writing - review & editing (equal, lead), supervision (equal, lead), funding acquisition (equal, lead); Xiaoyuan Yi: review & editing (equal, support), supervision (equal, lead); Ziang Xiao: conceptualization (equal, lead), methodology (equal, lead), validation (support), formal analysis (support), investigation (support), resources (equal, lead), data curation (support), writing - original draft (support), writing - review & editing (equal, lead), supervision (equal, lead), funding acquisition (equal, lead); Frederick L.
+Oswald: methodology (support), writing - review & editing (equal, lead).
+Appendix A Appendix: BFI-2 Scale and Behavioral Measures Scales
+A.1 BFI-2 Scale
 Instructions: Here are a number of characteristics that may or may not apply to you.
@@
 60 I am someone who is original, comes up with new ideas.
-Table 5: BFI-2 Scale
-Scoring: Reverse-keyed items are denoted by “R.”
-In psychological measurement and scale development, reverse coding is a common and essential technique.
-Its primary purpose is to ensure that the scoring direction of all items remains consistent, thereby improving the reliability of the measurement results and the accuracy of their interpretation.
-A balanced arrangement of positive and negative items helps reduce response biases, such as consistency effects or response patterns where participants select the same option repeatedly.
-These tendencies can mask the respondent’s true attitudes or behavioral traits, leading to distorted measurement outcomes.
-By alternating positive and negative statements and applying reverse coding to the appropriate items, this issue can be effectively mitigated.
-Reverse coding can also help reduce the influence of social desirability bias.
+Table 2: BFI-2 scale.
+Scoring: Reverse-keyed items appear as “R.”
+Reverse coding is a widely used technique in psychological measurement and scale development.
+Its primary purpose is to align the scoring direction of all items, ensuring consistency across the scale.
+This process enhances the reliability of the overall measurement and supports a more accurate interpretation of the results.
+A balanced inclusion of positively and negatively worded items is a well-established strategy to reduce response biases—such as acquiescence bias, consistency effects, or patterned responding (e.g., selecting the same scale point across items).
+These biases can obscure respondents’ true attitudes or behavioral tendencies, compromising measurement validity.
+Although it is not guaranteed, reverse coding can help reduce the influence of social desirability bias.
 When all items are presented in the same direction, participants may easily guess the purpose of the test and provide responses that align with perceived expectations.
-By mixing positive and negative statements and applying reverse coding, the scale can disrupt this pattern, making it harder for participants to determine the “correct” answers, thus providing a more genuine reflection of their inner states or attitudes.
+By mixing positive and negative statements and applying reverse coding, the scale can disrupt this pattern, ideally making participants reflect more carefully on their inner states or attitudes when providing responses.
 Domain Level Item Numbers
@@
 Openness 5R, 10, 15, 20, 25R, 30R, 35, 40, 45R, 50R, 55R, 60
-Table 6: BFI-2 Domain Level with Item Numbers
+Table 3: BFI-2 domain level with item numbers.
 Facet Level Item Numbers
@@
 Creative Imagination 15, 30R, 45R, 60
-Table 7: BFI-2 Facet Level with Item Numbers
-Appendix F Appendix: Prompts List
-Personality Scale Prompt Format
-Table 8 presents the prompts used to generate LLM responses for the selected personality test on a Likert scale, where personality_description denotes the personality prompt.
-In the PSI method, the personality prompt integrates both the interview question and the interviewee’s corresponding response, as collected in the PSI dataset (see Appendix D).
-For the Persona and Shape methods, more detailed descriptions of the personality prompts can be found in Appendix G.
-The test_item corresponds to each individual item in the BFI-2 scale.
-For the following task, respond in a way that matches this description: {personality_description}.
-Considering the statement, please indicate the extent to which you agree or disagree on a scale from 1 to 5 (where 1 = “disagree strongly”, 2 = “disagree a little”, 3 = “neither agree nor disagree”, 4 = “agree a little”, and 5 = “agree strongly”): {test_item}.
-Table 8: Prompt format for gathering LLMs’ responses to BFI-2 scale.
-Behavioral Question Prompt Format
-The complete prompt format for eliciting LLMs’ responses to personality-related behavioral questions is defined in Table 9.
-The questions_and_responses refers to the transcript of a structured human personality interview, while question_list comprises statements evaluating OCB and CWB, as detailed in Table 10 and Table 11 in Appendix G.
-The model is instructed to rate each statement using a frequency scale from one to five.
-Task: Simulate an individual’s behavior and predict their responses to a series of work-related questions.
-Data to Analyze: {questions_and_responses}
-Instructions:
-1. Simulate the Individual: Use the provided data to simulate the individual’s personality, behavior, and work habits.
-2. Predict Responses:
-- Based on the provided data, simulate the individual and answer the question.
-- Recall that in the past year, how often have you done the following things? (1 = Never, 2 = Once or twice, 3 = Once or twice per month, 4 = Once or twice per week, 5 = Every day)
-Questions: {question_list}
-Output Format:
-Question Score: [Provide the numerical score here.]
-Table 9: Prompt format for gathering LLMs’ responses to personality-related behavioral questions.
-Appendix G Appendix: Additional Experiment Settings
-Pearson Correlation Coefficient
-The correlation calculation formula is as follows:
-$r=\frac{\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}}$
-$X_{i}$ and $Y_{i}$ represent the data values for LLMs and human.
-The symbols $\overline{X}$ and $\overline{Y}$ represent the means of variables $X$ and $Y$, respectively.
-The numerator, $\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})$, represents the covariance between $X$ and $Y$.
-The denominator, $\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}$, standardizes the result, constraining the value of $r$ to range between -1 and 1.
-When $r$ is close to 1, it indicates a strong positive correlation between the two variables; when $r$ is close to -1, it indicates a strong negative correlation; and when $r$ is close to 0, it indicates no significant linear relationship between the two variables.
-Persona Method
-The Persona method is based on the Persona-Chat dataset constructed by Zhang et al.
-(2018).
-The dataset consists of persona descriptions, and each is made up of five short sentences containing demographic information collected through Amazon Mechanical Turk crowdsourcing.
-To avoid sentence similarity or repetition, these persona descriptions were required to be rewritten (e.g., changing “I am very shy” to “I am not a social person”).
-Zhang et al.
-(2018) demonstrated through machine learning model validation and human evaluations that such persona descriptions provide an effective method to enhance personalization.
-Currently, incorporating personal profiles into prompts is widely used in research related to LLM agents (Park et al., 2023; Wang et al., 2023; Xi et al., 2023).
-In the current study, we treated each persona description as an individual entity (i.e., a single subject) and randomly selected 300 persona descriptions from the dataset.
-One example is, “I wear a lot of leather.
-I have boots I always wear.
-I sleep in late during the day.
-I listen to metal music.
-I have black spiky hair.”
-Shape Method
-The Shape method is based on the work of (Serapio-García et al., 2023), who introduced a prompting approach to shape synthetic personality in LLMs along desired dimensions.
-The researchers expanded upon Goldberg (1990) lexical hypothesis, expanding his list of 70 bipolar adjectives (Goldberg, 1992) to include 104 trait adjectives.
-Additionally, they employed linguistic qualifiers commonly used in Likert-type scales (Likert, 1932), such as “a bit,” “very,” and “extremely,” to set target levels for each adjective.
-This resulted in a fine-grained prompting method with nine levels: 1.
-extremely low adjective; 2.
-very low adjective; 3.
-low adjective; 4.
-a bit low adjective; 5.
-neither low adjective nor high adjective; 6.
-a bit high adjective; 7.
-high adjective; 8.
-very high adjective; 9.
-extremely high adjective.
-In our study, each prompt involves five randomly selected adjective markers from a specific personality domain.
-These markers are positioned after a consistent linguistic qualifier to set the prompt at one of nine intensity levels.
-For example, one prompt is: “You are extremely friendly, extremely energetic, extremely assertive, extremely bold, and extremely active.” These are five positively-keyed adjectives targeting extraversion.
-In this case, the prompt seeks to create a highly sociable and dynamic personality profile, which might result in responses characterized by enthusiasm, confidence, and proactivity.
-We also randomly select 300 prompts here.
-CFA Model
-The basic form of the CFA model is:
-$y=\Lambda\eta+\epsilon$
-where $y$ represents the vector of observed variables; $\Lambda$ is the factor loading matrix (i.e., the loadings of each observed variable on the latent factors); $\eta$ is the vector of latent factors; $\epsilon$ is the vector of error terms, with the assumption that the error terms have a mean of zero and are mutually independent (Jöreskog, 1969).
-Latent factors are variables that are not directly observed but are inferred from other variables that are observed (measured).
-In the context of a CFA model, latent factors represent underlying constructs or traits that are believed to influence the observed variables.
-For example, in psychology, a latent factor might represent a construct like “intelligence” or “anxiety,” which cannot be measured directly but can be estimated through related observed behaviors or responses on a test.
-Model Fit Information
-Chi-Square Test, $\mathbf{\chi^{2}}$: The chi-square test is used to measure the difference between the observed covariance matrix and the factor model’s fitted covariance matrix.
-$\chi^{2}=(N-1)\times F_{\text{ML}}$
-where $N$ is the sample size; $F_{\text{ML}}$ is the value of the fit function under maximum likelihood estimation.
-A smaller chi-square value indicates a better model fit.
-However, with large samples, the chi-square value tends to be large, so other fit indices are usually the primary reference.
-Degrees of Freedom, df: The df represent the relationship between model parameters and observed variables:
-$df=\frac{p(p+1)}{2}-q$
-where $p$ is the number of observed variables, and $q$ is the number of model parameters.
-Degrees of freedom reflect the amount of independent information available in a statistical model to estimate its parameters.
-They are calculated as the number of information points provided by the data minus the number of model parameters, representing the extent to which the model can adjust freely.
-Therefore, a higher degree of freedom indicates fewer parameters in the model and fewer constraints on the data.
-With more degrees of freedom, the model has fewer restrictions, though the fitting difficulty may increase.
-Too few degrees of freedom may lead to overfitting, while too many can result in underfitting.
-Comparative Fit Index, CFI: The CFI is used to compare the goodness of fit of a model with a baseline model (usually an independent model).
-$\text{CFI}=1-\frac{\max(\chi^{2}-\textit{df},0)}{\max(\chi_{\text{null}}^{2}-\textit{df}_{\text{null}},0)}$
-where $\chi^{2}$ and df are the chi-square value and degrees of freedom of the target model; $\chi_{\textit{null}}^{2}$ and $\textit{df}_{\text{null}}$ are the chi-square value and degrees of freedom of the baseline (independent) model.
-Tucker-Lewis Index, TLI: TLI, also known as the Non-Normed Fit Index (NNFI), considers model complexity.
-$\text{TLI}=\frac{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-\left(\chi^{2}/\textit{df}\right)}{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-1}$
-This value ranges from 0 to 1, with a value typically greater than.90 indicating good model fit.
-Root Mean Square Error of Approximation, RMSEA: The RMSEA quantifies the error per degree of freedom in a model, with smaller values indicating better model fit.
-$\text{RMSEA}=\frac{\chi^{2}-df}{df(N-1)}$
-where $\chi^{2}$ is the chi-square value of the model; df is the degrees of freedom; N is the sample size.
-Standardized Root Mean Square Residual, SRMR: The SRMR measures the discrepancy between model-predicted values and actual observed values, calculated as:
-$\text{SRMR}=\sqrt{\frac{\sum_{i=1}^{p}\sum_{j=1}^{p}\left(s_{ij}-\hat{s}_{ij}\right)^{2}}{\frac{p(p+1)}{2}}}$
-where $s_{ij}$ is an element in the observed covariance matrix; $\hat{s}_{ij}$ is an element in the model-fitted covariance matrix.
-Based on Hu and Bentler (1999), CFI/TLI $\geq$ .95, RMSEA $\leq$ .06, and SRMR $\leq$ .08 are considered good fit thresholds.
-Tucker’s Congruence Coefficient
-Tucker’s congruence coefficient, also known as the coefficient of congruence, is typically used to assess the similarity between two-factor structures in factor analysis (Tucker, 1951).
-The formula is given by:
-$\phi=\frac{\sum_{i=1}^{n}a_{i}b_{i}}{\sqrt{\sum_{i=1}^{n}a_{i}^{2}\cdot\sum_{i=1}^{n}b_{i}^{2}}}$
-where $a_{i}$ and $b_{i}$ are the loadings of the $i$-th factor for two different factor solutions (or different samples or methods); $n$ is the total number of factors.
-The coefficient ranges from -1 to 1, where values close to 1 indicate high similarity (congruence) between the factor solutions, values close to 0 indicate low similarity, and negative values indicate a dissimilar or inverse relationship.
-According to Lorenzo-Seva and Ten Berge (2006), a TCC above .95 indicates good similarity, while a TCC of .85 to .94 suggests fair similarity.
-However, this is a relatively lenient criterion; specific differences still need to be determined based on the factor loadings.
-Behavior Variable Measures
+Table 4: BFI-2 facet level with item numbers.
+A.2 Behavioral Measures Scales
 Organizational Citizenship Behavior (OCB): OCB was measured using ten items from Spector and Fox (2010) to assess extra-role behaviors.
 Items were rated on a frequency scale ranging from 1 (never) to 5 (every day).
-Example item: “In the past year, how often have you helped new employees get oriented to the job?”.
-Internal consistency was Cronbach’s alpha = .83.
+Example item: ‘‘In the past year, how often have you helped new employees get oriented to the job?’’.
+Internal consistency was Cronbach’s alpha = 0.83.
 Counterproductive Work Behavior (CWB): CWB was measured using ten items from Spector and Fox (2010), designed to assess harmful workplace behaviors.
 Items were rated on a frequency scale ranging from 1 (never) to 5 (every day).
-Example item: “In the past year, how often have you ignored someone at work?”.
-Internal consistency was Cronbach’s alpha = .86.
+Example item: ‘‘In the past year, how often have you ignored someone at work?’’.
+Internal consistency was Cronbach’s alpha = 0.86.
 The detailed scale information for both can be found in Table 5 and Table 6 below.
-Instructions: Recall that in the past year, how often have you done the following things (1 = Never, 2 = Once or twice, 3 = Once or twice per month, 4 = Once or twice per week, 5 = Every day)?
 # Statement
@@
 Table 6: CWB Scale
+Instructions: Recall that in the past year, how often have you done the following things (1=Never, 2=Once or twice, 3=Once or twice per month, 4=Once or twice per week, 5=Every day)?
+Appendix B Appendix: Development and Validation of Personality-Structured Interview Questions
+B.1 Questions and Development Process
+Table 7 presents the final set of 32 questions that form the basis of our personality structured interview.
+This framework was developed by adapting and modifying McAdams’s life history interview and narrative identity approach (McAdams, 1995, 1996, 2001), while also incorporating components from the Structured Interview of the Five-Factor Model (SIFFM; Trull et al., 1998).
+We adapted and modified the life narrative identity approach proposed by McAdams (1995, 1996, 2001), incorporating elements from the Structured Interview for the Five-Factor Model (SIFFM; Trull et al., 1998).
+The initial pool of interview questions was drafted and refined through collaborative discussions among SMEs, including two doctoral students, one postdoctoral researcher, and one professor, all specializing in Personality Psychology.
+Following this, we conducted pilot testing with six undergraduate research assistants affiliated with a personality research laboratory.
+Based on their feedback, the interview questions were further revised by SMEs.
+Once the final set of questions was determined, field testing was conducted with a convenience sample of university participants.
+During field testing, the other three doctoral students in Industrial/Organizational and Personality Psychology served as independent observers: they reviewed the interview transcripts and provided observer ratings of participants’ personalities to validate the reliability and validity of the structured personality interview questions.
+If the evaluated metrics do not meet the predefined thresholds, the questions are further revised.
+This iterative process of development and feedback led to the construction of the personality structured interview.
+The development process is similar to the development of psychological tests or scales. We have provided more details on psychometrics and the development framework in Appendix G.
+# Questions
+1 To get us started, where are you from? Where did you grow up and what was the place like?
+2 Thinking back, what kind of student were you in school?
+3 Did you have a teacher or teachers that were influential? If so, why? What were they like?
+4 What was your favorite subject in school, and why?
+5 What was your least favorite subject in school, and why?
+6 Still thinking back, who were your heroes when you were young and why?
+7 When you were little, what did you want to be when you grew up? And why?
+8 What were your dreams and plans when you graduated from high school? What made you have those dreams or plans?
+9 If you had complete freedom, what would your dream job be, and why?
+10 How have your dreams and goals changed throughout your life?
+11 Shifting gears to your childhood, how would you describe the personalities of people in the family you grew up in? For example, what were your parents and/or siblings like?
+12 How are you similar or different from your parents and/or siblings?
+13 How do you think your similarities and/or differences influenced your relationship with them?
+14 What was the best part of your childhood?
+15 What do you think were the worst parts of your childhood?
+16 Switching gears a little bit, what was your first paid job? How old were you then? (If this is not applicable to you, then please put ‘NA’)
+17 What other jobs have you had? (If this is not applicable to you, then please put ‘NA’)
+18 What do you do now for a living? And why did you choose it?
+19 Please describe your typical work day.
+20 What is the best and worst part of your current work?
+21 Did you serve in the military? Please tell us about that experience, what was the best and worst part of it?
+22 Moving on, what are your adult friendships like?
+23 How are your adult friendships different from your childhood friendships?
+24 What are your strongest qualities as a friend? In other words, what makes you a great friend to have?
+25 What about your weakest qualities in friendships? In other words, what do you struggle with when you are trying to be a friend to someone?
+26 Moving onto more general questions, when thinking about your life in general, what are you most proud of?
+27 What hobbies or other interests do you have?
+28 What things frighten you now?
+29 What were some things that frightened you most as a child?
+30 What are the three biggest news events that have occurred in your lifetime?
+31 If you had the power to solve one and only one problem in the world, what would it be, and why?
+32 Tell me about a time when you did not know if you would make it. How did you overcome that challenge?
+Table 7: Personality Structured Interview (PSI) questions.
+B.2 Validation Results
+As previously described, three doctoral students specializing in Industrial/Organizational and Personality Psychology served as observers.
+They read the interview transcripts and rated participants’ personalities using an other-report format of the 15-item Extra Short Form of the Big Five Inventory–2 (BFI-2-XS;  Soto and John, 2017a).
+To adapt the BFI-2-XS from self-report to observer-report, raters evaluated the extent to which each item (e.g., ‘‘The participant worries a lot’’) appeared characteristic of the participant based on the transcript.
+The BFI-2-XS was selected due to its brevity and efficiency, which reduced the time burden on raters.
+The BFI-2-XS has demonstrated strong convergent validity with the full BFI-2 in both self- and observer-report formats, with trait-level correlations exceeding r = 0.85 (Soto and John, 2017a).
+These findings support its use as a brief yet psychometrically sound instrument for personality assessment.
+Observer ratings were averaged across the three raters, and interrater reliability was assessed.
+All five traits demonstrated adequate agreement (see Table 8), with more behaviorally salient traits such as conscientiousness and neuroticism showing higher intraclass correlations (ICCs), and less observable traits such as openness showing lower ICCs, consistent with prior findings (Allik et al., 2010).
+The average interrater reliability (ICC) for observer ratings was 0.76, exceeding the commonly accepted threshold of 0.70.
+As shown in Table 8, all convergent correlations between self- and observer-rated personality traits were positive and statistically significant, averaging 0.36 across the Big Five traits.
+These findings align with meta-analytic estimates of self–observer convergence (Connolly et al., 2007).
+Discriminant correlations among the five traits averaged 0.25 for self-reports and 0.18 for observer-reports, suggesting appropriate trait differentiation.
+Self-report Observer-rated
+$\mu$ $\sigma$ Ext Agr Con Neu Ope Ext Agr Con Neu Ope
+Self-report Ext 3.22 0.69 0.84
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 8: Correlation matrix of self- and observer-scored big five personality.
+Ext = extroversion.
+Agr = agreeableness.
+Con = conscientiousness.
+Neu = neuroticism.
+Ope = openness.
+Gray values indicate non-significant correlations.
+The diagonal reports reliabilities in italics, using Cronbach’s alpha for self-ratings and intraclass correlations for observer ratings.
+Red highlights indicate convergent correlations between self- and observer-reported scores.
+Light yellow and dark yellow highlights represent discriminant correlations for self-report and observer-rated personality scores, respectively.
+Appendix C Appendix: Personality Structured Interview Dataset and Data Collection
+The data were collected through an online questionnaire, followed by an online structured interview.
+We will share the dataset that we have obtained permission to share, which also excludes any personally identifiable information.
+This portion of the data is available upon request and may not be used for any commercial purposes.
+Academic use is permitted only with prior approval.
+Each example is composed of the following characteristics:
+  1. 1.
+    Gender: The gender of the participant.
+  2. 2.
+    Race: The racial background of the participant.
+  3. 3.
+    English: Whether English is the participant’s first language.
+  4. 4.
+    Age: The participant’s age.
+  5. 5.
+    Weight: The participant’s weight.
+  6. 6.
+    Height: The participant’s height.
+  7. 7.
+    OCB1–OCB10: Self-reported Organizational Citizenship Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 5 for specific item details.
+  8. 8.
+    CWB1–CWB10: Self-reported Counterproductive Work Behavior data, measured using the scale from  Spector and Fox (2010), and see Table 6 for specific item details.
+  9. 9.
+    Q1–Q32: Participant responses to each personality structured interview questions, see Table 7 for specific question details.
+  10. 10.
+    Item1–Item60: Responses to each item of the BFI-2 (already reverse coded). Refer to Appendix A for item descriptions and scoring guidelines.
+Appendix D Appendix: Prompts List
+D.1 Personality Scale Responses Prompt Format
+To minimize the interference of the prompt template on model behavior, we use a standardized prompt structure, defined as: $\text{Prompt}(d)=T_{\text{base}}+d$, where $T_{\text{base}}$ is a fixed base template: $T_{\text{base}}$ = ‘‘For the following task, respond in a way that matches:’’ The variable $d$ represents the personality description content, determined by the method used.
+Specifically: $d\in\{d_{\text{Persona}},d_{\text{Shape}},d_{\text{PSI}}\}$.
+The construction of $d$ varies across methods.
+For the Persona method, $d_{\text{Persona}}=\textstyle\sum_{i=1}^{5}s_{i}$, where each $s_{i}$ is a short sentence from the Persona-Chat dataset, such as: ‘‘I wear a lot of leather.’’ Each $d_{\text{Persona}}$ represents a unique individual profile.
+For the Shape method, $d_{\text{Shape}}=\textstyle\sum_{i=1}^{5}q(a_{i})$, where $a_{i}$ is an adjective representing a personality trait (e.g., friendly, energetic), and $q(\cdot)$ is a linguistic intensity modifier (e.g., extremely, very).
+For the PSI method, $d_{\text{PSI}}=\textstyle\sum_{i=1}^{32}(\texttt{[Q}_{i}\texttt{]}+\texttt{[A}_{i}\texttt{]})$, where $\texttt{Q}_{i}$ denotes the $i$-th interview question and $\texttt{A}_{i}$ is the corresponding response.
+Table 9 presents the prompts used to generate LLM-simulated responses for the selected personality test (BFI-2), formatted on a Likert scale.
+In each prompt, personality_description denotes the personality-specific framing, while test_item refers to an individual item from the BFI-2 scale.
+Example prompts for each method are also included in Table 9.
+PSI Method: In the PSI method, the personality prompt integrates both the interview question and the interviewee’s corresponding response, as collected in the PSI dataset (see Appendix C).
+Persona Method: The Persona method is based on the Persona-Chat dataset constructed by Zhang et al.
+(2018).
+The dataset consists of persona descriptions, and each is made up of five short sentences containing demographic information collected through Amazon Mechanical Turk crowdsourcing.
+To avoid sentence similarity or repetition, these persona descriptions were required to be paraphrased (e.g., changing ‘‘I am very shy’’ to ‘‘I am not a social person’’).
+Zhang et al.
+(2018) demonstrated through machine learning model validation and human evaluations that such persona descriptions provide an effective method to enhance personalization.
+Currently, incorporating personal profiles into prompts is widely used in research related to LLM agents (Park et al., 2023; Wang et al., 2023; Xi et al., 2023).
+In our study, we treated each persona description as an individual entity (i.e., a single subject) and randomly selected 300 persona descriptions from the dataset.
+One example is, ‘‘I wear a lot of leather.
+I have boots I always wear.
+I sleep in late during the day.
+I listen to metal music.
+I have black spiky hair.’’
+Shape Method: The Shape method is based on the work of Serapio-García et al.
+(2023), who introduced a prompting approach to shape synthetic personality in LLMs along desired dimensions.
+The researchers expanded upon Goldberg (1990), expanding his list of 70 bipolar adjectives (Goldberg, 1992) to include 104 trait adjectives.
+Additionally, they employed linguistic qualifiers commonly used in Likert-type scales (Likert, 1932), such as ‘‘a bit’’, ‘‘very’’, and ‘‘extremely’’ to set target levels for each adjective.
+This resulted in a fine-grained prompting method with nine levels: 1.
+extremely low adjective; 2.
+very low adjective; 3.
+low adjective; 4.
+a bit low adjective; 5.
+neither low adjective nor high adjective; 6.
+a bit high adjective; 7.
+high adjective; 8.
+very high adjective; 9.
+extremely high adjective.
+In our study, each prompt involves five randomly selected adjective markers from a specific personality domain.
+These markers are positioned after a consistent linguistic qualifier to set the prompt at one of nine intensity levels.
+For example, one prompt is: ‘‘You are extremely friendly, extremely energetic, extremely assertive, extremely bold, and extremely active.’’ These are five positively keyed adjectives targeting the factor (trait) of Extraversion.
+In this case, the prompt seeks to create a highly sociable and dynamic personality profile, which might result in responses characterized by narrower traits (facets) of enthusiasm, confidence, and proactivity.
+We also randomly select 300 prompts.
+D.2 Behavioral Question Responses Prompt Format
+The complete prompt format for eliciting LLMs’ responses to personality-related behavioral questions is defined in Table 10.
+The questions_and_responses refers to the transcript of a structured human personality interview, while question_list comprises statements evaluating OCB and CWB, as detailed in Table 5 and Table 6 in Appendix A.
+The model is instructed to rate each statement using a frequency scale from one to five.
+Table 9: Prompt format and examples for gathering LLMs’ responses to BFI-2 scale.
+Table 10: Prompt format for gathering LLMs’ responses to personality-related behavioral questions.
+D.3 Compute Resources
+All LLMs were conducted using the OpenAI and TogetherAI platforms, running a total of seven models: Mistral-7B (Jiang et al., 2023), Gemma-2-9B, Gemma-2-27B (GemmaTeam, 2024), Llama3-8B, Llama3-70B (AI@Meta, 2024), GPT-4o-mini (OpenAI, 2024a), and GPT-4o (OpenAI, 2024b).
+Using the GPT-4o tokenizer as an illustrative example, the approximate input token counts for generating BFI-2 LLM responses are as follows: Persona method: 2,555,820 tokens; Shape method: 2,316,240 tokens; and PSI method: 34,510,812 tokens.
+The total number of output tokens for the BFI-2 LLM responses is approximately 25,280.
+For the PSI method, the input tokens used to generate responses to the OCB and CWB scales were approximately 690,497 and 681,215, respectively, with output tokens totaling around 153,887 and 139,207.
+Token counts may vary slightly across models due to differences in tokenizers, formatting, or prompt structures, but they generally fall within a comparable range.
+Appendix E Appendix: Evaluation Metrics and Formulas
+E.1 Pearson Correlation Coefficient
+The formula for correlation is as follows:
+$r=\frac{\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})}{\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}}$
+$X_{i}$ and $Y_{i}$ represent the data values for LLMs and human.
+The symbols $\overline{X}$ and $\overline{Y}$ represent the means of variables $X$ and $Y$, respectively.
+The numerator, $\sum(X_{i}-\overline{X})(Y_{i}-\overline{Y})$, represents the covariance between $X$ and $Y$.
+The denominator, $\sqrt{\sum(X_{i}-\overline{X})^{2}\sum(Y_{i}-\overline{Y})^{2}}$, standardizes the result, let the value of $r$ range between -1 and 1.
+When $r$ is close to 1, it indicates a strong positive correlation between the two variables; when $r$ is close to -1, it indicates a strong negative correlation; and when $r$ is close to 0, it indicates no significant linear relationship between the two variables.
+E.2 CFA Model
+The basic form of the CFA model is:
+$y=\Lambda\eta+\epsilon$
+where $y$ represents the vector of observed variables; $\Lambda$ is the factor loading matrix (i.e., the loadings of each observed variable on the latent factors); $\eta$ is the vector of latent factors; $\epsilon$ is the vector of error terms, with the assumption that the error terms have a mean of zero and are mutually independent (Jöreskog, 1969).
+Latent factors are variables that are not directly observed but are inferred from other variables that are observed (measured).
+In the context of a CFA model, latent factors represent underlying constructs or traits that are believed to influence the observed variables.
+For example, in psychology, a latent factor might represent a construct like “intelligence” or “anxiety,” which cannot be measured directly but can be estimated through related observed behaviors or responses on a test.
+E.3 Model Fit Information
+Chi-Square Test, $\mathbf{\chi^{2}}$: The chi-square test is used to measure the difference between the observed covariance matrix and the factor model’s fitted covariance matrix.
+$\chi^{2}=(N-1)\times F_{\text{ML}}$
+where $N$ is the sample size; $F_{\text{ML}}$ is the value of the fit function under maximum likelihood estimation.
+A smaller chi-square value indicates a better model fit.
+However, with large samples, the chi-square value tends to be large, so other fit indices are usually the primary reference.
+Degrees of Freedom, df: The df represents the relationship between model parameters and observed variables:
+$df=\frac{p(p+1)}{2}-q$
+where $p$ is the number of observed variables, and $q$ is the number of model parameters.
+Degrees of freedom reflect the amount of independent information available in a statistical model to estimate its parameters.
+They are calculated as the number of information points provided by the data minus the number of model parameters, representing the extent to which the model can adjust freely.
+Therefore, a higher degree of freedom indicates fewer parameters in the model and fewer constraints on the data.
+With more degrees of freedom, the model has fewer restrictions, though the fitting difficulty may increase.
+Too few degrees of freedom may lead to overfitting, while too many can result in underfitting.
+Comparative Fit Index, CFI: The CFI is used to compare the goodness of fit of a model with a baseline model (usually an independent model).
+$\text{CFI}=1-\frac{\max(\chi^{2}-\textit{df},0)}{\max(\chi_{\text{null}}^{2}-\textit{df}_{\text{null}},0)}$
+where $\chi^{2}$ and df are the chi-square value and degrees of freedom of the target model; $\chi_{\textit{null}}^{2}$ and $\textit{df}_{\text{null}}$ are the chi-square value and degrees of freedom of the baseline (independent) model.
+Tucker-Lewis Index, TLI: TLI, also known as the Non-Normed Fit Index (NNFI), considers model complexity.
+$\text{TLI}=\frac{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-\left(\chi^{2}/\textit{df}\right)}{\left(\chi^{2}_{\text{null}}/\textit{df}_{\text{null}}\right)-1}$
+This value ranges from 0 to 1, with a value typically greater than.90 indicating good model fit.
+Root Mean Square Error of Approximation, RMSEA: The RMSEA quantifies the error per degree of freedom in a model, with smaller values indicating better model fit.
+$\text{RMSEA}=\frac{\chi^{2}-df}{df(N-1)}$
+where $\chi^{2}$ is the chi-square value of the model; df is the degrees of freedom; N is the sample size.
+Standardized Root Mean Square Residual, SRMR: The SRMR measures the discrepancy between model-predicted values and actual observed values, calculated as:
+$\text{SRMR}=\sqrt{\frac{\sum_{i=1}^{p}\sum_{j=1}^{p}\left(s_{ij}-\hat{s}_{ij}\right)^{2}}{\frac{p(p+1)}{2}}}$
+where $s_{ij}$ is an element in the observed covariance matrix; $\hat{s}_{ij}$ is an element in the model-fitted covariance matrix.
+Based on Hu and Bentler (1999), CFI/TLI $\geq$ 0.95, RMSEA $\leq$ 0.06, and SRMR $\leq$ 0.08 are considered good fit thresholds.
+E.4 Tucker’s Congruence Coefficient
+Tucker’s congruence coefficient, also known as the coefficient of congruence, is typically used to assess the similarity between two-factor structures in factor analysis (Tucker, 1951).
+The formula is given by:
+$\phi=\frac{\sum_{i=1}^{n}a_{i}b_{i}}{\sqrt{\sum_{i=1}^{n}a_{i}^{2}\cdot\sum_{i=1}^{n}b_{i}^{2}}}$
+where $a_{i}$ and $b_{i}$ are the loadings of the $i$-th factor for two different factor solutions (or different samples or methods); $n$ is the total number of factors.
+The coefficient ranges from -1 to 1, where values close to 1 indicate high similarity (congruence) between the factor solutions, values close to 0 indicate low similarity, and negative values indicate a dissimilar or inverse relationship.
+According to Lorenzo-Seva and Ten Berge (2006), a TCC above 0.95 indicates good similarity, while a TCC of 0.85 to 0.94 suggests fair similarity.
+However, this is a relatively lenient criterion; specific differences still need to be determined based on the factor loadings.
+Appendix F Appendix: The Structure of Personality
+A key question in personality-related LLM research pertains to personality structure: What is the nature and breadth of the personality traits we want to simulate the human personality distributions?
+In the research literature, personality structures often emerge from applying factor analysis to individuals’ responses to a large number of personality-relevant items.
+This approach is what has been used to identify the five personality factors in the FFM.
+Moreover, personality is better understood in terms of one’s continuous standing on each of multiple dimensions rather than as static types or profiles.
+Research data clearly supports this view (Wilmot, 2015; Wilmot et al., 2019).
+Dividing individuals into limited categories (e.g., 16 types in MBTI) artificially segments continuous dimensions into discrete units, which may overlook important individual differences (Ones and Wiernik, 2018).
+Another question worth paying attention to is that, although the FFM occupies a dominant position in personality research, it remains debatable whether focusing solely on the FFM is sufficient in light of the many newly proposed personality variables and alternative structures.
+It turns out that there can be conceptual overlap among these models (e.g., Hough et al., 2015).
+For example, a meta-analysis by Joseph and Newman (2010) revealed that emotional intelligence (EI) shows statistically and practically significant relationships with neuroticism and extroversion within the FFM.
+In fact, when controlling for personality variables, the unique contribution of EI almost disappears.
+Similarly, Credé et al.
+(2017) conducted a meta-analysis on grit and found that its core components can largely be explained by conscientiousness, with little added predictive validity beyond that.
+Moreover, although the HEXACO structure introduces an honesty-humility dimension, the remaining five dimensions align closely with the FFM structure (Lee and Ashton, 2004, 2006).
+Research by Cutler and Condon (2023) further indicates that nearly all personality semantic information can be classified within the FFM structure.
+However, this does not mean that the FFM has exhausted the depiction of personality.
+On the contrary, although the FFM performs well in terms of structure and predictive power, it may still miss some more nuanced aspects of personality.
+For instance, measurement tools such as the Facet MAP (Irwing et al., 2024) attempt to capture “micro-level differences” in personality by focusing on the lower-level structure of each dimension.
+Therefore, while the FFM is “good,” there remains room for further refinement in exploring the complexity and diversity of personality.
+Rather than pursuing a wide range of personality frameworks, this paper will delve deeper into the FFM structure, which is widely accepted among personality psychologists.
+Appendix G Appendix: Psychometrics and Structured Interview Development Framework
+Psychometrics is a field of psychology dedicated to the theory and practice of psychological measurement.
+It primarily focuses on quantifying psychological traits, behaviors, and abilities through systematic testing and analysis.
+Psychological traits, such as personality dimensions, cognitive abilities, and emotional states, are inherently abstract constructs that cannot be measured directly.
+Therefore, psychometricians rely on tools such as surveys, questionnaires, scales, or interviews to infer these traits through observable indicators or responses.
+G.1 Measuring Psychological Traits
+To measure a psychological trait, psychometricians typically operationalize the trait by identifying observable behaviors or self-identities that correlate with the underlying construct.
+For example, extraversion can be measured by assessing behaviors such as sociability, assertiveness, and enjoyment of social interactions, or by examining identities such as seeing oneself as an outgoing person and believing that one thrives in social situations.
+These behaviors are translated into measurable items (for scale; e.g., ‘‘I enjoy being the center of attention’’) or questions (for interview; e.g., ‘‘What are your strongest qualities as a friend?
+In other words, what makes you a great friend to have?’’).
+The challenge lies in ensuring that these items/questions accurately and consistently capture the construct across different populations and contexts.
+G.2 Theory-Informed Structured Interview Transcripts for LLM Data Simulation
+Theory-informed structured interviews are the most suitable method for enabling LLMs to simulate psychometric data.
+These interviews are specifically designed to capture the constructs underlying the targeted psychometric measures, ensuring that the simulated data aligns with the intended psychological construct.
+By extracting textual information that directly reflects the target construct, theory-informed structured interviews facilitate the representation of heterogeneous data while preserving a high degree of human diversity, thereby enhancing the validity and applicability of the simulated psychometric data.
+Moreover, since the information is extracted based on theoretical foundations, it also provides a certain level of interpretability for the LLM’s simulation.
+This not only allows the generated data to be compared with theoretical expectations but also increases its potential for practical applications.
+G.3 The Potential for Advancing Research
+A theory-informed structured interview transcript-based simulation can generate data more effectively by focusing on the target construct.
+Ideally, the simulated data should reproduce the same constructs reflected in real-world data and simulate behaviors associated with these constructs.
+Take personality as an example—if simulated data can accurately replicate real-world personality constructs, it enables research that would be difficult to conduct in reality, such as developing contextualized personality assessment tools and exploring new personality theories through multi-agent simulations.
+Developing Contextualized Personality Assessment Tools: Traditional personality assessments mainly rely on standardized questionnaires or laboratory tasks, which often fail to adequately simulate real-world social contexts.
+By using theory-informed structured interview transcript-based simulations, we can generate more fine-grained and context-sensitive individual response data.
+For instance, we can simulate various occupational scenarios (such as crisis management, teamwork, or remote work) and analyze how different personality traits manifest in these contexts.
+This approach not only aids in developing measurement tools tailored to specific applications but also enhances ecological validity, allowing for more accurate assessments of personality across different situations.
+Exploring New Personality Theories through Multi-Agent Simulation: If simulated data can accurately reflect real-world personality constructs, we can leverage multi-agent interactive systems to simulate individuals’ behavioral patterns and observe how different personality traits evolve in group dynamics.
+For example, virtual agents with distinct personality traits can be placed in cooperative tasks, competitive environments, or social interactions, enabling researchers to test whether existing personality theories effectively predict these interaction patterns.
+Additionally, this approach can uncover new personality dynamics, such as whether certain personality trait combinations produce unexpected group effects or whether behavior in specific situations deviates from traditional theoretical predictions.
+However, these assumptions are based on an ideal premise—that simulated data can successfully reflect the same constructs as real-world data.
+A theory-informed structured interview undoubtedly offers a promising pathway in this regard, warranting further in-depth exploration.
+G.4 Structured Interview Development Framework
+Developing a structured interview for LLM to simulate data involves a series of carefully designed steps to ensure that the resulting test reliably and validly measures the construct of interest.
+Table 11 outlines the framework.
+# Steps
+1 Identify behaviors/perceptions that represent the construct or define the domain.
+2 Prepare a set of structured interview specifications, structured interview blueprint.
+3 Build an initial question pool.
+4 Have questions reviewed by substantive experts (and revise as necessary).
+5 Hold preliminary question tryouts.
+6 Determine statistical properties of questions (and eliminate poor questions or revise as necessary).
+7 Field-test the structured interview on a large representative sample of the intended examinee population.
+8 Design and conduct reliability and validity studies for the final form of the structured interview.
+9 Develop guidelines for administration and interpretation of the structured interview.
+Table 11: Structured interview development framework.
+Identify behaviors/perceptions that represent the construct or define the domain: Clearly defining the construct is essential to developing relevant structured interview questions.
+The construct should be operationalized by identifying specific behaviors or attributes that indicate its presence.
+In other words, you need to find a theory to guide you on how to measure the target constructs.
+This step may involve reviewing literature, conducting expert interviews, or organizing focus groups to understand the various dimensions and observable characteristics of the construct.
+Prepare a set of structured interview specifications, structured interview blueprint: A blueprint outlines the structured interview’s structure and content, specifying how questions are distributed across the construct’s dimensions.
+It typically includes the number of questions per domain and the content of the questions.
+Build an initial question pool: In this step, an extensive list of questions is created to cover the full range of the construct.
+Question wording should be clear, concise, and relevant to the target population.
+It is common practice to generate more questions than needed to ensure that poorly performing questions can be removed later without compromising the structured interview.
+Have questions reviewed by substantive experts (and revised as necessary): SMEs review the question pool for content accuracy, relevance, clarity, and bias.
+Experts assess whether the questions align with the construct’s definition and whether any important aspects are missing.
+Feedback from experts helps refine the wording, remove ambiguous questions, and identify questions with potential cultural or gender biases.
+Hold preliminary question tryouts: Before large-scale testing, questions are piloted on a small group of individuals representative of the target population.
+This stage helps identify any immediate issues with question comprehension, response format, or instructions.
+It can also include cognitive interviews where participants are asked to explain their thought processes when answering questions.
+The feedback from this stage informs further revisions, ensuring that questions are clear and easy to understand.
+Determine statistical properties of questions (and eliminate poor questions or revise as necessary): Question performance is assessed through statistical analyses to evaluate difficulty, discrimination, and internal consistency.
+Standardized scoring criteria, such as Behaviorally Anchored Rating Scales (BARS), can be used for obtaining question scores.
+Then, methods like correlations, factor analysis, and item response theory (IRT) can help determine how effectively each question measures the intended construct.
+Questions that demonstrate poor psychometric properties—such as low discrimination or high measurement error—are either revised or removed.
+For example, questions with low correlations with the overall score or incorrect factor loadings may be eliminated.
+Field-test the structured interview on a large representative sample of the intended examinee population: The revised item set is administered to a large, representative sample to gather comprehensive data on the scale’s performance.
+This step ensures that the sample reflects the population for which the test is intended, which is critical for generalizability.
+Statistical analyses are conducted to refine the test further.
+This process may involve removing redundant items, assessing dimensionality, and ensuring that items work well across demographic subgroups.
+Design and conduct reliability and validity studies for the final form of the structured interview: To ensure the structured interview is psychometrically sound, various reliability and validity studies are conducted: Reliability studies measure the structured interview’s consistency and stability, including internal consistency (e.g., Cronbach’s alpha), test-retest reliability, and inter-rater reliability (if applicable).
+Validity studies assess whether the structured interview measures what it is intended to measure.
+This includes: Content validity (the extent to which items cover the construct); Construct validity (e.g., convergent and discriminant validity); Criterion-related validity (e.g., predictive or concurrent validity).
+These studies provide evidence that the structured interview is both reliable and valid for its intended purpose.
+Develop guidelines for administration and interpretation of the structured interview: The final step involves creating a comprehensive, structured interview manual that includes instructions for structured interview administration, scoring procedures, and guidelines for interpreting results.
+This manual ensures consistency in the structured interview’s use across different settings and helps minimize errors in administration and scoring.
+Guidelines for interpreting scores may include norms, cutoff points, and descriptions of what various score ranges indicate.
 Appendix H Appendix: Psychometric Data Evaluation Framework
@@
 Below, we outline each component and the rationale for its inclusion.
-Descriptive Statistics
+H.1 Descriptive Statistics
 Descriptive statistics are primarily used to summarize, outline, and present the basic features of data.
 They help researchers understand the distribution and fundamental trends of the data without interpreting or measuring specific psychological constructs.
-Do not forget that the evaluation of psychometric data is hierarchical, we need to evaluate on item and domain level.
-Mean (M): The mean represents the central tendency of the responses, reflecting the average score across individuals.
-Standard Deviation (SD): The standard deviation captures the dispersion of responses, indicating how much variation exists within the data.
-Both M and SD can be quantified for similarity to human distribution using MAE and $r$.
+Do not forget that the evaluation of psychometric data is hierarchical; we need to evaluate it at both the Item and Domain levels.
+Mean ($\mu$): The mean represents the central tendency of the responses, reflecting the average score across individuals.
+Standard Deviation ($\sigma$): The standard deviation captures the dispersion of responses, indicating how much variation exists within the data.
+Both $\mu$ and $\sigma$ can be quantified for similarity to human distribution using MAE and $r$.
 However, these metrics provide a more summarized level of comparison; we also need to examine performance on specific items and the domain.
@@
 High kurtosis (leptokurtic) suggests heavy tails with more extreme values, while low kurtosis (platykurtic) indicates a flatter distribution with fewer extreme values.
-Psychometric Performance
+Figure 6: Psychometric data evaluation framework.
+H.2 Psychometric Performance
 Psychometric performance is primarily used to evaluate the measurement quality of psychological constructs, ensuring that assessments accurately and reliably capture the intended traits.
@@
 Discriminant Validity: Discriminant validity ensures that distinct constructs are not excessively correlated. It can be examined by calculating the mean absolute correlation.
-Figure 2: Psychometric Data Evaluation Framework
 Appendix I Appendix: Additional Analyses and Results
-I.1 Additional Method Comparison Results
-Additional Descriptive Statistics Results
-Here we show the detailed means and standard deviations for human responses and LLM responses at the item, facet, and domain levels, see Tables 1213141516, and 17.
-No. Item Content Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-1 I am someone who is outgoing, sociable 3.03 3.41 2.77 2.76 2.45 2.67 2.14
-2 I am someone who is compassionate, has a soft heart 4.34 3.51 3.87 3.10 3.17 4.05 3.91
-3 I am someone who tends to be disorganized 3.62 3.39 3.77 3.32 3.31 3.34 3.24
-4 I am someone who is relaxed, handles stress well 2.56 3.20 2.94 3.47 3.32 2.91 2.96
-5 I am someone who has few artistic interests 3.68 3.72 3.80 3.27 2.92 2.75 2.73
-6 I am someone who has an assertive personality 2.96 3.08 3.07 2.82 2.81 2.57 2.25
-7 I am someone who is respectful, treats others with respect 4.62 3.38 4.43 3.48 3.62 4.16 4.22
-8 I am someone who tends to be lazy 3.62 4.00 4.23 3.51 3.44 3.54 3.75
-9 I am someone who stays optimistic after a setback 2.44 2.74 2.29 2.95 2.92 2.18 2.29
-10 I am someone who is curious about many different things 4.43 3.89 3.71 3.02 3.15 3.60 3.05
-11 I am someone who rarely feels excited or eager 3.54 4.30 4.18 3.21 2.99 2.98 2.40
-12 I am someone who tends to find fault with others 3.34 3.29 4.31 3.47 3.46 3.58 3.71
-13 I am someone who is dependable, steady 4.33 3.29 3.70 2.93 3.11 4.12 3.96
-14 I am someone who is moody, has up and down mood swings 2.56 3.01 2.52 2.72 2.60 3.04 2.31
-15 I am someone who is inventive, finds clever ways to do things 3.90 3.44 3.24 3.06 2.91 3.05 2.37
-16 I am someone who tends to be quiet 2.35 3.26 3.44 3.02 3.25 2.17 2.51
-17 I am someone who feels little sympathy for others 3.71 3.73 4.47 3.72 3.54 4.22 2.88
-18 I am someone who is systematic, likes to keep things in order 3.96 3.00 2.46 2.94 2.59 3.18 2.44
-19 I am someone who can be tense 3.29 3.23 3.18 3.12 3.04 3.66 2.73
-20 I am someone who is fascinated by art, music, or literature 4.06 3.61 2.99 3.15 2.90 3.43 2.81
-21 I am someone who has an assertive personality (dominant) 2.90 2.94 2.46 2.77 2.53 2.42 2.07
-22 I am someone who starts arguments with others 4.23 3.65 4.56 3.65 3.70 4.44 4.35
-23 I am someone who has difficulty getting started on tasks 3.31 3.29 3.85 3.16 3.23 3.10 2.79
-24 I am someone who feels secure, comfortable with self 2.26 2.75 2.24 3.14 2.65 2.65 2.70
-25 I am someone who avoids intellectual, philosophical discussions 3.93 3.40 3.36 3.23 3.13 2.85 2.57
-26 I am someone who is less active than other people 3.35 3.80 3.86 3.36 3.14 2.65 2.67
-27 I am someone who has a forgiving nature 3.78 3.05 3.48 3.05 3.28 3.51 3.59
-28 I am someone who can be somewhat careless 3.56 3.30 3.19 3.16 3.00 3.06 3.15
-29 I am someone who is emotionally stable, not easily upset 2.44 3.16 2.66 3.17 2.87 2.70 2.67
-30 I am someone who has little creativity 3.95 4.10 4.36 3.59 3.42 3.37 2.85
-31 I am someone who is sometimes shy, introverted 2.35 3.28 3.43 3.11 3.34 2.14 2.70
-32 I am someone who is helpful and unselfish with others 4.18 3.32 3.88 3.27 3.24 4.06 4.12
-33 I am someone who keeps things neat and tidy 3.72 2.99 2.47 3.07 2.63 3.04 2.70
-34 I am someone who worries a lot 3.30 2.96 2.59 2.79 2.50 3.64 3.04
-35 I am someone who values art and beauty 4.16 3.65 3.23 3.14 3.06 3.38 2.87
-36 I am someone who finds it hard to influence people 3.09 3.32 3.44 3.25 2.99 2.74 2.25
-37 I am someone who is sometimes rude to others 3.79 3.24 4.43 3.50 3.63 3.70 4.10
-38 I am someone who is efficient, gets things done 4.23 3.31 3.45 3.07 3.12 3.69 3.07
-39 I am someone who often feels sad 2.71 2.86 2.27 2.74 2.42 3.24 2.69
-40 I am someone who is complex, a deep thinker 4.02 3.64 3.59 3.19 3.14 3.34 2.74
-41 I am someone who is full of energy 3.14 3.45 3.39 2.99 2.86 2.81 2.41
-42 I am someone who is suspicious of others’ intentions 2.75 3.14 3.91 3.27 3.40 3.42 3.60
-43 I am someone who is reliable, can always be counted on 4.32 3.28 3.76 3.12 3.18 4.06 3.75
-44 I am someone who keeps their emotions under control 2.22 3.10 2.92 2.98 2.90 2.50 2.34
-45 I am someone who has difficulty imagining things 4.17 4.06 4.08 3.57 3.43 3.21 2.14
-46 I am someone who is talkative 2.88 3.11 2.53 2.87 2.50 3.31 2.36
-47 I am someone who can be cold and uncaring 3.84 3.95 4.44 3.60 3.47 4.08 4.19
-48 I am someone who leaves a mess, doesn’t clean up 4.15 3.47 4.25 3.39 3.25 3.32 4.03
-49 I am someone who rarely feels anxious or afraid 3.44 3.20 3.12 2.85 2.71 3.48 3.66
-50 I am someone who thinks poetry and plays are boring 3.66 3.54 3.85 3.44 3.15 2.90 2.29
-51 I am someone who prefers to have others take charge 2.96 3.54 4.33 3.39 3.20 3.52 3.87
-52 I am someone who is polite, courteous to others 4.50 3.28 4.13 3.33 3.56 3.97 4.38
-53 I am someone who is persistent, works until the task is finished 4.27 3.56 3.87 3.01 3.13 4.07 3.75
-54 I am someone who tends to feel depressed, blue 2.65 2.70 2.01 2.60 2.36 3.25 2.78
-55 I am someone who has little interest in abstract ideas 3.83 3.25 3.40 3.08 2.93 2.53 2.15
-56 I am someone who shows a lot of enthusiasm 3.35 3.86 4.08 2.81 2.81 2.70 2.58
-57 I am someone who assumes the best about people 3.37 3.12 3.40 3.04 3.13 3.46 3.64
-58 I am someone who sometimes behaves irresponsibly 3.64 3.11 3.44 2.90 3.22 3.13 3.55
-59 I am someone who is temperamental, gets emotional easily 2.53 3.02 2.33 2.80 2.66 2.96 2.04
-60 I am someone who is original, comes up with new ideas 3.84 3.28 3.25 2.96 2.88 3.05 2.40
-Table 12: Item Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-No. Item Content Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-1 I am someone who is outgoing, sociable 1.40 0.87 1.31 1.22 1.37 0.98 0.78
-2 I am someone who is compassionate, has a soft heart 0.90 0.84 1.12 1.22 1.44 0.93 1.03
-3 I am someone who tends to be disorganized 1.36 0.78 0.95 1.17 1.46 0.76 0.96
-4 I am someone who is relaxed, handles stress well 1.30 0.79 1.16 1.16 1.39 0.92 0.96
-5 I am someone who has few artistic interests 1.33 1.04 1.31 1.25 1.32 1.39 1.11
-6 I am someone who has an assertive personality 1.38 0.79 1.18 1.18 1.37 0.76 0.73
-7 I am someone who is respectful, treats others with respect 0.61 0.89 0.92 1.23 1.42 0.75 0.78
-8 I am someone who tends to be lazy 1.31 0.87 1.04 1.28 1.50 0.81 0.97
-9 I am someone who stays optimistic after experiencing a setback 1.24 0.84 0.98 1.34 1.54 0.89 0.93
-10 I am someone who is curious about many different things 0.81 0.89 1.12 1.37 1.55 0.98 0.91
-11 I am someone who rarely feels excited or eager 1.26 0.84 0.75 1.33 1.41 0.93 0.82
-12 I am someone who tends to find fault with others 1.32 0.59 0.69 1.26 1.48 0.83 0.70
-13 I am someone who is dependable, steady 0.86 0.85 1.12 1.33 1.59 0.79 0.75
-14 I am someone who is moody, has up and down mood swings 1.36 0.56 1.09 1.27 1.41 0.76 0.93
-15 I am someone who is inventive, finds clever ways to do things 1.03 0.69 1.14 1.21 1.44 0.63 0.74
-16 I am someone who tends to be quiet 1.34 0.84 1.12 1.12 1.27 1.02 0.97
-17 I am someone who feels little sympathy for others 1.45 0.93 0.79 1.27 1.54 0.68 1.00
-18 I am someone who is systematic, likes to keep things in order 1.06 0.88 1.09 1.16 1.37 0.94 0.81
-19 I am someone who can be tense 1.27 0.75 1.08 1.22 1.20 0.95 0.95
-20 I am someone who is fascinated by art, music, or literature 1.17 0.89 1.44 1.18 1.36 1.18 1.11
-21 I am someone who is dominant, acts as a leader 1.35 0.74 1.15 1.13 1.49 0.70 0.61
-22 I am someone who starts arguments with others 1.05 0.72 0.77 1.24 1.54 0.72 0.60
-23 I am someone who has difficulty getting started on tasks 1.39 0.69 0.86 1.26 1.49 0.74 0.69
-24 I am someone who feels secure, comfortable with self 1.27 0.88 1.08 1.42 1.49 0.96 0.92
-25 I am someone who avoids intellectual, philosophical discussions 1.21 0.79 1.36 1.25 1.56 1.10 0.99
-26 I am someone who is less active than other people 1.31 0.99 1.04 1.24 1.31 0.98 0.83
-27 I am someone who has a forgiving nature 1.20 0.57 0.90 1.20 1.37 0.85 0.78
-28 I am someone who can be somewhat careless 1.25 0.83 1.11 1.18 1.38 0.68 1.00
-29 I am someone who is emotionally stable, not easily upset 1.27 0.70 1.01 1.31 1.47 1.06 1.09
-30 I am someone who has little creativity 1.20 0.76 0.69 1.22 1.32 0.83 1.02
-31 I am someone who is sometimes shy, introverted 1.36 0.87 1.11 1.07 1.30 1.04 0.99
-32 I am someone who is helpful and unselfish with others 0.88 0.80 1.05 1.21 1.50 0.73 0.72
-33 I am someone who keeps things neat and tidy 1.24 0.64 1.00 1.09 1.35 0.73 0.71
-34 I am someone who worries a lot 1.47 0.70 1.08 1.23 1.33 1.18 1.07
-35 I am someone who values art and beauty 1.07 0.82 1.30 1.18 1.35 1.00 1.06
-36 I am someone who finds it hard to influence people 1.20 0.69 1.04 1.26 1.37 0.67 0.67
-37 I am someone who is sometimes rude to others 1.25 0.66 0.91 1.31 1.49 0.77 1.00
-38 I am someone who is efficient, gets things done 0.88 0.82 1.12 1.25 1.46 0.84 0.91
-39 I am someone who often feels sad 1.43 0.68 0.98 1.27 1.25 0.99 0.91
-40 I am someone who is complex, a deep thinker 1.08 0.72 1.06 1.06 1.26 0.83 0.99
-41 I am someone who is full of energy 1.30 0.89 1.24 1.23 1.35 0.85 0.76
-42 I am someone who is suspicious of others’ intentions 1.30 0.58 0.83 1.26 1.49 1.00 0.68
-43 I am someone who is reliable, can always be counted on 0.88 0.93 1.09 1.32 1.59 0.80 0.78
-44 I am someone who keeps their emotions under control 1.14 0.64 1.08 1.22 1.44 0.77 1.04
-45 I am someone who has difficulty imagining things 1.06 0.77 0.53 1.25 1.27 0.95 0.54
-46 I am someone who is talkative 1.38 0.70 1.11 1.04 1.28 1.09 0.92
-47 I am someone who can be cold and uncaring 1.23 0.80 0.87 1.35 1.49 0.78 0.98
-48 I am someone who leaves a mess, doesn’t clean up 1.12 0.71 0.95 1.26 1.64 0.55 0.85
-49 I am someone who rarely feels anxious or afraid 1.37 0.83 1.12 1.15 1.37 1.02 0.89
-50 I am someone who thinks poetry and plays are boring 1.37 0.75 1.31 1.38 1.53 0.87 0.68
-51 I am someone who prefers to have others take charge 1.28 0.69 0.76 1.27 1.57 0.70 0.51
-52 I am someone who is polite, courteous to others 0.70 0.65 0.97 1.14 1.44 0.75 0.79
-53 I am someone who is persistent, works until the task is finished 0.91 0.87 1.13 1.26 1.48 0.80 0.94
-54 I am someone who tends to feel depressed, blue 1.46 0.79 1.07 1.31 1.37 1.08 1.00
-55 I am someone who has little interest in abstract ideas 1.22 0.96 1.08 1.22 1.36 1.17 0.63
-56 I am someone who shows a lot of enthusiasm 1.22 0.85 1.09 1.34 1.51 0.83 0.94
-57 I am someone who assumes the best about people 1.28 0.78 0.98 1.27 1.47 0.92 0.74
-58 I am someone who sometimes behaves irresponsibly 1.30 0.74 1.13 1.14 1.40 0.70 0.95
-59 I am someone who is temperamental, gets emotional easily 1.33 0.64 1.08 1.24 1.52 1.00 0.87
-60 I am someone who is original, comes up with new ideas 1.07 0.65 1.13 1.22 1.41 0.68 0.75
-Table 13: Item Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Facet Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Sociability 2.65 3.26 3.04 2.94 2.89 2.57 2.43
-Assertiveness 2.97 3.22 3.33 3.06 2.88 2.81 2.61
-Energy Level 3.35 3.85 3.88 3.09 2.95 2.79 2.52
-Compassion 4.02 3.63 4.16 3.42 3.36 4.10 3.78
-Respectfulness 4.29 3.39 4.39 3.49 3.63 4.07 4.26
-Trust 3.31 3.15 3.78 3.21 3.32 3.49 3.64
-Organization 3.86 3.21 3.24 3.18 2.94 3.22 3.10
-Productiveness 3.86 3.54 3.85 3.19 3.23 3.60 3.34
-Responsibility 3.96 3.25 3.52 3.03 3.13 3.59 3.60
-Anxiety 3.15 3.15 2.96 3.06 2.89 3.42 3.10
-Depression 2.51 2.76 2.20 2.86 2.59 2.83 2.61
-Emotional Volatility 2.44 3.07 2.61 2.92 2.76 2.80 2.34
-Intellectual Curiosity 4.05 3.55 3.51 3.13 3.09 3.08 2.63
-Aesthetic Sensitivity 3.89 3.63 3.47 3.25 3.01 3.12 2.67
-Creative Imagination 3.96 3.72 3.73 3.29 3.16 3.17 2.44
-Table 14: Facet Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Facet Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Sociability 1.17 0.69 0.98 0.95 1.00 0.91 0.74
-Assertiveness 1.04 0.55 0.82 1.10 1.22 0.57 0.42
-Energy Level 0.95 0.75 0.88 1.12 1.19 0.77 0.64
-Compassion 0.81 0.69 0.81 1.18 1.36 0.69 0.76
-Respectfulness 0.70 0.58 0.74 1.13 1.33 0.61 0.68
-Trust 1.00 0.44 0.67 1.17 1.34 0.77 0.62
-Organization 1.01 0.57 0.79 0.98 1.18 0.63 0.68
-Productiveness 0.91 0.65 0.87 1.12 1.29 0.68 0.68
-Responsibility 0.84 0.59 0.95 1.10 1.26 0.64 0.74
-Anxiety 1.11 0.54 0.89 0.95 1.04 0.88 0.75
-Depression 1.14 0.56 0.89 1.16 1.26 0.87 0.81
-Emotional Volatility 1.10 0.43 0.92 1.06 1.25 0.76 0.87
-Intellectual Curiosity 0.83 0.68 0.96 1.04 1.24 0.89 0.70
-Aesthetic Sensitivity 1.01 0.75 1.09 1.12 1.26 0.96 0.83
-Creative Imagination 0.88 0.57 0.74 1.05 1.17 0.67 0.60
-Table 15: Facet Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Domain Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion 2.99 3.44 3.42 3.03 2.90 2.72 2.52
-Agreeableness 3.87 3.39 4.11 3.37 3.43 3.89 3.89
-Conscientiousness 3.89 3.33 3.54 3.13 3.10 3.47 3.35
-Neuroticism 2.70 2.99 2.59 2.94 2.75 3.02 2.68
-Openness 3.97 3.63 3.57 3.22 3.09 3.12 2.58
-Table 16: Domain Level Mean for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Domain Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion 0.85 0.57 0.81 0.95 1.00 0.67 0.52
-Agreeableness 0.69 0.51 0.70 1.11 1.28 0.62 0.62
-Conscientiousness 0.80 0.53 0.79 1.00 1.15 0.59 0.64
-Neuroticism 1.02 0.43 0.81 0.97 1.09 0.77 0.72
-Openness 0.76 0.58 0.78 1.03 1.15 0.72 0.59
-Table 17: Domain Level Standard Deviation for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Additional Psychometric Performance Results
-Here we show the TFM fit information in Table 18; FFM fit information in Table 19; TCC results for the TFM of each BFI-2 domain in Table 20; TCC results for the FFM in Table 21; specific standardized factor loading results in Tables 22 and 23; specific inter-factor correlation results in Tables 24 and 25.
-Domain Model $\mathbf{\chi^{2}}$ df CFI TLI RMSEA SRMR
-Extraversion Human 993.189 51 .892 .861 .109 .057
-Persona
-GPT-4o 261.099 51 .886 .852 .117 .058
-Llama3 316.715 51 .885 .851 .132 .060
-Shape
-GPT-4o 663.947 51 .823 .771 .200 .101
-Llama3 934.002 51 .684 .592 .241 .132
-PSI
-GPT-4o 329.341 51 .892 .860 .135 .057
-Llama3 332.944 51 .810 .755 .136 .092
-Agreeableness Human 904.640 51 .875 .838 .104 .060
-Persona
-GPT-4o 210.863 51 .900 .870 .102 .062
-Llama3 616.199 51 .798 .739 .192 .088
-Shape
-GPT-4o 525.799 51 .909 .882 .176 .039
-Llama3 874.914 51 .840 .793 .233 .070
-PSI
-GPT-4o 177.494 51 .952 .937 .091 .049
-Llama3 255.684 51 .923 .900 .116 .050
-Conscientiousness Human 1041.784 51 .897 .867 .112 .058
-Persona
-GPT-4o 385.339 51 .780 .716 .148 .088
-Llama3 705.359 51 .768 .699 .207 .109
-Shape
-GPT-4o 705.340 51 .826 .775 .207 .069
-Llama3 1502.901 51 .642 .536 .310 .196
-PSI
-GPT-4o 348.206 51 .887 .854 .139 .052
-Llama3 427.908 51 .834 .785 .157 .069
-Neuroticism Human 929.871 51 .931 .911 .105 .052
-Persona
-GPT-4o 304.611 51 .756 .684 .129 .091
-Llama3 346.207 51 .884 .850 .139 .059
-Shape
-GPT-4o 960.437 51 .720 .638 .244 .099
-Llama3 808.305 51 .772 .705 .224 .101
-PSI
-GPT-4o 448.933 51 .876 .840 .161 .069
-Llama3 448.871 51 .854 .811 .161 .067
-Openness Human 909.210 51 .899 .870 .104 .064
-Persona
-GPT-4o 226.649 51 .910 .883 .107 .057
-Llama3 256.032 51 .904 .876 .116 .077
-Shape
-GPT-4o 502.104 51 .878 .842 .172 .055
-Llama3 569.466 51 .861 .820 .185 .057
-PSI
-GPT-4o 310.479 51 .904 .876 .130 .071
-Llama3 352.224 51 .854 .811 .140 .114
-Table 18: Model Fits for BFI-2 Three-Factor Models of Each Domain
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Model $\mathbf{\chi^{2}}$ df CFI TLI RMSEA SRMR
-Human 1747.433 80 .852 .806 .116 .080
-Persona
-GPT-4o 551.459 80 .819 .762 .140 .097
-Llama3 954.848 80 .780 .711 .191 .120
-Shape
-GPT-4o 2132.208 80 .711 .621 .292 .143
-Llama3 1988.584 80 .717 .629 .283 .149
-PSI
-GPT-4o 899.057 80 .769 .697 .185 .119
-Llama3 802.352 80 .774 .703 .173 .114
-Table 19: Model Fits for BFI-2 Five-Factor Model
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Domain Facet Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion Sociability 1.00 .99 1.00 .99 1.00 .99
-Assertiveness .97 .98 .99 .97 .98 .97
-Energy Level .98 .98 .97 .96 .98 .98
-Agreeableness Compassion .95 .96 .97 .95 .97 1.00
-Respectfulness 1.00 1.00 1.00 1.00 .98 .99
-Trust .98 .99 1.00 .99 1.00 .99
-Conscientiousness Organization .99 1.00 .99 .95 .99 1.00
-Productiveness .99 .99 .99 .99 1.00 .97
-Responsibility .99 .98 .99 .96 1.00 .99
-Neuroticism Anxiety .99 .99 .98 .99 1.00 1.00
-Depression .97 .99 .99 .99 1.00 1.00
-Emotional Volatility .97 1.00 .99 .99 1.00 1.00
-Openness Intellectual Curiosity 1.00 1.00 .99 .98 1.00 .98
-Aesthetic Sensitivity .99 .98 .99 .98 .99 .97
-Creative Imagination 1.00 .98 1.00 1.00 .99 .96
-Table 20: Tucker’s Congruence Coefficient for BFI-2 Three-Factor Models of Each Domain
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Domain Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion 1.00 .99 1.00 1.00 1.00 1.00
-Agreeableness 1.00 1.00 .99 .99 1.00 1.00
-Conscientiousness .99 .98 .99 .99 .99 .99
-Neuroticism .99 1.00 .99 1.00 1.00 1.00
-Openness 1.00 .99 1.00 .99 1.00 .99
-Table 21: Tucker’s Congruence Coefficient for BFI-2 Five-Factor Model
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Domain Facet=~Item Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion Sociability=~item1 .80 .77 .72 .85 .53 .87 .63
-Sociability=~item16 .82 .79 .88 .74 .75 .84 .88
-Sociability=~item31 .82 .79 .87 .73 .80 .91 .76
-Sociability=~item46 .76 .74 .64 .85 .57 .76 .67
-Assertiveness=~item6 .76 .66 .84 .94 .85 .81 .69
-Assertiveness=~item21 .89 .61 .73 .90 .69 .80 .69
-Assertiveness=~item36 .53 .73 .72 .84 .88 .76 .44
-Assertiveness=~item51 .69 .59 .54 .80 .73 .55 .27
-Energy Level=~item11 .45 .82 .73 .91 .89 .82 .66
-Energy Level=~item26 .54 .69 .75 .72 .78 .74 .47
-Energy Level=~item41 .83 .81 .87 .82 .76 .87 .71
-Energy Level=~item56 .77 .82 .78 .88 .77 .80 .84
-Agreeableness Compassion=~item2 .73 .76 .75 .90 .74 .84 .83
-Compassion=~item17 .34 .84 .86 .93 .97 .79 .46
-Compassion=~item32 .62 .74 .74 .87 .86 .80 .79
-Compassion=~item47 .75 .67 .85 .96 .98 .89 .86
-Respectfulness=~item7 .71 .82 .80 .93 .84 .90 .91
-Respectfulness=~item22 .60 .54 .68 .80 .87 .43 .48
-Respectfulness=~item37 .70 .74 .79 .94 .94 .80 .82
-Respectfulness=~item52 .70 .70 .80 .83 .81 .85 .86
-Trust=~item12 .69 .63 .35 .57 .95 .95 .78
-Trust=~item27 .68 .60 .80 .84 .84 .79 .84
-Trust=~item42 .67 .35 .63 .95 .95 .78 .83
-Trust=~item57 .78 .64 .82 .90 .85 .88 .89
-Conscientiousness Organization=~item3 .86 .64 .75 .75 .48 .83 .83
-Organization=~item18 .70 .62 .66 .73 .90 .83 .65
-Organization=~item33 .86 .63 .68 .95 .55 .74 .74
-Organization=~item48 .74 .72 .72 .68 .55 .80 .79
-Productiveness=~item8 .66 .67 .68 .64 .67 .73 .43
-Productiveness=~item23 .86 .71 .65 .81 .70 .78 .43
-Productiveness=~item38 .87 .72 .88 .82 .85 .85 .52
-Productiveness=~item53 .65 .47 .66 .80 .63 .81 .82
-Responsibility=~item13 .72 .67 .94 .92 .97 .79 .83
-Responsibility=~item28 .77 .58 .58 .68 .46 .59 .71
-Responsibility=~item43 .66 .51 .74 .75 .70 .68 .57
-Responsibility=~item58 .80 .59 .80 .75 .87 .71 .74
-Neuroticism Anxiety=~item4 .60 .71 .72 .68 .62 .80 .79
-Anxiety=~item19 .63 .65 .61 .65 .72 .78 .43
-Anxiety=~item34 .65 .47 .81 .73 .79 .59 .65
-Anxiety=~item49 .74 .71 .84 .85 .70 .92 .89
-Depression=~item9 .66 .67 .60 .81 .77 .88 .44
-Depression=~item24 .86 .82 .59 .81 .78 .92 .91
-Depression=~item39 .87 .86 .65 .47 .86 .95 .95
-Depression=~item54 .65 .69 .88 .92 .95 .79 .85
-Emotional Volatility=~item14 .72 .77 .72 .59 .88 .84 .86
-Emotional Volatility=~item29 .77 .66 .51 .74 .75 .70 .68
-Emotional Volatility=~item44 .66 .64 .27 .68 .75 .60 .67
-Emotional Volatility=~item59 .80 .67 .72 .83 .79 .65 .85
-Openness Intellectual Curiosity=~item10 .60 .90 .81 .84 .91 .92 .89
-Intellectual Curiosity=~item25 .63 .91 .67 .88 .89 .95 .94
-Intellectual Curiosity=~item40 .65 .80 .59 .70 .57 .85 .85
-Intellectual Curiosity=~item55 .74 .77 .81 .84 .89 .88 .56
-Aesthetic Sensitivity=~item5 .66 .66 .84 .85 .91 .71 .74
-Aesthetic Sensitivity=~item20 .86 .77 .81 .65 .83 .79 .85
-Aesthetic Sensitivity=~item35 .87 .59 .70 .84 .81 .92 .88
-Aesthetic Sensitivity=~item50 .65 .63 .69 .41 .91 .63 .32
-Creative Imagination=~item15 .72 .77 .40 .71 .81 .95 .79
-Creative Imagination=~item30 .77 .80 .59 .70 .81 .86 .82
-Creative Imagination=~item45 .66 .84 .68 .49 .77 .92 .35
-Creative Imagination=~item60 .80 .86 .70 .95 .81 .84 .95
-Table 22: Standardized Factor Loadings for BFI-2 Three-Factor Models of Each Domain with Human Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
-Domain=~Facet Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-E=~Sociability .66 .76 .85 .91 .87 .81 .78
-E=~Assertiveness .59 .67 .86 .77 .73 .71 .68
-E=~Energy Level .77 .88 .86 .89 .85 .92 .89
-A=~Compassion .70 .82 .90 .96 .99 .82 .79
-A=~Respectfulness .80 .90 .94 .90 .85 .88 .94
-A=~Trust .65 .75 .87 .93 .95 .84 .84
-C=~Organization .70 .87 .87 .94 .97 .87 .86
-C=~Productiveness .89 .77 .75 .84 .81 .84 .82
-C=~Responsibility .79 .82 .95 .93 .92 .88 .91
-N=~Anxiety .84 .69 .75 .80 .86 .77 .77
-N=~Depression .88 .91 .96 1.02 .96 .97 .87
-N=~Emotional Volatility .85 .63 .79 .71 .80 .78 .85
-O=~Intellectual Curiosity .74 .77 .66 .97 .93 .74 .85
-O=~Aesthetic Sensitivity .71 .73 .65 .95 .96 .69 .59
-O=~Creative Imagination .81 .91 .89 .90 .82 .96 .76
-Table 23: Standardized Factor Loadings for BFI-2 Five-Factor Model
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
-Domain Facet~~Facet Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion Sociability~~Assertiveness .59 .72 .89 .77 .91 .69 .70
-Sociability~~Energy Level .64 .79 .84 .91 .80 .84 .85
-Assertiveness~~Energy Level .47 .66 .82 .65 .61 .72 .65
-Agreeableness Compassion~~Respectfulness .81 .90 .98 .92 .87 .85 .90
-Compassion~~Trust .70 .83 .94 .94 .96 .79 .84
-Respectfulness~~Trust .58 .95 .96 .93 .84 .85 .86
-Conscientiousness Organization~~Productiveness .75 .86 .81 .97 .70 .86 .86
-Organization~~Responsibility .70 1.01 .85 .99 .86 .89 .87
-Productiveness~~Responsibility .86 .82 .87 .87 .89 .82 .87
-Neuroticism Anxiety~~Depression .81 .88 .88 .95 .96 .86 .79
-Anxiety~~Emotional Volatility .90 .87 .79 .87 .87 .92 .83
-Depression~~Emotional Volatility .79 .84 .85 .89 .83 .86 .82
-Openness Intellectual Curiosity~~Aesthetic Sensitivity .66 .68 .59 1.02 1.00 .51 .51
-Intellectual Curiosity~~Creative Imagination .74 .85 .67 .99 .89 .78 .72
-Aesthetic Sensitivity~~Creative Imagination .63 .79 .60 .93 .86 .77 .50
-Table 24: Inter-factor Correlations for BFI-2 Three-Factor Models of Each Domain with Human Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
-Domain~~Domain Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Extraversion~~Agreeableness .28 .26 .10 .59 .54 .41 .24
-Extraversion~~Conscientiousness .56 .43 .26 .36 .31 .44 .34
-Extraversion~~Neuroticism -.61 -.61 -.59 -.80 -.72 -.62 -.31
-Extraversion~~Openness .35 .37 .51 .72 .71 .29 .48
-Agreeableness~~Conscientiousness .47 .56 .77 .65 .65 .59 .63
-Agreeableness~~Neuroticism -.43 -.40 -.54 -.66 -.73 -.44 -.47
-Agreeableness~~Openness .27 .31 .25 .79 .89 .14 .14
-Conscientiousness~~Neuroticism -.61 -.52 -.63 -.67 -.82 -.58 -.60
-Conscientiousness~~Openness .17 .10 .16 .27 .45 .06 .18
-Neuroticism~~Openness -.17 -.06 -.23 -.48 -.65 .14 .18
-Table 25: Inter-factor Correlations for BFI-2 Five-Factor Models with Human Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
-Scale Reliability: Facet level and domain level Cronbach’s alpha for different method LLM responses on BFI-2 and human responses are shown in Table 26.
-It can be observed that the PSI method, compared to the Persona and Shape methods, performs closer to the results of the human sample in terms of Cronbach‘s alpha (with the number of data marked in italics and bold being the smallest).
+I.1 Additional Response Similarity Results
+I.2 Additional Human Personality Distribution Simulation Results
+I.2.1 Additional Descriptive Statistics Results
+Here we show the detailed $\mu$ and $\sigma$ for human responses and LLM responses at the Item, Facet, and Domain levels, see Tables LABEL:tab:Mistral_d-LABEL:tab:GPT-4o.
+We also present detail MAE results in Table 20.
+I.2.2 Additional Psychometric Performance Results
+Table LABEL:tab:modelfit presents the model fit information for both the TFMs and FFM.
+TCC results are reported in Tables LABEL:tcc_tfm_e-27, while factor loading MAE results appear in Tables LABEL:tab:mae_extraversion-33.
+Specific standardized factor loadings for all TFMs are shown in Tables LABEL:tab:fc_sociability-LABEL:tab:fc_creative_imagination, while standardized factor loadings for the FFM are shown in Tables LABEL:tab:ffm_extraversion-LABEL:tab:ffm_openness.
+Inter-factor correlations are provided in Tables LABEL:tab:inter_extraversion-59.
+Scale Reliability: Facet level and domain level Cronbach’s alpha for different method LLM responses on BFI-2 and human responses are shown in Tables LABEL:tab:alpha_extraversion-65.
+It can be observed that the PSI method, compared to the Persona and Shape methods, performs closer to the results of the human sample in terms of Cronbach’s alpha (with the number of data marked in italics and bold being the smallest).
 This indicates that the LLM personality data generated by the PSI method holds an advantage in consistency and reliability, enabling it to more accurately simulate the statistical characteristics of human samples.
-Facet / Domain Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-Sociability .87 .85 .86 .87 .77 .91 .82
-Assertiveness .81 .74 .80 .93 .86 .82 .57
-Energy Level .73 .86 .86 .90 .87 .88 .76
-Compassion .67 .84 .87 .95 .93 .90 .82
-Respectfulness .74 .79 .85 .94 .93 .84 .86
-Trust .80 .64 .79 .96 .94 .87 .88
-Organization .87 .73 .80 .86 .82 .86 .82
-Productiveness .80 .80 .85 .91 .90 .88 .76
-Responsibility .77 .67 .88 .91 .86 .88 .86
-Anxiety .84 .65 .81 .81 .79 .88 .78
-Depression .86 .66 .88 .90 .91 .91 .88
-Emotional Volatility .89 .60 .89 .87 .88 .89 .91
-Intellectual Curiosity .75 .82 .85 .87 .89 .89 .79
-Aesthetic Sensitivity .83 .87 .83 .92 .93 .87 .84
-Creative Imagination .82 .80 .82 .88 .89 .88 .77
-Extraversion .87 .90 .93 .95 .92 .93 .87
-Agreeableness .85 .90 .94 .98 .97 .93 .93
-Conscientiousness .90 .88 .93 .95 .94 .94 .92
-Neuroticism .94 .83 .93 .94 .94 .95 .93
-Openness .89 .91 .89 .96 .96 .92 .88
-Table 26: Cronbach’s alpha for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Italics for absolute differences compared to the human responses of .100 to .199, and boldface for differences of .200 or higher.
 Discriminant Validity: The results for discriminant validity are shown in Table 66.
@@
 Therefore, our focus here is on identifying the approach that most closely aligns with the performance of human samples.
-Dimensions Human Persona Shape PSI
-GPT-4o Llama3 GPT-4o Llama3 GPT-4o Llama3
-E, A .17 .21 .12 .48 .40 .33 .21
-E, C .35 .40 .33 .46 .41 .38 .30
-E, N -.45 -.43 -.42 -.58 -.57 -.43 -.21
-E, O .22 .23 .28 .66 .68 .13 .33
-A, C .33 .47 .66 .67 .70 .49 .52
-A, N -.36 -.30 -.54 -.64 -.71 -.29 -.37
-A, O .20 .30 .31 .71 .78 .14 .14
-C, N -.49 -.43 -.60 -.74 -.87 -.47 -.52
-C, O .09 .09 .20 .29 .48 .02 .08
-N, O -.11 .03 -.08 -.31 -.57 .32 .23
-Mean of Absolute Values .28 .29 .35 .55 .62 .30 .29
-Table 27: Domain Level Correlation Analysis for BFI-2 Human Responses and Different Methods LLM Responses
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-E = Extraversion; A = Agreeableness; C = Conscientiousness; N = Neuroticism; O = Openness.
-I.2 Additional Personality-Related Behavioral Performance Results
-Table 28 presents the correlations between the personality dimensions and OCB/CWB reported by PSI GPT-4o, with a comparison to human self-reported data.
-The relevant results for PSI Llama3 are provided in Table 29, while Figures 3 and Figure 4 display the complete correlation matrices for GPT-4o and Llama3, respectively.
-Domain OCB CWB
-Human PSI Human PSI
-Ext .42 .54 -.01 -.08
-Agr .18 .36 -.30 -.54
-Con .12 .45 -.35 -.47
-Neu -.21 -.35 .23 .36
-Ope .17 .02 -.17 .01
-Table 28: Comparison of OCB and CWB Correlations with Personality Domains: Human vs. PSI GPT-4o
-Note: n = 357 for human and PSI method.
-Domain OCB CWB
-Human PSI Human PSI
-Ext .42 .48 -.01 .05
-Agr .18 .46 -.30 -.45
-Con .12 .47 -.35 -.41
-Neu -.21 -.38 .23 .33
-Ope .17 .11 -.17 .00
-Table 29: Comparison of OCB and CWB Correlations with Personality Domains: Human vs. PSI Llama3
-Note: n = 357 for human and PSI method.
-Figure 3: Complete Correlation Matrices for Human and PSI GPT-4o
-Note: n = 357 for human and PSI method.
-Figure 4: Complete Correlation Matrices for Human and PSI Llama3
-Note: n = 357 for human and PSI method.
-I.3 The Influence of Social Desirability
-In previous studies, researchers have observed some intriguing phenomena, specifically that LLMs tend to exhibit human-like social desirability bias when simulating human samples.
-This bias is reflected in the LLMs’ responses, which lean toward behaviors and traits that are socially approved.
-Numerous studies support this finding.
-For example, Hilliard et al.
-(2024) noted that newer and larger-parameter LLMs display more diverse personality traits, including higher levels of agreeableness, emotional stability, and openness.
-Similarly, Salecha et al.
-(2024) found that LLMs exhibit human-like social desirability bias when generating simulated data.
-In this section, using the data we have collected, we directly compare the results from human samples and LLM-simulated data.
-We examined the self-reported personality traits of human samples alongside the simulated personality traits from LLMs, analyzing how both were influenced by social desirability ratings.
-Specifically, we used social desirability ratings to predict and test the correlations and differences between the mean scores of human and LLM data across various items.
-By analyzing the predictive relationships and correlations between item means and social desirability ratings, we aim to uncover how social desirability influences the characteristics and patterns in both human and LLM data.
-Furthermore, by comparing the differences in item means between human data and LLM-simulated data, we explore whether social desirability ratings affect human samples and model-simulated data in similar ways.
-Social Desirability Rating
-Social desirability ratings of the BFI-2 items were obtained from another ongoing study where 142 human resource practitioners were asked to rate how desirable each item was in general (1 = “Very undesirable,” 2 = “Undesirable,” 3 = “Slightly undesirable,” 4 = “Neither desirable nor undesirable,” 5 = “Slightly desirable,” 6 = “Desirable,” 7 = “Very desirable”).
-If a certain measurement item receives a high score (6 or 7), it indicates that the trait is generally viewed as desirable or positive in society.
-On the other hand, a low score (1 or 2) suggests that the trait may be considered less ideal or inconsistent with social expectations.
-Some desirable traits may lead participants to lean toward choosing higher scores (due to social desirability bias) rather than responding based on their true circumstances.
-Results
-Table 30 presents the regression results predicting the self-reported personality in the human sample, the LLM-simulated personality, and the difference between human and LLM personalities, all based on the social desirability rating.
-Additionally, Figure 5 visually illustrates the trends through a regression plot.
-It can be observed that when an item’s social desirability rating is neutral (4), most LLMs tend to generate neutral responses (rated 3 on the scale).
-This result indicates that LLMs have learned to associate neutral social expectations with neutral opinions, consistent with previous research findings (Salecha et al., 2024).
-Regarding the mean difference between the human responses and the LLM responses, most differences show a positive correlation with the social desirability rating, and when the social desirability rating is neutral (4), the mean difference in most data is approximately 0.
-In Figure 5, the line graph clearly illustrates that as the social desirability rating of the item increases, the average scores of both human responses and LLM responses rise accordingly.
-Moreover, as the social desirability rating increases, the average difference between human and LLM responses gradually widens, whereas this difference narrows when the social desirability scores are lower.
-This trend indicates that when social desirability scores are high, the mean score of human responses tends to be higher than that of LLM responses.
-Conversely, when social desirability scores are low, the average score of human responses falls below that of LLM responses.
-This suggests that human responses are more strongly influenced by social desirability than those of LLMs.
-Variable Human Persona GPT-4o Shape GPT-4o PSI GPT-4o Persona Llama3 Shape Llama3 PSI Llama3
-Intercept 1.15 (0.23) 1.90 (0.17) 2.33 (0.09) 2.13 (0.23) 0.71 (0.21) 2.25 (0.11) 2.08 (0.28)
-Social desirability 0.46 (0.05) 0.23 (0.04) 0.12 (0.02) 0.23 (0.05) 0.47 (0.05) 0.14 (0.02) 0.20 (0.06)
-$R^{2}$59 .39 .38 .27 .65 .37 .15
-Predicted score at neutral point 2.99 2.82 2.81 3.05 2.59 2.81 2.88
-Correlation .77 .63 .62 .52 .80 .61 .39
-Mean Score Difference
-Intercept - -0.76 (0.19) -1.18 (0.19) -0.99 (0.20) 0.44 (0.21) -1.11 (0.23) -0.93 (0.32)
-Social desirability - 0.23 (0.04) 0.34 (0.04) 0.23 (0.04) -0.01 (0.05) 0.33 (0.05) 0.26 (0.07)
-$R^{2}$34 .53 .32 .48 .42 .19
-Predicted score at neutral point - 0.16 0.18 -0.07 0.40 0.21 0.11
-Correlation - .58 .73 .56 -.03 .65 .43
-Table 30: Regression Analysis and Correlation of Social Desirability Ratings
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
-Figure 5: Regression Line Chart of Social Desirability Ratings
-Note: n = 1,559 for human responses, n = 297 for Shape Llama3, and n = 300 for other LLM responses.
-Some sample sizes are below 300, because certain generated data exceeded reasonable thresholds (1-5) for specific items and were excluded from the analysis.
+I.3 Additional Personality-Related Behavioral Performance Results
+Table 67 and 68 present the correlations between the personality dimensions and OCB/CWB reported by all model+PSI simulations, with a comparison to human self-reported data.
+I.4 Additional Abalation Test Results
+Figure 7-10 present the relative importance of individual questions for different prediction domains.
+MAE $r$
+Model Ext Agr Neu Con Ope Ext Agr Neu Con Ope
+Mistral-7B+PSI 0.57 0.66 0.60 0.57 0.58 0.56 0.34 0.56 0.42 0.36
+Gemma-2-9B+PSI 0.62 0.65 0.79 0.73 0.95 0.55 0.37 0.50 0.39 0.34
+Gemma-2-27B+PSI 0.76 0.50 0.71 0.65 1.09 0.48 0.35 0.54 0.40 0.40
+Llama3-8B+PSI 0.62 0.51 0.61 0.72 0.93 0.60 0.40 0.62 0.42 0.44
+Llama3-70B+PSI 0.75 0.56 0.59 0.66 1.27 0.43 0.36 0.57 0.40 0.39
+GPT-4o-mini+PSI 0.60 0.55 0.63 0.56 0.81 0.63 0.41 0.63 0.48 0.43
+GPT-4o+PSI 0.58 0.53 0.63 0.59 0.80 0.64 0.41 0.63 0.46 0.43
+GPT-4o+Life Interview (Park et al., 2024) 0.72 0.60 0.75 0.63 0.62 0.45 0.35 0.68 0.52 0.39
+Table 12: MAE and $r$ of human and LLM-simulated data across seven Models+PSI and GPT-4o+Life Interview (see Park et al., 2024, page 35 Table 3) for BFI-2 each Personality domain.
+Table 13: Mistral-7B descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $3.23$ $3.30$ $2.69$ $1.40$ $1.09$ $1.30$ $1.10$
+2 I am someone who is compassionate, has a soft heart $4.34$ $3.86$ $3.74$ $4.17$ $0.90$ $1.04$ $1.29$ $1.03$
+3 I am someone who tends to be disorganized $3.62$ $2.89$ $3.13$ $3.90$ $1.36$ $0.96$ $1.22$ $0.68$
+4 I am someone who is relaxed, handles stress well $2.56$ $2.28$ $2.38$ $2.73$ $1.30$ $0.69$ $0.85$ $0.93$
+5 I am someone who has few artistic interests $3.68$ $3.73$ $2.96$ $4.07$ $1.33$ $0.54$ $0.78$ $0.26$
+6 I am someone who has an assertive personality $2.96$ $3.61$ $3.26$ $3.34$ $1.38$ $0.71$ $0.89$ $1.02$
+7 I am someone who is respectful, treats others with respect $4.62$ $4.65$ $4.64$ $4.50$ $0.61$ $0.77$ $0.89$ $0.80$
+8 I am someone who tends to be lazy $3.62$ $3.96$ $3.28$ $4.38$ $1.31$ $0.59$ $1.52$ $0.77$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.17$ $2.12$ $1.99$ $1.24$ $0.52$ $0.87$ $0.78$
+10 I am someone who is curious about many different things $4.43$ $4.28$ $4.37$ $3.61$ $0.81$ $0.65$ $0.96$ $0.94$
+11 I am someone who rarely feels excited or eager $3.54$ $3.67$ $3.01$ $3.91$ $1.26$ $0.65$ $1.07$ $0.44$
+12 I am someone who tends to find fault with others $3.34$ $3.63$ $3.28$ $4.45$ $1.32$ $0.60$ $1.47$ $0.58$
+13 I am someone who is dependable, steady $4.33$ $3.85$ $3.77$ $4.04$ $0.86$ $0.90$ $1.09$ $0.92$
+14 I am someone who is moody, has up and down mood swings $2.56$ $3.19$ $2.82$ $2.69$ $1.36$ $0.60$ $1.22$ $0.95$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.68$ $3.48$ $2.82$ $1.03$ $0.83$ $1.09$ $0.98$
+16 I am someone who tends to be quiet $2.35$ $2.83$ $2.47$ $3.21$ $1.34$ $0.68$ $1.15$ $1.06$
+17 I am someone who feels little sympathy for others $3.71$ $3.58$ $3.62$ $4.54$ $1.45$ $0.94$ $1.51$ $0.50$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $2.84$ $3.34$ $2.49$ $1.06$ $0.91$ $0.82$ $0.97$
+19 I am someone who can be tense $3.29$ $3.04$ $3.00$ $3.63$ $1.27$ $0.87$ $0.95$ $0.78$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $3.44$ $4.34$ $2.03$ $1.16$ $1.02$ $1.16$ $1.04$
+21 I am someone who is dominant, acts as a leader $2.90$ $2.43$ $2.46$ $1.99$ $1.35$ $0.80$ $1.05$ $0.68$
+22 I am someone who starts arguments with others $4.23$ $4.54$ $3.72$ $4.96$ $1.05$ $0.87$ $1.75$ $0.23$
+23 I am someone who has difficulty getting started on tasks $3.31$ $3.43$ $2.78$ $3.74$ $1.39$ $0.77$ $1.22$ $0.70$
+24 I am someone who feels secure, comfortable with self $2.26$ $1.90$ $1.97$ $2.06$ $1.27$ $0.70$ $0.97$ $0.90$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $3.11$ $2.72$ $3.96$ $1.21$ $0.74$ $1.14$ $0.42$
+26 I am someone who is less active than other people $3.35$ $3.37$ $2.79$ $3.47$ $1.31$ $0.97$ $1.16$ $0.98$
+27 I am someone who has a forgiving nature $3.79$ $3.36$ $3.51$ $3.72$ $1.20$ $0.49$ $0.94$ $0.75$
+28 I am someone who can be somewhat careless $3.56$ $2.85$ $3.12$ $3.40$ $1.25$ $0.92$ $0.98$ $0.95$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $2.62$ $2.27$ $2.93$ $1.27$ $0.56$ $0.98$ $0.85$
+30 I am someone who has little creativity $3.95$ $3.73$ $3.12$ $4.05$ $1.20$ $0.45$ $0.82$ $0.22$
+31 I am someone who is sometimes shy, introverted $2.35$ $2.15$ $2.56$ $3.04$ $1.36$ $0.45$ $0.66$ $0.94$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.69$ $4.15$ $4.03$ $0.88$ $0.76$ $1.14$ $0.92$
+33 I am someone who keeps things neat and tidy $3.71$ $3.20$ $3.50$ $2.32$ $1.24$ $0.84$ $1.05$ $0.60$
+34 I am someone who worries a lot $3.30$ $3.68$ $3.30$ $3.60$ $1.47$ $0.65$ $1.07$ $1.00$
+35 I am someone who values art and beauty $4.16$ $3.94$ $4.45$ $3.11$ $1.07$ $0.93$ $0.89$ $1.17$
+36 I am someone who finds it hard to influence people $3.09$ $3.38$ $2.79$ $3.94$ $1.20$ $0.55$ $0.85$ $0.32$
+37 I am someone who is sometimes rude to others $3.79$ $3.65$ $3.52$ $4.27$ $1.25$ $0.78$ $1.48$ $0.66$
+38 I am someone who is efficient, gets things done $4.23$ $3.52$ $3.70$ $3.40$ $0.88$ $0.91$ $1.08$ $1.10$
+39 I am someone who often feels sad $2.71$ $3.46$ $3.19$ $3.07$ $1.43$ $0.83$ $1.14$ $1.08$
+40 I am someone who is complex, a deep thinker $4.02$ $4.34$ $3.88$ $4.17$ $1.08$ $0.70$ $0.93$ $0.73$
+41 I am someone who is full of energy $3.14$ $3.75$ $3.50$ $3.18$ $1.30$ $0.97$ $1.36$ $1.20$
+42 I am someone who is suspicious of others’ intentions $2.75$ $3.19$ $3.28$ $4.25$ $1.30$ $0.49$ $1.21$ $0.71$
+43 I am someone who is reliable, can always be counted on $4.32$ $3.79$ $3.84$ $4.13$ $0.88$ $0.94$ $1.15$ $0.95$
+44 I am someone who keeps their emotions under control $2.22$ $2.88$ $2.35$ $3.31$ $1.14$ $0.42$ $0.90$ $0.79$
+45 I am someone who has difficulty imagining things $4.17$ $3.09$ $2.86$ $3.89$ $1.06$ $0.37$ $0.72$ $0.37$
+46 I am someone who is talkative $2.87$ $3.04$ $2.94$ $3.20$ $1.38$ $0.96$ $1.01$ $1.19$
+47 I am someone who can be cold and uncaring $3.84$ $3.66$ $3.47$ $4.26$ $1.23$ $0.63$ $1.27$ $0.72$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $3.44$ $3.07$ $4.79$ $1.12$ $1.24$ $1.90$ $0.45$
+49 I am someone who rarely feels anxious or afraid $3.44$ $3.00$ $2.27$ $3.34$ $1.37$ $0.29$ $1.17$ $0.73$
+50 I am someone who thinks poetry and plays are boring $3.66$ $2.65$ $2.66$ $3.46$ $1.37$ $1.34$ $1.39$ $1.68$
+51 I am someone who prefers to have others take charge $2.96$ $3.26$ $2.86$ $4.11$ $1.28$ $0.75$ $1.49$ $0.55$
+52 I am someone who is polite, courteous to others $4.50$ $4.21$ $4.55$ $4.30$ $0.70$ $0.89$ $0.88$ $0.87$
+53 I am someone who is persistent, works until the task is finished $4.28$ $4.36$ $4.09$ $4.55$ $0.91$ $0.68$ $1.03$ $0.66$
+54 I am someone who tends to feel depressed, blue $2.65$ $3.37$ $2.85$ $2.85$ $1.46$ $0.84$ $1.36$ $1.14$
+55 I am someone who has little interest in abstract ideas $3.83$ $3.32$ $3.03$ $3.97$ $1.22$ $0.55$ $0.61$ $0.25$
+56 I am someone who shows a lot of enthusiasm $3.35$ $4.39$ $3.70$ $3.71$ $1.22$ $0.75$ $1.30$ $1.18$
+57 I am someone who assumes the best about people $3.37$ $3.39$ $3.67$ $3.44$ $1.28$ $0.53$ $0.87$ $0.89$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $3.05$ $3.24$ $3.81$ $1.30$ $0.99$ $0.99$ $0.76$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $3.17$ $2.79$ $2.15$ $1.33$ $0.96$ $1.42$ $0.92$
+60 I am someone who is original, comes up with new ideas $3.84$ $3.67$ $3.36$ $2.96$ $1.07$ $0.91$ $1.17$ $1.11$
+Facet Level
+1 Sociability $2.65$ $2.81$ $2.82$ $3.04$ $1.17$ $0.59$ $0.85$ $0.89$
+2 Assertiveness $2.98$ $3.17$ $2.84$ $3.34$ $1.04$ $0.46$ $0.86$ $0.47$
+3 Energy Level $3.35$ $3.79$ $3.25$ $3.57$ $0.95$ $0.66$ $1.05$ $0.73$
+4 Compassion $4.02$ $3.70$ $3.74$ $4.25$ $0.81$ $0.56$ $1.03$ $0.65$
+5 Respectfulness $4.29$ $4.26$ $4.11$ $4.51$ $0.70$ $0.64$ $1.01$ $0.55$
+6 Trust $3.31$ $3.39$ $3.44$ $3.97$ $1.00$ $0.36$ $0.93$ $0.57$
+7 Organization $3.86$ $3.09$ $3.26$ $3.37$ $1.01$ $0.67$ $0.94$ $0.50$
+8 Productiveness $3.86$ $3.82$ $3.46$ $4.02$ $0.91$ $0.57$ $1.05$ $0.64$
+9 Responsibility $3.96$ $3.39$ $3.49$ $3.85$ $0.84$ $0.76$ $0.87$ $0.75$
+10 Anxiety $3.15$ $3.00$ $2.74$ $3.32$ $1.11$ $0.43$ $0.70$ $0.66$
+11 Depression $2.51$ $2.72$ $2.53$ $2.49$ $1.14$ $0.51$ $0.91$ $0.81$
+12 Emotional Volatility $2.44$ $2.96$ $2.56$ $2.77$ $1.10$ $0.45$ $0.90$ $0.72$
+13 Intellectual Curiosity $4.05$ $3.76$ $3.50$ $3.93$ $0.83$ $0.48$ $0.65$ $0.41$
+14 Aesthetic Sensitivity $3.89$ $3.44$ $3.60$ $3.17$ $1.01$ $0.67$ $0.86$ $0.73$
+15 Creative Imagination $3.96$ $3.54$ $3.20$ $3.43$ $0.88$ $0.47$ $0.77$ $0.53$
+Domain Level
+1 Extraversion $2.99$ $3.26$ $2.97$ $3.32$ $0.85$ $0.47$ $0.84$ $0.62$
+2 Agreeableness $3.87$ $3.78$ $3.76$ $4.24$ $0.69$ $0.45$ $0.92$ $0.54$
+3 Conscientiousness $3.89$ $3.43$ $3.41$ $3.75$ $0.80$ $0.59$ $0.90$ $0.56$
+4 Neuroticism $2.70$ $2.90$ $2.61$ $2.86$ $1.02$ $0.38$ $0.75$ $0.65$
+5 Openness $3.97$ $3.58$ $3.43$ $3.51$ $0.76$ $0.44$ $0.70$ $0.45$
+Table 14: Gemma-2-9B descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $2.98$ $2.46$ $2.26$ $1.40$ $0.79$ $1.22$ $0.80$
+2 I am someone who is compassionate, has a soft heart $4.34$ $3.50$ $3.09$ $3.44$ $0.90$ $0.82$ $1.31$ $0.93$
+3 I am someone who tends to be disorganized $3.62$ $2.97$ $3.05$ $2.95$ $1.36$ $0.39$ $1.03$ $0.37$
+4 I am someone who is relaxed, handles stress well $2.56$ $3.09$ $3.33$ $3.59$ $1.30$ $0.62$ $1.09$ $0.65$
+5 I am someone who has few artistic interests $3.68$ $3.15$ $3.11$ $2.76$ $1.33$ $0.72$ $1.01$ $0.62$
+6 I am someone who has an assertive personality $2.96$ $3.10$ $2.93$ $2.38$ $1.38$ $0.62$ $1.15$ $0.72$
+7 I am someone who is respectful, treats others with respect $4.62$ $3.99$ $3.71$ $3.51$ $0.61$ $1.07$ $1.42$ $0.72$
+8 I am someone who tends to be lazy $3.62$ $3.38$ $3.37$ $3.36$ $1.31$ $0.67$ $1.27$ $0.80$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.67$ $2.72$ $2.82$ $1.24$ $0.74$ $1.48$ $0.94$
+10 I am someone who is curious about many different things $4.43$ $3.65$ $3.27$ $3.01$ $0.81$ $0.69$ $1.40$ $0.62$
+11 I am someone who rarely feels excited or eager $3.54$ $3.57$ $3.02$ $2.85$ $1.26$ $0.72$ $1.57$ $0.96$
+12 I am someone who tends to find fault with others $3.34$ $3.05$ $3.31$ $3.15$ $1.32$ $0.32$ $1.29$ $0.52$
+13 I am someone who is dependable, steady $4.33$ $3.58$ $3.10$ $3.55$ $0.86$ $0.78$ $1.40$ $0.73$
+14 I am someone who is moody, has up and down mood swings $2.56$ $3.16$ $2.85$ $3.10$ $1.36$ $0.49$ $1.31$ $0.71$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.29$ $2.89$ $2.95$ $1.03$ $0.52$ $1.13$ $0.50$
+16 I am someone who tends to be quiet $2.35$ $2.99$ $2.97$ $2.56$ $1.34$ $0.52$ $1.08$ $0.98$
+17 I am someone who feels little sympathy for others $3.71$ $3.64$ $3.64$ $3.41$ $1.45$ $0.90$ $1.45$ $1.00$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $3.27$ $2.98$ $2.82$ $1.06$ $0.66$ $1.14$ $0.69$
+19 I am someone who can be tense $3.29$ $3.37$ $3.36$ $3.42$ $1.27$ $0.61$ $1.18$ $0.68$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $3.49$ $2.95$ $2.94$ $1.16$ $0.82$ $1.16$ $0.83$
+21 I am someone who is dominant, acts as a leader $2.90$ $3.03$ $3.08$ $2.46$ $1.35$ $0.57$ $1.13$ $0.73$
+22 I am someone who starts arguments with others $4.23$ $3.61$ $3.74$ $3.52$ $1.05$ $0.86$ $1.40$ $0.84$
+23 I am someone who has difficulty getting started on tasks $3.31$ $3.00$ $3.11$ $2.81$ $1.39$ $0.26$ $1.21$ $0.47$
+24 I am someone who feels secure, comfortable with self $2.26$ $2.73$ $3.00$ $3.45$ $1.27$ $0.82$ $1.40$ $0.74$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $3.02$ $2.93$ $2.85$ $1.21$ $0.28$ $1.19$ $0.63$
+26 I am someone who is less active than other people $3.35$ $3.21$ $3.20$ $2.74$ $1.31$ $0.70$ $1.12$ $0.52$
+27 I am someone who has a forgiving nature $3.79$ $3.03$ $3.20$ $3.03$ $1.20$ $0.36$ $1.16$ $0.59$
+28 I am someone who can be somewhat careless $3.56$ $2.90$ $2.66$ $3.10$ $1.25$ $0.42$ $1.00$ $0.64$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $3.11$ $2.87$ $3.31$ $1.27$ $0.53$ $1.28$ $0.78$
+30 I am someone who has little creativity $3.95$ $3.01$ $3.28$ $3.01$ $1.20$ $0.49$ $1.13$ $0.49$
+31 I am someone who is sometimes shy, introverted $2.35$ $2.88$ $3.04$ $2.89$ $1.36$ $0.57$ $1.06$ $0.87$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.24$ $3.28$ $3.51$ $0.88$ $0.76$ $1.50$ $0.76$
+33 I am someone who keeps things neat and tidy $3.71$ $2.97$ $2.89$ $3.03$ $1.24$ $0.44$ $1.02$ $0.36$
+34 I am someone who worries a lot $3.30$ $3.20$ $2.80$ $3.52$ $1.47$ $0.61$ $1.31$ $1.06$
+35 I am someone who values art and beauty $4.16$ $3.56$ $3.16$ $3.22$ $1.07$ $0.81$ $1.18$ $0.64$
+36 I am someone who finds it hard to influence people $3.09$ $3.01$ $3.04$ $2.89$ $1.20$ $0.32$ $1.06$ $0.47$
+37 I am someone who is sometimes rude to others $3.79$ $3.09$ $3.49$ $3.29$ $1.25$ $0.50$ $1.32$ $0.79$
+38 I am someone who is efficient, gets things done $4.23$ $3.39$ $3.25$ $3.33$ $0.88$ $0.69$ $1.24$ $0.66$
+39 I am someone who often feels sad $2.71$ $3.09$ $2.65$ $3.38$ $1.43$ $0.52$ $1.21$ $0.72$
+40 I am someone who is complex, a deep thinker $4.02$ $3.40$ $3.00$ $2.72$ $1.08$ $0.67$ $1.04$ $0.87$
+41 I am someone who is full of energy $3.14$ $3.47$ $2.79$ $2.61$ $1.30$ $0.85$ $1.14$ $0.87$
+42 I am someone who is suspicious of others’ intentions $2.75$ $2.93$ $3.15$ $2.90$ $1.30$ $0.39$ $1.28$ $0.52$
+43 I am someone who is reliable, can always be counted on $4.32$ $3.49$ $3.37$ $3.47$ $0.88$ $0.86$ $1.49$ $0.75$
+44 I am someone who keeps their emotions under control $2.22$ $3.12$ $2.61$ $3.26$ $1.14$ $0.47$ $1.37$ $0.66$
+45 I am someone who has difficulty imagining things $4.17$ $3.10$ $3.45$ $3.06$ $1.06$ $0.38$ $1.09$ $0.42$
+46 I am someone who is talkative $2.87$ $3.02$ $3.03$ $2.50$ $1.38$ $0.38$ $0.89$ $0.88$
+47 I am someone who can be cold and uncaring $3.84$ $3.42$ $3.26$ $3.45$ $1.23$ $0.73$ $1.48$ $1.10$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $3.27$ $3.06$ $3.16$ $1.12$ $0.62$ $1.44$ $0.51$
+49 I am someone who rarely feels anxious or afraid $3.44$ $3.18$ $2.77$ $3.64$ $1.37$ $0.48$ $1.26$ $0.83$
+50 I am someone who thinks poetry and plays are boring $3.66$ $3.53$ $3.34$ $3.12$ $1.37$ $0.57$ $1.45$ $0.43$
+51 I am someone who prefers to have others take charge $2.96$ $3.29$ $2.84$ $3.05$ $1.28$ $0.62$ $1.46$ $0.78$
+52 I am someone who is polite, courteous to others $4.50$ $3.55$ $3.65$ $3.43$ $0.70$ $0.93$ $1.40$ $0.75$
+53 I am someone who is persistent, works until finished $4.28$ $3.65$ $3.49$ $3.63$ $0.91$ $0.71$ $1.22$ $0.65$
+54 I am someone who tends to feel depressed, blue $2.65$ $3.09$ $2.63$ $3.40$ $1.46$ $0.46$ $1.30$ $0.72$
+55 I am someone who has little interest in abstract ideas $3.83$ $3.04$ $2.87$ $2.92$ $1.22$ $0.40$ $1.14$ $0.63$
+56 I am someone who shows a lot of enthusiasm $3.35$ $3.67$ $2.71$ $2.93$ $1.22$ $0.84$ $1.34$ $0.94$
+57 I am someone who assumes the best about people $3.37$ $3.21$ $3.39$ $2.93$ $1.28$ $0.64$ $1.40$ $0.75$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $2.88$ $2.81$ $2.82$ $1.30$ $0.40$ $1.07$ $0.62$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $2.90$ $2.72$ $2.61$ $1.33$ $0.52$ $1.34$ $0.76$
+60 I am someone who is original, comes up with new ideas $3.84$ $3.10$ $2.77$ $2.72$ $1.07$ $0.52$ $1.23$ $0.55$
+Facet Level
+1 Sociability $2.65$ $2.97$ $2.88$ $2.55$ $1.17$ $0.47$ $0.84$ $0.74$
+2 Assertiveness $2.98$ $3.11$ $2.97$ $2.69$ $1.04$ $0.40$ $0.90$ $0.49$
+3 Energy Level $3.35$ $3.48$ $2.93$ $2.78$ $0.95$ $0.61$ $1.17$ $0.52$
+4 Compassion $4.02$ $3.45$ $3.32$ $3.46$ $0.81$ $0.68$ $1.35$ $0.79$
+5 Respectfulness $4.29$ $3.56$ $3.65$ $3.44$ $0.70$ $0.67$ $1.30$ $0.55$
+6 Trust $3.31$ $3.06$ $3.26$ $3.00$ $1.00$ $0.32$ $1.21$ $0.41$
+7 Organization $3.86$ $3.12$ $3.00$ $2.99$ $1.01$ $0.41$ $0.92$ $0.30$
+8 Productiveness $3.86$ $3.36$ $3.30$ $3.28$ $0.91$ $0.48$ $1.14$ $0.48$
+9 Responsibility $3.96$ $3.21$ $2.98$ $3.23$ $0.84$ $0.52$ $1.06$ $0.55$
+10 Anxiety $3.15$ $3.21$ $3.07$ $3.54$ $1.11$ $0.49$ $0.89$ $0.65$
+11 Depression $2.51$ $2.89$ $2.75$ $3.26$ $1.14$ $0.52$ $1.17$ $0.65$
+12 Emotional Volatility $2.44$ $3.07$ $2.76$ $3.07$ $1.10$ $0.40$ $1.03$ $0.55$
+13 Intellectual Curiosity $4.05$ $3.28$ $3.02$ $2.87$ $0.83$ $0.36$ $1.05$ $0.51$
+14 Aesthetic Sensitivity $3.89$ $3.43$ $3.14$ $3.01$ $1.01$ $0.53$ $1.09$ $0.33$
+15 Creative Imagination $3.96$ $3.13$ $3.10$ $2.93$ $0.88$ $0.32$ $1.02$ $0.30$
+Domain Level
+1 Extraversion $2.99$ $3.19$ $2.93$ $2.68$ $0.85$ $0.42$ $0.89$ $0.50$
+2 Agreeableness $3.87$ $3.35$ $3.41$ $3.30$ $0.69$ $0.51$ $1.25$ $0.52$
+3 Conscientiousness $3.89$ $3.23$ $3.09$ $3.17$ $0.80$ $0.41$ $0.96$ $0.39$
+4 Neuroticism $2.70$ $3.06$ $2.86$ $3.29$ $1.02$ $0.42$ $0.95$ $0.55$
+5 Openness $3.97$ $3.28$ $3.08$ $2.94$ $0.76$ $0.33$ $1.00$ $0.30$
+Table 15: Gemma-2-27B descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who Is outgoing, sociable $3.02$ $2.92$ $2.50$ $2.31$ $1.40$ $0.73$ $1.15$ $0.70$
+2 I am someone who Is compassionate, has a soft heart $4.34$ $3.62$ $3.16$ $3.71$ $0.90$ $0.75$ $1.30$ $0.82$
+3 I am someone who Tends to be disorganized $3.62$ $3.03$ $3.23$ $3.38$ $1.36$ $0.42$ $1.02$ $0.82$
+4 I am someone who Is relaxed, handles stress well $2.56$ $3.27$ $3.40$ $3.67$ $1.30$ $0.61$ $1.09$ $0.55$
+5 I am someone who Has few artistic interests $3.68$ $3.40$ $2.93$ $2.70$ $1.33$ $0.70$ $0.93$ $0.98$
+6 I am someone who Has an assertive personality $2.96$ $3.06$ $2.80$ $2.31$ $1.38$ $0.55$ $1.11$ $0.60$
+7 I am someone who Is respectful, treats others with respect $4.62$ $4.77$ $4.00$ $4.30$ $0.61$ $0.58$ $1.29$ $0.68$
+8 I am someone who Tends to be lazy $3.62$ $3.56$ $3.27$ $3.60$ $1.31$ $0.64$ $1.24$ $0.79$
+9 I am someone who Stays optimistic after experiencing a setback $2.43$ $2.33$ $2.71$ $2.31$ $1.24$ $0.64$ $1.34$ $0.65$
+10 I am someone who Is curious about many different things $4.43$ $4.08$ $3.58$ $3.53$ $0.81$ $0.54$ $1.21$ $0.64$
+11 I am someone who Rarely feels excited or eager $3.54$ $3.62$ $2.91$ $2.63$ $1.26$ $0.72$ $1.27$ $0.92$
+12 I am someone who Tends to find fault with others $3.34$ $3.67$ $3.15$ $3.63$ $1.32$ $0.54$ $1.25$ $0.71$
+13 I am someone who Is dependable, steady $4.33$ $3.60$ $3.33$ $3.79$ $0.86$ $0.70$ $1.30$ $0.72$
+14 I am someone who Is moody, has up and down mood swings $2.56$ $3.13$ $2.73$ $2.98$ $1.36$ $0.47$ $1.18$ $0.84$
+15 I am someone who Is inventive, finds clever ways to do things $3.90$ $3.28$ $3.21$ $2.41$ $1.03$ $0.51$ $1.09$ $0.63$
+16 I am someone who Tends to be quiet $2.35$ $3.03$ $3.04$ $2.29$ $1.34$ $0.58$ $0.99$ $0.96$
+17 I am someone who Feels little sympathy for others $3.71$ $3.97$ $3.55$ $3.76$ $1.45$ $0.49$ $1.42$ $0.61$
+18 I am someone who Is systematic, likes to keep things in order $3.96$ $3.05$ $2.92$ $2.65$ $1.06$ $0.74$ $1.06$ $0.75$
+19 I am someone who Can be tense $3.29$ $3.30$ $3.18$ $3.53$ $1.27$ $0.52$ $1.01$ $0.58$
+20 I am someone who Is fascinated by art, music, or literature $4.06$ $3.39$ $3.21$ $2.74$ $1.16$ $0.88$ $1.13$ $0.97$
+21 I am someone who Is dominant, acts as a leader $2.90$ $2.83$ $2.66$ $2.14$ $1.35$ $0.57$ $1.09$ $0.43$
+22 I am someone who Starts arguments with others $4.23$ $3.96$ $3.73$ $3.94$ $1.05$ $0.38$ $1.39$ $0.37$
+23 I am someone who Has difficulty getting started on tasks $3.31$ $2.95$ $3.00$ $2.58$ $1.39$ $0.48$ $1.24$ $0.66$
+24 I am someone who Feels secure, comfortable with self $2.26$ $2.87$ $2.73$ $3.43$ $1.27$ $0.72$ $1.35$ $0.71$
+25 I am someone who Avoids intellectual, philosophical discussions $3.94$ $3.18$ $2.92$ $2.55$ $1.21$ $0.82$ $1.22$ $0.85$
+26 I am someone who Is less active than other people $3.35$ $3.26$ $2.98$ $2.32$ $1.31$ $0.78$ $1.01$ $0.76$
+27 I am someone who Has a forgiving nature $3.79$ $3.52$ $3.48$ $3.71$ $1.20$ $0.56$ $1.23$ $0.58$
+28 I am someone who Can be somewhat careless $3.56$ $2.91$ $2.99$ $3.04$ $1.25$ $0.41$ $0.93$ $0.59$
+29 I am someone who Is emotionally stable, not easily upset $2.44$ $3.18$ $3.08$ $3.37$ $1.27$ $0.58$ $1.15$ $0.76$
+30 I am someone who Has little creativity $3.95$ $3.59$ $3.14$ $3.16$ $1.20$ $0.59$ $1.16$ $0.93$
+31 I am someone who Is sometimes shy, introverted $2.35$ $2.75$ $2.96$ $2.36$ $1.36$ $0.56$ $0.83$ $0.87$
+32 I am someone who Is helpful and unselfish with others $4.18$ $3.73$ $3.51$ $3.94$ $0.88$ $0.66$ $1.43$ $0.62$
+33 I am someone who Keeps things neat and tidy $3.71$ $2.90$ $2.86$ $2.87$ $1.24$ $0.62$ $1.07$ $0.64$
+34 I am someone who Worries a lot $3.30$ $3.12$ $2.89$ $3.70$ $1.47$ $0.60$ $1.17$ $0.76$
+35 I am someone who Values art and beauty $4.16$ $3.71$ $3.37$ $2.97$ $1.07$ $0.78$ $1.09$ $1.03$
+36 I am someone who Finds it hard to influence people $3.09$ $2.97$ $2.86$ $2.29$ $1.20$ $0.48$ $1.12$ $0.59$
+37 I am someone who Is sometimes rude to others $3.79$ $3.57$ $3.52$ $3.74$ $1.25$ $0.54$ $1.23$ $0.58$
+38 I am someone who Is efficient, gets things done $4.23$ $3.56$ $3.49$ $3.48$ $0.88$ $0.66$ $1.16$ $0.68$
+39 I am someone who Often feels sad $2.71$ $3.00$ $2.69$ $2.91$ $1.43$ $0.51$ $1.20$ $1.00$
+40 I am someone who Is complex, a deep thinker $4.02$ $3.55$ $3.20$ $2.92$ $1.08$ $0.63$ $1.00$ $0.93$
+41 I am someone who Is full of energy $3.14$ $3.36$ $2.98$ $2.45$ $1.30$ $0.74$ $1.20$ $0.69$
+42 I am someone who Is suspicious of others’ intentions $2.75$ $3.16$ $3.03$ $3.33$ $1.30$ $0.63$ $1.23$ $0.85$
+43 I am someone who Is reliable, can always be counted on $4.32$ $3.63$ $3.48$ $3.86$ $0.88$ $0.74$ $1.37$ $0.78$
+44 I am someone who Keeps their emotions under control $2.22$ $3.06$ $2.68$ $2.62$ $1.14$ $0.56$ $1.16$ $0.78$
+45 I am someone who Has difficulty imagining things $4.17$ $3.41$ $3.19$ $2.78$ $1.06$ $0.54$ $1.02$ $0.61$
+46 I am someone who Is talkative $2.87$ $2.96$ $2.83$ $2.62$ $1.38$ $0.46$ $1.04$ $0.82$
+47 I am someone who Can be cold and uncaring $3.84$ $3.73$ $3.20$ $3.58$ $1.23$ $0.59$ $1.33$ $0.79$
+48 I am someone who Leaves a mess, doesn’t clean up $4.15$ $3.55$ $3.18$ $3.90$ $1.12$ $0.60$ $1.33$ $0.34$
+49 I am someone who Rarely feels anxious or afraid $3.44$ $3.78$ $2.85$ $3.72$ $1.37$ $0.44$ $1.24$ $0.73$
+50 I am someone who Thinks poetry and plays are boring $3.66$ $3.65$ $3.30$ $2.97$ $1.37$ $0.72$ $1.37$ $0.75$
+51 I am someone who Prefers to have others take charge $2.96$ $3.69$ $2.97$ $3.55$ $1.28$ $0.61$ $1.18$ $0.83$
+52 I am someone who Is polite, courteous to others $4.50$ $4.48$ $3.87$ $4.34$ $0.70$ $0.72$ $1.31$ $0.64$
+53 I am someone who Is persistent, works until the task is finished $4.28$ $3.78$ $3.60$ $3.76$ $0.91$ $0.62$ $1.16$ $0.67$
+54 I am someone who Tends to feel depressed, blue $2.65$ $2.93$ $2.66$ $3.03$ $1.46$ $0.54$ $1.27$ $0.98$
+55 I am someone who Has little interest in abstract Ideas $3.83$ $2.84$ $2.74$ $2.12$ $1.22$ $0.74$ $1.01$ $0.46$
+56 I am someone who Shows a lot of Enthusiasm $3.35$ $3.69$ $3.04$ $2.30$ $1.22$ $0.76$ $1.37$ $0.61$
+57 I am someone who Assumes the best about people $3.37$ $3.60$ $3.42$ $3.37$ $1.28$ $0.68$ $1.29$ $0.77$
+58 I am someone who Sometimes behaves irresponsibly $3.64$ $2.86$ $2.88$ $2.84$ $1.30$ $0.44$ $0.91$ $0.65$
+59 I am someone who Is temperamental, gets emotional easily $2.53$ $2.89$ $2.69$ $2.44$ $1.33$ $0.56$ $1.15$ $0.81$
+60 I am someone who Is original, comes up with new Ideas $3.84$ $3.24$ $3.23$ $2.46$ $1.07$ $0.49$ $1.04$ $0.63$
+Facet Level
+1 Sociability $2.65$ $2.91$ $2.83$ $2.40$ $1.17$ $0.48$ $0.87$ $0.72$
+2 Assertiveness $2.98$ $3.14$ $2.82$ $2.57$ $1.04$ $0.43$ $0.95$ $0.43$
+3 Energy Level $3.35$ $3.48$ $2.98$ $2.43$ $0.95$ $0.64$ $1.11$ $0.61$
+4 Compassion $4.02$ $3.76$ $3.36$ $3.75$ $0.81$ $0.51$ $1.28$ $0.59$
+5 Respectfulness $4.29$ $4.19$ $3.78$ $4.08$ $0.70$ $0.45$ $1.22$ $0.47$
+6 Trust $3.31$ $3.49$ $3.27$ $3.51$ $1.00$ $0.49$ $1.16$ $0.62$
+7 Organization $3.86$ $3.13$ $3.05$ $3.20$ $1.01$ $0.47$ $0.93$ $0.48$
+8 Productiveness $3.86$ $3.46$ $3.34$ $3.35$ $0.91$ $0.50$ $1.10$ $0.52$
+9 Responsibility $3.96$ $3.25$ $3.17$ $3.38$ $0.84$ $0.50$ $1.02$ $0.57$
+10 Anxiety $3.15$ $3.37$ $3.08$ $3.66$ $1.11$ $0.41$ $0.77$ $0.49$
+11 Depression $2.51$ $2.78$ $2.70$ $2.92$ $1.14$ $0.49$ $1.22$ $0.69$
+12 Emotional Volatility $2.44$ $3.06$ $2.80$ $2.85$ $1.10$ $0.44$ $1.02$ $0.66$
+13 Intellectual Curiosity $4.05$ $3.41$ $3.11$ $2.78$ $0.83$ $0.52$ $0.92$ $0.57$
+14 Aesthetic Sensitivity $3.89$ $3.54$ $3.20$ $2.84$ $1.01$ $0.63$ $1.03$ $0.79$
+15 Creative Imagination $3.96$ $3.38$ $3.19$ $2.71$ $0.88$ $0.42$ $0.99$ $0.54$
+Domain Level
+1 Extraversion $2.99$ $3.18$ $2.88$ $2.47$ $0.85$ $0.46$ $0.91$ $0.53$
+2 Agreeableness $3.87$ $3.81$ $3.47$ $3.78$ $0.69$ $0.45$ $1.17$ $0.51$
+3 Conscientiousness $3.89$ $3.28$ $3.18$ $3.31$ $0.80$ $0.44$ $0.95$ $0.47$
+4 Neuroticism $2.70$ $3.07$ $2.86$ $3.14$ $1.02$ $0.39$ $0.92$ $0.54$
+5 Openness $3.97$ $3.44$ $3.17$ $2.78$ $0.76$ $0.45$ $0.94$ $0.54$
+Table 16: Llama3-8B descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $2.31$ $2.62$ $2.80$ $1.40$ $0.70$ $1.32$ $0.92$
+2 I am someone who is compassionate, has a soft heart $4.34$ $2.86$ $2.56$ $4.08$ $0.90$ $0.89$ $1.16$ $0.87$
+3 I am someone who tends to be disorganized $3.62$ $3.46$ $2.87$ $2.84$ $1.36$ $0.66$ $1.12$ $0.61$
+4 I am someone who is relaxed, handles stress well $2.56$ $3.29$ $3.34$ $2.32$ $1.30$ $0.88$ $1.24$ $0.58$
+5 I am someone who has few artistic interests $3.68$ $2.24$ $2.11$ $2.17$ $1.33$ $0.62$ $0.94$ $0.84$
+6 I am someone who has an assertive personality $2.96$ $3.25$ $3.12$ $3.11$ $1.38$ $0.82$ $1.16$ $0.74$
+7 I am someone who is respectful, treats others with respect $4.62$ $3.58$ $3.17$ $4.25$ $0.61$ $0.79$ $1.23$ $0.61$
+8 I am someone who tends to be lazy $3.62$ $3.60$ $2.88$ $2.69$ $1.31$ $0.68$ $1.26$ $0.62$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.63$ $2.69$ $2.06$ $1.24$ $0.85$ $1.17$ $0.77$
+10 I am someone who is curious about many different things $4.43$ $3.63$ $3.04$ $3.88$ $0.81$ $0.73$ $1.17$ $0.83$
+11 I am someone who rarely feels excited or eager $3.54$ $2.19$ $2.10$ $2.28$ $1.26$ $0.58$ $0.75$ $0.58$
+12 I am someone who tends to find fault with others $3.34$ $3.64$ $2.97$ $3.23$ $1.32$ $0.69$ $1.02$ $0.67$
+13 I am someone who is dependable, steady $4.33$ $2.90$ $2.76$ $4.13$ $0.86$ $0.88$ $1.20$ $0.72$
+14 I am someone who is moody, has up and down mood swings $2.56$ $2.99$ $2.95$ $3.60$ $1.36$ $0.91$ $1.11$ $0.61$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.03$ $2.68$ $3.22$ $1.03$ $0.69$ $1.01$ $0.58$
+16 I am someone who tends to be quiet $2.35$ $3.71$ $3.17$ $2.20$ $1.34$ $0.63$ $1.10$ $0.88$
+17 I am someone who feels little sympathy for others $3.71$ $3.15$ $2.80$ $2.93$ $1.45$ $0.98$ $1.11$ $0.32$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $2.35$ $2.59$ $3.34$ $1.06$ $0.59$ $1.05$ $0.77$
+19 I am someone who can be tense $3.29$ $3.45$ $3.03$ $3.99$ $1.27$ $0.76$ $1.12$ $0.64$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $2.79$ $2.77$ $3.40$ $1.16$ $0.90$ $1.12$ $0.94$
+21 I am someone who is dominant, acts as a leader $2.90$ $2.74$ $2.65$ $3.05$ $1.35$ $0.75$ $1.02$ $0.61$
+22 I am someone who starts arguments with others $4.23$ $3.94$ $3.50$ $3.79$ $1.05$ $0.37$ $1.12$ $0.47$
+23 I am someone who has difficulty getting started on tasks $3.31$ $2.90$ $2.36$ $2.26$ $1.39$ $0.81$ $0.94$ $0.45$
+24 I am someone who feels secure, comfortable with self $2.26$ $2.83$ $3.32$ $2.26$ $1.27$ $0.92$ $1.20$ $0.71$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $2.11$ $1.98$ $2.30$ $1.21$ $0.44$ $0.63$ $0.60$
+26 I am someone who is less active than other people $3.35$ $2.98$ $2.22$ $2.15$ $1.31$ $0.94$ $0.96$ $0.58$
+27 I am someone who has a forgiving nature $3.79$ $2.91$ $2.65$ $3.74$ $1.20$ $0.90$ $1.08$ $0.73$
+28 I am someone who can be somewhat careless $3.56$ $2.98$ $3.23$ $2.27$ $1.25$ $0.92$ $1.06$ $0.48$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $2.94$ $2.55$ $2.22$ $1.27$ $0.93$ $1.16$ $0.63$
+30 I am someone who has little creativity $3.95$ $2.25$ $2.23$ $2.50$ $1.20$ $0.57$ $0.99$ $0.50$
+31 I am someone who is sometimes shy, introverted $2.35$ $3.63$ $3.13$ $1.97$ $1.36$ $0.64$ $1.07$ $0.85$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.35$ $3.23$ $4.20$ $0.88$ $0.84$ $1.19$ $0.58$
+33 I am someone who keeps things neat and tidy $3.71$ $2.35$ $2.81$ $3.34$ $1.24$ $0.60$ $1.09$ $0.75$
+34 I am someone who worries a lot $3.30$ $2.47$ $3.06$ $4.01$ $1.47$ $0.81$ $1.23$ $0.79$
+35 I am someone who values art and beauty $4.16$ $3.24$ $3.16$ $3.48$ $1.07$ $0.86$ $1.10$ $0.91$
+36 I am someone who finds it hard to influence people $3.09$ $2.20$ $2.01$ $2.06$ $1.20$ $0.58$ $0.60$ $0.36$
+37 I am someone who is sometimes rude to others $3.79$ $3.49$ $2.87$ $2.74$ $1.25$ $0.82$ $1.17$ $0.64$
+38 I am someone who is efficient, gets things done $4.23$ $2.48$ $2.76$ $3.89$ $0.88$ $0.74$ $1.19$ $0.63$
+39 I am someone who often feels sad $2.71$ $2.68$ $2.71$ $3.68$ $1.43$ $0.86$ $1.20$ $0.77$
+40 I am someone who is complex, a deep thinker $4.02$ $3.15$ $2.82$ $3.75$ $1.08$ $0.87$ $1.02$ $0.67$
+41 I am someone who is full of energy $3.14$ $2.78$ $2.91$ $3.15$ $1.30$ $0.89$ $1.28$ $0.86$
+42 I am someone who is suspicious of others’ intentions $2.75$ $3.43$ $2.96$ $2.74$ $1.30$ $0.68$ $1.04$ $0.65$
+43 I am someone who is reliable, can always be counted on $4.32$ $2.82$ $2.84$ $4.17$ $0.88$ $0.83$ $1.21$ $0.66$
+44 I am someone who keeps their emotions under control $2.22$ $2.92$ $2.46$ $2.08$ $1.14$ $0.92$ $1.04$ $0.42$
+45 I am someone who has difficulty imagining things $4.17$ $2.14$ $2.18$ $2.42$ $1.06$ $0.46$ $0.81$ $0.52$
+46 I am someone who is talkative $2.87$ $2.73$ $2.58$ $3.43$ $1.38$ $0.93$ $1.16$ $1.00$
+47 I am someone who can be cold and uncaring $3.84$ $3.64$ $2.93$ $3.01$ $1.23$ $0.70$ $1.19$ $0.61$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $3.91$ $3.46$ $2.59$ $1.12$ $0.44$ $1.14$ $0.53$
+49 I am someone who rarely feels anxious or afraid $3.44$ $2.69$ $2.16$ $2.67$ $1.37$ $0.94$ $0.91$ $0.61$
+50 I am someone who thinks poetry and plays are boring $3.66$ $3.31$ $2.63$ $2.44$ $1.37$ $0.89$ $1.17$ $0.61$
+51 I am someone who prefers to have others take charge $2.96$ $3.17$ $2.57$ $2.58$ $1.28$ $0.94$ $1.03$ $0.51$
+52 I am someone who is polite, courteous to others $4.50$ $2.98$ $2.88$ $4.11$ $0.70$ $0.97$ $1.17$ $0.59$
+53 I am someone who is persistent, works until the task is finished $4.28$ $3.58$ $3.10$ $4.20$ $0.91$ $0.76$ $1.27$ $0.67$
+54 I am someone who tends to feel depressed, blue $2.65$ $2.41$ $2.39$ $3.52$ $1.46$ $0.74$ $1.27$ $0.89$
+55 I am someone who has little interest in abstract ideas $3.83$ $2.09$ $2.05$ $2.15$ $1.22$ $0.38$ $0.56$ $0.68$
+56 I am someone who shows a lot of enthusiasm $3.35$ $3.58$ $2.96$ $2.99$ $1.22$ $0.75$ $1.31$ $0.83$
+57 I am someone who assumes the best about people $3.37$ $3.01$ $2.89$ $3.48$ $1.28$ $0.95$ $1.09$ $0.85$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $2.63$ $2.81$ $2.33$ $1.30$ $0.78$ $1.13$ $0.49$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $2.65$ $2.86$ $3.34$ $1.33$ $0.88$ $1.19$ $0.77$
+60 I am someone who is original, comes up with new ideas $3.84$ $2.93$ $2.91$ $3.62$ $1.07$ $0.83$ $1.08$ $0.60$
+Facet Level
+1 Sociability $2.65$ $3.10$ $2.88$ $2.60$ $1.17$ $0.45$ $0.77$ $0.76$
+2 Assertiveness $2.98$ $2.84$ $2.59$ $2.70$ $1.04$ $0.39$ $0.61$ $0.42$
+3 Energy Level $3.35$ $2.88$ $2.55$ $2.64$ $0.95$ $0.42$ $0.63$ $0.59$
+4 Compassion $4.02$ $3.25$ $2.88$ $3.55$ $0.81$ $0.56$ $0.83$ $0.47$
+5 Respectfulness $4.29$ $3.50$ $3.10$ $3.72$ $0.70$ $0.47$ $0.88$ $0.43$
+6 Trust $3.31$ $3.25$ $2.87$ $3.30$ $1.00$ $0.48$ $0.76$ $0.55$
+7 Organization $3.86$ $3.02$ $2.93$ $3.03$ $1.01$ $0.29$ $0.75$ $0.47$
+8 Productiveness $3.86$ $3.14$ $2.78$ $3.26$ $0.91$ $0.38$ $0.79$ $0.47$
+9 Responsibility $3.96$ $2.83$ $2.91$ $3.22$ $0.84$ $0.49$ $0.82$ $0.47$
+10 Anxiety $3.15$ $2.98$ $2.90$ $3.25$ $1.11$ $0.40$ $0.47$ $0.53$
+11 Depression $2.51$ $2.64$ $2.78$ $2.88$ $1.14$ $0.51$ $0.80$ $0.64$
+12 Emotional Volatility $2.44$ $2.88$ $2.70$ $2.81$ $1.10$ $0.49$ $0.55$ $0.48$
+13 Intellectual Curiosity $4.05$ $2.74$ $2.47$ $3.02$ $0.83$ $0.34$ $0.57$ $0.57$
+14 Aesthetic Sensitivity $3.89$ $2.90$ $2.67$ $2.87$ $1.01$ $0.46$ $0.76$ $0.69$
+15 Creative Imagination $3.96$ $2.59$ $2.50$ $2.94$ $0.88$ $0.35$ $0.62$ $0.43$
+Domain Level
+1 Extraversion $2.99$ $2.94$ $2.67$ $2.65$ $0.85$ $0.33$ $0.57$ $0.52$
+2 Agreeableness $3.87$ $3.33$ $2.95$ $3.53$ $0.69$ $0.44$ $0.75$ $0.43$
+3 Conscientiousness $3.89$ $3.00$ $2.87$ $3.17$ $0.80$ $0.30$ $0.70$ $0.41$
+4 Neuroticism $2.70$ $2.83$ $2.79$ $2.98$ $1.02$ $0.37$ $0.49$ $0.49$
+5 Openness $3.97$ $2.74$ $2.55$ $2.94$ $0.76$ $0.29$ $0.59$ $0.48$
+Table 17: Llama3-70B descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $2.77$ $2.43$ $2.15$ $1.40$ $1.31$ $1.37$ $0.78$
+2 I am someone who is compassionate, has a soft heart $4.34$ $3.87$ $3.15$ $3.93$ $0.90$ $1.12$ $1.44$ $1.01$
+3 I am someone who tends to be disorganized $3.62$ $3.77$ $3.29$ $3.27$ $1.36$ $0.95$ $1.47$ $0.96$
+4 I am someone who is relaxed, handles stress well $2.56$ $2.94$ $3.34$ $2.97$ $1.30$ $1.16$ $1.39$ $0.95$
+5 I am someone who has few artistic interests $3.68$ $3.80$ $2.91$ $2.71$ $1.33$ $1.31$ $1.32$ $1.12$
+6 I am someone who has an assertive personality $2.96$ $3.07$ $2.79$ $2.24$ $1.38$ $1.18$ $1.37$ $0.70$
+7 I am someone who is respectful, treats others with respect $4.62$ $4.43$ $3.60$ $4.23$ $0.61$ $0.92$ $1.43$ $0.78$
+8 I am someone who tends to be lazy $3.62$ $4.23$ $3.44$ $3.77$ $1.31$ $1.04$ $1.49$ $0.96$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.29$ $2.94$ $2.28$ $1.24$ $0.98$ $1.54$ $0.92$
+10 I am someone who is curious about many different things $4.43$ $3.71$ $3.14$ $3.06$ $0.81$ $1.12$ $1.55$ $0.92$
+11 I am someone who rarely feels excited or eager $3.54$ $4.18$ $2.98$ $2.40$ $1.26$ $0.75$ $1.42$ $0.81$
+12 I am someone who tends to find fault with others $3.34$ $4.31$ $3.44$ $3.72$ $1.32$ $0.69$ $1.48$ $0.68$
+13 I am someone who is dependable, steady $4.33$ $3.70$ $3.09$ $3.98$ $0.86$ $1.12$ $1.59$ $0.75$
+14 I am someone who is moody, has up and down mood swings $2.56$ $2.52$ $2.62$ $2.31$ $1.36$ $1.09$ $1.42$ $0.93$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.24$ $2.90$ $2.35$ $1.03$ $1.14$ $1.44$ $0.71$
+16 I am someone who tends to be quiet $2.35$ $3.44$ $3.26$ $2.53$ $1.34$ $1.12$ $1.27$ $0.96$
+17 I am someone who feels little sympathy for others $3.71$ $4.47$ $3.52$ $2.93$ $1.45$ $0.79$ $1.54$ $1.01$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $2.46$ $2.57$ $2.45$ $1.06$ $1.09$ $1.37$ $0.83$
+19 I am someone who can be tense $3.29$ $3.18$ $3.06$ $2.73$ $1.27$ $1.08$ $1.21$ $0.94$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $2.99$ $2.89$ $2.81$ $1.16$ $1.44$ $1.36$ $1.10$
+21 I am someone who is dominant, acts as a leader $2.90$ $2.46$ $2.51$ $2.06$ $1.35$ $1.15$ $1.49$ $0.59$
+22 I am someone who starts arguments with others $4.23$ $4.56$ $3.68$ $4.35$ $1.05$ $0.77$ $1.54$ $0.59$
+23 I am someone who has difficulty getting started on tasks $3.31$ $3.85$ $3.21$ $2.81$ $1.39$ $0.86$ $1.50$ $0.70$
+24 I am someone who feels secure, comfortable with self $2.26$ $2.24$ $2.67$ $2.70$ $1.27$ $1.08$ $1.50$ $0.92$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $3.36$ $3.12$ $2.62$ $1.21$ $1.36$ $1.56$ $1.01$
+26 I am someone who is less active than other people $3.35$ $3.86$ $3.14$ $2.68$ $1.31$ $1.04$ $1.31$ $0.83$
+27 I am someone who has a forgiving nature $3.79$ $3.48$ $3.26$ $3.61$ $1.20$ $0.90$ $1.38$ $0.77$
+28 I am someone who can be somewhat careless $3.56$ $3.19$ $2.99$ $3.18$ $1.25$ $1.11$ $1.38$ $1.01$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $2.66$ $2.89$ $2.68$ $1.27$ $1.01$ $1.48$ $1.09$
+30 I am someone who has little creativity $3.95$ $4.36$ $3.42$ $2.85$ $1.20$ $0.69$ $1.32$ $1.02$
+31 I am someone who is sometimes shy, introverted $2.35$ $3.43$ $3.33$ $2.72$ $1.36$ $1.11$ $1.31$ $0.98$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.88$ $3.22$ $4.12$ $0.88$ $1.05$ $1.51$ $0.73$
+33 I am someone who keeps things neat and tidy $3.71$ $2.47$ $2.61$ $2.71$ $1.24$ $1.00$ $1.35$ $0.70$
+34 I am someone who worries a lot $3.30$ $2.59$ $2.53$ $3.08$ $1.47$ $1.08$ $1.34$ $1.05$
+35 I am someone who values art and beauty $4.16$ $3.23$ $3.05$ $2.88$ $1.07$ $1.30$ $1.35$ $1.06$
+36 I am someone who finds it hard to influence people $3.09$ $3.44$ $2.97$ $2.27$ $1.20$ $1.04$ $1.37$ $0.70$
+37 I am someone who is sometimes rude to others $3.79$ $4.43$ $3.61$ $4.13$ $1.25$ $0.91$ $1.50$ $0.99$
+38 I am someone who is efficient, gets things done $4.23$ $3.45$ $3.10$ $3.09$ $0.88$ $1.12$ $1.47$ $0.92$
+39 I am someone who often feels sad $2.71$ $2.27$ $2.44$ $2.68$ $1.43$ $0.98$ $1.27$ $0.90$
+40 I am someone who is complex, a deep thinker $4.02$ $3.59$ $3.15$ $2.75$ $1.08$ $1.06$ $1.26$ $0.99$
+41 I am someone who is full of energy $3.14$ $3.39$ $2.86$ $2.39$ $1.30$ $1.24$ $1.35$ $0.74$
+42 I am someone who is suspicious of others’ intentions $2.75$ $3.91$ $3.38$ $3.61$ $1.30$ $0.83$ $1.50$ $0.68$
+43 I am someone who is reliable, can always be counted on $4.32$ $3.76$ $3.15$ $3.76$ $0.88$ $1.09$ $1.60$ $0.77$
+44 I am someone who keeps their emotions under control $2.22$ $2.92$ $2.92$ $2.36$ $1.14$ $1.08$ $1.45$ $1.05$
+45 I am someone who has difficulty imagining things $4.17$ $4.08$ $3.43$ $2.13$ $1.06$ $0.53$ $1.26$ $0.52$
+46 I am someone who is talkative $2.87$ $2.53$ $2.49$ $2.36$ $1.38$ $1.11$ $1.28$ $0.92$
+47 I am someone who can be cold and uncaring $3.84$ $4.44$ $3.46$ $4.21$ $1.23$ $0.87$ $1.49$ $0.96$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $4.25$ $3.23$ $4.06$ $1.12$ $0.95$ $1.65$ $0.83$
+49 I am someone who rarely feels anxious or afraid $3.44$ $3.12$ $2.73$ $3.66$ $1.37$ $1.12$ $1.38$ $0.88$
+50 I am someone who thinks poetry and plays are boring $3.66$ $3.85$ $3.14$ $2.31$ $1.37$ $1.31$ $1.54$ $0.70$
+51 I am someone who prefers to have others take charge $2.96$ $4.33$ $3.21$ $3.86$ $1.28$ $0.76$ $1.57$ $0.51$
+52 I am someone who is polite, courteous to others $4.50$ $4.13$ $3.54$ $4.38$ $0.70$ $0.97$ $1.45$ $0.79$
+53 I am someone who is persistent, works until the task is finished $4.28$ $3.87$ $3.12$ $3.76$ $0.91$ $1.13$ $1.48$ $0.95$
+54 I am someone who tends to feel depressed, blue $2.65$ $2.01$ $2.39$ $2.76$ $1.46$ $1.07$ $1.38$ $0.98$
+55 I am someone who has little interest in abstract ideas $3.83$ $3.40$ $2.92$ $2.17$ $1.22$ $1.08$ $1.36$ $0.65$
+56 I am someone who shows a lot of enthusiasm $3.35$ $4.08$ $2.79$ $2.58$ $1.22$ $1.09$ $1.51$ $0.93$
+57 I am someone who assumes the best about people $3.37$ $3.40$ $3.11$ $3.63$ $1.28$ $0.98$ $1.47$ $0.75$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $3.44$ $3.20$ $3.57$ $1.30$ $1.13$ $1.40$ $0.96$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $2.33$ $2.66$ $2.04$ $1.33$ $1.08$ $1.52$ $0.89$
+60 I am someone who is original, comes up with new ideas $3.84$ $3.25$ $2.88$ $2.39$ $1.07$ $1.13$ $1.40$ $0.72$
+Facet Level
+1 Sociability $2.65$ $3.04$ $2.88$ $2.44$ $1.17$ $0.98$ $1.00$ $0.73$
+2 Assertiveness $2.98$ $3.33$ $2.87$ $2.61$ $1.04$ $0.82$ $1.22$ $0.41$
+3 Energy Level $3.35$ $3.88$ $2.94$ $2.51$ $0.95$ $0.88$ $1.19$ $0.64$
+4 Compassion $4.02$ $4.16$ $3.34$ $3.80$ $0.81$ $0.81$ $1.37$ $0.76$
+5 Respectfulness $4.29$ $4.39$ $3.61$ $4.27$ $0.70$ $0.74$ $1.34$ $0.68$
+6 Trust $3.31$ $3.78$ $3.30$ $3.64$ $1.00$ $0.67$ $1.35$ $0.61$
+7 Organization $3.86$ $3.24$ $2.93$ $3.13$ $1.01$ $0.79$ $1.19$ $0.68$
+8 Productiveness $3.86$ $3.85$ $3.22$ $3.36$ $0.91$ $0.87$ $1.29$ $0.69$
+9 Responsibility $3.96$ $3.52$ $3.11$ $3.62$ $0.84$ $0.95$ $1.26$ $0.74$
+10 Anxiety $3.15$ $2.96$ $2.91$ $3.11$ $1.11$ $0.89$ $1.06$ $0.75$
+11 Depression $2.51$ $2.20$ $2.61$ $2.61$ $1.14$ $0.89$ $1.28$ $0.80$
+12 Emotional Volatility $2.44$ $2.61$ $2.77$ $2.35$ $1.10$ $0.92$ $1.26$ $0.88$
+13 Intellectual Curiosity $4.05$ $3.51$ $3.08$ $2.65$ $0.83$ $0.96$ $1.24$ $0.71$
+14 Aesthetic Sensitivity $3.89$ $3.47$ $3.00$ $2.68$ $1.01$ $1.09$ $1.26$ $0.83$
+15 Creative Imagination $3.96$ $3.73$ $3.16$ $2.43$ $0.88$ $0.74$ $1.17$ $0.59$
+Domain Level
+1 Extraversion $2.99$ $3.42$ $2.90$ $2.52$ $0.85$ $0.81$ $1.00$ $0.52$
+2 Agreeableness $3.87$ $4.11$ $3.41$ $3.90$ $0.69$ $0.70$ $1.29$ $0.62$
+3 Conscientiousness $3.89$ $3.54$ $3.08$ $3.37$ $0.80$ $0.79$ $1.16$ $0.65$
+4 Neuroticism $2.70$ $2.59$ $2.77$ $2.69$ $1.02$ $0.81$ $1.10$ $0.72$
+5 Openness $3.97$ $3.57$ $3.08$ $2.59$ $0.76$ $0.78$ $1.15$ $0.59$
+Table 18: GPT-4o-mini descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $2.90$ $3.15$ $2.64$ $1.40$ $0.98$ $1.31$ $0.96$
+2 I am someone who is compassionate, has a soft heart $4.34$ $3.94$ $3.58$ $4.11$ $0.90$ $0.86$ $1.23$ $0.96$
+3 I am someone who tends to be disorganized $3.62$ $3.42$ $3.05$ $3.12$ $1.36$ $0.83$ $1.34$ $0.69$
+4 I am someone who is relaxed, handles stress well $2.56$ $2.82$ $2.63$ $2.96$ $1.30$ $0.78$ $1.25$ $0.90$
+5 I am someone who has few artistic interests $3.68$ $4.05$ $3.14$ $2.94$ $1.33$ $0.89$ $1.34$ $1.10$
+6 I am someone who has an assertive personality $2.96$ $2.98$ $3.12$ $2.55$ $1.38$ $0.76$ $1.22$ $0.74$
+7 I am someone who is respectful, treats others with respect $4.62$ $4.70$ $4.06$ $4.56$ $0.61$ $0.59$ $1.21$ $0.73$
+8 I am someone who tends to be lazy $3.62$ $3.94$ $3.29$ $3.56$ $1.31$ $0.72$ $1.36$ $0.73$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.34$ $2.36$ $2.26$ $1.24$ $0.70$ $1.25$ $0.92$
+10 I am someone who is curious about many different things $4.43$ $4.43$ $3.86$ $3.70$ $0.81$ $0.74$ $1.26$ $0.96$
+11 I am someone who rarely feels excited or eager $3.54$ $4.06$ $3.09$ $3.20$ $1.26$ $0.73$ $1.46$ $0.80$
+12 I am someone who tends to find fault with others $3.34$ $4.00$ $3.45$ $3.79$ $1.32$ $0.44$ $1.37$ $0.57$
+13 I am someone who is dependable, steady $4.33$ $3.78$ $3.62$ $4.27$ $0.86$ $0.84$ $1.36$ $0.78$
+14 I am someone who is moody, has up and down mood swings $2.56$ $2.80$ $2.63$ $3.02$ $1.36$ $0.66$ $1.38$ $0.78$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.62$ $3.45$ $3.07$ $1.03$ $0.67$ $1.20$ $0.63$
+16 I am someone who tends to be quiet $2.35$ $3.21$ $2.84$ $2.22$ $1.34$ $0.89$ $1.15$ $0.97$
+17 I am someone who feels little sympathy for others $3.71$ $4.68$ $3.51$ $4.23$ $1.45$ $0.60$ $1.52$ $0.63$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $3.06$ $3.21$ $3.16$ $1.06$ $0.96$ $1.33$ $0.94$
+19 I am someone who can be tense $3.29$ $3.49$ $3.30$ $3.78$ $1.27$ $0.60$ $1.00$ $0.77$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $3.49$ $3.53$ $3.38$ $1.16$ $1.04$ $1.23$ $1.18$
+21 I am someone who is dominant, acts as a leader $2.90$ $2.58$ $2.93$ $2.39$ $1.35$ $0.80$ $1.25$ $0.66$
+22 I am someone who starts arguments with others $4.23$ $4.87$ $4.05$ $4.48$ $1.05$ $0.38$ $1.34$ $0.57$
+23 I am someone who has difficulty getting started on tasks $3.31$ $3.04$ $2.85$ $2.85$ $1.39$ $0.70$ $1.25$ $0.50$
+24 I am someone who feels secure, comfortable with self $2.26$ $2.63$ $2.37$ $2.65$ $1.27$ $0.78$ $1.20$ $0.92$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $3.61$ $2.95$ $2.69$ $1.21$ $0.87$ $1.45$ $0.94$
+26 I am someone who is less active than other people $3.35$ $3.36$ $2.79$ $2.55$ $1.31$ $0.95$ $1.20$ $0.83$
+27 I am someone who has a forgiving nature $3.79$ $3.58$ $3.51$ $3.56$ $1.20$ $0.63$ $1.11$ $0.82$
+28 I am someone who can be somewhat careless $3.56$ $3.32$ $2.87$ $3.07$ $1.25$ $0.72$ $1.11$ $0.65$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $2.69$ $2.31$ $2.71$ $1.27$ $0.81$ $1.22$ $1.03$
+30 I am someone who has little creativity $3.95$ $4.38$ $3.29$ $3.39$ $1.20$ $0.69$ $1.43$ $0.75$
+31 I am someone who is sometimes shy, introverted $2.35$ $2.85$ $2.86$ $2.11$ $1.36$ $0.81$ $1.11$ $0.98$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.96$ $3.75$ $4.34$ $0.88$ $0.72$ $1.24$ $0.77$
+33 I am someone who keeps things neat and tidy $3.71$ $3.03$ $3.42$ $3.10$ $1.24$ $0.91$ $1.26$ $0.71$
+34 I am someone who worries a lot $3.30$ $2.97$ $2.93$ $3.72$ $1.47$ $0.78$ $1.25$ $1.07$
+35 I am someone who values art and beauty $4.16$ $3.64$ $3.61$ $3.36$ $1.07$ $0.99$ $1.25$ $0.94$
+36 I am someone who finds it hard to influence people $3.09$ $3.06$ $2.68$ $2.68$ $1.20$ $0.68$ $1.08$ $0.58$
+37 I am someone who is sometimes rude to others $3.79$ $4.26$ $3.68$ $3.91$ $1.25$ $0.66$ $1.39$ $0.66$
+38 I am someone who is efficient, gets things done $4.23$ $3.67$ $3.68$ $3.82$ $0.88$ $0.84$ $1.17$ $0.83$
+39 I am someone who often feels sad $2.71$ $2.56$ $2.67$ $3.39$ $1.43$ $0.71$ $1.14$ $0.82$
+40 I am someone who is complex, a deep thinker $4.02$ $3.66$ $3.53$ $3.45$ $1.08$ $0.69$ $0.97$ $0.87$
+41 I am someone who is full of energy $3.14$ $3.31$ $3.09$ $2.75$ $1.30$ $0.94$ $1.29$ $0.80$
+42 I am someone who is suspicious of others’ intentions $2.75$ $3.54$ $3.19$ $3.35$ $1.30$ $0.68$ $1.32$ $0.93$
+43 I am someone who is reliable, can always be counted on $4.32$ $3.84$ $3.85$ $4.21$ $0.88$ $0.83$ $1.20$ $0.81$
+44 I am someone who keeps their emotions under control $2.22$ $2.55$ $2.45$ $2.46$ $1.14$ $0.70$ $1.24$ $0.78$
+45 I am someone who has difficulty imagining things $4.17$ $3.96$ $3.20$ $3.10$ $1.06$ $0.59$ $1.32$ $0.73$
+46 I am someone who is talkative $2.87$ $2.89$ $3.03$ $3.07$ $1.38$ $0.78$ $1.21$ $0.95$
+47 I am someone who can be cold and uncaring $3.84$ $3.93$ $3.33$ $3.98$ $1.23$ $0.64$ $1.41$ $0.84$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $4.37$ $3.18$ $3.63$ $1.12$ $0.91$ $1.65$ $0.53$
+49 I am someone who rarely feels anxious or afraid $3.44$ $3.47$ $2.41$ $3.58$ $1.37$ $0.73$ $1.16$ $0.89$
+50 I am someone who thinks poetry and plays are boring $3.66$ $3.77$ $3.17$ $2.76$ $1.37$ $0.79$ $1.41$ $0.61$
+51 I am someone who prefers to have others take charge $2.96$ $3.34$ $2.49$ $3.21$ $1.28$ $0.79$ $1.27$ $0.72$
+52 I am someone who is polite, courteous to others $4.50$ $4.31$ $3.92$ $4.20$ $0.70$ $0.81$ $1.24$ $0.81$
+53 I am someone who is persistent, works until the task is finished $4.28$ $4.15$ $3.67$ $4.24$ $0.91$ $0.83$ $1.18$ $0.84$
+54 I am someone who tends to feel depressed, blue $2.65$ $2.45$ $2.56$ $3.40$ $1.46$ $0.72$ $1.26$ $0.89$
+55 I am someone who has little interest in abstract ideas $3.83$ $3.31$ $2.89$ $2.63$ $1.22$ $1.10$ $1.39$ $1.12$
+56 I am someone who shows a lot of enthusiasm $3.35$ $3.77$ $3.42$ $2.81$ $1.22$ $0.93$ $1.35$ $0.92$
+57 I am someone who assumes the best about people $3.37$ $3.54$ $3.72$ $3.48$ $1.28$ $0.75$ $1.23$ $0.95$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $3.27$ $3.00$ $3.12$ $1.30$ $0.73$ $1.14$ $0.66$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $2.86$ $2.71$ $2.82$ $1.33$ $0.71$ $1.32$ $0.84$
+60 I am someone who is original, comes up with new ideas $3.84$ $3.55$ $3.56$ $3.07$ $1.07$ $0.73$ $1.19$ $0.68$
+Facet Level
+1 Sociability $2.65$ $2.96$ $2.97$ $2.51$ $1.17$ $0.71$ $0.90$ $0.87$
+2 Assertiveness $2.98$ $2.99$ $2.81$ $2.71$ $1.04$ $0.63$ $0.93$ $0.54$
+3 Energy Level $3.35$ $3.62$ $3.10$ $2.83$ $0.95$ $0.77$ $1.09$ $0.72$
+4 Compassion $4.02$ $4.13$ $3.54$ $4.17$ $0.81$ $0.59$ $1.16$ $0.70$
+5 Respectfulness $4.29$ $4.54$ $3.93$ $4.29$ $0.70$ $0.50$ $1.10$ $0.56$
+6 Trust $3.31$ $3.66$ $3.47$ $3.55$ $1.00$ $0.51$ $1.05$ $0.70$
+7 Organization $3.86$ $3.47$ $3.21$ $3.25$ $1.01$ $0.74$ $0.99$ $0.61$
+8 Productiveness $3.86$ $3.70$ $3.37$ $3.62$ $0.91$ $0.66$ $1.01$ $0.61$
+9 Responsibility $3.96$ $3.55$ $3.34$ $3.67$ $0.84$ $0.67$ $0.94$ $0.62$
+10 Anxiety $3.15$ $3.19$ $2.82$ $3.51$ $1.11$ $0.55$ $0.79$ $0.78$
+11 Depression $2.51$ $2.49$ $2.49$ $2.92$ $1.14$ $0.61$ $0.97$ $0.78$
+12 Emotional Volatility $2.44$ $2.73$ $2.53$ $2.75$ $1.10$ $0.59$ $0.98$ $0.74$
+13 Intellectual Curiosity $4.05$ $3.75$ $3.31$ $3.12$ $0.83$ $0.67$ $0.97$ $0.84$
+14 Aesthetic Sensitivity $3.89$ $3.74$ $3.36$ $3.11$ $1.01$ $0.79$ $1.13$ $0.83$
+15 Creative Imagination $3.96$ $3.88$ $3.38$ $3.16$ $0.88$ $0.54$ $1.01$ $0.59$
+Domain Level
+1 Extraversion $2.99$ $3.19$ $2.96$ $2.68$ $0.85$ $0.64$ $0.86$ $0.63$
+2 Agreeableness $3.87$ $4.11$ $3.65$ $4.00$ $0.69$ $0.50$ $1.05$ $0.60$
+3 Conscientiousness $3.89$ $3.58$ $3.31$ $3.51$ $0.80$ $0.64$ $0.92$ $0.56$
+4 Neuroticism $2.70$ $2.80$ $2.61$ $3.06$ $1.02$ $0.52$ $0.83$ $0.70$
+5 Openness $3.97$ $3.79$ $3.35$ $3.13$ $0.76$ $0.59$ $1.00$ $0.65$
+Table 19: GPT-4o descriptive statistics results.
+$\mu$ $\sigma$
+No. Content Human Persona Shape PSI Human Persona Shape PSI
+Item Level
+1 I am someone who is outgoing, sociable $3.02$ $3.41$ $2.76$ $2.67$ $1.40$ $0.87$ $1.22$ $0.96$
+2 I am someone who is compassionate, has a soft heart $4.34$ $3.51$ $3.10$ $4.07$ $0.90$ $0.84$ $1.22$ $0.93$
+3 I am someone who tends to be disorganized $3.62$ $3.39$ $3.32$ $3.34$ $1.36$ $0.78$ $1.17$ $0.76$
+4 I am someone who is relaxed, handles stress well $2.56$ $3.20$ $3.47$ $2.92$ $1.30$ $0.79$ $1.16$ $0.94$
+5 I am someone who has few artistic interests $3.68$ $3.72$ $3.27$ $2.79$ $1.33$ $1.04$ $1.25$ $1.38$
+6 I am someone who has an assertive personality $2.96$ $3.08$ $2.82$ $2.57$ $1.38$ $0.79$ $1.18$ $0.77$
+7 I am someone who is respectful, treats others with respect $4.62$ $3.38$ $3.48$ $4.15$ $0.61$ $0.89$ $1.23$ $0.76$
+8 I am someone who tends to be lazy $3.62$ $4.00$ $3.51$ $3.56$ $1.31$ $0.87$ $1.28$ $0.79$
+9 I am someone who stays optimistic after experiencing a setback $2.43$ $2.74$ $2.95$ $2.19$ $1.24$ $0.84$ $1.34$ $0.91$
+10 I am someone who is curious about many different things $4.43$ $4.43$ $3.86$ $3.70$ $0.81$ $0.74$ $1.26$ $0.96$
+11 I am someone who rarely feels excited or eager $3.54$ $4.06$ $3.09$ $3.20$ $1.26$ $0.73$ $1.46$ $0.80$
+12 I am someone who tends to find fault with others $3.34$ $4.00$ $3.45$ $3.79$ $1.32$ $0.44$ $1.37$ $0.57$
+13 I am someone who is dependable, steady $4.33$ $3.78$ $3.62$ $4.27$ $0.86$ $0.84$ $1.36$ $0.78$
+14 I am someone who is moody, has up and down mood swings $2.56$ $2.80$ $2.63$ $3.02$ $1.36$ $0.66$ $1.38$ $0.78$
+15 I am someone who is inventive, finds clever ways to do things $3.90$ $3.62$ $3.45$ $3.07$ $1.03$ $0.67$ $1.20$ $0.63$
+16 I am someone who tends to be quiet $2.35$ $3.21$ $2.84$ $2.22$ $1.34$ $0.89$ $1.15$ $0.97$
+17 I am someone who feels little sympathy for others $3.71$ $4.68$ $3.51$ $4.24$ $1.45$ $0.60$ $1.52$ $0.63$
+18 I am someone who is systematic, likes to keep things in order $3.96$ $3.06$ $3.21$ $3.16$ $1.06$ $0.96$ $1.33$ $0.94$
+19 I am someone who can be tense $3.29$ $3.49$ $3.30$ $3.78$ $1.27$ $0.60$ $1.00$ $0.77$
+20 I am someone who is fascinated by art, music, or literature $4.06$ $3.49$ $3.53$ $3.38$ $1.16$ $1.04$ $1.23$ $1.18$
+21 I am someone who is dominant, acts as a leader $2.90$ $2.94$ $2.77$ $2.42$ $1.35$ $0.74$ $1.13$ $0.70$
+22 I am someone who starts arguments with others $4.23$ $3.65$ $3.65$ $4.43$ $1.05$ $0.72$ $1.24$ $0.71$
+23 I am someone who has difficulty getting started on tasks $3.31$ $3.29$ $3.16$ $3.11$ $1.39$ $0.69$ $1.26$ $0.73$
+24 I am someone who feels secure, comfortable with self $2.26$ $2.75$ $3.14$ $2.67$ $1.27$ $0.88$ $1.42$ $0.98$
+25 I am someone who avoids intellectual, philosophical discussions $3.94$ $3.40$ $3.23$ $2.83$ $1.21$ $0.79$ $1.25$ $1.11$
+26 I am someone who is less active than other people $3.35$ $3.80$ $3.36$ $2.65$ $1.31$ $0.99$ $1.24$ $0.97$
+27 I am someone who has a forgiving nature $3.79$ $3.05$ $3.05$ $3.51$ $1.20$ $0.57$ $1.20$ $0.86$
+28 I am someone who can be somewhat careless $3.56$ $3.30$ $3.16$ $3.06$ $1.25$ $0.83$ $1.18$ $0.67$
+29 I am someone who is emotionally stable, not easily upset $2.44$ $3.16$ $3.17$ $2.72$ $1.27$ $0.70$ $1.31$ $1.06$
+30 I am someone who has little creativity $3.95$ $4.10$ $3.59$ $3.38$ $1.20$ $0.76$ $1.22$ $0.84$
+31 I am someone who is sometimes shy, introverted $2.35$ $2.85$ $2.86$ $2.13$ $1.36$ $0.87$ $1.07$ $1.04$
+32 I am someone who is helpful and unselfish with others $4.18$ $3.32$ $3.27$ $4.07$ $0.88$ $0.80$ $1.21$ $0.74$
+33 I am someone who keeps things neat and tidy $3.71$ $2.99$ $3.07$ $3.06$ $1.24$ $0.64$ $1.09$ $0.74$
+34 I am someone who worries a lot $3.30$ $2.96$ $2.79$ $3.65$ $1.47$ $0.70$ $1.23$ $1.18$
+35 I am someone who values art and beauty $4.16$ $3.65$ $3.61$ $3.36$ $1.07$ $0.82$ $1.18$ $1.00$
+36 I am someone who finds it hard to influence people $3.09$ $3.32$ $3.25$ $2.74$ $1.20$ $0.68$ $1.26$ $0.69$
+37 I am someone who is sometimes rude to others $3.79$ $3.24$ $3.50$ $3.68$ $1.25$ $0.66$ $1.31$ $0.78$
+38 I am someone who is efficient, gets things done $4.23$ $3.31$ $3.07$ $3.70$ $0.88$ $0.82$ $1.25$ $0.82$
+39 I am someone who often feels sad $2.71$ $2.56$ $2.67$ $3.39$ $1.43$ $0.71$ $1.14$ $0.82$
+40 I am someone who is complex, a deep thinker $4.02$ $3.66$ $3.53$ $3.45$ $1.08$ $0.69$ $0.97$ $0.87$
+41 I am someone who is full of energy $3.14$ $3.45$ $2.99$ $2.83$ $1.30$ $0.94$ $1.23$ $0.87$
+42 I am someone who is suspicious of others’ intentions $2.75$ $3.14$ $3.27$ $3.42$ $1.30$ $0.58$ $1.26$ $0.99$
+43 I am someone who is reliable, can always be counted on $4.32$ $3.84$ $3.85$ $4.21$ $0.88$ $0.83$ $1.20$ $0.81$
+44 I am someone who keeps their emotions under control $2.22$ $2.55$ $2.45$ $2.46$ $1.14$ $0.70$ $1.24$ $0.78$
+45 I am someone who has difficulty imagining things $4.17$ $4.06$ $3.57$ $3.23$ $1.06$ $0.77$ $1.25$ $0.95$
+46 I am someone who is talkative $2.87$ $3.11$ $2.87$ $3.31$ $1.38$ $0.70$ $1.04$ $1.09$
+47 I am someone who can be cold and uncaring $3.84$ $3.95$ $3.60$ $4.08$ $1.23$ $0.80$ $1.35$ $0.79$
+48 I am someone who leaves a mess, doesn’t clean up $4.15$ $3.47$ $3.39$ $3.31$ $1.12$ $0.71$ $1.26$ $0.55$
+49 I am someone who rarely feels anxious or afraid $3.44$ $3.20$ $2.85$ $3.50$ $1.37$ $0.83$ $1.15$ $1.02$
+50 I am someone who thinks poetry and plays are boring $3.66$ $3.54$ $3.44$ $2.93$ $1.37$ $0.75$ $1.38$ $0.89$
+51 I am someone who prefers to have others take charge $2.96$ $3.54$ $3.39$ $3.51$ $1.28$ $0.69$ $1.27$ $0.71$
+52 I am someone who is polite, courteous to others $4.50$ $3.28$ $3.33$ $3.96$ $0.70$ $0.65$ $1.14$ $0.75$
+53 I am someone who is persistent, works until the task is finished $4.28$ $3.56$ $3.01$ $4.09$ $0.91$ $0.87$ $1.26$ $0.80$
+54 I am someone who tends to feel depressed, blue $2.65$ $2.70$ $2.60$ $3.28$ $1.46$ $0.79$ $1.31$ $1.07$
+55 I am someone who has little interest in abstract ideas $3.83$ $3.25$ $3.08$ $2.55$ $1.22$ $0.96$ $1.22$ $1.17$
+56 I am someone who shows a lot of enthusiasm $3.35$ $3.86$ $2.81$ $2.70$ $1.22$ $0.85$ $1.34$ $0.85$
+57 I am someone who assumes the best about people $3.37$ $3.54$ $3.72$ $3.48$ $1.28$ $0.75$ $1.23$ $0.95$
+58 I am someone who sometimes behaves irresponsibly $3.64$ $3.27$ $3.00$ $3.12$ $1.30$ $0.74$ $1.14$ $0.66$
+59 I am someone who is temperamental, gets emotional easily $2.53$ $2.86$ $2.71$ $2.82$ $1.33$ $0.71$ $1.32$ $0.84$
+60 I am someone who is original, comes up with new ideas $3.84$ $3.28$ $2.96$ $3.07$ $1.07$ $0.73$ $1.19$ $0.68$
+Facet Level
+1 Sociability $2.65$ $3.26$ $2.94$ $2.57$ $1.17$ $0.69$ $0.95$ $0.91$
+2 Assertiveness $2.98$ $3.22$ $3.06$ $2.81$ $1.04$ $0.55$ $1.10$ $0.57$
+3 Energy Level $3.35$ $3.85$ $3.09$ $2.79$ $0.95$ $0.75$ $1.12$ $0.78$
+4 Compassion $4.02$ $3.63$ $3.42$ $4.11$ $0.81$ $0.69$ $1.18$ $0.68$
+5 Respectfulness $4.29$ $3.39$ $3.49$ $4.05$ $0.70$ $0.58$ $1.13$ $0.62$
+6 Trust $3.31$ $3.15$ $3.21$ $3.49$ $1.00$ $0.44$ $1.17$ $0.78$
+7 Organization $3.86$ $3.21$ $3.18$ $3.22$ $1.01$ $0.57$ $0.98$ $0.63$
+8 Productiveness $3.86$ $3.54$ $3.19$ $3.62$ $0.91$ $0.65$ $1.12$ $0.67$
+9 Responsibility $3.96$ $3.24$ $3.03$ $3.59$ $0.84$ $0.59$ $1.10$ $0.63$
+10 Anxiety $3.15$ $3.15$ $3.06$ $3.44$ $1.11$ $0.54$ $0.95$ $0.89$
+11 Depression $2.51$ $2.76$ $2.86$ $2.85$ $1.14$ $0.56$ $1.16$ $0.88$
+12 Emotional Volatility $2.44$ $3.07$ $2.92$ $2.82$ $1.10$ $0.43$ $1.06$ $0.78$
+13 Intellectual Curiosity $4.05$ $3.55$ $3.13$ $3.09$ $0.83$ $0.68$ $1.04$ $0.89$
+14 Aesthetic Sensitivity $3.89$ $3.63$ $3.25$ $3.15$ $1.01$ $0.75$ $1.12$ $0.97$
+15 Creative Imagination $3.96$ $3.72$ $3.29$ $3.19$ $0.88$ $0.57$ $1.05$ $0.67$
+Domain Level
+1 Extraversion $2.99$ $3.44$ $3.03$ $2.72$ $0.85$ $0.57$ $0.95$ $0.67$
+2 Agreeableness $3.87$ $3.39$ $3.37$ $3.89$ $0.69$ $0.51$ $1.11$ $0.62$
+3 Conscientiousness $3.89$ $3.33$ $3.13$ $3.47$ $0.80$ $0.53$ $1.00$ $0.58$
+4 Neuroticism $2.70$ $2.99$ $2.94$ $3.04$ $1.02$ $0.43$ $0.97$ $0.77$
+5 Openness $3.97$ $3.63$ $3.22$ $3.14$ $0.76$ $0.58$ $1.03$ $0.73$
+Model Method Item Facet Domain
+[…]
+Mistral-7B Persona $0.42$ $0.46$ $0.31$ $0.41$ $0.28$ $0.36$
+Shape $0.40$ $0.25$ $0.31$ $0.18$ $0.25$ $0.13$
+PSI $0.51$ $0.43$ $0.32$ $0.33$ $0.29$ $0.26$
+Gemma-2-9B Persona $0.51$ $0.62$ $0.48$ $0.49$ $0.49$ $0.41$
+Shape $0.58$ $0.21$ $0.52$ $0.22$ $0.47$ $0.21$
+PSI $0.68$ $0.50$ $0.65$ $0.45$ $0.64$ $0.37$
+Gemma-2-27B Persona $0.42$ $0.59$ $0.37$ $0.47$ $0.35$ $0.39$
+Shape $0.53$ $0.20$ $0.47$ $0.19$ $0.44$ $0.19$
+PSI $0.63$ $0.48$ $0.59$ $0.38$ $0.56$ $0.31$
+Llama3-8B Persona $0.78$ $0.45$ $0.65$ $0.53$ $0.57$ $0.48$
+Shape $0.89$ $0.23$ $0.82$ $0.29$ $0.75$ $0.23$
+PSI $0.64$ $0.53$ $0.54$ $0.44$ $0.54$ $0.36$
+Llama3-70B Persona $0.51$ $0.26$ $0.33$ $0.15$ $0.31$ $0.06$
+Shape $0.63$ $0.25$ $0.52$ $0.29$ $0.46$ $0.32$
+PSI $0.66$ $0.36$ $0.53$ $0.27$ $0.48$ $0.20$
+GPT-4o-mini Persona $0.33$ $0.44$ $0.21$ $0.33$ $0.21$ $0.25$
+Shape $0.42$ $0.18$ $0.39$ $0.17$ $0.31$ $0.18$
+PSI $0.45$ $0.40$ $0.40$ $0.27$ $0.40$ $0.20$
+GPT-4o Persona $0.48$ $0.44$ $0.42$ $0.36$ $0.42$ $0.30$
+Shape $0.58$ $0.16$ $0.50$ $0.18$ $0.46$ $0.21$
+PSI $0.47$ $0.34$ $0.40$ $0.22$ $0.38$ $0.15$
+Table 20: MAE results of $\mu$ and $\sigma$ for Item, Facet, and Domain across different Backbone Model + Method combinations.
+Table 21: Model fit indices for BFI-2 TFMs of each domain and FFM.
+Model Chi‑square df CFI TLI RMSEA SRMR
+Extraversion TFM
+Mistral-7B+Persona 188.882 51 0.873 0.836 0.095 0.066
+Mistral-7B+Shape 676.095 51 0.781 0.717 0.202 0.095
+Mistral-7B+PSI 395.981 51 0.828 0.778 0.138 0.079
+Gemma-2-9B+Persona 254.304 51 0.880 0.845 0.115 0.074
+Gemma-2-9B+Shape 1084.518 51 0.695 0.605 0.260 0.113
+Gemma-2-9B+PSI 267.081 51 0.884 0.850 0.109 0.080
+Gemma-2-27B+Persona 353.266 51 0.860 0.818 0.141 0.073
+Gemma-2-27B+Shape 1101.801 51 0.744 0.668 0.262 0.154
+Gemma-2-27B+PSI 290.205 51 0.900 0.871 0.115 0.057
+Llama3-8B+Persona 430.846 51 0.556 0.425 0.158 0.166
+Llama3-8B+Shape 553.196 51 0.607 0.492 0.181 0.151
+Llama3-8B+PSI 275.953 51 0.899 0.869 0.111 0.055
+Llama3-70B+Persona 316.715 51 0.885 0.851 0.132 0.060
+Llama3-70B+Shape 944.452 51 0.681 0.588 0.242 0.130
+Llama3-70B+PSI 367.869 51 0.815 0.761 0.132 0.074
+GPT-4o-mini+Persona 362.974 51 0.874 0.837 0.143 0.065
+GPT-4o-mini+Shape 1674.008 51 0.504 0.359 0.326 0.264
+GPT-4o-mini+PSI 369.695 51 0.898 0.868 0.132 0.063
+GPT-4o+Persona 261.099 51 0.886 0.852 0.117 0.058
+GPT-4o+Shape 663.947 51 0.823 0.771 0.200 0.101
+GPT-4o+PSI 376.669 51 0.895 0.864 0.134 0.055
+Human 994.226 51 0.892 0.861 0.109 0.057
+Agreeableness TFM
+Mistral-7B+Persona 294.949 51 0.818 0.764 0.126 0.091
+Mistral-7B+Shape 578.669 51 0.802 0.744 0.186 0.102
+Mistral-7B+PSI 298.885 51 0.908 0.881 0.117 0.061
+Gemma-2-9B+Persona 355.402 51 0.853 0.810 0.141 0.073
+Gemma-2-9B+Shape 771.602 51 0.881 0.845 0.217 0.042
+Gemma-2-9B+PSI 257.858 51 0.914 0.889 0.107 0.067
+Gemma-2-27B+Persona 377.813 51 0.863 0.823 0.146 0.068
+Gemma-2-27B+Shape 865.382 51 0.863 0.823 0.231 0.058
+Gemma-2-27B+PSI 347.315 51 0.900 0.870 0.128 0.065
+Llama3-8B+Persona 588.254 51 0.655 0.554 0.187 0.179
+Llama3-8B+Shape 1183.880 51 0.557 0.426 0.272 0.222
+Llama3-8B+PSI 241.972 51 0.898 0.868 0.102 0.071
+Llama3-70B+Persona 616.199 51 0.798 0.739 0.192 0.088
+Llama3-70B+Shape 884.948 51 0.841 0.794 0.233 0.069
+Llama3-70B+PSI 295.615 51 0.924 0.901 0.116 0.048
+GPT-4o-mini+Persona 340.628 51 0.884 0.849 0.138 0.075
+GPT-4o-mini+Shape 1213.078 51 0.731 0.652 0.276 0.154
+GPT-4o-mini+PSI 212.855 51 0.948 0.933 0.094 0.046
+GPT-4o+Persona 210.863 51 0.900 0.870 0.102 0.062
+GPT-4o+Shape 525.799 51 0.909 0.882 0.176 0.039
+GPT-4o+PSI 182.956 51 0.957 0.945 0.085 0.045
+Human 903.288 51 0.875 0.838 0.104 0.060
+Conscientiousness TFM
+Mistral-7B+Persona 415.371 51 0.809 0.753 0.154 0.103
+Mistral-7B+Shape 979.589 51 0.730 0.650 0.246 0.144
+Mistral-7B+PSI 501.680 51 0.811 0.755 0.157 0.092
+Gemma-2-9B+Persona 580.705 51 0.751 0.678 0.186 0.114
+Gemma-2-9B+Shape 1592.597 51 0.653 0.551 0.317 0.122
+Gemma-2-9B+PSI 330.247 51 0.848 0.804 0.124 0.087
+Gemma-2-27B+Persona 445.890 51 0.837 0.789 0.161 0.080
+Gemma-2-27B+Shape 1666.873 51 0.678 0.584 0.325 0.114
+Gemma-2-27B+PSI 354.743 51 0.861 0.821 0.129 0.075
+Llama3-8B+Persona 357.786 51 0.783 0.719 0.142 0.117
+Llama3-8B+Shape 987.900 51 0.568 0.441 0.247 0.215
+Llama3-8B+PSI 437.070 51 0.805 0.748 0.146 0.114
+Llama3-70B+Persona 705.359 51 0.768 0.699 0.207 0.109
+Llama3-70B+Shape 1498.095 51 0.645 0.541 0.308 0.192
+Llama3-70B+PSI 467.824 51 0.850 0.806 0.151 0.062
+GPT-4o-mini+Persona 625.986 51 0.810 0.755 0.194 0.080
+GPT-4o-mini+Shape 1215.114 51 0.614 0.500 0.276 0.200
+GPT-4o-mini+PSI 396.471 51 0.882 0.848 0.138 0.056
+GPT-4o+Persona 385.339 51 0.780 0.716 0.148 0.088
+GPT-4o+Shape 705.340 51 0.826 0.775 0.207 0.069
+GPT-4o+PSI 411.654 51 0.883 0.849 0.141 0.057
+Human 1045.113 51 0.897 0.867 0.112 0.058
+Neuroticism TFM
+Mistral-7B+Persona 530.516 51 0.585 0.462 0.177 0.158
+Mistral-7B+Shape 557.616 51 0.810 0.754 0.182 0.095
+Mistral-7B+PSI 431.937 51 0.841 0.794 0.145 0.066
+Gemma-2-9B+Persona 469.967 51 0.816 0.762 0.165 0.096
+Gemma-2-9B+Shape 1072.848 51 0.718 0.635 0.258 0.134
+Gemma-2-9B+PSI 611.373 51 0.787 0.724 0.175 0.094
+Gemma-2-27B+Persona 598.476 51 0.743 0.668 0.189 0.118
+Gemma-2-27B+Shape 759.167 51 0.839 0.791 0.215 0.085
+Gemma-2-27B+PSI 478.337 51 0.836 0.787 0.153 0.082
+Llama3-8B+Persona 540.289 51 0.586 0.464 0.179 0.150
+Llama3-8B+Shape 956.346 51 0.545 0.412 0.243 0.257
+Llama3-8B+PSI 522.249 51 0.814 0.760 0.161 0.081
+Llama3-70B+Persona 346.207 51 0.884 0.850 0.139 0.059
+Llama3-70B+Shape 808.502 51 0.776 0.710 0.223 0.099
+Llama3-70B+PSI 525.423 51 0.855 0.812 0.161 0.065
+GPT-4o-mini+Persona 477.940 51 0.803 0.745 0.167 0.076
+GPT-4o-mini+Shape 856.807 51 0.667 0.569 0.229 0.151
+GPT-4o-mini+PSI 539.536 51 0.872 0.834 0.164 0.070
+GPT-4o+Persona 304.611 51 0.756 0.684 0.129 0.091
+GPT-4o+Shape 960.437 51 0.720 0.638 0.244 0.099
+GPT-4o+PSI 541.928 51 0.874 0.837 0.164 0.069
+Human 927.971 51 0.931 0.911 0.105 0.052
+Openness TFM
+Mistral-7B+Persona 316.958 51 0.748 0.674 0.132 0.119
+Mistral-7B+Shape 962.242 51 0.653 0.551 0.244 0.149
+Mistral-7B+PSI 455.750 51 0.722 0.640 0.149 0.114
+Gemma-2-9B+Persona 313.753 51 0.772 0.704 0.131 0.103
+Gemma-2-9B+Shape 737.946 51 0.837 0.789 0.212 0.055
+Gemma-2-9B+PSI 380.816 51 0.748 0.674 0.135 0.105
+Gemma-2-27B+Persona 381.298 51 0.823 0.771 0.147 0.103
+Gemma-2-27B+Shape 702.555 51 0.862 0.821 0.206 0.051
+Gemma-2-27B+PSI 471.557 51 0.831 0.781 0.152 0.116
+Llama3-8B+Persona 211.861 51 0.806 0.749 0.103 0.087
+Llama3-8B+Shape 825.535 51 0.637 0.530 0.225 0.209
+Llama3-8B+PSI 318.777 51 0.889 0.856 0.121 0.070
+Llama3-70B+Persona 256.032 51 0.904 0.876 0.116 0.077
+Llama3-70B+Shape 562.061 51 0.864 0.823 0.183 0.057
+Llama3-70B+PSI 408.648 51 0.853 0.810 0.140 0.117
+GPT-4o-mini+Persona 257.955 51 0.899 0.869 0.116 0.068
+GPT-4o-mini+Shape 1464.309 51 0.656 0.555 0.304 0.201
+GPT-4o-mini+PSI 307.850 51 0.914 0.889 0.119 0.067
+GPT-4o+Persona 226.649 51 0.910 0.883 0.107 0.057
+GPT-4o+Shape 502.104 51 0.878 0.842 0.172 0.055
+GPT-4o+PSI 350.639 51 0.909 0.882 0.128 0.069
+Human 904.843 51 0.900 0.870 0.104 0.064
+FFM
+Mistral-7B+Persona 560.227 80 0.802 0.740 0.141 0.088
+Mistral-7B+Shape 1457.055 80 0.756 0.680 0.240 0.115
+Mistral-7B+PSI 1067.765 80 0.773 0.701 0.186 0.111
+Gemma-2-9B+Persona 731.270 80 0.790 0.725 0.165 0.107
+Gemma-2-9B+Shape 2336.944 80 0.696 0.601 0.307 0.144
+Gemma-2-9B+PSI 627.697 80 0.836 0.784 0.138 0.084
+Gemma-2-27B+Persona 756.777 80 0.821 0.765 0.168 0.093
+Gemma-2-27B+Shape 2325.784 80 0.714 0.624 0.306 0.139
+Gemma-2-27B+PSI 706.399 80 0.836 0.784 0.148 0.101
+Llama3-8B+Persona 302.351 80 0.849 0.802 0.096 0.078
+Llama3-8B+Shape 839.663 80 0.784 0.716 0.178 0.113
+Llama3-8B+PSI 868.954 80 0.782 0.714 0.166 0.117
+Llama3-70B+Persona 954.848 80 0.780 0.711 0.191 0.120
+Llama3-70B+Shape 1996.472 80 0.720 0.633 0.283 0.148
+Llama3-70B+PSI 901.061 80 0.788 0.722 0.170 0.112
+GPT-4o-mini+Persona 1079.430 80 0.779 0.710 0.204 0.108
+GPT-4o-mini+Shape 1350.813 80 0.797 0.734 0.230 0.104
+GPT-4o-mini+PSI 1144.935 80 0.760 0.685 0.193 0.120
+GPT-4o+Persona 551.459 80 0.819 0.762 0.140 0.097
+GPT-4o+Shape 2132.208 80 0.711 0.621 0.292 0.143
+GPT-4o+PSI 1029.123 80 0.776 0.706 0.182 0.120
+Human 1747.433 80 0.852 0.806 0.116 0.080
+Table 22: Tucker’s congruence coefficient for BFI-2 TFM of Extraversion.
+Model Sociability Assertiveness Energy Level
+Mistral-7B+Persona 0.91 0.92 0.98
+Mistral-7B+Shape 0.99 0.94 0.95
+Mistral-7B+PSI 1.00 0.96 0.99
+Gemma-2-9B+Persona 1.00 0.97 1.00
+Gemma-2-9B+Shape 0.97 0.86 0.97
+Gemma-2-9B+PSI 1.00 0.97 0.86
+Gemma-2-27B+Persona 1.00 0.97 0.98
+Gemma-2-27B+Shape 1.00 0.97 0.97
+Gemma-2-27B+PSI 1.00 0.92 0.98
+Llama3-8B+Persona -0.78 0.55 0.71
+Llama3-8B+Shape -0.81 0.79 -0.80
+Llama3-8B+PSI 1.00 0.98 0.98
+Llama3-70B+Persona 0.99 0.98 0.98
+Llama3-70B+Shape 0.99 0.97 0.96
+Llama3-70B+PSI 1.00 0.95 0.99
+GPT-4o-mini+Persona 1.00 0.99 0.99
+GPT-4o-mini+Shape 0.91 0.95 1.00
+GPT-4o-mini+PSI 1.00 0.99 0.98
+GPT-4o+Persona 1.00 0.97 0.98
+GPT-4o+Shape 1.00 0.99 0.97
+GPT-4o+PSI 1.00 0.98 0.98
+Table 23: Tucker’s congruence coefficient for BFI-2 TFM of Agreeableness.
+Model Compassion Respectfulness Trust
+Mistral-7B+Persona 0.95 0.98 0.92
+Mistral-7B+Shape 0.92 0.97 0.96
+Mistral-7B+PSI 0.99 0.97 0.98
+Gemma-2-9B+Persona 0.97 0.99 0.97
+Gemma-2-9B+Shape 0.96 1.00 1.00
+Gemma-2-9B+PSI 0.99 0.90 -0.87
+Gemma-2-27B+Persona 0.97 0.99 1.00
+Gemma-2-27B+Shape 0.97 1.00 0.99
+Gemma-2-27B+PSI 0.96 0.98 1.00
+Llama3-8B+Persona 0.86 0.71 -0.60
+Llama3-8B+Shape 0.96 0.93 0.94
+Llama3-8B+PSI 0.99 0.94 0.97
+Llama3-70B+Persona 0.96 1.00 0.99
+Llama3-70B+Shape 0.95 1.00 0.99
+Llama3-70B+PSI 1.00 0.99 1.00
+GPT-4o-mini+Persona 0.98 0.99 0.98
+GPT-4o-mini+Shape 0.91 0.98 0.94
+GPT-4o-mini+PSI 0.97 0.97 1.00
+GPT-4o+Persona 0.95 1.00 0.98
+GPT-4o+Shape 0.97 1.00 1.00
+GPT-4o+PSI 0.98 0.98 1.00
+Table 24: Tucker’s congruence coefficient for BFI-2 TFM of Conscientiousness.
+Model Organization Productiveness Responsibility
+Mistral-7B+Persona 0.98 0.99 0.95
+Mistral-7B+Shape 0.93 0.99 0.97
+Mistral-7B+PSI 0.98 0.99 0.98
+Gemma-2-9B+Persona 0.98 0.95 0.98
+Gemma-2-9B+Shape 0.99 1.00 0.99
+Gemma-2-9B+PSI 0.81 0.98 0.94
+Gemma-2-27B+Persona 0.99 0.99 0.99
+Gemma-2-27B+Shape 0.98 1.00 0.99
+Gemma-2-27B+PSI 0.99 1.00 0.97
+Llama3-8B+Persona -0.43 0.33 0.51
+Llama3-8B+Shape 0.81 0.82 0.86
+Llama3-8B+PSI 0.89 0.99 0.97
+Llama3-70B+Persona 1.00 0.99 0.97
+Llama3-70B+Shape 0.96 0.99 0.96
+Llama3-70B+PSI 1.00 0.97 1.00
+GPT-4o-mini+Persona 1.00 0.99 0.99
+GPT-4o-mini+Shape 0.90 0.98 0.96
+GPT-4o-mini+PSI 1.00 1.00 1.00
+GPT-4o+Persona 0.99 0.99 0.99
+GPT-4o+Shape 0.99 0.99 0.99
+GPT-4o+PSI 0.99 1.00 1.00
+Table 25: Tucker’s congruence coefficient for BFI-2 TFM of Neuroticism.
+Model Anxiety Depression Emotional Volatility
+Mistral-7B+Persona 0.91 0.93 0.88
+Mistral-7B+Shape 0.86 1.00 0.93
+Mistral-7B+PSI 0.99 0.99 1.00
+Gemma-2-9B+Persona 0.99 1.00 0.99
+Gemma-2-9B+Shape 0.92 1.00 0.91
+Gemma-2-9B+PSI 0.99 1.00 0.98
+Gemma-2-27B+Persona 0.97 1.00 1.00
+Gemma-2-27B+Shape 0.85 0.99 0.97
+Gemma-2-27B+PSI 0.99 0.99 0.99
+Llama3-8B+Persona 0.17 0.03 -0.42
+Llama3-8B+Shape -0.15 0.60 0.32
+Llama3-8B+PSI 0.99 1.00 0.99
+Llama3-70B+Persona 0.99 0.99 1.00
+Llama3-70B+Shape 0.99 0.99 0.99
+Llama3-70B+PSI 1.00 1.00 1.00
+GPT-4o-mini+Persona 1.00 0.99 0.99
+GPT-4o-mini+Shape 0.85 0.99 0.96
+GPT-4o-mini+PSI 1.00 1.00 1.00
+GPT-4o+Persona 0.99 0.97 0.97
+GPT-4o+Shape 0.98 0.99 0.99
+GPT-4o+PSI 1.00 1.00 1.00
+Table 26: Tucker’s congruence coefficient for BFI-2 TFM of Openness.
+Model Intellectual Curiosity Aesthetic Sensitivity Creative Imagination
+Mistral-7B+Persona 1.00 0.95 0.94
+Mistral-7B+Shape 0.96 0.99 0.98
+Mistral-7B+PSI 0.81 0.94 0.86
+Gemma-2-9B+Persona 0.96 0.94 0.87
+Gemma-2-9B+Shape 0.99 0.99 0.99
+Gemma-2-9B+PSI 0.96 -0.62 0.70
+Gemma-2-27B+Persona 0.98 0.98 0.99
+Gemma-2-27B+Shape 0.95 0.99 0.99
+Gemma-2-27B+PSI 0.98 0.98 1.00
+Llama3-8B+Persona 0.46 -0.70 0.34
+Llama3-8B+Shape 0.69 0.94 0.72
+Llama3-8B+PSI 1.00 0.98 0.98
+Llama3-70B+Persona 1.00 0.98 0.98
+Llama3-70B+Shape 0.98 0.98 1.00
+Llama3-70B+PSI 0.98 0.97 0.96
+GPT-4o-mini+Persona 0.99 1.00 0.99
+GPT-4o-mini+Shape 0.93 0.93 0.91
+GPT-4o-mini+PSI 1.00 0.98 1.00
+GPT-4o+Persona 1.00 0.99 1.00
+GPT-4o+Shape 0.99 0.99 1.00
+GPT-4o+PSI 1.00 0.99 0.99
+Table 27: Tucker’s congruence coefficient for BFI-2 FFM.
+Model Extraversion Agreeableness Conscientiousness Neuroticism Openness
+Mistral-7B+Persona 0.99 1.00 1.00 0.99 1.00
+Mistral-7B+Shape 0.99 0.99 0.99 0.99 0.99
+Mistral-7B+PSI 1.00 1.00 1.00 0.99 0.99
+Gemma-2-9B+Persona 0.99 1.00 0.98 0.99 0.99
+Gemma-2-9B+Shape 1.00 0.99 0.98 0.99 1.00
+Gemma-2-9B+PSI 0.99 1.00 0.99 0.99 0.96
+Gemma-2-27B+Persona 1.00 1.00 0.99 1.00 0.99
+Gemma-2-27B+Shape 1.00 0.99 0.99 0.99 1.00
+Gemma-2-27B+PSI 1.00 1.00 1.00 0.99 1.00
+Llama3-8B+Persona 0.99 1.00 0.97 0.96 0.99
+Llama3-8B+Shape 0.98 0.99 0.99 0.96 0.99
+Llama3-8B+PSI 1.00 0.99 1.00 1.00 1.00
+Llama3-70B+Persona 0.99 1.00 0.98 1.00 0.99
+Llama3-70B+Shape 1.00 0.99 0.99 1.00 0.99
+Llama3-70B+PSI 1.00 1.00 0.99 1.00 0.99
+GPT-4o-mini+Persona 1.00 1.00 0.99 0.99 1.00
+GPT-4o-mini+Shape 0.99 0.99 0.99 1.00 1.00
+GPT-4o-mini+PSI 1.00 1.00 0.99 0.99 1.00
+GPT-4o+Persona 1.00 0.99 0.99 0.99 1.00
+GPT-4o+Shape 1.00 1.00 0.99 0.99 1.00
+GPT-4o+PSI 1.00 1.00 0.99 0.99 1.00
+Table 28: Factor loading MAEs for BFI-2 TFM of Extraversion.
+Model Sociability Assertiveness Energy Level
+Mistral-7B+Persona 0.09 0.08 0.04
+Mistral-7B+Shape 0.03 0.07 0.07
+Mistral-7B+PSI 0.02 0.07 0.02
+Gemma-2-9B+Persona 0.01 0.05 0.02
+Gemma-2-9B+Shape 0.04 0.10 0.07
+Gemma-2-9B+PSI 0.01 0.05 0.09
+Gemma-2-27B+Persona 0.02 0.06 0.05
+Gemma-2-27B+Shape 0.02 0.06 0.08
+Gemma-2-27B+PSI 0.02 0.09 0.04
+Llama3-8B+Persona 0.41 0.16 0.14
+Llama3-8B+Shape 0.40 0.11 0.33
+Llama3-8B+PSI 0.02 0.04 0.04
+Llama3-70B+Persona 0.03 0.05 0.05
+Llama3-70B+Shape 0.05 0.05 0.06
+Llama3-70B+PSI 0.03 0.09 0.03
+GPT-4o-mini+Persona 0.01 0.04 0.05
+GPT-4o-mini+Shape 0.09 0.07 0.03
+GPT-4o-mini+PSI 0.02 0.03 0.05
+GPT-4o+Persona 0.01 0.06 0.05
+GPT-4o+Shape 0.02 0.05 0.06
+GPT-4o+PSI 0.02 0.04 0.06
+Table 29: Factor loading MAEs for BFI-2 TFM of Agreeableness.
+Model Compassion Respectfulness Trust
+Mistral-7B+Persona 0.06 0.04 0.07
+Mistral-7B+Shape 0.08 0.06 0.07
+Mistral-7B+PSI 0.03 0.06 0.04
+Gemma-2-9B+Persona 0.06 0.02 0.05
+Gemma-2-9B+Shape 0.10 0.08 0.07
+Gemma-2-9B+PSI 0.05 0.09 0.42
+Gemma-2-27B+Persona 0.05 0.04 0.01
+Gemma-2-27B+Shape 0.10 0.08 0.06
+Gemma-2-27B+PSI 0.05 0.05 0.03
+Llama3-8B+Persona 0.09 0.12 0.35
+Llama3-8B+Shape 0.04 0.07 0.07
+Llama3-8B+PSI 0.04 0.07 0.05
+Llama3-70B+Persona 0.06 0.03 0.03
+Llama3-70B+Shape 0.09 0.07 0.06
+Llama3-70B+PSI 0.04 0.05 0.03
+GPT-4o-mini+Persona 0.06 0.04 0.04
+GPT-4o-mini+Shape 0.09 0.06 0.08
+GPT-4o-mini+PSI 0.07 0.05 0.03
+GPT-4o+Persona 0.06 0.02 0.05
+GPT-4o+Shape 0.10 0.07 0.07
+GPT-4o+PSI 0.07 0.05 0.04
+Table 30: Factor loading MAEs for BFI-2 TFM of Conscientiousness.
+Model Organization Productiveness Responsibility
+Mistral-7B+Persona 0.08 0.04 0.08
+Mistral-7B+Shape 0.08 0.04 0.06
+Mistral-7B+PSI 0.07 0.02 0.06
+Gemma-2-9B+Persona 0.06 0.06 0.05
+Gemma-2-9B+Shape 0.03 0.06 0.04
+Gemma-2-9B+PSI 0.15 0.04 0.08
+Gemma-2-27B+Persona 0.03 0.03 0.05
+Gemma-2-27B+Shape 0.05 0.06 0.06
+Gemma-2-27B+PSI 0.06 0.03 0.06
+Llama3-8B+Persona 0.35 0.19 0.19
+Llama3-8B+Shape 0.13 0.11 0.11
+Llama3-8B+PSI 0.11 0.03 0.06
+Llama3-70B+Persona 0.03 0.04 0.06
+Llama3-70B+Shape 0.07 0.04 0.07
+Llama3-70B+PSI 0.02 0.04 0.04
+GPT-4o-mini+Persona 0.02 0.03 0.04
+GPT-4o-mini+Shape 0.09 0.04 0.06
+GPT-4o-mini+PSI 0.02 0.03 0.03
+GPT-4o+Persona 0.05 0.02 0.03
+GPT-4o+Shape 0.03 0.04 0.06
+GPT-4o+PSI 0.02 0.02 0.04
+Table 31: Factor loading MAEs for BFI-2 TFM of Neuroticism.
+Model Anxiety Depression Emotional Volatility
+Mistral-7B+Persona 0.10 0.08 0.12
+Mistral-7B+Shape 0.11 0.01 0.09
+Mistral-7B+PSI 0.04 0.03 0.02
+Gemma-2-9B+Persona 0.02 0.01 0.04
+Gemma-2-9B+Shape 0.08 0.03 0.09
+Gemma-2-9B+PSI 0.02 0.01 0.05
+Gemma-2-27B+Persona 0.06 0.01 0.02
+Gemma-2-27B+Shape 0.10 0.05 0.06
+Gemma-2-27B+PSI 0.05 0.03 0.03
+Llama3-8B+Persona 0.23 0.23 0.34
+Llama3-8B+Shape 0.28 0.17 0.23
+Llama3-8B+PSI 0.02 0.02 0.04
+Llama3-70B+Persona 0.03 0.03 0.01
+Llama3-70B+Shape 0.03 0.03 0.03
+Llama3-70B+PSI 0.02 0.02 0.02
+GPT-4o-mini+Persona 0.03 0.04 0.03
+GPT-4o-mini+Shape 0.11 0.02 0.07
+GPT-4o-mini+PSI 0.02 0.02 0.01
+GPT-4o+Persona 0.06 0.07 0.09
+GPT-4o+Shape 0.04 0.03 0.04
+GPT-4o+PSI 0.02 0.02 0.01
+Table 32: Factor loading MAEs for BFI-2 TFM of Openness.
+Model Intellectual Curiosity Aesthetic Sensitivity Creative Imagination
+Mistral-7B+Persona 0.02 0.07 0.07
+Mistral-7B+Shape 0.05 0.02 0.05
+Mistral-7B+PSI 0.10 0.07 0.11
+Gemma-2-9B+Persona 0.06 0.06 0.09
+Gemma-2-9B+Shape 0.06 0.04 0.04
+Gemma-2-9B+PSI 0.05 0.36 0.13
+Gemma-2-27B+Persona 0.03 0.04 0.02
+Gemma-2-27B+Shape 0.09 0.04 0.06
+Gemma-2-27B+PSI 0.04 0.05 0.03
+Llama3-8B+Persona 0.15 0.37 0.19
+Llama3-8B+Shape 0.14 0.07 0.14
+Llama3-8B+PSI 0.03 0.04 0.04
+Llama3-70B+Persona 0.04 0.04 0.05
+Llama3-70B+Shape 0.07 0.05 0.02
+Llama3-70B+PSI 0.04 0.05 0.06
+GPT-4o-mini+Persona 0.02 0.01 0.03
+GPT-4o-mini+Shape 0.07 0.09 0.10
+GPT-4o-mini+PSI 0.05 0.03 0.02
+GPT-4o+Persona 0.02 0.02 0.02
+GPT-4o+Shape 0.05 0.04 0.02
+GPT-4o+PSI 0.06 0.02 0.03
+Table 33: Factor loading MAEs for BFI-2 FFM.
+Model Extraversion Agreeableness Conscientiousness Neuroticism Openness
+Mistral-7B+Persona 0.02 0.01 0.01 0.03 0.01
+Mistral-7B+Shape 0.04 0.03 0.03 0.03 0.03
+Mistral-7B+PSI 0.03 0.03 0.01 0.02 0.01
+Gemma-2-9B+Persona 0.02 0.03 0.03 0.02 0.02
+Gemma-2-9B+Shape 0.05 0.03 0.02 0.03 0.03
+Gemma-2-9B+PSI 0.02 0.02 0.02 0.02 0.04
+Gemma-2-27B+Persona 0.03 0.04 0.03 0.01 0.02
+Gemma-2-27B+Shape 0.04 0.03 0.02 0.03 0.04
+Gemma-2-27B+PSI 0.03 0.02 0.02 0.01 0.02
+Llama3-8B+Persona 0.02 0.03 0.05 0.03 0.03
+Llama3-8B+Shape 0.03 0.04 0.03 0.04 0.02
+Llama3-8B+PSI 0.03 0.02 0.02 0.01 0.02
+Llama3-70B+Persona 0.04 0.04 0.03 0.02 0.02
+Llama3-70B+Shape 0.03 0.04 0.03 0.01 0.03
+Llama3-70B+PSI 0.03 0.03 0.02 0.01 0.02
+GPT-4o-mini+Persona 0.03 0.04 0.03 0.02 0.01
+GPT-4o-mini+Shape 0.02 0.04 0.03 0.02 0.04
+GPT-4o-mini+PSI 0.03 0.03 0.02 0.02 0.01
+GPT-4o+Persona 0.02 0.02 0.02 0.03 0.01
+GPT-4o+Shape 0.04 0.03 0.04 0.03 0.02
+GPT-4o+PSI 0.03 0.02 0.02 0.02 0.01
+Table 34: Standardized factor loadings for Sociability items in BFI-2 TFM of Extraversion.
+Model Item 1 Item 16 Item 31 Item 46
+Mistral-7B+Persona 0.88 0.46 0.22 0.77
+Mistral-7B+Shape 0.64 0.87 0.65 0.78
+Mistral-7B+PSI 0.77 0.73 0.72 0.81
+Gemma-2-9B+Persona 0.75 0.81 0.79 0.71
+Gemma-2-9B+Shape 0.71 0.90 0.80 0.41
+Gemma-2-9B+PSI 0.79 0.83 0.71 0.77
+Gemma-2-27B+Persona 0.75 0.80 0.70 0.76
+Gemma-2-27B+Shape 0.84 0.87 0.77 0.80
+Gemma-2-27B+PSI 0.70 0.91 0.84 0.80
+Llama3-8B+Persona -0.15 -0.78 -0.75 0.01
+Llama3-8B+Shape -0.71 -0.18 -0.08 -0.67
+Llama3-8B+PSI 0.82 0.71 0.77 0.78
+Llama3-70B+Persona 0.72 0.88 0.87 0.64
+Llama3-70B+Shape 0.55 0.73 0.79 0.58
+Llama3-70B+PSI 0.65 0.86 0.74 0.69
+GPT-4o-mini+Persona 0.79 0.78 0.72 0.74
+GPT-4o-mini+Shape 0.88 0.33 0.37 0.86
+GPT-4o-mini+PSI 0.85 0.88 0.88 0.82
+GPT-4o+Persona 0.77 0.79 0.79 0.74
+GPT-4o+Shape 0.85 0.74 0.73 0.85
+GPT-4o+PSI 0.87 0.84 0.90 0.76
+Human 0.80 0.82 0.82 0.76
+Table 35: Standardized factor loadings for Assertiveness items in BFI-2 TFM of Extraversion.
+Model Item 6 Item 21 Item 36 Item 51
+Mistral-7B+Persona 0.71 0.33 0.59 0.35
+Mistral-7B+Shape 0.76 0.54 0.86 0.80
+Mistral-7B+PSI 0.85 0.63 0.34 0.35
+Gemma-2-9B+Persona 0.90 0.84 0.33 0.44
+Gemma-2-9B+Shape 0.70 0.20 0.81 0.81
+Gemma-2-9B+PSI 0.71 0.86 0.37 0.35
+Gemma-2-27B+Persona 0.83 0.67 0.76 0.52
+Gemma-2-27B+Shape 0.99 0.84 0.80 0.52
+Gemma-2-27B+PSI 0.58 0.41 0.74 0.45
+Llama3-8B+Persona 0.78 0.74 -0.32 -0.23
+Llama3-8B+Shape 0.79 0.83 -0.28 0.31
+Llama3-8B+PSI 0.80 0.78 0.50 0.43
+Llama3-70B+Persona 0.84 0.73 0.72 0.54
+Llama3-70B+Shape 0.86 0.70 0.88 0.72
+Llama3-70B+PSI 0.49 0.53 0.59 0.28
+GPT-4o-mini+Persona 0.87 0.80 0.74 0.65
+GPT-4o-mini+Shape 0.99 0.86 0.28 0.37
+GPT-4o-mini+PSI 0.78 0.76 0.72 0.63
+GPT-4o+Persona 0.66 0.61 0.73 0.59
+GPT-4o+Shape 0.94 0.90 0.84 0.80
+GPT-4o+PSI 0.76 0.79 0.78 0.56
+Human 0.76 0.89 0.53 0.69
+Table 36: Standardized factor loadings for Energy Level items in BFI-2 TFM of Extraversion.
+Model Item 11 Item 26 Item 41 Item 56
+Mistral-7B+Persona 0.60 0.74 0.79 0.64
+Mistral-7B+Shape 0.87 0.88 0.73 0.73
+Mistral-7B+PSI 0.31 0.54 0.87 0.79
+Gemma-2-9B+Persona 0.47 0.64 0.88 0.77
+Gemma-2-9B+Shape 0.84 0.92 0.84 0.86
+Gemma-2-9B+PSI 0.06 -0.00 0.88 0.82
+Gemma-2-27B+Persona 0.72 0.83 0.86 0.79
+Gemma-2-27B+Shape 0.91 0.84 0.90 0.90
+Gemma-2-27B+PSI 0.80 0.63 0.85 0.75
+Llama3-8B+Persona -0.13 -0.14 0.57 0.65
+Llama3-8B+Shape -0.04 0.14 -0.69 -0.76
+Llama3-8B+PSI 0.77 0.62 0.83 0.80
+Llama3-70B+Persona 0.73 0.75 0.87 0.78
+Llama3-70B+Shape 0.88 0.77 0.76 0.77
+Llama3-70B+PSI 0.64 0.50 0.72 0.84
+GPT-4o-mini+Persona 0.70 0.78 0.88 0.88
+GPT-4o-mini+Shape 0.55 0.48 0.89 0.87
+GPT-4o-mini+PSI 0.76 0.75 0.89 0.82
+GPT-4o+Persona 0.82 0.69 0.81 0.82
+GPT-4o+Shape 0.91 0.72 0.82 0.88
+GPT-4o+PSI 0.82 0.75 0.88 0.82
+Human 0.45 0.54 0.83 0.77
+Table 37: Standardized factor loadings for Compassion items in BFI-2 TFM of Agreeableness.
+Model Item 2 Item 17 Item 32 Item 47
+Mistral-7B+Persona 0.68 0.11 0.76 0.48
+Mistral-7B+Shape 0.46 0.82 0.69 0.89
+Mistral-7B+PSI 0.78 0.44 0.86 0.73
+Gemma-2-9B+Persona 0.80 0.76 0.80 0.77
+Gemma-2-9B+Shape 0.86 0.92 0.93 0.96
+Gemma-2-9B+PSI 0.83 0.55 0.87 0.80
+Gemma-2-27B+Persona 0.70 0.68 0.86 0.79
+Gemma-2-27B+Shape 0.85 0.90 0.93 0.96
+Gemma-2-27B+PSI 0.74 0.81 0.73 0.81
+Llama3-8B+Persona 0.71 0.04 0.88 0.23
+Llama3-8B+Shape 0.77 0.25 0.71 0.43
+Llama3-8B+PSI 0.80 0.46 0.75 0.62
+Llama3-70B+Persona 0.75 0.86 0.74 0.85
+Llama3-70B+Shape 0.74 0.97 0.86 0.98
+Llama3-70B+PSI 0.84 0.48 0.79 0.86
+GPT-4o-mini+Persona 0.82 0.70 0.84 0.76
+GPT-4o-mini+Shape 0.52 0.96 0.69 0.98
+GPT-4o-mini+PSI 0.84 0.79 0.81 0.89
+GPT-4o+Persona 0.76 0.84 0.74 0.67
+GPT-4o+Shape 0.90 0.93 0.87 0.96
+GPT-4o+PSI 0.83 0.77 0.79 0.89
+Human 0.73 0.34 0.62 0.75
+Table 38: Standardized factor loadings for Respectfulness items in BFI-2 TFM of Agreeableness.
+Model Item 7 Item 22 Item 37 Item 52
+Mistral-7B+Persona 0.77 0.53 0.54 0.88
+Mistral-7B+Shape 0.60 0.77 0.88 0.48
+Mistral-7B+PSI 0.92 0.36 0.70 0.93
+Gemma-2-9B+Persona 0.77 0.60 0.63 0.86
+Gemma-2-9B+Shape 0.98 0.83 0.91 0.97
+Gemma-2-9B+PSI 0.90 0.10 0.50 0.89
+Gemma-2-27B+Persona 0.82 0.50 0.64 0.87
+Gemma-2-27B+Shape 0.96 0.83 0.87 0.99
+Gemma-2-27B+PSI 0.93 0.42 0.70 0.89
+Llama3-8B+Persona 0.73 0.12 -0.13 0.77
+Llama3-8B+Shape 0.94 0.50 0.28 0.75
+Llama3-8B+PSI 0.90 0.26 0.51 0.82
+Llama3-70B+Persona 0.80 0.68 0.79 0.80
+Llama3-70B+Shape 0.85 0.87 0.94 0.84
+Llama3-70B+PSI 0.92 0.48 0.84 0.90
+GPT-4o-mini+Persona 0.86 0.45 0.67 0.85
+GPT-4o-mini+Shape 0.64 0.84 0.97 0.62
+GPT-4o-mini+PSI 0.85 0.35 0.76 0.89
+GPT-4o+Persona 0.82 0.54 0.74 0.70
+GPT-4o+Shape 0.93 0.80 0.94 0.88
+GPT-4o+PSI 0.91 0.44 0.78 0.87
+Human 0.71 0.60 0.70 0.70
+Table 39: Standardized factor loadings for Trust items in BFI-2 TFM of Agreeableness.
+Model Item 12 Item 27 Item 42 Item 57
+Mistral-7B+Persona 0.34 0.75 0.21 0.78
+Mistral-7B+Shape 0.94 0.50 0.90 0.58
+Mistral-7B+PSI 0.40 0.78 0.67 0.84
+Gemma-2-9B+Persona 0.35 0.62 0.56 0.88
+Gemma-2-9B+Shape 0.95 0.87 0.95 0.90
+Gemma-2-9B+PSI -0.01 -0.81 -0.51 -0.85
+Gemma-2-27B+Persona 0.66 0.70 0.68 0.88
+Gemma-2-27B+Shape 0.96 0.80 0.96 0.86
+Gemma-2-27B+PSI 0.80 0.71 0.81 0.88
+Llama3-8B+Persona 0.18 -0.79 0.12 -0.83
+Llama3-8B+Shape 0.46 0.80 0.27 0.85
+Llama3-8B+PSI 0.38 0.79 0.56 0.86
+Llama3-70B+Persona 0.57 0.80 0.63 0.82
+Llama3-70B+Shape 0.94 0.84 0.95 0.85
+Llama3-70B+PSI 0.66 0.85 0.80 0.87
+GPT-4o-mini+Persona 0.49 0.85 0.66 0.90
+GPT-4o-mini+Shape 0.94 0.61 0.94 0.44
+GPT-4o-mini+PSI 0.69 0.80 0.81 0.91
+GPT-4o+Persona 0.63 0.60 0.35 0.64
+GPT-4o+Shape 0.96 0.84 0.95 0.90
+GPT-4o+PSI 0.76 0.80 0.80 0.89
+Human 0.69 0.68 0.67 0.78
+Table 40: Standardized factor loadings for Organization items in BFI-2 TFM of Conscientiousness.
+Model Item 3 Item 18 Item 33 Item 48
+Mistral-7B+Persona 0.47 0.59 0.67 0.41
+Mistral-7B+Shape 0.81 0.41 0.36 0.87
+Mistral-7B+PSI 0.73 0.58 0.42 0.60
+Gemma-2-9B+Persona 0.71 0.83 0.69 0.50
+Gemma-2-9B+Shape 0.77 0.65 0.66 0.78
+Gemma-2-9B+PSI 0.25 0.82 0.39 0.13
+Gemma-2-27B+Persona 0.63 0.70 0.73 0.69
+Gemma-2-27B+Shape 0.93 0.54 0.64 0.88
+Gemma-2-27B+PSI 0.80 0.60 0.63 0.42
+Llama3-8B+Persona 0.41 -0.72 -0.78 0.06
+Llama3-8B+Shape 0.11 0.89 0.81 0.22
+Llama3-8B+PSI 0.33 0.89 0.73 0.26
+Llama3-70B+Persona 0.75 0.66 0.68 0.74
+Llama3-70B+Shape 0.49 0.90 0.95 0.56
+Llama3-70B+PSI 0.83 0.65 0.75 0.76
+GPT-4o-mini+Persona 0.80 0.72 0.76 0.76
+GPT-4o-mini+Shape 0.85 0.27 0.32 0.85
+GPT-4o-mini+PSI 0.84 0.79 0.75 0.73
+GPT-4o+Persona 0.64 0.62 0.63 0.70
+GPT-4o+Shape 0.75 0.73 0.73 0.88
+GPT-4o+PSI 0.84 0.84 0.75 0.72
+Human 0.85 0.70 0.86 0.74
+Table 41: Standardized factor loadings for Productiveness items in BFI-2 TFM of Conscientiousness.
+Model Item 8 Item 23 Item 38 Item 53
+Mistral-7B+Persona 0.54 0.59 0.82 0.71
+Mistral-7B+Shape 0.96 0.89 0.60 0.73
+Mistral-7B+PSI 0.74 0.60 0.70 0.74
+Gemma-2-9B+Persona 0.68 0.36 0.85 0.86
+Gemma-2-9B+Shape 0.84 0.95 0.90 0.91
+Gemma-2-9B+PSI 0.47 0.57 0.80 0.72
+Gemma-2-27B+Persona 0.70 0.67 0.86 0.81
+Gemma-2-27B+Shape 0.82 0.92 0.91 0.92
+Gemma-2-27B+PSI 0.64 0.59 0.75 0.60
+Llama3-8B+Persona -0.13 -0.42 0.71 0.52
+Llama3-8B+Shape 0.16 0.13 0.78 0.73
+Llama3-8B+PSI 0.65 0.62 0.79 0.73
+Llama3-70B+Persona 0.68 0.65 0.88 0.82
+Llama3-70B+Shape 0.67 0.72 0.83 0.95
+Llama3-70B+PSI 0.79 0.43 0.74 0.78
+GPT-4o-mini+Persona 0.75 0.67 0.89 0.84
+GPT-4o-mini+Shape 0.89 0.81 0.48 0.64
+GPT-4o-mini+PSI 0.82 0.68 0.86 0.75
+GPT-4o+Persona 0.72 0.60 0.71 0.81
+GPT-4o+Shape 0.80 0.73 0.86 0.95
+GPT-4o+PSI 0.79 0.79 0.87 0.71
+Human 0.72 0.75 0.73 0.70
+Table 42: Standardized factor loadings for Responsibility items in BFI-2 TFM of Conscientiousness.
+Model Item 13 Item 28 Item 43 Item 58
+Mistral-7B+Persona 0.93 0.46 0.97 0.51
+Mistral-7B+Shape 0.50 0.90 0.60 0.88
+Mistral-7B+PSI 0.93 0.60 0.93 0.61
+Gemma-2-9B+Persona 0.88 0.58 0.88 0.60
+Gemma-2-9B+Shape 0.83 0.72 0.77 0.92
+Gemma-2-9B+PSI 0.95 0.35 0.94 0.63
+Gemma-2-27B+Persona 0.91 0.66 0.93 0.68
+Gemma-2-27B+Shape 0.94 0.74 0.92 0.85
+Gemma-2-27B+PSI 0.92 0.51 0.93 0.65
+Llama3-8B+Persona 0.90 -0.26 0.95 -0.20
+Llama3-8B+Shape 0.91 0.16 0.90 0.31
+Llama3-8B+PSI 0.88 0.52 0.89 0.50
+Llama3-70B+Persona 0.94 0.58 0.95 0.66
+Llama3-70B+Shape 0.92 0.48 0.96 0.64
+Llama3-70B+PSI 0.81 0.74 0.82 0.84
+GPT-4o-mini+Persona 0.89 0.72 0.92 0.74
+GPT-4o-mini+Shape 0.46 0.86 0.52 0.85
+GPT-4o-mini+PSI 0.81 0.77 0.79 0.78
+GPT-4o+Persona 0.66 0.58 0.64 0.47
+GPT-4o+Shape 0.92 0.68 0.94 0.80
+GPT-4o+PSI 0.83 0.78 0.84 0.78
+Human 0.68 0.73 0.70 0.65
+Table 43: Standardized factor loadings for Anxiety items in BFI-2 TFM of Neuroticism.
+Model Item 4 Item 19 Item 34 Item 49
+Mistral-7B+Persona 0.25 0.55 0.71 0.25
+Mistral-7B+Shape 0.35 0.96 0.79 0.14
+Mistral-7B+PSI 0.69 0.59 0.63 0.75
+Gemma-2-9B+Persona 0.76 0.72 0.82 0.80
+Gemma-2-9B+Shape 0.58 0.83 0.88 0.18
+Gemma-2-9B+PSI 0.69 0.71 0.74 0.73
+Gemma-2-27B+Persona 0.75 0.85 0.69 0.35
+Gemma-2-27B+Shape 0.66 0.91 0.89 -0.08
+Gemma-2-27B+PSI 0.68 0.78 0.63 0.46
+Llama3-8B+Persona 0.56 -0.48 -0.31 0.53
+Llama3-8B+Shape 0.51 -0.72 -0.71 0.52
+Llama3-8B+PSI 0.71 0.78 0.77 0.64
+Llama3-70B+Persona 0.72 0.61 0.81 0.74
+Llama3-70B+Shape 0.63 0.72 0.80 0.71
+Llama3-70B+PSI 0.79 0.64 0.76 0.59
+GPT-4o-mini+Persona 0.74 0.58 0.67 0.64
+GPT-4o-mini+Shape 0.27 0.79 0.92 0.15
+GPT-4o-mini+PSI 0.84 0.86 0.84 0.70
+GPT-4o+Persona 0.52 0.65 0.59 0.51
+GPT-4o+Shape 0.55 0.80 0.78 0.75
+GPT-4o+PSI 0.85 0.87 0.80 0.68
+Human 0.85 0.70 0.79 0.66
+Table 44: Standardized factor loadings for Depression items in BFI-2 TFM of Neuroticism.
+Model Item 9 Item 24 Item 39 Item 54
+Mistral-7B+Persona 0.19 0.27 0.79 0.91
+Mistral-7B+Shape 0.68 0.69 0.83 0.92
+Mistral-7B+PSI 0.64 0.79 0.76 0.82
+Gemma-2-9B+Persona 0.52 0.66 0.88 0.90
+Gemma-2-9B+Shape 0.74 0.80 0.85 0.95
+Gemma-2-9B+PSI 0.57 0.63 0.89 0.95
+Gemma-2-27B+Persona 0.67 0.63 0.88 0.86
+Gemma-2-27B+Shape 0.88 0.91 0.93 0.98
+Gemma-2-27B+PSI 0.42 0.63 0.92 0.95
+Llama3-8B+Persona 0.57 0.58 -0.50 -0.28
+Llama3-8B+Shape 0.67 0.93 -0.09 0.15
+Llama3-8B+PSI 0.56 0.67 0.84 0.86
+Llama3-70B+Persona 0.68 0.86 0.81 0.88
+Llama3-70B+Shape 0.78 0.86 0.87 0.95
+Llama3-70B+PSI 0.61 0.80 0.89 0.94
+GPT-4o-mini+Persona 0.71 0.82 0.76 0.80
+GPT-4o-mini+Shape 0.50 0.54 0.90 0.90
+GPT-4o-mini+PSI 0.63 0.81 0.92 0.95
+GPT-4o+Persona 0.27 0.72 0.70 0.67
+GPT-4o+Shape 0.75 0.83 0.84 0.89
+GPT-4o+PSI 0.69 0.83 0.93 0.95
+Human 0.64 0.67 0.90 0.91
+Table 45: Standardized factor loadings for Emotional Volatility items in BFI-2 TFM of Neuroticism.
+Model Item 14 Item 29 Item 44 Item 59
+Mistral-7B+Persona 0.64 0.30 0.23 0.90
+Mistral-7B+Shape 0.92 0.62 0.25 0.93
+Mistral-7B+PSI 0.77 0.87 0.63 0.69
+Gemma-2-9B+Persona 0.72 0.90 0.62 0.62
+Gemma-2-9B+Shape 0.96 0.57 0.20 0.89
+Gemma-2-9B+PSI 0.72 0.79 0.69 0.40
+Gemma-2-27B+Persona 0.76 0.84 0.74 0.63
+Gemma-2-27B+Shape 0.99 0.71 0.54 0.91
+Gemma-2-27B+PSI 0.88 0.81 0.71 0.63
+Llama3-8B+Persona 0.23 -0.75 -0.70 0.34
+Llama3-8B+Shape 0.90 -0.41 -0.44 0.90
+Llama3-8B+PSI 0.63 0.87 0.57 0.67
+Llama3-70B+Persona 0.86 0.86 0.73 0.80
+Llama3-70B+Shape 0.88 0.85 0.60 0.84
+Llama3-70B+PSI 0.86 0.91 0.83 0.74
+GPT-4o-mini+Persona 0.77 0.90 0.58 0.66
+GPT-4o-mini+Shape 0.85 0.57 0.38 0.74
+GPT-4o-mini+PSI 0.80 0.93 0.77 0.69
+GPT-4o+Persona 0.71 0.44 0.40 0.59
+GPT-4o+Shape 0.92 0.80 0.58 0.75
+GPT-4o+PSI 0.78 0.92 0.75 0.72
+Human 0.80 0.88 0.77 0.80
+Table 46: Standardized factor loadings for Intellectual Curiosity items in BFI-2 TFM of Openness.
+Model Item 10 Item 25 Item 40 Item 55
+Mistral-7B+Persona 0.60 0.63 0.54 0.65
+Mistral-7B+Shape 0.81 0.62 0.39 0.58
+Mistral-7B+PSI 0.78 0.19 0.68 0.14
+Gemma-2-9B+Persona 0.69 0.42 0.62 0.40
+Gemma-2-9B+Shape 0.90 0.90 0.67 0.89
+Gemma-2-9B+PSI 0.80 0.55 0.65 0.44
+Gemma-2-27B+Persona 0.65 0.80 0.48 0.77
+Gemma-2-27B+Shape 0.87 0.95 0.35 0.93
+Gemma-2-27B+PSI 0.81 0.67 0.73 0.54
+Llama3-8B+Persona 0.64 -0.10 0.55 -0.21
+Llama3-8B+Shape 0.93 -0.05 0.61 0.11
+Llama3-8B+PSI 0.72 0.77 0.73 0.79
+Llama3-70B+Persona 0.81 0.76 0.70 0.82
+Llama3-70B+Shape 0.91 0.88 0.56 0.89
+Llama3-70B+PSI 0.81 0.64 0.78 0.57
+GPT-4o-mini+Persona 0.79 0.63 0.68 0.68
+GPT-4o-mini+Shape 0.61 0.87 0.23 0.86
+GPT-4o-mini+PSI 0.80 0.77 0.81 0.86
+GPT-4o+Persona 0.71 0.68 0.78 0.75
+GPT-4o+Shape 0.90 0.87 0.62 0.82
+GPT-4o+PSI 0.76 0.84 0.81 0.89
+Human 0.60 0.63 0.65 0.74
+Table 47: Standardized factor loadings for Aesthetic Sensitivity items in BFI-2 TFM of Openness.
+Model Item 5 Item 20 Item 35 Item 50
+Mistral-7B+Persona 0.35 0.76 0.95 0.25
+Mistral-7B+Shape 0.65 0.85 0.73 0.73
+Mistral-7B+PSI 0.41 0.89 0.79 0.18
+Gemma-2-9B+Persona 0.24 0.87 0.91 0.34
+Gemma-2-9B+Shape 0.85 0.86 0.88 0.89
+Gemma-2-9B+PSI 0.29 -1.01 -0.63 0.10
+Gemma-2-27B+Persona 0.85 0.83 0.89 0.44
+Gemma-2-27B+Shape 0.85 0.83 0.85 0.93
+Gemma-2-27B+PSI 0.88 0.88 0.95 0.40
+Llama3-8B+Persona 0.17 -0.79 -0.86 0.07
+Llama3-8B+Shape 0.19 0.86 0.80 0.38
+Llama3-8B+PSI 0.82 0.88 0.91 0.41
+Llama3-70B+Persona 0.85 0.85 0.88 0.41
+Llama3-70B+Shape 0.91 0.83 0.82 0.91
+Llama3-70B+PSI 0.87 0.84 0.94 0.33
+GPT-4o-mini+Persona 0.79 0.87 0.88 0.64
+GPT-4o-mini+Shape 0.91 0.56 0.60 0.96
+GPT-4o-mini+PSI 0.90 0.85 0.89 0.51
+GPT-4o+Persona 0.84 0.82 0.86 0.69
+GPT-4o+Shape 0.81 0.91 0.85 0.89
+GPT-4o+PSI 0.87 0.85 0.90 0.64
+Human 0.66 0.86 0.87 0.65
+Table 48: Standardized factor loadings for Creative Imagination items in BFI-2 TFM of Openness.
+Model Item 15 Item 30 Item 45 Item 60
+Mistral-7B+Persona 0.76 0.54 0.18 0.72
+Mistral-7B+Shape 0.91 0.55 0.57 0.84
+Mistral-7B+PSI 0.85 0.34 0.06 0.91
+Gemma-2-9B+Persona 0.74 0.11 0.34 0.78
+Gemma-2-9B+Shape 0.82 0.93 0.86 0.78
+Gemma-2-9B+PSI 0.78 -0.08 0.04 0.81
+Gemma-2-27B+Persona 0.80 0.70 0.55 0.80
+Gemma-2-27B+Shape 0.87 0.97 0.95 0.75
+Gemma-2-27B+PSI 0.60 0.74 0.55 0.76
+Llama3-8B+Persona 0.71 -0.40 -0.34 0.76
+Llama3-8B+Shape 0.93 -0.04 0.05 0.80
+Llama3-8B+PSI 0.62 0.78 0.80 0.59
+Llama3-70B+Persona 0.89 0.63 0.49 0.95
+Llama3-70B+Shape 0.86 0.81 0.76 0.81
+Llama3-70B+PSI 0.85 0.60 0.37 0.95
+GPT-4o-mini+Persona 0.80 0.74 0.50 0.90
+GPT-4o-mini+Shape 0.47 0.97 0.87 0.32
+GPT-4o-mini+PSI 0.70 0.86 0.78 0.81
+GPT-4o+Persona 0.65 0.81 0.68 0.70
+GPT-4o+Shape 0.79 0.84 0.79 0.79
+GPT-4o+PSI 0.66 0.88 0.88 0.79
+Human 0.72 0.77 0.66 0.80
+Table 49: Standardized factor loadings for five‐factor model: Extraversion subdimension.
+Model Sociability Assertiveness Energy Level
+Mistral-7B+Persona 0.59 0.69 0.87
+Mistral-7B+Shape 0.74 0.94 0.91
+Mistral-7B+PSI 0.76 0.76 0.95
+Gemma-2-9B+Persona 0.66 0.83 0.81
+Gemma-2-9B+Shape 0.86 0.84 0.92
+Gemma-2-9B+PSI 0.65 0.81 0.84
+Gemma-2-27B+Persona 0.76 0.82 0.90
+Gemma-2-27B+Shape 0.88 0.83 0.95
+Gemma-2-27B+PSI 0.84 0.78 0.89
+Llama3-8B+Persona 0.63 0.66 0.65
+Llama3-8B+Shape 0.70 0.87 0.69
+Llama3-8B+PSI 0.76 0.74 0.94
+Llama3-70B+Persona 0.85 0.86 0.86
+Llama3-70B+Shape 0.87 0.74 0.85
+Llama3-70B+PSI 0.77 0.68 0.90
+GPT-4o-mini+Persona 0.82 0.78 0.93
+GPT-4o-mini+Shape 0.76 0.58 1.03
+GPT-4o-mini+PSI 0.82 0.69 0.94
+GPT-4o+Persona 0.76 0.67 0.88
+GPT-4o+Shape 0.91 0.77 0.89
+GPT-4o+PSI 0.82 0.71 0.93
+Human 0.66 0.59 0.77
+Table 50: Standardized factor loadings for five‐factor model: Agreeableness subdimension.
+Model Compassion Respectfulness Trust
+Mistral-7B+Persona 0.76 0.85 0.67
+Mistral-7B+Shape 0.93 0.83 0.90
+Mistral-7B+PSI 0.86 0.88 0.90
+Gemma-2-9B+Persona 0.86 0.92 0.81
+Gemma-2-9B+Shape 0.98 0.92 0.95
+Gemma-2-9B+PSI 0.85 0.82 0.81
+Gemma-2-27B+Persona 0.88 0.90 0.91
+Gemma-2-27B+Shape 1.00 0.87 0.95
+Gemma-2-27B+PSI 0.80 0.91 0.85
+Llama3-8B+Persona 0.80 0.86 0.76
+Llama3-8B+Shape 0.94 0.79 0.90
+Llama3-8B+PSI 0.83 0.79 0.84
+Llama3-70B+Persona 0.90 0.94 0.87
+Llama3-70B+Shape 0.99 0.86 0.95
+Llama3-70B+PSI 0.79 0.95 0.84
+GPT-4o-mini+Persona 0.89 0.94 0.88
+GPT-4o-mini+Shape 0.97 0.86 0.94
+GPT-4o-mini+PSI 0.85 0.89 0.86
+GPT-4o+Persona 0.82 0.90 0.75
+GPT-4o+Shape 0.96 0.90 0.93
+GPT-4o+PSI 0.83 0.88 0.84
+Human 0.70 0.80 0.65
+Table 51: Standardized factor loadings for five‐factor model: Conscientiousness subdimension.
+Model Organization Productiveness Responsibility
+Mistral-7B+Persona 0.75 0.85 0.85
+Mistral-7B+Shape 0.96 0.88 0.90
+Mistral-7B+PSI 0.78 0.85 0.86
+Gemma-2-9B+Persona 0.77 0.68 0.93
+Gemma-2-9B+Shape 0.92 0.76 0.95
+Gemma-2-9B+PSI 0.66 0.79 0.89
+Gemma-2-27B+Persona 0.86 0.80 0.92
+Gemma-2-27B+Shape 0.91 0.80 0.97
+Gemma-2-27B+PSI 0.85 0.84 0.86
+Llama3-8B+Persona 0.62 0.50 0.77
+Llama3-8B+Shape 0.83 0.75 0.92
+Llama3-8B+PSI 0.79 0.80 0.86
+Llama3-70B+Persona 0.87 0.75 0.95
+Llama3-70B+Shape 0.97 0.80 0.93
+Llama3-70B+PSI 0.87 0.83 0.92
+GPT-4o-mini+Persona 0.91 0.83 0.94
+GPT-4o-mini+Shape 0.96 0.86 0.90
+GPT-4o-mini+PSI 0.86 0.81 0.90
+GPT-4o+Persona 0.87 0.77 0.82
+GPT-4o+Shape 0.94 0.84 0.93
+GPT-4o+PSI 0.87 0.83 0.88
+Human 0.70 0.89 0.79
+Table 52: Standardized factor loadings for five‐factor model: Neuroticism subdimension.
+Model Anxiety Depression Emotional Volatility
+Mistral-7B+Persona 0.51 0.80 0.75
+Mistral-7B+Shape 0.67 1.00 0.76
+Mistral-7B+PSI 0.66 0.96 0.81
+Gemma-2-9B+Persona 0.76 0.96 0.70
+Gemma-2-9B+Shape 0.82 1.01 0.73
+Gemma-2-9B+PSI 0.72 0.96 0.78
+Gemma-2-27B+Persona 0.70 0.85 0.82
+Gemma-2-27B+Shape 0.79 1.01 0.72
+Gemma-2-27B+PSI 0.62 0.90 0.82
+Llama3-8B+Persona 0.44 0.91 0.60
+Llama3-8B+Shape 0.42 0.91 0.65
+Llama3-8B+PSI 0.80 0.87 0.81
+Llama3-70B+Persona 0.75 0.96 0.79
+Llama3-70B+Shape 0.87 0.96 0.81
+Llama3-70B+PSI 0.75 0.88 0.85
+GPT-4o-mini+Persona 0.73 0.97 0.74
+GPT-4o-mini+Shape 0.78 0.97 0.78
+GPT-4o-mini+PSI 0.74 0.97 0.78
+GPT-4o+Persona 0.69 0.91 0.63
+GPT-4o+Shape 0.80 1.02 0.71
+GPT-4o+PSI 0.74 0.99 0.76
+Human 0.84 0.88 0.85
+Table 53: Standardized factor loadings for five‐factor model: Openness subdimension.
+Model Intellectual Curiosity Aesthetic Sensitivity Creative Imagination
+Mistral-7B+Persona 0.74 0.64 0.70
+Mistral-7B+Shape 0.89 0.94 0.79
+Mistral-7B+PSI 0.74 0.53 0.77
+Gemma-2-9B+Persona 0.85 0.58 0.74
+Gemma-2-9B+Shape 0.96 0.94 0.87
+Gemma-2-9B+PSI 0.86 0.40 0.62
+Gemma-2-27B+Persona 0.82 0.81 0.71
+Gemma-2-27B+Shape 0.94 0.96 0.89
+Gemma-2-27B+PSI 0.74 0.67 0.91
+Llama3-8B+Persona 0.69 0.50 0.63
+Llama3-8B+Shape 0.78 0.96 0.78
+Llama3-8B+PSI 0.80 0.63 0.91
+Llama3-70B+Persona 0.66 0.65 0.89
+Llama3-70B+Shape 0.93 0.96 0.82
+Llama3-70B+PSI 0.85 0.58 0.78
+GPT-4o-mini+Persona 0.86 0.77 0.80
+GPT-4o-mini+Shape 0.94 0.96 0.93
+GPT-4o-mini+PSI 0.79 0.70 0.88
+GPT-4o+Persona 0.77 0.73 0.91
+GPT-4o+Shape 0.97 0.95 0.90
+GPT-4o+PSI 0.75 0.70 0.96
+Human 0.74 0.71 0.81
+Table 54: Inter‐factor correlations for BFI-2 TFM of Extraversion.
+Model Soc–Ass Soc–Ene Ass–Ene
+Mistral-7B+Persona 0.72 0.67 0.87
+Mistral-7B+Shape 0.81 0.80 1.01
+Mistral-7B+PSI 0.79 0.89 0.88
+Gemma-2-9B+Persona 0.73 0.66 0.64
+Gemma-2-9B+Shape 0.91 0.88 0.97
+Gemma-2-9B+PSI 0.72 0.88 0.81
+Gemma-2-27B+Persona 0.79 0.83 0.79
+Gemma-2-27B+Shape 0.71 0.92 0.68
+Gemma-2-27B+PSI 0.89 0.87 0.95
+Llama3-8B+Persona -0.15 -0.32 0.96
+Llama3-8B+Shape -0.79 1.15 -0.69
+Llama3-8B+PSI 0.73 0.85 0.84
+Llama3-70B+Persona 0.89 0.84 0.82
+Llama3-70B+Shape 0.91 0.82 0.61
+Llama3-70B+PSI 0.80 0.85 0.88
+GPT-4o-mini+Persona 0.81 0.90 0.78
+GPT-4o-mini+Shape 0.60 0.95 0.53
+GPT-4o-mini+PSI 0.73 0.87 0.73
+GPT-4o+Persona 0.72 0.79 0.66
+GPT-4o+Shape 0.77 0.91 0.65
+GPT-4o+PSI 0.71 0.86 0.74
+Human 0.59 0.64 0.47
+Table 55: Inter‐factor correlations for BFI-2 TFM of Agreeableness.
+Model Com–Res Com–Tru Res–Tru
+Mistral-7B+Persona 1.00 0.89 0.71
+Mistral-7B+Shape 0.93 0.96 0.91
+Mistral-7B+PSI 0.91 0.98 0.92
+Gemma-2-9B+Persona 0.93 0.88 0.94
+Gemma-2-9B+Shape 0.90 0.98 0.91
+Gemma-2-9B+PSI 0.93 -0.93 -0.90
+Gemma-2-27B+Persona 0.96 0.96 0.97
+Gemma-2-27B+Shape 0.88 1.00 0.84
+Gemma-2-27B+PSI 0.79 0.79 0.84
+Llama3-8B+Persona 1.05 -0.85 -0.97
+Llama3-8B+Shape 0.99 1.09 0.86
+Llama3-8B+PSI 0.87 0.88 0.83
+Llama3-70B+Persona 0.98 0.94 0.96
+Llama3-70B+Shape 0.87 0.96 0.85
+Llama3-70B+PSI 0.90 0.85 0.88
+GPT-4o-mini+Persona 0.99 0.91 0.94
+GPT-4o-mini+Shape 0.88 0.99 0.90
+GPT-4o-mini+PSI 0.90 0.82 0.86
+GPT-4o+Persona 0.90 0.83 0.95
+GPT-4o+Shape 0.92 0.94 0.93
+GPT-4o+PSI 0.86 0.80 0.85
+Human 0.81 0.70 0.58
+Table 56: Inter‐factor correlations for BFI-2 TFM of Conscientiousness.
+Model Org–Pro Org–Res Pro–Res
+Mistral-7B+Persona 0.93 0.85 0.87
+Mistral-7B+Shape 1.01 1.04 0.90
+Mistral-7B+PSI 0.96 0.83 0.81
+Gemma-2-9B+Persona 0.62 0.81 0.77
+Gemma-2-9B+Shape 0.89 1.00 0.66
+Gemma-2-9B+PSI 0.77 0.73 0.82
+Gemma-2-27B+Persona 0.84 0.90 0.83
+Gemma-2-27B+Shape 0.94 0.88 0.75
+Gemma-2-27B+PSI 0.97 0.85 0.84
+Llama3-8B+Persona -0.98 -0.75 0.99
+Llama3-8B+Shape 0.87 0.87 0.91
+Llama3-8B+PSI 0.76 0.76 0.74
+Llama3-70B+Persona 0.81 0.85 0.87
+Llama3-70B+Shape 0.69 0.86 0.87
+Llama3-70B+PSI 0.88 0.91 0.88
+GPT-4o-mini+Persona 0.88 0.93 0.90
+GPT-4o-mini+Shape 1.07 0.92 0.85
+GPT-4o-mini+PSI 0.85 0.90 0.83
+GPT-4o+Persona 0.86 1.01 0.82
+GPT-4o+Shape 0.97 0.99 0.87
+GPT-4o+PSI 0.86 0.86 0.80
+Human 0.75 0.70 0.86
+Table 57: Inter‐factor correlations for BFI-2 TFM of Neuroticism.
+Model Anx–Dep Anx–Emo Dep–Emo
+Mistral-7B+Persona 0.78 0.51 0.60
+Mistral-7B+Shape 0.93 0.82 0.91
+Mistral-7B+PSI 0.81 0.83 0.93
+Gemma-2-9B+Persona 0.84 0.79 0.68
+Gemma-2-9B+Shape 1.01 0.94 0.92
+Gemma-2-9B+PSI 0.77 0.85 0.87
+Gemma-2-27B+Persona 0.66 0.77 0.63
+Gemma-2-27B+Shape 0.97 0.89 0.91
+Gemma-2-27B+PSI 0.70 0.82 0.84
+Llama3-8B+Persona 1.13 -1.10 -1.04
+Llama3-8B+Shape 0.18 -0.90 0.23
+Llama3-8B+PSI 0.85 0.95 0.87
+Llama3-70B+Persona 0.88 0.79 0.85
+Llama3-70B+Shape 0.96 0.87 0.84
+Llama3-70B+PSI 0.77 0.83 0.83
+GPT-4o-mini+Persona 0.88 0.83 0.91
+GPT-4o-mini+Shape 0.90 0.78 0.93
+GPT-4o-mini+PSI 0.80 0.89 0.86
+GPT-4o+Persona 0.88 0.87 0.84
+GPT-4o+Shape 0.95 0.87 0.89
+GPT-4o+PSI 0.85 0.91 0.86
+Human 0.81 0.90 0.79
+Table 58: Inter‐factor correlations for BFI-2 TFM of Openness.
+Model Int–Aes Int–Cre Aes–Cre
+Mistral-7B+Persona 0.35 0.65 0.58
+Mistral-7B+Shape 1.09 0.83 0.77
+Mistral-7B+PSI 0.51 0.67 0.63
+Gemma-2-9B+Persona 0.61 0.96 0.64
+Gemma-2-9B+Shape 0.97 0.94 0.90
+Gemma-2-9B+PSI -0.47 0.89 -0.54
+Gemma-2-27B+Persona 0.65 0.66 0.67
+Gemma-2-27B+Shape 0.98 0.94 0.93
+Gemma-2-27B+PSI 0.44 0.87 0.71
+Llama3-8B+Persona -0.70 1.06 -0.64
+Llama3-8B+Shape 0.97 0.96 0.98
+Llama3-8B+PSI 0.54 0.85 0.74
+Llama3-70B+Persona 0.59 0.67 0.60
+Llama3-70B+Shape 1.00 0.89 0.85
+Llama3-70B+PSI 0.48 0.74 0.51
+GPT-4o-mini+Persona 0.81 0.82 0.64
+GPT-4o-mini+Shape 1.00 0.98 1.00
+GPT-4o-mini+PSI 0.57 0.79 0.75
+GPT-4o+Persona 0.68 0.85 0.79
+GPT-4o+Shape 1.02 0.99 0.93
+GPT-4o+PSI 0.54 0.80 0.78
+Human 0.66 0.74 0.63
+Model E~A E~C E~N E~O A~C A~N A~O C~N C~O N~O
+Mistral-7B+Persona 0.57 0.74 -0.68 0.42 0.86 -0.59 0.49 -0.78 0.33 -0.07
+Mistral-7B+Shape 0.58 0.67 -0.80 0.64 0.79 -0.78 0.75 -0.82 0.51 -0.61
+Mistral-7B+PSI 0.57 0.60 -0.64 0.64 0.73 -0.68 0.43 -0.78 0.19 -0.18
+Gemma-2-9B+Persona 0.25 0.45 -0.63 0.53 0.74 -0.60 0.42 -0.67 0.32 -0.30
+Gemma-2-9B+Shape 0.66 0.42 -0.77 0.83 0.66 -0.80 0.82 -0.73 0.43 -0.62
+Gemma-2-9B+PSI 0.48 0.64 -0.70 0.71 0.69 -0.64 0.52 -0.76 0.45 -0.34
+Gemma-2-27B+Persona 0.48 0.57 -0.67 0.53 0.79 -0.72 0.62 -0.84 0.41 -0.36
+Gemma-2-27B+Shape 0.72 0.47 -0.83 0.82 0.77 -0.82 0.85 -0.80 0.48 -0.71
+Gemma-2-27B+PSI 0.34 0.41 -0.43 0.47 0.67 -0.52 0.13 -0.76 0.08 0.17
+Llama3-8B+Persona 0.30 0.30 -0.64 0.42 0.62 -0.63 0.40 -0.59 0.03 -0.25
+Llama3-8B+Shape 0.40 0.27 -0.62 0.68 0.58 -0.69 0.76 -0.55 0.27 -0.62
+Llama3-8B+PSI 0.53 0.56 -0.46 0.36 0.63 -0.35 0.22 -0.55 0.16 0.16
+Llama3-70B+Persona 0.10 0.26 -0.59 0.51 0.77 -0.54 0.25 -0.63 0.16 -0.23
+Llama3-70B+Shape 0.54 0.31 -0.71 0.70 0.65 -0.74 0.89 -0.82 0.45 -0.65
+Llama3-70B+PSI 0.28 0.36 -0.33 0.51 0.64 -0.50 0.15 -0.63 0.20 0.12
+GPT-4o-mini+Persona 0.51 0.55 -0.76 0.52 0.79 -0.73 0.54 -0.69 0.36 -0.39
+GPT-4o-mini+Shape 0.81 0.59 -0.61 0.87 0.75 -0.79 0.91 -0.75 0.58 -0.62
+GPT-4o-mini+PSI 0.50 0.45 -0.62 0.33 0.65 -0.52 0.19 -0.64 0.07 0.10
+GPT-4o+Persona 0.26 0.43 -0.61 0.37 0.56 -0.40 0.31 -0.52 0.10 -0.06
+GPT-4o+Shape 0.59 0.36 -0.80 0.72 0.65 -0.66 0.79 -0.67 0.27 -0.48
+GPT-4o+PSI 0.44 0.45 -0.63 0.30 0.60 -0.47 0.15 -0.58 0.06 0.09
+Human 0.28 0.56 -0.61 0.35 0.47 -0.43 0.27 -0.61 0.17 -0.17
+Table 59: Inter-factor correlations for BFI-2 FFM. E = Extraversion; A = Agreeableness; C = Conscientiousness; N = Neuroticism; O = Openness.
+Table 60: Cronbach’s alpha for Extraversion facets.
+Model Sociability Assertiveness Energy Level
+Mistral-7B+Persona 0.68 0.55 0.78
+Mistral-7B+Shape 0.81 0.79 0.88
+Mistral-7B+PSI 0.84 0.61 0.72
+Gemma-2-9B+Persona 0.82 0.71 0.79
+Gemma-2-9B+Shape 0.79 0.74 0.92
+Gemma-2-9B+PSI 0.86 0.67 0.47
+Gemma-2-27B+Persona 0.83 0.77 0.88
+Gemma-2-27B+Shape 0.89 0.86 0.93
+Gemma-2-27B+PSI 0.88 0.62 0.83
+Llama3-8B+Persona 0.44 0.01 0.13
+Llama3-8B+Shape 0.58 0.49 0.32
+Llama3-8B+PSI 0.86 0.71 0.83
+Llama3-70B+Persona 0.86 0.80 0.86
+Llama3-70B+Shape 0.76 0.86 0.87
+Llama3-70B+PSI 0.82 0.55 0.77
+GPT-4o-mini+Persona 0.84 0.85 0.88
+GPT-4o-mini+Shape 0.74 0.77 0.84
+GPT-4o-mini+PSI 0.92 0.81 0.88
+GPT-4o+Persona 0.85 0.74 0.86
+GPT-4o+Shape 0.87 0.93 0.90
+GPT-4o+PSI 0.91 0.81 0.89
+Human 0.87 0.81 0.73
+Table 61: Cronbach’s alpha for Agreeableness facets.
+Model Compassion Respectfulness Trust
+Mistral-7B+Persona 0.56 0.78 0.61
+Mistral-7B+Shape 0.79 0.78 0.83
+Mistral-7B+PSI 0.80 0.81 0.77
+Gemma-2-9B+Persona 0.86 0.78 0.69
+Gemma-2-9B+Shape 0.95 0.96 0.96
+Gemma-2-9B+PSI 0.84 0.66 0.64
+Gemma-2-27B+Persona 0.84 0.79 0.83
+Gemma-2-27B+Shape 0.95 0.95 0.95
+Gemma-2-27B+PSI 0.85 0.83 0.87
+Llama3-8B+Persona 0.54 0.44 0.36
+Llama3-8B+Shape 0.68 0.74 0.69
+Llama3-8B+PSI 0.74 0.72 0.75
+Llama3-70B+Persona 0.87 0.85 0.79
+Llama3-70B+Shape 0.93 0.93 0.94
+Llama3-70B+PSI 0.83 0.87 0.87
+GPT-4o-mini+Persona 0.86 0.80 0.81
+GPT-4o-mini+Shape 0.88 0.87 0.85
+GPT-4o-mini+PSI 0.89 0.81 0.87
+GPT-4o+Persona 0.84 0.79 0.64
+GPT-4o+Shape 0.95 0.94 0.96
+GPT-4o+PSI 0.89 0.84 0.88
+Human 0.67 0.74 0.80
+Table 62: Cronbach’s alpha for Conscientiousness facets.
+Model Organization Productiveness Responsibility
+Mistral-7B+Persona 0.58 0.77 0.83
+Mistral-7B+Shape 0.68 0.88 0.84
+Mistral-7B+PSI 0.67 0.77 0.86
+Gemma-2-9B+Persona 0.75 0.78 0.80
+Gemma-2-9B+Shape 0.80 0.94 0.87
+Gemma-2-9B+PSI 0.38 0.72 0.82
+Gemma-2-27B+Persona 0.77 0.84 0.86
+Gemma-2-27B+Shape 0.84 0.94 0.91
+Gemma-2-27B+PSI 0.70 0.74 0.85
+Llama3-8B+Persona 0.04 0.03 0.31
+Llama3-8B+Shape 0.61 0.60 0.68
+Llama3-8B+PSI 0.64 0.79 0.79
+Llama3-70B+Persona 0.80 0.85 0.88
+Llama3-70B+Shape 0.83 0.89 0.87
+Llama3-70B+PSI 0.83 0.77 0.87
+GPT-4o-mini+Persona 0.84 0.87 0.89
+GPT-4o-mini+Shape 0.67 0.82 0.79
+GPT-4o-mini+PSI 0.85 0.85 0.86
+GPT-4o+Persona 0.73 0.80 0.67
+GPT-4o+Shape 0.86 0.91 0.91
+GPT-4o+PSI 0.86 0.87 0.88
+Human 0.87 0.80 0.77
+Table 63: Cronbach’s alpha for Neuroticism facets.
+Model Anxiety Depression Emotional Volatility
+Mistral-7B+Persona 0.54 0.64 0.61
+Mistral-7B+Shape 0.64 0.85 0.79
+Mistral-7B+PSI 0.75 0.84 0.83
+Gemma-2-9B+Persona 0.85 0.80 0.80
+Gemma-2-9B+Shape 0.71 0.89 0.78
+Gemma-2-9B+PSI 0.79 0.84 0.75
+Gemma-2-27B+Persona 0.74 0.83 0.83
+Gemma-2-27B+Shape 0.61 0.96 0.90
+Gemma-2-27B+PSI 0.73 0.89 0.90
+Llama3-8B+Persona -0.17 0.40 0.19
+Llama3-8B+Shape -0.57 0.57 -0.04
+Llama3-8B+PSI 0.82 0.83 0.78
+Llama3-70B+Persona 0.81 0.88 0.89
+Llama3-70B+Shape 0.80 0.92 0.88
+Llama3-70B+PSI 0.82 0.89 0.90
+GPT-4o-mini+Persona 0.76 0.86 0.84
+GPT-4o-mini+Shape 0.60 0.82 0.76
+GPT-4o-mini+PSI 0.88 0.90 0.88
+GPT-4o+Persona 0.65 0.66 0.60
+GPT-4o+Shape 0.81 0.90 0.87
+GPT-4o+PSI 0.88 0.91 0.87
+Human 0.84 0.86 0.89
+Table 64: Cronbach’s alpha for Openness facets.
+Model Intellectual Curiosity Aesthetic Sensitivity Creative Imagination
+Mistral-7B+Persona 0.69 0.60 0.62
+Mistral-7B+Shape 0.66 0.80 0.81
+Mistral-7B+PSI 0.52 0.50 0.62
+Gemma-2-9B+Persona 0.60 0.68 0.56
+Gemma-2-9B+Shape 0.90 0.92 0.92
+Gemma-2-9B+PSI 0.70 0.04 0.44
+Gemma-2-27B+Persona 0.75 0.83 0.80
+Gemma-2-27B+Shape 0.84 0.92 0.94
+Gemma-2-27B+PSI 0.77 0.86 0.74
+Llama3-8B+Persona 0.19 0.27 0.16
+Llama3-8B+Shape 0.54 0.65 0.51
+Llama3-8B+PSI 0.83 0.84 0.79
+Llama3-70B+Persona 0.85 0.83 0.82
+Llama3-70B+Shape 0.89 0.93 0.88
+Llama3-70B+PSI 0.80 0.84 0.77
+GPT-4o-mini+Persona 0.77 0.87 0.82
+GPT-4o-mini+Shape 0.76 0.88 0.79
+GPT-4o-mini+PSI 0.88 0.86 0.87
+GPT-4o+Persona 0.82 0.87 0.80
+GPT-4o+Shape 0.87 0.92 0.88
+GPT-4o+PSI 0.89 0.88 0.88
+Human 0.75 0.83 0.82
+Table 65: Cronbach’s alpha for BFI-2 each domain.
+Model Extraversion Agreeableness Conscientiousness Neuroticism Openness
+Mistral-7B+Persona 0.83 0.83 0.88 0.79 0.79
+Mistral-7B+Shape 0.93 0.92 0.93 0.90 0.91
+Mistral-7B+PSI 0.89 0.92 0.90 0.91 0.74
+Gemma-2-9B+Persona 0.88 0.91 0.89 0.91 0.80
+Gemma-2-9B+Shape 0.93 0.98 0.94 0.92 0.96
+Gemma-2-9B+PSI 0.85 0.88 0.86 0.91 0.70
+Gemma-2-27B+Persona 0.92 0.93 0.92 0.90 0.89
+Gemma-2-27B+Shape 0.95 0.98 0.95 0.94 0.96
+Gemma-2-27B+PSI 0.91 0.93 0.90 0.90 0.89
+Llama3-8B+Persona 0.59 0.78 0.55 0.59 0.55
+Llama3-8B+Shape 0.76 0.89 0.85 0.59 0.84
+Llama3-8B+PSI 0.90 0.88 0.88 0.91 0.90
+Llama3-70B+Persona 0.93 0.94 0.93 0.93 0.89
+Llama3-70B+Shape 0.91 0.97 0.94 0.94 0.96
+Llama3-70B+PSI 0.87 0.93 0.93 0.93 0.88
+GPT-4o-mini+Persona 0.93 0.93 0.94 0.91 0.91
+GPT-4o-mini+Shape 0.90 0.95 0.91 0.89 0.94
+GPT-4o-mini+PSI 0.93 0.94 0.93 0.94 0.92
+GPT-4o+Persona 0.90 0.90 0.88 0.83 0.91
+GPT-4o+Shape 0.95 0.98 0.95 0.94 0.96
+GPT-4o+PSI 0.93 0.94 0.94 0.95 0.92
+Human 0.87 0.85 0.90 0.94 0.89
+Model E–A E–C E–N E–O A–C A–N A–O C–N C–O N–O Mean of abs. values
+Mistral-7B+Persona 0.42 0.55 –0.49 0.30 0.71 –0.42 0.37 –0.58 0.23 –0.04 0.41
+Mistral-7B+Shape 0.48 0.59 –0.56 0.63 0.75 –0.74 0.63 –0.80 0.46 –0.46 0.61
+Mistral-7B+PSI 0.45 0.46 –0.45 0.43 0.64 –0.54 0.29 –0.66 0.06 –0.03 0.40
+Gemma-2-9B+Persona 0.23 0.43 –0.46 0.38 0.62 –0.53 0.39 –0.56 0.30 –0.19 0.41
+Gemma-2-9B+Shape 0.57 0.54 –0.54 0.78 0.67 –0.72 0.75 –0.78 0.49 –0.42 0.63
+Gemma-2-9B+PSI 0.37 0.50 –0.49 0.46 0.57 –0.53 0.40 –0.64 0.36 –0.18 0.45
+Gemma-2-27B+Persona 0.41 0.54 –0.56 0.44 0.70 –0.61 0.53 –0.71 0.37 –0.27 0.51
+Gemma-2-27B+Shape 0.57 0.56 –0.62 0.78 0.78 –0.78 0.74 –0.87 0.54 –0.55 0.68
+Gemma-2-27B+PSI 0.32 0.36 –0.31 0.34 0.57 –0.39 0.16 –0.63 0.05 0.24 0.34
+Llama3-8B+Persona 0.24 0.26 –0.39 0.26 0.45 –0.51 0.32 –0.40 0.02 –0.13 0.30
+Llama3-8B+Shape 0.28 0.24 –0.39 0.59 0.55 –0.60 0.61 –0.46 0.24 –0.47 0.44
+Llama3-8B+PSI 0.38 0.44 –0.40 0.19 0.53 –0.29 0.19 –0.46 0.07 0.26 0.32
+Llama3-70B+Persona 0.12 0.33 –0.42 0.28 0.66 –0.54 0.31 –0.60 0.20 –0.08 0.35
+Llama3-70B+Shape 0.41 0.42 –0.57 0.68 0.70 –0.72 0.78 –0.87 0.48 –0.57 0.62
+Llama3-70B+PSI 0.24 0.31 –0.23 0.35 0.53 –0.40 0.15 –0.54 0.09 0.20 0.31
+GPT-4o-mini+Persona 0.42 0.54 –0.59 0.40 0.71 –0.64 0.50 –0.66 0.31 –0.21 0.50
+GPT-4o-mini+Shape 0.63 0.50 –0.52 0.75 0.77 –0.71 0.85 –0.72 0.60 –0.48 0.65
+GPT-4o-mini+PSI 0.39 0.40 –0.43 0.23 0.55 –0.34 0.20 –0.53 0.03 0.26 0.34
+GPT-4o+Persona 0.21 0.40 –0.43 0.23 0.47 –0.30 0.30 –0.43 0.09 0.03 0.29
+GPT-4o+Shape 0.48 0.46 –0.58 0.66 0.67 –0.64 0.71 –0.74 0.29 –0.31 0.55
+GPT-4o+PSI 0.35 0.39 –0.43 0.16 0.51 –0.30 0.14 –0.47 0.02 0.30 0.31
+Human 0.17 0.35 –0.45 0.22 0.33 –0.36 0.20 –0.49 0.09 –0.11 0.28
+Table 66: Domain level correlation analysis. E = Extraversion; A = Agreeableness; C = Conscientiousness; N = Neuroticism; O = Openness.
+Model + Method EXT~OCB AGR~OCB NEU~OCB CON~OCB OPE~OCB
+Mistral-7B+PSI 0.58 (+0.16) 0.63 (+0.45) -0.44 (-0.23) 0.60 (+0.48) 0.25 (+0.08)
+Gemma-2-9B+PSI 0.59 (+0.17) 0.60 (+0.42) -0.52 (-0.31) 0.58 (+0.46) 0.35 (+0.18)
+Gemma-2-27B+PSI 0.53 (+0.11) 0.63 (+0.45) -0.48 (-0.27) 0.62 (+0.50) 0.17 (+0.00)
+Llama3-8B+PSI 0.59 (+0.17) 0.54 (+0.36) -0.32 (-0.11) 0.55 (+0.43) 0.15 (-0.02)
+Llama3-70B+PSI 0.48 (+0.06) 0.46 (+0.28) -0.38 (-0.17) 0.47 (+0.35) 0.11 (-0.06)
+GPT4o-mini+PSI 0.56 (+0.14) 0.43 (+0.25) -0.40 (-0.19) 0.52 (+0.40) 0.08 (-0.09)
+GPT4o+PSI 0.54 (+0.12) 0.36 (+0.18) -0.35 (-0.14) 0.45 (+0.33) 0.02 (-0.15)
+Human 0.42 0.18 -0.21 0.12 0.17
+Table 67: OCB correlation with five personality domains. Values in parentheses show the difference from human, bold for positive, underline for negative.
+Model + Method EXT~CWB AGR~CWB NEU~CWB CON~CWB OPE~CWB
+Mistral-7B+PSI -0.05 (-0.04) -0.29 (+0.01) 0.26 (+0.03) -0.27 (+0.08) -0.07 (+0.10)
+Gemma-2-9B+PSI -0.04 (-0.03) -0.58 (-0.28) 0.45 (+0.22) -0.45 (-0.10) -0.11 (+0.06)
+Gemma-2-27B+PSI -0.02 (-0.01) -0.49 (-0.19) 0.40 (+0.17) -0.48 (-0.13) 0.10 (+0.27)
+Llama3-8B+PSI 0.04 (+0.05) -0.20 (+0.10) 0.07 (-0.16) -0.18 (+0.17) 0.00 (+0.17)
+Llama3-70B+PSI 0.05 (+0.06) -0.45 (-0.15) 0.33 (+0.10) -0.41 (-0.06) 0.00 (+0.17)
+GPT4o-mini+PSI -0.21 (-0.20) -0.56 (-0.26) 0.43 (+0.20) -0.52 (-0.17) 0.04 (+0.21)
+GPT4o+PSI -0.08 (-0.07) -0.54 (-0.24) 0.36 (+0.13) -0.47 (-0.12) 0.01 (+0.18)
+Human -0.01 -0.30 0.23 -0.35 -0.17
+Table 68: CWB correlation with five personality domains. Values in parentheses show the difference from human, bold for positive, underline for negative.
+Figure 7: Ablation test results for predicting Openness.
+The x-axis represents each removed question, and its corresponding response, and the y-axis shows the $R^{2}$ value.
+The red dashed line indicates the baseline $R^{2}$ when all questions are included.
+Figure 8: Ablation test results for predicting Extraversion.
+The x-axis represents each removed question, and its corresponding response, and the y-axis shows the $R^{2}$ value.
+The red dashed line indicates the baseline $R^{2}$ when all questions are included.
+Figure 9: Ablation test results for predicting Agreeableness.
+The x-axis represents each removed question, and its corresponding response, and the y-axis shows the $R^{2}$ value.
+The red dashed line indicates the baseline $R^{2}$ when all questions are included.
+Figure 10: Ablation test results for predicting Neuroticism.
+The x-axis represents each removed question, and its corresponding response, and the y-axis shows the $R^{2}$ value.
+The red dashed line indicates the baseline $R^{2}$ when all questions are included.